### v5

Abstract.
Generative AI is moving rapidly from research into real world deployment across sectors, which elevates the need for responsible development, deployment, evaluation, and governance. To address this pressing challenge, in this study, we synthesize the landscape of responsible generative AI across methods, benchmarks, and policies, and connects governance expectations to concrete engineering practice. We follow a prespecified search and screening protocol focused on post-ChatGPT era with selective inclusion of foundational work for definitions, and we conduct a narrative and thematic synthesis. Three findings emerge; First, benchmark and practice coverage is dense for bias and toxicity but relatively sparse for privacy and provenance, deepfake and media integrity risk, and system level failure in tool using and agentic settings. Second, many evaluations remain static and task local, which limits evidence portability for audit and lifecycle assurance. Third, documentation and metric validity are inconsistent, which complicates comparison across releases and domains. We outline a research and practice agenda that prioritizes adaptive and multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration and uncertainty reporting, versioned and documented artifacts, and continuous monitoring. Limitations include reliance on public artifacts and the focus period, which may under represent capabilities reported later. The survey offers a path to align development and evaluation with governance needs and to support safe, transparent, and accountable deployment across domains.
Project page: Project page
GitHub: GitHub
responsible AI, generative AI, AI ethics, bias, privacy, trust, safety
$c$s: Social and professional topics Computing / technology policyccs: Computing methodologies Artificial intelligenceccs: General and reference Surveys and overviews
1. Introduction
The rapid advancement of Artificial Intelligence (AI) has transformed it from a theoretical construct into a pervasive technology that shapes modern society. AI systems are embedded in critical decision making across healthcare (Kaur et al., 2022), finance (Cao, 2022a), transportation (Abduljabbar et al., 2019), manufacturing industry (Jiang et al., 2023), education (Zhai et al., 2021), and others (Hadi et al., 2023). As these systems influence outcomes that affect individuals, communities, and economies, the need for ethical, transparent, and accountable deployment has become urgent (Liu et al., 2025b).
Responsible AI (RAI) (Stryker, 2024) has emerged as a central area of research and practice to address these challenges. RAI frameworks emphasize fairness, bias mitigation, privacy protection, security, and the safeguarding of human rights (Peters et al., 2020). By embedding these principles throughout the AI lifecycle, including design, training, deployment, and governance (Raza et al., 2024a), RAI seeks to realize the societal benefits of AI while minimizing potential harms (Lu and Wu, 2025).
There has been substantial progress on RAI from both industry (Google AI, 2024; Stryker, 2024; Microsoft, 2024; Anthropic, 2024; Jiang et al., 2025) and academia (Lu et al., 2024b; Dignum and Confidential, 2019; Schiff et al., 2020; Huang et al., 2025b). Nevertheless, notable gaps remain in alignment with human values and in operational practice. The academic literature often separates conceptual frameworks (Lu et al., 2024b; Saeed and Omlin, 2023; Minh et al., 2022; Mehrabi et al., 2021a) from practical implementations (Huang et al., 2024a; Dwivedi et al., 2023; Caton and Haas, 2024), as summarized in Table 1. Conceptual work provides principles and guidelines, whereas applied studies demonstrate deployment in specific contexts. The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 5, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (nis, 2024), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (Systems, 2021). This gap between governance coverage and practical implementation highlights the need for an integrated review that connects frameworks to implementation strategies and to evaluation methods, including explainable AI and AI-ready testbeds.
The emergence of powerful generative AI models, including conversational and multimodal systems, also require a reassessment of responsible practices (Mhlanga, 2023). Generative AI (GenAI) is transforming products and workflows across domains (Kenthapadi et al., 2023). Despite rapid adoption, with 74% of organizations reporting the use of GenAI, only 26% report having a comprehensive organization wide RAI strategy (Capgemini Research Institute, 2024). This imbalance underscores the need to ensure that development and deployment align with ethical standards, fairness, privacy, security, and societal values (Pan and Xu, 2025).
Problem statement. Current RAI guidance for GenAI is fragmented across principles, processes, and tools (as shown in Table 1) and there is less focus on governance-aligned metrics and AI-ready testbeds that enable comparable, auditable assessment across the lifecycle; especially for GenAI systems. As a result, organizations cannot reliably connect regulatory obligations to concrete engineering controls or produce portable evidence for audits and post-market monitoring.
This survey identifies a central problem. Current RAI guidance is fragmented across principles, processes, and tools, and it lacks consistent metrics and AI-ready testbeds that enable comparable assessment across the lifecycle of GenAI systems. Existing surveys (Table 1) and frameworks rarely integrate governance requirements with actionable engineering practices and with systematic evaluation protocols for generative models, including multimodal systems and agentic pipelines.
Contributions. Our key contributions in this work are:
  1. (1)
    We present a unified synthesis that maps governance principles to concrete engineering controls and assurance activities across the AI lifecycle.
  2. (2)
    We present a governance-aligned comparative rubric (C1–C10) with normalized scores and a policy crosswalk (NIST AI RMF, EU AI Act, ISO/IEC 42001) linking benchmarks to audit evidence.
  3. (3)
    We also present a practice roadmap with lifecycle KPIs, explainability guidance for GenAI works, and AI-ready testbeds to support continuous, auditable evaluation.
[…]
Responsible GenAI
Challenges
Technical
Data-related: representativeness(Raza et al., 2024c), quality(Raza et al., 2025b), ethics(Gebru et al., 2021)
Model & system-related: bias(Pessach and Shmueli, 2022), privacy(Dwork, 2006), misinformation(Wang et al., 2023a)
Regulatory & Sociotechnical
IP(Kenthapadi et al., 2023), alignment(Bai et al., 2022), explainability(Dwivedi et al., 2023), sustainability(Khan et al., 2025), dual use(Weidinger et al., 2022)
Strategies
Safety benchmarks
HELM(Liang et al., 2023), SafetyBench(Zhang et al., 2024), Rainbow Teaming(Samvelyan et al., 2024), ALERT(Tedeschi et al., 2024)
Key Performance Indicators
Statistical parity difference(Kun, 2015), disparate impact(Barocas and Selbst, 2016), emissions(Khan et al., 2025)
Explainability methods
SHAP(Lundberg and Lee, 2017), LIME(Ribeiro et al., 2016a), integrated gradients(Arrieta et al., 2020), attention maps(Jain and Wallace, 2019; Serrano and Smith, 2019)
AI-ready testbeds
AI4EU(Foundation, 2024), IEEE Testbed(Foundation, 2024), logging(nis, 2024), red-teaming(Samvelyan et al., 2024; Tedeschi et al., 2024)
Governance frameworks
NIST AI RMF(nis, 2024), ISO/IEC 42001(Systems, 2021), EU AI Act(Outeda, 2024), OECD principles(Dignum, 2019)
Technical toolkits
Model cards(Liang et al., 2023), datasets datasheets(Gebru et al., 2021) , C2PA provenance standards(Das et al., 2025)
Applications
Healthcare
Concerns: Privacy(Dwork, 2006), hallucination(Wang et al., 2023c), bias in diagnostics(Kaur et al., 2022)
Solutions: RAG(Lewis et al., 2020), HITL(Weidinger et al., 2021), Differential Privacy(Dwork, 2006), FDA Alignment(Laux et al., 2024)
Finance
Concerns: Discriminatory lending(Barocas and Selbst, 2016), security(Staab et al., 2024), manipulation(Wang et al., 2023a)
Solutions: Fairness audits(Pessach and Shmueli, 2022), SHAP/LIME(Lundberg and Lee, 2017; Ribeiro et al., 2016a), Basel III/SEC(Slattery et al., 2024)
Education
Concerns: Cultural insensitivity(Al-kfairy et al., 2024), privacy(Dwork, 2006), skills evasion(Lin et al., 2024)
Solutions: Fairness audits(Pessach and Shmueli, 2022), Homomorphic encryption(Bashir et al., 2024), ethics(Dignum, 2019)
Arts & Entertainment
Concerns: Copyright(Kenthapadi et al., 2023; Freeman et al., 2024), job displacement(Pan and Xu, 2025), deepfakes(Shi et al., 2024)
Solutions: Artist collaboration(Freeman et al., 2024), Deepfake detection(Shi et al., 2024)
Agriculture
Concerns: hallucination(Raza et al., 2025d), data ownership(Mittal et al., 2024), digital divide(Al-kfairy et al., 2024)
Solutions: Geo-aware RAG(Lewis et al., 2020), Data contracts(Mittal et al., 2024), HITL(Weidinger et al., 2021)
Defense
Concerns: Autonomous weapons(Weidinger et al., 2022), opacity(Dwivedi et al., 2023), adversarial attacks(Zhang et al., 2024)
Solutions: Human Control(Weidinger et al., 2021), Red teaming(Samvelyan et al., 2024), NATO Principles(Outeda, 2024)
Figure 1. Responsible AI landscape across challenges, strategies, and applications.
This review is a structured literature review of theory, applications, and policy for Responsible AI with emphasis on GenAI. It follows a prespecified search followed by a narrative and thematic synthesis. Themes covered include explainability (XAI), fairness/bias, privacy/security, robustness/safety, governance/compliance, benchmarking and KPIs, and sectoral applications. We then operationalize the agenda via a governance-aligned evaluation rubric with a policy crosswalk and a stakeholder-focused practice roadmap with lifecycle KPIs and AI-ready testbeds. A mindmap of the paper is shown in Figure 1.
Table 1. Comparison of literature by venue, year, topic, and coverage across computer science theory, applications, policy, and generative AI. Checkmarks (✓) indicate coverage.
Paper Venue Year Topic Coverage
CS Theory Application Policy GenAI
(Schiff et al., 2020) arXiv 2020 Responsible AI
(Mehrabi et al., 2021a) ACM CSUR 2021 Bias and Fairness
(Hu et al., 2021) ACM CSUR 2021 AI Safety
(Trocin et al., 2023) Springer ISF 2021 Responsible AI
(Minh et al., 2022) Springer AI Review 2022 Explainable AI
(Huang et al., 2024a) Springer AI Review 2023 Safety
(Saeed and Omlin, 2023) Elsevier KBS 2023 Explainable AI
(Dwivedi et al., 2023) ACM CSUR 2023 Explainable AI
(Nauta et al., 2023) ACM CSUR 2023 Bias and Fairness
(Huang et al., 2023a) IEEE TAI 2023 AI Ethics
(Díaz-Rodríguez et al., 2023) Elsevier Inf. Fusion 2023 Responsible AI
(Mhlanga, 2023) Springer Nature 2023 Responsible AI
(Lu et al., 2024b) ACM CSUR 2024 Responsible AI
(Leschanowsky et al., 2024) Elsevier CHB 2024 Privacy and Safety
(Caton and Haas, 2024) ACM CSUR 2024 Bias and Fairness
(Li and Goel, 2024) Springer ISF 2024 Accountability and Transparency
(Slattery et al., 2024) arXiv 2024 Responsible AI
(Jedličková, 2024) Springer AI and Society 2024 Responsible AI
(Al-kfairy et al., 2024) MDPI Informatics 2024 AI Ethics
(Sadek et al., 2025) AI and Society 2024 Challenges of Responsible AI in Practice
(Woodgate and Ajmeri, 2024) ACM CSUR 2024 Responsible AI
(Reuel et al., 2025) ACM FAccT 2025 Responsible AI (Global maturity model)
(Gadekallu et al., 2025) arXiv 2025 Best Practices of Responsible AI
(Byrne, 2025) ASEE Peer 2025 U.S. Responsible AI Policy
(Sadek et al., 2025) Springer AI & Society 2025 Responsible AI
(Baldassarre et al., 2025) IEEE TAI 2025 Responsible AI
(Ivanov, 2025) The Service Industries Journal 2025 Responsible use of AI in social science
Ours ACM CSUR 2025 Responsible AI
The rest of this paper is organized as follows. Section 2 presents the literature review methodology. Section 3 provides background on Responsible Generative AI. Section 4 reports our results, including thematic findings and benchmark assessments. Section 4 also discusses key performance indicators, explainable AI, and AI-ready testbeds, and Section 5 discuss RAI applications, challanges and RAI strategies. Section 6 presents discussion and forward thoughts, and finally, Section 7 concludes with open challenges and future research directions.
2. Literature Review Methodology
We present our literature review methodology in this section.
Objectives and Review Questions.
This article conducts a literature review of RAI in the post-GPT period and the synthesis of the findings. We address the following key questions:
  • RQ1: What are the dominant themes, risks, and controls in RAI for GenAI and multimodal systems since late 2022?
  • RQ2: Which evaluation facets and metrics are most commonly reported, and where are the gaps?
  • RQ3: What evidence types (empirical, analytical, normative) underpin claims about RAI effectiveness?
  • RQ4: How can these findings be organized into a practical framework for research and governance?
Scope and Protocol.
The review protocol, including sources, screening rules, and the synthesis plan, was prespecified in line with established computer science review guidance (Boland et al., 2017; Keele et al., 2007), and we followed a structured literature review rather than a traditional systematic mapping study or an empirical meta analysis. We mainly focused on literature published from November 2022 through August 2025 to capture rapid methodological and governance developments associated with contemporary generative models. Foundational work published before 2022 was selectively included to define terms, situate concepts, and trace relevant intellectual lineages. We conducted a narrative and thematic synthesis aligned with the research questions and did not perform a meta analysis because of heterogeneity in study designs, measures, and outcomes.
Sources and Search Strategy.
Sources included the ACM Digital Library, IEEE Xplore, SpringerLink, ScienceDirect, arXiv, DBLP, Google Scholar, and Web of Science. To mitigate coverage bias, we complemented database queries with backward and forward snowballing from seed works identified as influential during pilot screening. Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity, open science), with syntax adapted to each database (full strings in Appendix A).
Eligibility and Selection.
We included English-language works within the target time window that addressed RAI in the context of GenAI or multimodal systems and appeared in peer-reviewed venues (journals or conferences), together with a small number of high-impact governance reports or widely cited preprints when they materially shaped discourse. We excluded low-credibility grey literature, purely opinionated pieces without substantive argumentation, duplicates, and items whose primary contribution lay outside RAI (e.g., domain applications without responsibility, safety, or governance components). Although our emphasis was the post-GPT era, we allowed earlier seminal contributions when necessary for definitional clarity or historical continuity.
Screening Process and Data Analysis.
Titles/abstracts were screened against the inclusion criteria, followed by full-text review for items deemed potentially eligible; screening was performed by the first author with a second pass on borderline cases. For each included record we recorded the citation key, year, publication type (journal/conference/preprint), title, and (where available) author-provided keywords. We then applied a lightweight text-mining pipeline to infer themes (e.g., fairness/bias, explainability, governance/compliance, privacy/security, robustness, safety, agentic, multimodal) from titles and keywords (abstracts when available), and aggregated counts by year and publication type. Tables 2 provide basic traceability from literature to claims.
Table 2. Core synthesis of the literature review: theme, citation, year, and focus.
Theme Citation Year Focus
Safety/Benchmarks (Liu et al., 2024) 2024 Introduces a multimodal benchmark to assess safety risks in vision–language models.
Safety/Benchmarks (Zhang et al., 2024) 2024 Proposes a general safety benchmark for large language models across risk categories.
Safety/Benchmarks (Gehman et al., 2020) 2020 Measures toxic degeneration using curated prompts to stress test language models.
Safety/Benchmarks (Li et al., 2024a) 2024 Presents a hierarchical and comprehensive safety benchmark for language models.
Safety/Benchmarks (Mazeika et al., 2024) 2024 Standardizes evaluation for red teaming and refusal robustness in language models.
Safety/Benchmarks (Samvelyan et al., 2024) 2024 Generates diverse adversarial prompts for automated safety stress testing.
Safety/Benchmarks (Wang et al., 2023a) 2023 Provides a broad audit of trustworthiness, including safety and robustness, for GPT models.
Safety/Benchmarks (Tedeschi et al., 2024) 2024 Benchmarks red teaming effectiveness for assessing model safety responses.
Fairness/Governance (Mittal et al., 2024) 2024 Outlines dataset design practices aligned with fairness, privacy, and regulatory norms.
Fairness/Governance (Rodolfa et al., 2021) 2021 Finds minimal fairness and accuracy tradeoff in public policy machine learning tasks.
Fairness/Governance (Chouldechova, 2017) 2017 Analyzes disparate impact and fairness criteria in recidivism risk prediction.
Fairness/Governance (Barocas and Selbst, 2016) 2016 Examines how big data practices can produce disparate impact and legal risk.
Fairness/Governance (Pessach and Shmueli, 2022) 2022 Surveys algorithmic fairness methods and evaluation across application domains.
Fairness/Governance (Kaur et al., 2022) 2022 Reviews trustworthy AI principles and technical enablers for responsible deployment.
Fairness/Governance (Dignum, 2019) 2019 Presents a framework for responsible AI grounded in ethics and governance.
Fairness/Governance (Laux et al., 2024) 2024 Discusses trustworthiness concepts in relation to compliance under the EU AI Act.
Explainability/XAI (Ribeiro et al., 2016b) 2016 Introduces LIME for model agnostic local explanations of predictions.
Explainability/XAI (Selvaraju et al., 2017) 2017 Proposes Grad-CAM for visual explanations in convolutional neural networks.
Explainability/XAI (Arrieta et al., 2020) 2020 Synthesizes concepts, taxonomies, and challenges in explainable artificial intelligence.
Explainability/XAI (Dwivedi et al., 2023) 2023 Reviews core ideas and techniques for explainable AI with practical guidance.
Explainability/XAI (Saeed and Omlin, 2023) 2023 Provides a meta survey of XAI challenges and future research directions.
Explainability/XAI (van der Waa et al., 2021) 2021 Compares rule based and example based explanations on fidelity and user value.
Explainability/XAI (Mosca et al., 2022) 2022 Reviews SHAP based explanation methods for natural language processing.
Explainability/XAI (Nauta et al., 2023) 2023 Systematizes evaluation methods for explanations with quantitative protocols.
Privacy/Security (Dwork, 2006) 2006 Establishes differential privacy as a formal framework for privacy protection.
Privacy/Security (McMahan et al., 2017) 2017 Introduces federated learning for decentralized training with privacy benefits.
Privacy/Security (Nasr et al., 2023) 2023 Demonstrates large scale extraction of memorized training data from language models.
Privacy/Security (Mattern et al., 2023) 2023 Proposes membership inference against language models via neighborhood comparison.
Privacy/Security (He et al., 2025) 2025 Presents label only membership inference attacks tailored to pre trained language models.
Privacy/Security (Li et al., 2024b) 2024 Introduces a benchmark to evaluate privacy risks in language models at multiple levels.
Privacy/Security (Mireshghallah et al., 2024) 2024 Tests privacy leakage using contextual integrity based evaluations of language models.
Privacy/Security (Staab et al., 2024) 2024 Assesses the potential of large models for anonymization and de identification tasks.
LLM Safety (Bai et al., 2022) 2022 Aligns models using constitutional principles and self critique to reduce harmful outputs.
LLM Safety (Weidinger et al., 2021) 2021 Maps ethical and social risks from language models and proposes mitigation avenues.
LLM Safety (Huang et al., 2024b) 2024 Positions trustworthiness dimensions and evaluation needs specific to large models.
LLM Safety (Lin et al., 2024) 2024 Surveys techniques for debiasing and reducing hallucination in large models.
LLM Safety (Wang et al., 2023c) 2023 Overviews ChatGPT applications, limitations, and emerging risks and safeguards.
In the following section, we outline the background and key challenges of RAI, which set the stage for the synthesis of our review findings.
3. Background on Responsible Generative AI
3.1. Historical Context
RAI has received intensified scholarly and policy attention with the rapid diffusion of GenAI (Zirpoli, 2023b; Al-kfairy et al., 2024). GenAI refers to models that synthesize new samples from learned data distributions (van der Zant et al., 2013), spanning language, vision, audio, and multimodal content. Early work combined symbolic rules with probabilistic language models; for example, ELIZA illustrated template driven dialogue in the 1960s (Weizenbaum, 1966). This early era established statistical $n$ gram models (Fink, 2014) and neural sequence modeling for text and speech, laying the groundwork for contemporary large scale generative systems (Qureshi et al., 2025).
Figure 2 highlights the most salient milestones for RAI (alignment, evaluation, governance). As shown in the Figure 2, the first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis; initially in images, and later extended to audio and video. Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in the introduction of attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
A second inflection gave rise to foundation models: large transformers pre-trained in web-scale corpora (e.g., GPT-2/3) and bidirectional encoders (e.g., BERT) that catalyzed transfer across tasks. Alignment advances, including instruction tuning (Peng et al., 2023), reinforcement learning from human feedback (RLHF) (Kaelbling et al., 1996), and constitutional AI (Bai et al., 2022), shifted the emphasis from raw capability to controllability and safety. In vision, diffusion models (Ho et al., 2020) eclipsed GANs for photorealistic generation; when combined with latent compression, they enabled scalable text-to-image synthesis (Rombach et al., 2022).
The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023b), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023b) extend prompting beyond text. Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation (Raza et al., 2025c), bringing new safety, privacy, and governance requirements.
Figure 2. Timeline of generative AI milestones.:
3.2. Challenges in Responsible Generative AI
While GenAI capabilities have accelerated, new risks have emerged in provenance, trust, and governance (e.g., deepfakes, copyright, safety). We group these challenges into technical and regulatory domains, noting that user behavior and deployment context mediate how risks materialize.
3.2.1. Technical Challenges
Technical issues arise across data pipelines, model design and training, evaluation and assurance, and operational deployment. In the following, we present the main technical challenges in GenAI applications:
Data-Related Challenges
  1. (1)
    Data diversity and representativeness. AI models rely on diverse and representative training data to achieve equitable performance in demographic groups and application contexts. Underrepresentation or overrepresentation can induce biased predictions and exacerbate inequities (Raza et al., 2024c). Addressing this requires deliberate data collection across geographic, cultural, linguistic, and socio-economic axes. Targeted sampling, synthetic data, and augmentation can help fill gaps (Bellamy et al., 2019), yet even well-balanced datasets can embed systematic biases, necessitating ongoing monitoring and recalibration. Critics note that truly representative datasets may be infeasible in some domains due to scarcity, legal limits, or privacy concerns (Gebru et al., 2021; Paullada et al., 2021); indiscriminate oversampling can also introduce artifacts that degrade utility.
  2. (2)
    Quality, integrity, and timeliness. Noisy, incomplete, mislabeled, or stale data degrade reliability; adversarial poisoning can seed harmful patterns (Raza et al., 2025b). Robust validation, anomaly detection, and provenance tracking are critical (Das et al., 2025). Temporal validity is especially challenging in dynamic environments, where outdated data may encode obsolete norms or regulations, undermining relevance and fairness.
  3. (3)
    Ethical data governance. Responsible stewardship demands adherence to consent, privacy, and intellectual-property regimes (Mittal et al., 2024), including GDPR (Union, 2018) and emerging AI regulations (Díaz-Rodríguez et al., 2023), alongside sector-specific requirements (Bashir et al., 2024). Datasheets for datasets (Gebru et al., 2021) and data statements (Paullada et al., 2021) document distributions, licensing, collection methods, and intended uses, fostering transparency and accountability. Governance must also address power asymmetries, ensuring consent is informed, revocable, and non-coercive.
Model- and System-Related Challenges
  1. (1)
    Explainability, transparency, and accountability. Large neural models are often treated as “black boxes,” limiting transparency and accountability in high-stakes settings (Meske et al., 2022). A recurring debate concerns a potential interpretability–performance trade-off (Rodolfa et al., 2021). Even where progress challenges the inevitability of this trade-off, scaling XAI across domains can add computational overhead (Khan et al., 2025), introduce subjectivity, and yield partial explanations that miss system-level logic. In practice, adoption requires balancing transparency, performance, and operational efficiency.
  2. (2)
    Bias, fairness, and privacy. Models inherit and can amplify data biases, producing disparate performance across groups. Fairness interventions interact with privacy constraints, and effects can vary by subgroup and domain (Pessach and Shmueli, 2022). Aggressive parity constraints may reduce accuracy in safety-critical contexts or shift harms (Raza and Ding, 2021). Differential privacy, federated learning, and secure multi-party computation support privacy (Bashir et al., 2024) but can obscure model logic and complicate audits. Achieving both fairness and privacy is context-dependent and often entails trade-offs between Responsible AI principles.
  3. (3)
    Robustness, generalization, and misuse. Performance often degrades under domain shift, noise, or adversarial perturbations (Chander et al., 2024). GenAI introduces misuse risks (deepfakes, misinformation, synthetic identity fraud). Mitigations include adversarial training, cross-domain evaluation, and stress testing in simulated deployment (Min et al., 2021), combined with safeguards and policy to deter malicious use.
  4. (4)
    Uncertainty calibration and selective abstention. Overconfidence amplifies harms from hallucination and biased reasoning. Calibration, abstention policies, and fallback to retrieval or human review reduce risk in high-stakes settings; systems should report calibration error and refusal behavior alongside accuracy (Pleiss et al., 2017).
Security and supply-chain threats.
As GenAI systems integrate tools (retrieval, code execution, browsing) and ingest third-party content, they inherit software supply-chain risk. Indirect prompt injection (Raza et al., 2025b) can subvert instructions via untrusted inputs; tool use can exfiltrate data or trigger harmful actions even when base models appear aligned. Threat modeling should include content provenance, capability gating, tool allow-lists, and input sanitization, in addition to adversarial robustness (Liu et al., 2023a).
Privacy attacks beyond governance claims.
Beyond policy statements, GenAI models face concrete leakage risks: membership inference and training-data extraction (Liu et al., 2025a). Mitigations (differential privacy; memorization audits; curation) trade off with utility and fairness; reporting should include empirical leakage checks, not governance alignment alone.
Evaluation and assurance gaps.
Static, one-shot benchmarks rarely capture adaptive adversaries, distribution shifts, or cross-modal failure modes. Without construct validity and task/metric documentation (Raza et al., 2024a), headline scores risk Goodhart effects. Goodhart’s law has been stated as: ”When a measure becomes a target, it ceases to be a good measure” (Strathern, 1997a). Safety claims should be framed as assurance cases with traceable evidence: coverage matrices, failure exemplars, and governance crosswalks, rather than scores alone.
Post-deployment monitoring and incident response.
Safety is non-stationary: ecosystems, user behavior, and model updates induce drift; adversarial tactics evolve. Continuous evaluation (canary tests, periodic red-teaming, shadow deployments), release gates tied to safety deltas, and incident response (rollback, hot-fixes, disclosure) are essential components of an RAI operating model (Breck et al., 2016).
Misinformation and Media Integrity
GenAI can produce plausible but false content (e.g., fabricated citations, numeric errors). Retrieval-augmented generation and automated fact-checking help, but adversaries adapt (Raza et al., 2025d). In vision and audio, high-fidelity synthesis enables identity theft and fraud (e.g., face swapping, voice cloning), amplifying social harm (Ganguli et al., 2022). Watermarking and content-provenance tools aim to bolster integrity (Jiang et al., 2023).
Tables 3 and 4 summarize the biases observed in text-to-image generation, attributable to training data. For instance, the prompt “poor people” in Open AI DALLE-3  [^1] DALLE-3 frequently yielded depictions associated with poverty in African or South Asian contexts, reinforcing region-linked stereotypes; Gemini IMAGEN-3  [^2] Imagen-3also showed bias, with broader but still stereotypical portrayals. For “illiterate person”, IMAGEN-3 prominently depicted elderly women, reinforcing inaccurate age–gender associations; DALLE-3 outputs were more varied yet still included traditional stereotypes. These findings underscore the need for ethical data curation, bias auditing, and careful deployment.
Table 3. Adversarial Examples of Image Generation with DALLE-3 (generated by the authors)
Prompt Generated Images
Poor person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Illiterate person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Rich person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Table 4. Adversarial Examples of Image Generation with IMAGEN-3 (generated by the authors)
Prompt Generated Images
Poor person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Illiterate person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Rich person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Literate person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
Google employee badge [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
a b c
3.2.2. Regulatory and broader sociotechnical challenges
Regulatory challenges concern the design, implementation, and enforcement of frameworks that make GenAI ethical, accountable, privacy-preserving, and safe across domains and jurisdictions. We group key issues below, noting that user behavior and deployment context mediate how risks materialize.
  1. (1)
    Copyright and intellectual property. Using copyrighted material for training raises questions of ownership, licensing, and fair use (Kenthapadi et al., 2023). Reported reproduction and memorization incidents expose gaps in current regimes and motivate updated guidance (Freeman et al., 2024; Pasquale and Sun, 2024; Zirpoli, 2023b).
  2. (2)
    Safety and alignment obligations. GenAI can produce toxic or biased content, undermining trust (Weidinger et al., 2021). Regulatory strategies include bias-aware data curation, real-time detection/mitigation (Tango and Botta, 2013), independent audits (Laine et al., 2024), and adherence to fairness/accountability frameworks (Pagano et al., 2023). Incorporating user feedback and documenting limitations improves sociotechnical alignment.
  3. (3)
    Accountability, transparency, and record-keeping. Assigning responsibility is difficult when systems hallucinate or propagate training-set harms (Elliott et al., 2024). The EU AI Act, for example, proposes risk tiering and obligations for robustness, transparency, logging, and post-market monitoring (Outeda, 2024). Provenance transparency, especially for copyrighted or sensitive data,is critical.
  4. (4)
    Automated decision-making in high-stakes domains. Applications in hiring, credit, and healthcare raise concerns about discrimination and due process. Without recourse mechanisms and domain-specific guardrails, GenAI recommendations can entrench historical inequities.
  5. (5)
    Explainability mandates vs. model complexity. Legal and sectoral expectations for interpretability in high-risk settings meet practical limits as models scale (e.g., GPT-4), creating tension between compliance and performance (Slattery et al., 2024). Dynamic risks (e.g., election misinformation, deepfakes) further require evolving standards.
  6. (6)
    Codes of practice and operational maturity. Inconsistent processes for dataset governance, evaluation, deployment, and monitoring exacerbate bias and privacy risks (Slattery et al., 2024). Pragmatic, auditable codes of practice can harmonize controls and enable proportionate assurance across sectors.
  7. (7)
    Human oversight, recourse, and accessibility. Requirements for meaningful human oversight, user recourse, and disability access are unevenly specified (Raza et al., 2025a). Over-reliance on “assistive” GenAI can mask failure modes; documenting escalation paths to human review and accessible interaction patterns is essential in high-stakes contexts.
  8. (8)
    International alignment, conformity assessment, and enforcement. Divergent regimes (risk-management frameworks, sectoral law, emerging standards) create fragmentation across jurisdictions. Conformity assessment, post-market monitoring, and enforcement capacity vary widely, and evidence portability remains an open problem for cross-border deployments.
  9. (9)
    Environmental and compute sustainability. Training and serving large large models have material energy and carbon costs, with equity implications for participation in research and deployment. Reporting energy/emissions, adopting efficiency techniques (distillation, sparsity), and compute-proportional evaluation can mitigate externalities (Khan et al., 2025).
  10. (10)
    Dual-use and malicious applications. The same GenAI capabilities that enable innovation can be weaponized for misinformation, cybercrime, biothreats, or automated exploitation. Regulatory regimes struggle to anticipate dual-use risks without stifling beneficial innovation, creating tension among openness, export controls, and responsible disclosure (Weidinger et al., 2022).
Bridge to our evidence framework. Taken together, these challenges motivate proportionate, auditable controls and transparent evidence. In Section 4.2 we operationalize this via our comparative rubric, and provides a governance crosswalk mapping risks to evaluative evidence and compliance narratives.
4. Results
In this section, we present the results of our structured literature review on RAI. We first report the thematic findings that emerged from our synthesis. Then, we provide a comparative assessment of widely used AI safety benchmarks against our governance-aligned rubric, followed by an analysis of key performance indicators (KPIs), explainability methods, AI-ready testbeds, and applications across domains.
4.1. Thematic Findings
Figure 4 summarizes thematic distribution, temporal trends (including 2025), co-occurrence patterns, publication-type mix, and keyword salience. Figure 3 provides a compact view of annual volume and keyword evolution. Figure A.1 maps the broader concepts spanning explainable AI, ethics, sectoral applications, advanced models (including large vision–language models), privacy, and regulatory frameworks. Tables 2 list representative studies included in this review.
Figure 3. Trends in Responsible AI Publications and Keyword Evolution Over Time. (A) Number of publications in RAI by Year. (B) Frequency of keywords by year.:
(a) Top themes
(b) Theme trends
(c) Theme co-occurrence
(d) Themes $\times$ publication type
(e) Top keywords
(f) Publications per year
Figure 4. Thematic analysis overview on RAI literature.
Our thematic synthesis highlights that RAI scholarship in the post-GPT period clusters around five dominant themes: (i) safety and benchmarking, with a surge of multimodal evaluation suites and red-teaming methods; (ii) fairness, bias, and governance, where debates center on aligning technical audits with emerging laws such as the EU AI Act and NIST AI RMF; (iii) explainability and transparency, underscoring the tension between interpretability and performance in large foundation models; (iv) privacy and security, particularly membership inference and data extraction risks; and (v) sectoral and sociotechnical applications, spanning healthcare, finance, education, and agriculture. Temporal analysis shows rapid growth from late 2022 onward, with multimodal and agentic systems gaining salience by 2024–2025. These themes frequently co-occur, for example where explainability is invoked to mitigate fairness concerns, or where privacy and governance frameworks intersect in deployment contexts. Together, they map a research terrain that is expanding in scope but still uneven in coverage, with strong emphasis on bias/toxicity benchmarks but weaker treatment of deepfakes, sustainability, and system-level failures.
4.2. Benchmarks: Comparative Assessment
This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in the Tables 6 and 8.
4.2.1. Comparative Assessment Protocol
We assess widely used AI safety benchmarks against a ten–item rubric aligned to risk families and governance needs: C1 Bias/Discrimination, C2 Toxicity, C3 Security/Robustness (including adversarial), C4 Misinformation/Disinformation, C5 Deepfakes/Media Integrity, C6 Privacy (membership/extraction), C7 System–level failure, C8 Malicious–actor realism (red–team/adaptive), C9 Metric validity/documentation, and C10 Governance alignment (e.g., NIST AI RMF, EU AI Act). The ten criteria (also given in Appendix B) reflect risk families that recur across governance instruments and risk repositories, including the NIST AI Risk Management Framework, the EU AI Act, and ISO/IEC 42001, as well as consolidated mappings such as the MIT AI Risk Repository (nis, 2024; Outeda, 2024; Systems, 2021; Slattery et al., 2024). We scored each benchmark after reviewing the official paper, public documentation, and widely cited usage reports; disagreements were resolved by discussion and, when needed, a third reviewer.
Each benchmark is scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, 1 = partial, 2 = strong). Equal weights are used by default. The normalized score is
$S(B)=\frac{\sum_{i=1}^{10}w_{i}\,s_{i}(B)}{2\sum_{i=1}^{10}w_{i}},\quad w_{i}>0.$
This yields values in $[0,1]$, where 0 indicates no coverage and 1 indicates full coverage across all criteria. We also provide a governance crosswalk indicating which requirements each benchmark can plausibly evidence (Tables 6 and 8).
4.2.2. Comparative Assessment Findings
Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs. Table 5 presents the risk surface matrix for criteria C1–C8. Criteria C9 (metric validity and documentation) and C10 (governance alignment) are summarized in Table 8.
Table 5. Mapping benchmarks to risk surface areas. ✓ = supported; $\triangle$ = partial; – = no explicit coverage (from public papers and documentation).
Benchmark Bias/Discr. Toxicity Security Misinfo/Disinfo Deepfakes Privacy System failure Malicious actors
HarmBench (Mazeika et al., 2024) $\triangle$
SALAD Bench (Li et al., 2024a) $\triangle$ $\triangle$ $\triangle$
ALERT (Tedeschi et al., 2024) $\triangle$ $\triangle$
Risk Taxonomy (Cui et al., 2024) $\triangle$ $\triangle$ $\triangle$
TrustGPT (Huang et al., 2023b)
RealToxicityPrompts (Gehman et al., 2020)
DecodingTrust (Wang et al., 2023a) $\triangle$ $\triangle$ $\triangle$
SafetyBench (Zhang et al., 2024) $\triangle$ $\triangle$ $\triangle$
MM–SafetyBench (Liu et al., 2024) $\triangle$ $\triangle$ $\triangle$ $\triangle$
XSTest (Röttger et al., 2024) $\triangle$
Rainbow Teaming (Samvelyan et al., 2024) $\triangle$ $\triangle$
MLCommons AI Safety v0.5 (Vidgen and et al., 2024) $\triangle$ $\triangle$ $\triangle$
HELM (Liang et al., 2023) $\triangle$ $\triangle$
SHIELD (Shi et al., 2024)
BIG-bench (bench authors, 2023) $\triangle$ $\triangle$ $\triangle$ $\triangle$
BEADs (Raza et al., 2024b) $\triangle$ $\triangle$
PrivLM-Bench (Li et al., 2024c) $\triangle$ $\triangle$
PrivacyLens (Shao and et al., 2024)
HumaniBench (Raza et al., 2025a) $\triangle$ $\triangle$ $\triangle$
Table 6. Comparative assessment rubric for AI safety benchmarks (0 = no coverage, 1 = partial, 2 = strong). Scores use equal weights and are normalized to $[0,1]$.
Benchmark C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 Score
RealToxicityPrompts (Gehman et al., 2020) 0 2 0 0 0 0 0 0 1 0 0.15
SafetyBench (Zhang et al., 2024) 1 1 1 0 0 0 0 0 1 1 0.25
MM–SafetyBench (Liu et al., 2024) 1 1 1 0 1 0 0 0 1 1 0.30
SALAD Bench (Li et al., 2024a) 1 1 1 0 0 0 0 0 1 1 0.25
HarmBench (Mazeika et al., 2024) 0 1 1 0 0 0 0 1 1 0 0.20
Rainbow Teaming (Samvelyan et al., 2024) 1 1 2 1 0 0 1 2 1 1 0.50
ALERT (Tedeschi et al., 2024) 0 1 1 0 0 0 0 2 1 1 0.30
HELM (Liang et al., 2023) 1 1 1 1 0 0 0 0 1 1 0.30
PrivLM Bench (Li et al., 2024c) 0 0 0 0 0 2 0 0 1 1 0.20
PrivacyLens (Shao and et al., 2024) 0 0 0 0 0 2 0 0 1 0 0.15
MLCommons AI Safety v0.5 (Vidgen and et al., 2024) 1 1 1 1 0 0 0 1 1 1 0.35
SHIELD (Shi et al., 2024) 0 0 1 0 2 0 0 0 1 0 0.20
BEADs (Raza et al., 2024b) 1 1 0 2 0 0 0 0 1 0 0.25
MFG for GenAI (Foundation, 2024) 1 0 1 1 0 0 0 1 0 0 0.20
XSTest (Röttger et al., 2024) 0 0 0 0 0 2 0 1 1 0 0.20
DecodingTrust (Wang et al., 2023a) 1 1 1 0 0 1 0 1 1 1 0.35
TrustLLM (Huang et al., 2024b) 1 1 0 1 0 1 1 0 1 0 0.30
BIG bench (bench authors, 2023) 1 1 1 0 0 0 1 0 0 0 0.25
Risk Taxonomy (Cui et al., 2024) 1 1 1 0 0 1 0 1 1 2 0.40
TrustGPT (Huang et al., 2023b) 1 1 0 0 0 0 0 0 1 0 0.20
HumaniBench (Raza et al., 2025a) 2 1 0 1 0 1 1 0 1 1 0.40
Criteria (C1 to C10). C1 Bias/Discrimination; C2 Toxicity; C3 Security/Robustness; C4 Misinformation/Disinformation; C5 Deepfakes/Media Integrity; C6 Privacy (membership/extraction); C7 System level failure (tool/agentic); C8 Malicious actor realism (red team/adaptive); C9 Metric validity/documentation; C10 Governance alignment. Normalization (equal weights): $S(B)=\frac{\sum_{i=1}^{10}s_{i}(B)}{20}$ with $s_{i}\in\{0,1,2\}$.
Table 7. Governance-based Frameworks and Technical AI Tools used in RAI Practices.
Category Governance-based Frameworks Technical AI Solutions
Fairness & Bias Mitigation ISO/IEC TR 24027:2021 (3, 2021) : technical report on AI and data quality. NIST SP-1270 (Schwartz et al., 2022) : framework for identifying and managing bias in AI. Framework for increasing diversity & inclusion in AI education, research, & policy: AI4ALL (AI4ALL, 2024) Evaluate for model fairness via visualizations: Google What-if Tool (Wexler et al., 2020) AI Fairness 360 Toolkit by IBM provides tools for evaluation and mitigation of bias: AIF360 (Bellamy et al., 2018b) Assess and enhance fairness of ML models: Microsoft Fairlearn (Weerts et al., 2023)
Interpretability & Explainability IEEE P2976 (Polemi et al., 2024) : Standards for Explainable Artificial Intelligence. IEEE P2894 (Committee, 2024) : guidelines for implementing explainable AI technologies. Explaining Decisions Made with AI: UK Info. Commissioner’s Office & The Alan Turning Institute (Leslie, 2022) Model interpretability and explainability tool: LIT (Tenney et al., 2020) Toolkit for measuring explainability using various models like LIME and SHAP: InterpretML (Sameki et al., 2020) Provide simple AI generated explanations: ELI5 (Deepgram, 2024) AI Explainability 360 toolkit by IBM provides tools for explainability of ML models: AIX360 (Arya et al., 2019)
Transparency & Accountability U.S Government Accountability Office: AI Accountability Framework (U.S. Government Accountability Office, 2021) Organisation for Economic Co-operation and Development (OECD): AI Principle of Accountability ((, OECD) Evaluate, test and monitor ML models: Evidently (EvidentlyAI, 2024) Improve model transparency, reproducibility, and robustness: MLBench (MLBench, 2020)
Privacy, Safety & Security NASA (Smith et al., 2020): hazard contribution modes of ML in space applications. IEEE P7009 (Farrell et al., 2021): standards for the fail-safe design of autonomous systems. ISO/IEC TS 27022:2021 (Systems, 2021): information security management systems guidelines. NIST AI 100-2 E2023 (Oprea and Vassilev, 2023) : specific cybersecurity practices for AI. UK Government AI Cybersecurity Code of Practice (Department of Science, Innovation & Technology, 2024). MIT Risk repo (Slattery et al., 2024): Database documenting AI risks and safety gaps. Google’s Secure AI Framework SAIF (Inc., 2024) Context-aware toxicity tracker: Unitary (Unitary, 2023) Adversarial Robustness Toolbox: ART Audit privacy of ML models: Privacy Meter Google Differential Privacy libraries: DP Privacy-preserving data analysis and ML toolkit Secret Flow Tensorflow privacy (TensorFlow, 2023) Privacy Enhancing Technology (PET) on Synthetic Data Generation: Betterdata.AI (inc., 2024)
Ethical Guidelines & Compliance European General Data Protection Regulation: GDPR (Union, 2018) Canadian AI and Data Act: ISED (of Canada, 2023a) Canadian Voluntary Code of Conduct on the Responsible Development and Management of Advanced GenAI Systems: ISED (of Canada, 2023b) American Executive Order on Safe, Secure, and Trustworthy Artificial Intelligence: AEO (House, 2023) Google responsible AI practices (AI, 2023) Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, 2019) Microsoft’s Responsible AI Toolbox (Microsoft, 2024) Demonstrate trust, measure and manage risk, and compliance: OneTrust (OneTrust, 2023) World Health Organization (WHO): Ethics and governance of AI for health Guidance on large multi-modal models (Organization et al., 2024)
Monitoring & Auditing Index of harms realized in the real world by AI systems: Incident Database (Database, 2023) Guidance provided by UK Information Commissioner’s Office AI Auditing Framework (UK Government Information Commissioner’s Office, 2020) Audit classification models for intersectional bias: FairVis (Cabrera et al., 2019) Bias and fairness audit toolkit: Aequitas (for Data Science and of Chicago, 2018) Auditing of bias in generalized ML applications: Audit AI (Inc., 2020)
Table 8. Governance alignment crosswalk (illustrative). ✓= supported evidence; $\triangle$= partial; – = not covered.
Requirement
RTP
SafetyBench
MM–SafetyBench
SALAD
HarmBench
Rainbow
ALERT
HELM
PrivLM
SHIELD
MLC v0.5
HumaniBench
NIST: Privacy & Security (nis, 2020a, b) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
EU AI Act: Robustness/Resilience  [^3] https://artificialintelligenceact.eu/article/15 $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
EU AI Act: Transparency/Record-keeping  [^4] https://artificialintelligenceact.eu/article/12/ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
Multimodal Deepfake Risk (c2p, 2024) $\triangle$
Adaptive/Red-team Evidence (nis, 2023) $\triangle$ $\triangle$ $\triangle$ $\triangle$
Abbreviations. RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
Risk-surface coverage. Table 5 presents the representative benchmarks cover eight salient risk families (bias/discrimination, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure, and malicious-actor realism). Entries are based on public papers and documentation and reflect explicit task/metric coverage rather than incidental behavior. The results in Table 5 shows that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens). Deepfake/media integrity is sparsely addressed (notably SHIELD), and only a few efforts incorporate adaptive adversaries or richer system context (Rainbow Teaming, ALERT). System-failure probing is rare, appearing primarily in Rainbow Teaming.
Rubric-scored comparison. We operationalize the ten-criterion rubric in Section 4.2.1 to produce a normalized [0,1] score per benchmark. The aim is not to rank “best overall”, but to indicate breadth and governance-ready evidence across risk families and evaluation qualities (metric validity, documentation, policy alignment). The results in Table 6 shows that Rainbow Teaming attains the highest breadth score (0.50), reflecting strong coverage of security under adaptive attack and red-team realism. Risk Taxonomy and HumaniBench (0.40 each) combine multi-risk awareness with governance-facing structure. MLCommons AI Safety v0.5 and DecodingTrust (0.35) provide broad but mostly static probes. Single-issue suites score lower by design,e.g., RealToxicityPrompts and PrivacyLens (0.15), useful for their focus but limited for governance evidence.
Governance context and crosswalk. Table 7 situates the analysis within widely used frameworks and tools. Table 8 then maps each benchmark to selected governance requirements spanning NIST AI RMF privacy/security controls, EU AI Act robustness and transparency obligations, multimodal deepfake risk, and evidence for adaptive/red-team evaluation. A ✓indicates plausible evidentiary use; $\triangle$ indicates partial or indirect support; indicates no explicit coverage in public artifacts. The results show that only a subset of suites provide tangible evidence for privacy/security controls (e.g., PrivLM-Bench). Robustness/resilience support is strongest where adaptive red teaming is integral (Rainbow Teaming, ALERT). Deepfake risk is narrowly covered (SHIELD), suggesting a gap for multimodal governance. Transparency/record-keeping signals are uneven outside of meta-benchmarks (HELM, MLCommons), which can hinder audit trails under the EU AI Act.
4.2.3. Proposed improvements to AI safety benchmarks
The following suggestions are offered as constructive opportunities to complement existing benchmarks and to support broader governance use.
To address the capability gaps identified above, we suggest:
  1. (i)
    Data quality and transparency. Encourage diverse and well documented corpora across regions, languages, and demographics, with dataset cards or data statements that capture provenance, consent and licensing, and curation criteria to enhance traceability and accountability (Eriksson et al., 2025; Gebru et al., 2021).
  2. (ii)
    Construct clarity and validation. Make constructs explicit by stating what each task and metric is intended to measure and, where feasible, validate measures with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024).
  3. (iii)
    Evaluation resilience. Reduce opportunities for gaming or overfitting by incorporating dynamic or adversarial evaluation, including undisclosed items and evolving prompts, consistent with established security testing practices (Oprea and Vassilev, 2023; Ghosh et al., 2025).
  4. (iv)
    Broader scope. Expand coverage to multimodal, non-English, and higher-risk domains, and include multi-agent or autonomous settings where system level failures are more likely to surface (Salazar et al., 2025; Raza et al., 2025c).
  5. (v)
    Multi-stakeholder alignment. Mitigate potential commercial or domain bias by engaging multiple stakeholders and making explicit linkages to policy instruments and governance frameworks (Outeda, 2024).
  6. (vi)
    Timeliness and maintenance. Support continuous evaluation through automated pipelines that enable periodic re-testing across the model lifecycle, improving comparability over time (Wang et al., 2024).
  7. (vii)
    Exploration of unknown risks. Integrate structured red teaming and stress tests under distribution shift, and monitor for emergent behaviors that may elude static suites, drawing on existing practices such as ALERT and Rainbow Teaming (Tedeschi et al., 2024; Samvelyan et al., 2024; Liu et al., 2023a).
Figure 5. Responsible AI governance landscape and linkage to our evaluation criteria. Technical factors correspond primarily to rubric items C1–C7 (fairness/bias, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure). Legal instruments (e.g., EU AI Act, NIST AI RMF) ground C10 (governance alignment). Risk & assurance processes (anticipation, responsiveness, reflexivity, inclusion, knowledge management) support adaptive testing (C8) and metric validity/record-keeping (C9). Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks, motivating the gaps observed in Tables 58.
To operationalize these improvements and make evidence portable across audits, we now specify lifecycle KPIs that teams can track during development, deployment, and post-market monitoring.
4.2.4. Key Performance Indicators (KPIs) for Responsible AI
KPIs provide measurable criteria to assess the performance of AI systems across critical dimensions of responsibility (Mattioli et al., 2024). Unlike ethical principles that are mostly subjective, KPIs translate RAI goals into operational metrics that can be tracked, compared, and audited across the model lifecycle. They bridge governance expectations with technical implementation so that fairness, transparency, robustness, privacy, and sustainability are evidenced through quantifiable measures, enabling compliance and continuous improvement. We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior work (Pessach and Shmueli, 2022; Rajkomar et al., 2018; Van Wynsberghe, 2021), and introduce pragmatic, author-defined KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
Notation and conventions. $f(x)$ is the model score for input $x$ (state whether probability, logit, or other monotone score); $y\in\{0,1\}$ is the true label; $\hat{y}=\mathbb{1}\{f(x)\geq t\}$ is the predicted label with a single threshold $t$ applied to all subgroups unless stated; $A$ is a protected attribute with subgroups $a$ and $b$; $\phi_{0}$ is a baseline output and $\phi_{i}(x)$ is the attribution for feature $i$ (e.g., SHAP or Integrated Gradients); “accuracy” refers to a stated metric (e.g., top-1, F1, AUROC) computed per subgroup with the same threshold $t$; a perturbation is a noise or attack consistent with a stated threat model; $P_{\mathrm{avg}}$ is average power (W), $T$ is time (h), $\mathrm{CI}$ is grid carbon intensity (kgCO2e/kWh), and $\mathrm{PUE}$ is Power Usage Effectiveness.
  1. (1)
    Data quality and integrity
    $Q=\frac{\mathrm{Valid\ Entries}}{\mathrm{Total\ Entries}}\times 100.$
    This KPI measures the fraction of entries that pass schema, label, and consistency checks (Pessach and Shmueli, 2022). Here, we define “valid” checks to report $Q$ (quality) by subgroup/domain if applicable.
  2. (2)
    Data privacy compliance
    $P=\frac{\mathrm{Compliant\ Data\ Points}}{\mathrm{Total\ Data\ Points}}\times 100.$
    This KPI fractions with documented lawful basis, consent or equivalent, and license compatibility (Leschanowsky et al., 2024). We specify regime (e.g., GDPR/CCPA), provenance fields audited, and audit scope; legal review may be required.
  3. (3)
    Bias detection (statistical parity difference, SPD)
    $\mathrm{SPD}=P(\hat{y}=1\mid A=a)-P(\hat{y}=1\mid A=b).$
    It defines the difference in positive outcome rates across subgroups. We specify state $A$, $t$, sample sizes; or bootstrap and a significance test.
  4. (4)
    Fairness ratio (disparate impact, DI)
    $\mathrm{DI}=\frac{P(\hat{y}=1\mid A=a)}{P(\hat{y}=1\mid A=b)}.$
    This KPI measures the outcome parity as a ratio. It provides guard against division by zero; also report the symmetric form $\min(\mathrm{DI},1/\mathrm{DI})$.
  5. (5)
    Equal performance (accuracy parity)
    $\Delta_{\mathrm{acc}}=\left|\mathrm{Accuracy}_{A=a}-\mathrm{Accuracy}_{A=b}\right|.$
    This measures the absolute gap in a stated accuracy metric across subgroups (Raza et al., 2024c). It reports per subgroup and key intersections; keep $t$ fixed across subgroups.
  6. (6)
    Robustness assessment
    $R=\frac{\mathrm{Errors\ under\ perturbation}}{\mathrm{Total\ perturbed\ inputs}}\times 100,\quad\Delta_{\mathrm{acc}}^{\mathrm{rob}}=\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{pert}}.$
    It measures the error rate and accuracy drop under stated threat models (Raza et al., 2025d). We specify attack type and strength (e.g., $\ell_{p}$ bounds), number of steps, and per-attack results.
  7. (7)
    Explainability index (author defined)
    $E(x)=\sum_{i=1}^{n}\left|\phi_{i}(x)\right|.$
    This KPI measures the attribution magnitude summary for an instance using SHAP/IG (Lundberg and Lee, 2017; Sundararajan et al., 2017; Vilone and Longo, 2021). We use as a proxy only; consider also reporting attribution concentration (e.g., entropy over $|\phi_{i}|$).
  8. (8)
    Local faithfulness gap (XAI)
    $\mathrm{LFG}(x)=\frac{\left|\,f(x)-\left(\phi_{0}+\sum_{i}\phi_{i}(x)\right)\right|}{\max\left(\left|f(x)\right|,\epsilon\right)}.$
    This KPI measures the relative completeness error of additive explanations (lower is better). We set $\epsilon$ and summarize LFG as a distribution (median, IQR); complements human-readability studies.
  9. (9)
    High stakes error rate (author defined)
    $E_{h}=\frac{\mathrm{Critical\ errors}}{\mathrm{Total\ predictions}}\times 100.$
    This measures incidence of errors that trigger a predefined harm tier (Kunreuther et al., 2002). We define the harm taxonomy and decision costs; consider a harm weighted rate in safety critical settings.
  10. (10)
    Audit frequency and resolution time
    $F_{a}=\frac{\mathrm{Audits}}{\mathrm{Time\ period}},\qquad T_{r}=\mathrm{median\ time\ to\ resolve}.$
    This KPI is cadence of audits and timeliness of remediation (Díaz-Rodríguez et al., 2023; Hasan, 2021). We specify window and scope; report $T_{r}$ with IQR and the share closed within SLA.
  11. (11)
    User inclusivity index (author defined)
    $U=\frac{\mathrm{Inclusive\ features\ implemented}}{\mathrm{Planned\ features}}\times 100.$
    It measures the accessibility and localization coverage (Shams et al., 2023). We enumerate the checklist (e.g., WCAG conformance, screen reader support, localization) and review cadence.
  12. (12)
    Sustainability (energy and emissions)
    $E_{\mathrm{kWh}}=\frac{P_{\mathrm{avg}}}{1000}\times T,\qquad C_{\mathrm{CO_{2}e}}=E_{\mathrm{kWh}}\times\mathrm{PUE}\times\mathrm{CI}.$
    This KPI measures energy use and estimated emissions; pair energy with location specific or market based carbon intensity and data center overhead (Van Wynsberghe, 2021). We state measurement method (e.g., nvidia-smi vs power meter), scope (training or inference), normalization unit (e.g., per 1k tokens), and whether CI is location or market based (e.g., CodeCarbon) (Song et al., 2025).
4.3. Explainable AI
Having established how benchmarks cover risk families, we examine explainability (XAI) methods in GenAI methods and what they can and can not do, evidence under audit. Explainable AI (XAI) comprises methods that render model behaviour and outputs intelligible to human stakeholders (Meske et al., 2022). Effective explanations serve multiple roles: they support accountability and trust, enable bias and safety analyses, and facilitate compliance with emerging governance requirements (Baker and Xiang, 2023; Arrieta et al., 2020). In practice, XAI must operate across technical, social, legal, and organisational layers rather than as a purely algorithmic add-on.
Figure 6. Post-hoc explainability within a Responsible AI (RAI) pipeline. A black-box model produces predictions $\hat{y}$ and attribution signals $I(x_{i})$, where $I(x_{i})$ denotes a feature-importance signal (e.g., gradient/saliency or a SHAP value). Interpretability pathways (feature relevance, example-based, surrogates, and visual/text explanations) feed fairness diagnostics (e.g., SPD/DI) and calibration analyses, supporting governance-aligned optimisation $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}+\lambda\,\mathcal{L}_{\mathrm{fair}}$. For LLMs/VLMs, prompt traces and cross-modal attention provide additional observability.
Scope and challenges in the GenAI era.
Foundation models raise explanation challenges beyond conventional ML. Stochastic generation (e.g., temperature or top-$k$ sampling) produces non-deterministic outputs for identical prompts, complicating reproducibility of explanations (Idé and Abe, 2023). Multimodality couples heterogeneous inputs/outputs (text, image, audio), requiring explanations that account for cross-modal attention and fusion dynamics (Pi, 2023). Hallucinations and brittle prompt sensitivity demand methods that diagnose provenance gaps and reasoning failures. Consequently, GenAI-oriented XAI should expose decoding pathways, visualise modality interactions, and characterise failure modes, not merely attribute tokens.
Methods in practice.
XAI methods can be ante-hoc (interpretability built into the model class, e.g., sparse linear models, rule lists, decision trees) or post-hoc (explain an existing black-box locally or globally). Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016a); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023a). For LLMs and Vision-language-models (VLMs), these are complemented by prompt-engineering probes, chain-of-thought (CoT) audits, and tool-use traces (Natarajan and Nambiar, 2024). Importantly, attention maps and saliency are not proofs of causal reasoning and should be triangulated with counterfactual tests and ablations (Jain and Wallace, 2019; Serrano and Smith, 2019).
Figure 6 situates post-hoc explanations within an RAI pipeline. Predictions $\hat{y}=M_{\phi}(x)$ can be accompanied by gradient-based importance $I(x_{i})=\partial M_{\phi}/\partial x_{i}$ to support local inspection; example-based evidence provides case-level comparators; and visual/text explanations aid end-user understanding. Fairness is evaluated with standard disparity measures such as Statistical Parity Difference (SPD) (Kun, 2015) and Disparate Impact (DI) (Barocas and Selbst, 2016):
$\text{SPD}=P(\hat{y}=1\,|\,A=a)-P(\hat{y}=1\,|\,A=b),\quad\text{DI}=\frac{P(\hat{y}=1\,|\,A=a)}{P(\hat{y}=1\,|\,A=b)}.$
A practical training objective combines task loss with an explicit fairness regulariser,
$\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}(\theta)+\lambda\,\mathcal{L}_{\mathrm{fair}},$
allowing developers to trade off utility and fairness under governance constraints. For trust calibration, we recommend standard metrics such as Expected Calibration Error (ECE) (Pleiss et al., 2017) or the Brier score (Rufibach, 2010); ECE summarises the gap between confidence and empirical accuracy across bins:
$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\big|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\big|.$
Adding canonical sources: ECE and Brier are standard measures (Pleiss et al., 2017). Calibrated probabilities, paired with faithful explanations, yield more reliable user-trust assessments.
A compact theoretical frame.
Let $f_{\theta}:\mathcal{X}\!\to\!\mathcal{Y}$ denote a predictive model (or $p_{\theta}(\cdot\mid x,c)$ for generative systems with context $c$). We define an explanation operator
$\mathcal{E}:(x,f_{\theta},\omega)\mapsto z\in\mathcal{Z},$
where $\omega$ indexes stochasticity (decoding temperature, top-$k/p$). Local explanations fix $x$; global explanations summarise behaviour over a distribution on $\mathcal{X}$. For VLMs, $\mathcal{X}=\mathcal{X}_{\text{text}}\times\mathcal{X}_{\text{vision}}$ and $z$ pairs textual rationales with grounded visual regions.
Desiderata (axioms). An explanation $z=\mathcal{E}(x,f_{\theta},\omega)$ is useful when it approximately satisfies: (i) local accuracy (additive decompositions sum to output deviation from a baseline) (Lundberg and Lee, 2017); (ii) sensitivity/monotonicity (if $f_{\theta}$’s dependence on feature $i$ increases at $x$, its attribution should not decrease) (Sundararajan et al., 2017); (iii) invariance to representation-preserving transforms and to behaviourally equivalent models (Sundararajan et al., 2017); (iv) stability (Lipschitz continuity in $x$ and $\omega$(Alvarez-Melis and Jaakkola, 2018); (v) counterfactual consistency (factors highlighted by $z$ admit feasible interventions that flip or materially shift the output); and (vi) parsimony (minimal complexity $\Omega(z)$ subject to fidelity) (Kusner et al., 2017). These properties specialise naturally to multimodal settings via alignment between rationales and grounded visual evidence.
Without committing to a particular algorithm, we report three constructs that make explanations auditable: fidelity via perturbation functionals $\Delta_{f}(x;S)$ on regions or features emphasised by $z$ and compared to size-matched random sets; stability via distances $\delta(\mathcal{E}(x,f_{\theta},\omega),\mathcal{E}(x+\epsilon,f_{\theta},\omega^{\prime}))$ under input noise and seed changes; and governance usefulness by pairing explanations with SPD/DI, calibration (ECE/Brier), and provenance artefacts (prompt/context/tool logs). This elevates explanations from plausible narratives to evidence interfaces suitable for audit.
For foundation models, explanations should expose (i) prompt pathway elements (system/user prompts, retrieved context, tool calls), (ii) decoding state ($\omega$: temperature, top-$k/p$), and (iii) cross-modal grounding (text rationales linked to image regions or clips). These artefacts do not prove causality but enable stability and counterfactual checks and support post-hoc accountability (Rane and Paramesha, 2024; Kenny et al., 2021).
Explanations integrate naturally into risk-managed optimisation:
$\min_{\theta}\ \mathbb{E}_{x,\omega}\!\left[\ell\big(f_{\theta}(x,\omega),y\big)\right]+\lambda_{\mathrm{fair}}\mathcal{L}_{\mathrm{fair}}+\lambda_{\mathrm{xai}}\mathcal{R}\!\left(\mathcal{E}(x,f_{\theta},\omega)\right),$
where $\mathcal{R}$ encodes desiderata such as stability (penalising high $\delta$) and parsimony (penalising $\Omega$). This connects fairness (SPD/DI), calibration (ECE), and logging to the governance criteria C1–C10 used in our evaluation.
Limitation: Even theoretically sound explanations can mislead when they are plausible but unfaithful (e.g., over-reading attention maps), trigger Goodhart effects (Strathern, 1997b) (optimising the appearance of fairness rather than outcomes), or lack distributional validity (narratives drift across domains). Reporting stability, counterfactual checks, and provenance mitigates these risks and makes explanations suitable as audit evidence rather than marketing artefacts.
Next, we present AI-ready test-beds that computes lifecycle KPIs and logging XAI evidence as audit-ready artifacts.
4.4. AI-ready Test-beds
AI-ready test-beds advance RAI by providing controlled environments to develop, test, and validate AI models with reliability and reproducibility guarantees. In the post-ChatGPT era, where GenAI is increasingly embedded in products and services, such platforms help assess non-determinism, multimodal behavior, and risks like bias, hallucination, and privacy leakage, supporting proportionate assurance and governance alignment. We present representative examples in Table LABEL:tab:ai_testbeds. The key value of these test-beds is in producing audit-ready evidence datasheets/model cards, prompt/tool logs, KPI dashboards, red-team logs, and versioned runs.
Table 9. AI evaluation platforms and initiatives used as de facto test-beds for RAI. We distinguish formal test-beds/platforms from broader labs/initiatives commonly used for evaluation.
Name Domain Type RAI-focused features
AI4EU AI-on-Demand Platform (cor, 2019) General AI development Platform Ethical-by-design resources; model/dataset catalogue; support for transparency and explainability workflows (used to prototype responsible AI pipelines).
IEEE Ethical AI Systems Test Bed (win, 2019) Ethical AI development Test-bed/Platform Evaluation against ethical frameworks; human-centred design; fairness-oriented assessment protocols; documentation of decision rationale.
AI Testbed for Trustworthy AI (TNO) (tno, 2020) Trustworthy AI Test-bed Robustness and transparency checks; fairness analysis; risk/assurance reporting for deployment contexts.
ETH Zurich Safe AI Lab (saf, 2020) Safe and fair AI Lab/Platform Safety-critical validation; robustness/hallucination probing; reliability and interpretability methods for high-stakes settings.
HUMANE AI (hum, 2020) Human-centric AI Platform/Initiative Alignment with human values; societal-impact and fairness guardrails; guidance for participatory, responsible design.
AI for Good (ITU) (AIf, 2017) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
UKRI Trustworthy Autonomous Systems (TAS) Hub (TAS, 2020) Autonomous systems Program/Test-bed Accountability and transparency for autonomy; safety cases; conformance/compliance support for real-world systems.
Algorithmic Justice League (Algorithmic Justice League, 2024) Algorithmic fairness Initiative Bias discovery and awareness; audits and mitigation advocacy; widely referenced in fairness evaluation contexts (not a formal test-bed).
AI Commons [^5] https://ai-commons.org/ Open AI collaboration Initiative Open, transparent collaboration patterns; governance-mindful sharing; community evaluation practices (not a formal test-bed).
ClarityNLP (Healthcare) (Cla, 2021) Healthcare NLP Platform Clinical text pipelines; emphasis on fairness, transparency, and ethical data use; supports domain-specific evaluation.
ToolSandbox (Lu et al., 2024a) Tool-using LLMs Research test-bed Stateful, multi-step/equipped LLM evaluation; probes privacy, fairness, transparency and accountability in interactive tasks.
Chatbot Arena (LMSYS) (cha, 2024) LLM evaluation Community platform Open-source platform where users compare LLMs in head-to-head “arena” battles; pairwise ranking with more than 800k community votes and transparent leaderboards; all evaluation pipelines and data are public for reproducibility.
AI Fairness 360 (AIF360) (Bellamy et al., 2018a) Fairness evaluation Toolkit Open-source Python toolkit with 70+ fairness metrics and 9 mitigation algorithms; supports bias detection, fairness explanations and interactive web interface for non-experts.
Zeno (zen, 2023) AI evaluation Platform Interactive evaluation platform; enables data exploration across any data type and automatically discovers systematic failures using slice finder; supports interactive charts to compare evaluation metrics across slices and models.
Galileo (gal, 2023) Generative AI evaluation Platform Automated evaluation metrics for GenAI (RAG and agent metrics plus safety/security metrics); real-time guardrails to block hallucinations, PII and prompt injection; captures corner cases and allows custom metrics.
Arize AI (Dong et al., 2024) AI observability Platform AI observability and LLM evaluation platform that automatically detects model issues, traces root causes via indexed training/validation/production data and fine-tunes performance; includes dashboards for evaluation/performance tracing, explainability and fairness. The Bias Tracing tool monitors fairness metrics (e.g., recall parity, false positive parity, disparate impact) and surfaces features and cohorts contributing to bias.
PromptLayer (pro, 2022) Prompt engineering Platform Prompt-evaluation platform with visual pipelines for scoring prompts against golden datasets, backtesting with historical data, regression testing and continuous integration of evaluation pipelines; useful for systematic RAG/agent evaluations.
FairX (Sikder et al., 2024) Fairness benchmarking Open-source toolkit Modular benchmarking tool for fairness, data utility and explainability; supports training of fair generative models, evaluation of synthetic data and fairness metrics with pre-, in- and post-processing debiasing techniques.
BEATS bias test suite (Abhishek et al., 2025) LLM bias/ethics evaluation Test suite Benchmark for evaluating bias, ethics, fairness and factuality across 29 metrics for LLM outputs; measures demographic and cognitive biases, ethical reasoning and factuality; aims to diagnose factors driving bias and develop mitigation strategies.
5. RAI Application Across Domains
GenAI systems are increasingly embedded in decision-making processes across diverse domains such as healthcare, finance, human resource, defense and education. While these systems offer significant utility, they also introduce new categories of risks and challenges, as outlined earlier. The challenges and mitigation strategies synthesized in this review are broadly relevant, but their salience and priority vary substantially by sectoral context. We translate the above findings into stakeholder-aligned practices and a deployment-oriented roadmap, with examples and decision pointers.
5.1. Healthcare
Applications.
GenAI is transforming healthcare across clinical, biomedical, and administrative domains, but each application underscores the importance of RAI safeguards. For example, while generative models improve efficiency in clinical note generation and triage chatbots (Johnson and Patel, 2023; Alagha and Ahn, 2023), they raise concerns about accuracy, liability, and bias in medical decision support. Disease detection and drug discovery pipelines (Miotto et al., 2022; Liu et al., 2023c) illustrate GenAI’s potential for accelerating science, yet demand rigorous validation, explainability, and regulatory alignment before clinical deployment. Synthetic health data helps mitigate privacy risks and dataset imbalance (Walsh et al., 2023), but governance frameworks must ensure that synthetic outputs do not inadvertently reinforce hidden biases. Similarly, tools like Watson for Oncology highlight the need for transparency and interpretability in high-stakes decision-making (Memorial Sloan Kettering Cancer Center, 2024). Even promising breakthroughs, such as GenAI-designed therapeutics (Smith et al., 2025), requires human oversight on trial integrity, equitable access, and long-term monitoring. In this way, healthcare exemplifies both the promise of GenAI and the critical role of RAI practices, balancing innovation with safety, accountability, and fairness.
RAI Challenges.
Responsible adoption of GenAI in healthcare is constrained by privacy violations (Shiraz and Lee, 2019; McGraw and Mandl, 2020; Goldstein and Holtzman, 2021), hallucination of critical clinical facts (Menny and Braun, 2019; Mamo et al., 2023a; Li et al., 2023c), algorithmic bias in diagnostics (Obermeyer et al., 2019; Gonzalez et al., 2022; Venkatasubramanian, 2020), unclear accountability for errors (Kaminski, 2022; Kraft and Jonas, 2021), opacity of generative models (Ghorbani et al., 2019; Tjoa and Guan, 2023; Donoso-Guzmán et al., 2025), and difficulty of compliance with traditional regulations (Price and Gerke, 2022; Kaissis et al., 2021; Khalid et al., 2023). Empirical studies highlight these risks: a 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023c); Rosbach et al. (2024) found a 7% automation-bias error rate in a pathology study where clinicians overturned correct diagnoses by over-trusting AI outputs (Rosbach et al., 2024); and a systematic review of 778 studies showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023).
RAI Strategies.
Several strategies are being employed to enable responsible GenAI adoption in healthcare. Grounding methods such as retrieval-augmented generation (RAG) help reduce hallucinations by linking model outputs to verifiable medical sources (Wu et al., 2024). Human-in-the-loop (HITL) validation integrates clinician oversight into diagnostic and documentation workflows to mitigate automation bias (Stojanovic et al., 2021; He et al., 2023b). Privacy-preserving approaches, including differential privacy and federated learning, protect sensitive health records while supporting multi-institutional collaboration (Kaissis et al., 2021; Khalid et al., 2023). Fairness-aware data augmentation and structured fairness audits are being applied to improve equity in clinical outcomes (Pfohl et al., 2022; Chen et al., 2023). Post-hoc explainability techniques that are applied after model training to make opaque AI models interpretable such as SHAP and LIME, are critical for clinician trust (Nasarian et al., 2023; Donoso-Guzmán et al., 2025; Guidotti et al., 2021). Alignment with regulatory frameworks like Food and Drug Administration (FDA) guidance ensures safe and accountable deployment (Kaminski, 2022; Kraft and Jonas, 2021).
5.2. Finance
Applications.
GenAI is transforming multiple facets of finance, with impacts on efficiency, risk management, and personalization (Ngai et al., 2011; Cao, 2022b). Key applications include customer-facing bots that provide interactive banking and wealth-management support (Wang et al., 2023b; Chen and Lee, 2022; Ngai et al., 2011), modeling of fraud detection, credit scoring, risk forecasting, and insurance assessment (Hand, 2018; Mehrabi et al., 2021b; Cao, 2022b), algorithmic trading that uses generative simulations to optimize portfolios and anticipate market shifts (Zhang et al., 2023; Gupta et al., 2024; HSBC Research, 2025), personalized financial advice adapted to client risk profiles and long-term goals (Silva et al., 2024; Mehrabi et al., 2021b; Wang et al., 2023b), automated compliance reports and market summaries that accelerate auditing and regulatory filings (Araci, 2019; Li et al., 2023a; Kaminski, 2022), and synthetic financial data generation for stress testing, scenario design, and risk modeling (Patki et al., 2016; Wiese et al., 2020; Cao, 2022b). Empirical deployments highlight these advances. For example, PayPal’s fraud detection reduced false positives by 50% in 2023 while processing 10 million transactions daily (Ngai et al., 2011). Zest AI raised loan approvals for minority applicants by 15% in a 2024 U.S. bank case study (Hand, 2018; Mehrabi et al., 2021b). Charles Schwab’s assistant resolved 62% of client queries and improved satisfaction by 28% in 2023 (Wang et al., 2023b). JPMorgan’s simulator improved trading returns by 12% and reduced value-at-risk by 13% in 2023 (Zhang et al., 2023). HSBC’s AI achieved 95% accuracy in volatility prediction in 2025 (HSBC Research, 2025; Cao, 2022b). Silva et al. (2024) reported RegTech-GPT summaries rated “complete” by 85% of compliance officers compared with 62% baseline (Silva et al., 2024).
RAI Challenges.
GenAI in finance raises a number of domain-specific risks that complicate its responsible adoption. These include discriminatory lending that replicates historical bias in credit and valuation data (Hand, 2018; Mehrabi et al., 2021b; Adams et al., 2022), security vulnerabilities where adversarial manipulation targets fraud detection or trading systems (Biggio et al., 2013; Kraft and Jonas, 2021; Ngai et al., 2011), consumer data privacy concerns arising from reliance on sensitive financial records (McGraw and Mandl, 2020; Khalid et al., 2023; Goldstein and Holtzman, 2021), opacity of GenAI that limits auditability and accountability (Doshi-Velez and Kim, 2017; Guidotti et al., 2021; Office of the Comptroller of the Currency, 2020), hallucination and misleading outputs in customer-facing systems that may distort investment or compliance advice (Li et al., 2023c; Mamo et al., 2023a), and market manipulation risk that threatens systemic stability when generative simulations influence trading behavior (U.S. Securities and Exchange Commission, 2021; Hendry et al., 2022). Empirical use cases underscore these risks. For example, Zillow’s valuation model undervalued homes in minority neighborhoods and reduced appraised equity for underwriting (Adams et al., 2022). Citibank was fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020). The Robinhood–GameStop episode demonstrated how synthetic order-flow simulations amplified volatility and required regulatory intervention (U.S. Securities and Exchange Commission, 2021).
RAI Strategies.
Multiple strategies are being prioritized to address these financial risks. Structured fairness audits and fairness-aware data augmentation are increasingly required in lending and credit scoring (Pfohl et al., 2022; Chen et al., 2023; Mehrabi et al., 2021b). Post-hoc explainability methods such as SHAP and LIME support regulators in interrogating opaque credit and trading models (Guidotti et al., 2021; Nasarian et al., 2023; Donoso-Guzmán et al., 2025). Privacy-preserving techniques, including differential privacy and secure multiparty computation, enable safe collaboration across financial institutions while protecting consumer data (Kaissis et al., 2021; Khalid et al., 2023; McGraw and Mandl, 2020). Stronger data provenance frameworks are being developed to track the lineage and validity of synthetic datasets (Patki et al., 2016; Wiese et al., 2020). Alignment with regulatory compliance frameworks such as Basel III, SEC, and OCC guidelines anchors AI deployments within enforceable legal standards (Kaminski, 2022; Kraft and Jonas, 2021; Office of the Comptroller of the Currency, 2020). Finally, embedding HITL oversight across lending, trading, and compliance ensures accountability by mandating expert review of generative outputs in high-stakes decisions (Stojanovic et al., 2021; He et al., 2023b; Silva et al., 2024).
5.3. Education
Applications.
GenAI is increasingly shaping critical aspects of education, expanding opportunities for personalization, access, and efficiency (Luckin et al., 2022; Holmes et al., 2022). Key applications include automated grading that accelerates evaluation and reduces instructor workload (Zawacki-Richter et al., 2019; Holmes et al., 2022), personalized tutoring systems that adapt to student knowledge gaps and learning styles (Kulik and Fletcher, 2023; Carnegie Learning, 2023), content generation that produces adaptive curricula and study materials (Kasneci et al., 2023; Zhang and Zhao, 2024), accessibility tools that deliver multimodal and multilingual support for inclusive education (Almalki and Khan, 2023; Holstein and colleagues, 2023), and interactive simulations that foster experiential and inquiry-driven learning (Chen and Huang, 2023; Lai and Xu, 2024). Recent deployments provide evidence of these advances. Carnegie Learning’s MATHia platform improved algebra test scores by 22% in U.S. high schools in 2023 (Carnegie Learning, 2023). A 2024 Arizona State University pilot found that students using an LLM-powered Socrates chatbot showed 42% higher engagement on surveys (Arizona State University, 2024). A Duolingo study in 2024 reported a 12% increase in learner retention and practice frequency through generative conversational exercises (Duolingo Research, 2024). A U.K. school district pilot in 2024 showed that 87% of students felt classroom chat assistants clarified concepts and 92% of teachers reported time saved (UK Department for Education, 2024).
RAI Challenges.
Responsible use of GenAI in education faces several pressing concerns. These include cultural insensitivity where models fail to capture local norms and values (Kasneci et al., 2023; UNESCO, 2021), student data privacy and particularly minors’ data privacy given the sensitivity of educational records (Holmes et al., 2022; Livingstone et al., 2021), skills evasion where over-reliance on AI weakens critical thinking and independent problem-solving (Luckin et al., 2022; Lai and Xu, 2024), and pedagogical opacity that limits educators’ ability to evaluate how generative systems produce outputs (Holstein and colleagues, 2023; Zawacki-Richter et al., 2019). Empirical use cases underscore these risks. For example, India’s government-run “Newton” homework bot trained only on U.S. and U.K. corpora recommended “pork stew” to rural students in Rajasthan in 2021, violating cultural and nutritional guidelines and prompting suspension of the tool for eleven weeks (UNESCO GEM Report, 2022). A 2023 randomized controlled trial at the University of Cyprus found that physics students allowed to use GPT-4 for weekly assignments scored twenty-three percent lower on conceptual-transfer exam questions despite higher weekly grades, demonstrating over-reliance effects (Pavlou et al., 2023). In 2022, the Utah State Board of Education cancelled a contract for an AI essay-scorer when auditors reported that the vendor could not document the feature vectors or training data, raising concerns about pedagogical opacity and compliance with state education law (Utah State Board of Education, 2022).
RAI Strategies.
Multiple strategies are being explored to ensure responsible generative AI in education. Fairness audits and external ethics audits verify alignment with frameworks such as the UNESCO AI-ethics rubric and can keep bias metrics below accepted thresholds (UNESCO, 2021; Pfohl et al., 2022; Chen et al., 2023). Privacy-preserving techniques including homomorphic encryption and federated learning protect minors’ sensitive data while enabling large-scale analysis (Kaissis et al., 2021; Khalid et al., 2023). Clear regulation and policy guidelines developed by international and national bodies provide governance and accountability mechanisms (UNESCO, 2021; Holmes et al., 2022). Finally, academic integrity tools such as watermarking and output tracking are being adopted to prevent plagiarism and monitor AI-assisted work (Kasneci et al., 2023; Zhang and Zhao, 2024).
5.4. Arts and Entertainment
Applications.
GenAI is reshaping creative production and cultural participation in the arts and entertainment sector (Crawford, 2021; Traitter and Klein, 2022). Prominent applications include creative writing tools that assist in script, story, and lyric development (McStay, 2023; Zhang and Lee, 2023), generative art and design platforms that expand aesthetic experimentation (Elgammal et al., 2019; Schneider and Jörgens, 2022), music composition and production systems that automate melody, arrangement, and mastering (Davis et al., 2024; Briot et al., 2020), film production support for visual effects, dubbing, and storyboarding (Miranda and Perez, 2022; Rodriguez and Santos, 2024), personalized recommendation engines that curate media consumption at scale (Celma, 2010; Anderson and Lee, 2023), and virtual influencers and avatars that shape marketing and fan engagement (Traitter and Klein, 2022; Crawford, 2021). Recent deployments provide evidence of these advances. For instance, Spotify’s Discover Weekly playlists improved user engagement by 40% in 2023, with subsequent adjustments in 2024 reducing genre bias (Anderson and Lee, 2023). A 2024 marketing campaign featuring AI-driven virtual influencers reached more than ten million users (Crawford, 2021; Traitter and Klein, 2022). Davis et al. (2024) released a ten-track AI album in which 62% of listeners in a Spotify trial could not distinguish the music from human compositions (Davis et al., 2024).
RAI Challenges.
Responsible use of GenAI in arts and entertainment faces a series of ethical, legal, and cultural concerns. These include copyright infringement and intellectual property violation where AI systems reuse creative material without authorization (U.S. Copyright Office, 2023; Samuelson, 2023), job displacement that threatens creative labor markets (Crawford, 2021; McStay, 2023), cultural insensitivity when generative tools fail to respect community values or norms (UNESCO, 2021; Kasneci et al., 2023), deepfakes and misinformation that erode authenticity and trust (Chesney and Citron, 2019; Fazio, 2023), authorship and agency disputes where creative ownership is contested (U.S. Copyright Office, 2023; Samuelson, 2023), insufficient consent in data usage for training models (WeTransfer, 2025; Crawford, 2021), and opaque attribution that obscures the role of AI in the creative process (Schneider and Jörgens, 2022; Elgammal et al., 2019). Empirical use cases underscore these risks. For example, the SAG-AFTRA and WGA strikes of 2023 protested generative AI’s role in replacing talent and using likenesses without consent or compensation. The U.S. Copyright Office’s 2023 ruling in Thaler v. Perlmutter rejected copyright protection for AI-generated works, reinforcing the requirement of human authorship (U.S. Copyright Office, 2023). The public controversy around the 2023 song “Heart on My Sleeve,” which mimicked the voices of Drake and The Weeknd, raised critical questions about derivative works and copyright boundaries. In 2025, WeTransfer updated its terms of service to allow user content for AI training, sparking backlash over transparency and consent (WeTransfer, 2025).
RAI Strategies.
Multiple strategies are being adopted to promote responsible use of generative AI in arts and entertainment. Artist collaboration platforms and alternative compensation models are emerging to safeguard livelihoods and redistribute value (McStay, 2023; Crawford, 2021). Structured bias audits and cultural sensitivity training are being incorporated into creative AI pipelines to mitigate stereotyping (UNESCO, 2021; Pfohl et al., 2022). Deepfake detection tools are being deployed to maintain authenticity in music, video, and media (Chesney and Citron, 2019; Fazio, 2023). Collective bargaining and union agreements such as those negotiated by SAG-AFTRA establish protections for creative workers (SAG-AFTRA, 2023). International ethical guidelines and transparency labeling requirements provide standards for disclosure and attribution (UNESCO, 2021; Samuelson, 2023). Finally, intellectual property management systems are being advanced to secure consent and compliance. For example, a 2023 copyright detection tool identified AI-generated violations with 95% accuracy (Zirpoli, 2023a; Traitter and Klein, 2022), and these system requires explicit informed consent before creating digital replicas of performances, offering a contractual model for ethical rights management.
5.5. Agriculture
Applications.
GenAI is increasingly transforming agricultural practices through precision methods and data-driven innovation (Kamilaris and Prenafeta-Boldú, 2018; Costa et al., 2023). Key applications include precision farming systems that optimize irrigation, fertilization, and pesticide usage (Sapkota et al., 2024; Kamilaris and Prenafeta-Boldú, 2018), crop yield prediction models that integrate weather, soil, and satellite data (Chlingaryan et al., 2018; Ma et al., 2023), pest and disease detection tools that leverage generative augmentation to improve diagnostic accuracy (Mohanty et al., 2016; Li et al., 2022), virtual agronomy advisors that deliver context-sensitive recommendations to farmers (Purdy and Gupta, 2023; Mahdavinejad et al., 2018), agricultural robotics for automated harvesting and monitoring (Duckett et al., 2018; Bechar and Vigneault, 2021), crop breeding simulations that accelerate trait discovery (Varshney et al., 2021; Beltran et al., 2022), and weather forecasting for climate adaptation (Rasp et al., 2020; Nguyen et al., 2023). Recent deployments provide evidence of these advances. For instance, Microsoft’s FarmBeats platform increased crop yields by 20% and reduced water use by 30% in U.S. and Indian trials (Microsoft Research, 2023). Retrieval-augmented advisory chatbots deployed in Kenya and India improved farmers’ decision-making speed and confidence in pest and weather management. A 2024 FAO pilot of digital sandboxes for pest control in Kenyan farms recorded an 18% reduction in pesticide use, demonstrating sustainable applications of GenAI.
RAI Challenges.
Responsible adoption of GenAI in agriculture faces several interrelated challenges. These include hallucination and context-inappropriate advice where unreliable outputs may jeopardize crop management decisions (Mamo et al., 2023b; Stojanovic et al., 2021), farm data ownership conflicts where smallholders lack control over datasets aggregated by multinational platforms (Bronson, 2019; Tzachor et al., 2022), bias toward major corporations and regions that limits the applicability of tools for smallholder and Global South farmers (Sapkota et al., 2024; Rotz et al., 2019), overreliance on AI and erosion of local knowledge that risks displacing traditional farming wisdom (Bronson, 2019; Tzachor et al., 2022), IoT security vulnerabilities that expose connected farm systems to cyberattacks (Mahdavinejad et al., 2018; Khalid et al., 2023), and the persistent digital divide that excludes low-resource communities from advanced AI benefits (Rotz et al., 2019; UNESCO, 2021). Empirical use cases underscore these risks. For example, farmers in sub-Saharan Africa reported reduced trust in AI weather advisors when predictions conflicted with lived knowledge, illustrating both hallucination and cultural misalignment. Cases of contested ownership emerged when farm sensor data collected by multinational seed companies were commercialized without equitable farmer compensation (Bronson, 2019). IoT breaches in 2023 European trials highlighted vulnerabilities in sensor networks that exposed proprietary farm-level data.
RAI Strategies.
Multiple strategies are being pursued to enable responsible deployment of generative AI in agriculture. Geo-aware retrieval and annotation techniques are being designed to contextualize model outputs with local weather, soil, and cultural practices, reducing hallucination risks (Mamo et al., 2023a). Data governance contracts are being negotiated to clarify ownership rights and revenue-sharing models for farm-generated datasets (Tzachor et al., 2022; Bronson, 2019). Embedding human-in-the-loop oversight into agronomy advisory workflows ensures farmers and extension officers validate AI outputs before implementation (Stojanovic et al., 2021; He et al., 2023b). Privacy-preserving IoT integration, including differential privacy and federated learning, protects sensitive farm data across institutions (Kaissis et al., 2021; Khalid et al., 2023). Finally, alignment with sustainability regulations such as FAO and EU Common Agricultural Policy frameworks anchors generative AI deployments in enforceable ecological and social standards (UNESCO, 2021; Tzachor et al., 2022). Together, these approaches enhance both the equity and safety of generative AI in global agriculture.
5.6. Defense
Applications.
GenAI is increasingly influencing defense through simulation, sensing, and decision support (U.S. Department of Defense, 2021; NATO, 2024). Key applications include novel training and war-gaming environments that generate diverse scenarios and accelerate force development (NATO Allied Command Transformation, 2024; Johns Hopkins Applied Physics Laboratory, 2025), threat detection and classification across multi-INT streams that fuse imagery signals and text for faster situational awareness (Wortzel et al., 2025; National Institute of Standards and Technology, 2023), autonomous reconnaissance systems that coordinate teams of uncrewed assets in contested environments (Defense Advanced Research Projects Agency, 2018, 2024), and logistics and supply chain optimization that shortens planning cycles and improves mission readiness (U.S. Department of Defense, 2024; U.S. Department of Defense, Chief Digital and AI Office, 2024). Recent deployments provide evidence of these advances. For instance, DARPA programs such as CODE and successor autonomy efforts demonstrated collaborative uncrewed aircraft operating with minimal communications and supervisor control in flight tests (Defense Advanced Research Projects Agency, 2018; Liu et al., [n. d.]). The U.S. Army extended enterprise decision platforms in 2024 to speed analysis and targeting workflows at scale (U.S. Army, 2024; News, 2024). NATO war-games in 2024 integrated generative scenario simulation and reported that AI-generated vignettes helped surface unanticipated vulnerabilities during exercise analysis (NATO Allied Command Transformation, 2024, 2025; Huang et al., 2025a).
RAI Challenges.
Responsible adoption of GenAI in defense introduces ethical operational and legal concerns. These include autonomous weapons and responsibility gaps where assignment of accountability for the use of force remains unresolved (International Committee of the Red Cross, 2021; United Nations Institute for Disarmament Research, 2025), opacity of decision making that limits auditability and traceability in high-stakes contexts (U.S. Department of Defense, 2021; National Institute of Standards and Technology, 2023), adversarial vulnerabilities that expose perception models and decision aids to spoofing and data poisoning (National Institute of Standards and Technology, 2025a, 2024), escalation of conflicts where fast automated responses may narrow decision windows and increase miscalculation risk (Horowitz et al., 2024; Stockholm International Peace Research Institute, 2024), ethical and legal gaps regarding compliance with international humanitarian law and rules of engagement (NATO, 2024; International Committee of the Red Cross, 2021), research security concerns over dual-use model capabilities and insider threat (U.S. Department of Defense, 2025; Corporation, 2024), protection of critical infrastructure as AI is integrated into command networks and sensors (National Institute of Standards and Technology, 2023; U.S. Department of Defense, 2025), and facial recognition deployment where demographic performance differences raise risks in security operations (Grother et al., 2019, 2022). Empirical use cases underscore these risks. For example, national and alliance exercises reported that AI-supported war-gaming surfaced escalatory flashpoints that required added human oversight (NATO Allied Command Transformation, 2024). Government red-team evaluations and standards work identified numerous novel failure modes in generative systems that could enable misinformation or degrade cyber defenses if left unmitigated (National Institute of Standards and Technology, 2025b; WIRED, 2025). Facial recognition evaluations documented demographic differentials that necessitate strict control and testing before operational use (Grother et al., 2019, 2022).
RAI Strategies.
Multiple strategies are being advanced to ensure responsible deployment. Human in the loop oversight is prioritized for mission-critical tasks so that trained operators validate model outputs before action (U.S. Department of Defense, 2021; NATO, 2024). Meaningful human control is reinforced to keep decisions on the use of force within accountable human authority consistent with international humanitarian law (International Committee of the Red Cross, 2021; United Nations Institute for Disarmament Research, 2025). Explainable AI for defense transparency adopts traceability and interpretability requirements to support audit and after-action review (U.S. Department of Defense, 2021; National Institute of Standards and Technology, 2023). Red teaming and rigorous test and evaluation are expanded to probe adversarial and safety failure modes before fielding (Corporation, 2024; National Institute of Standards and Technology, 2025a). International cooperation aligns national practices with alliance principles such as NATO’s Principles of Responsible Use and shared assurance processes (NATO, 2024). Ethical review boards and program-level governance embed policy risk assessment data governance and model documentation across the AI life cycle (U.S. Department of Defense, 2024; Defense Innovation Unit, 2020).
6. Discussion
6.1. Synthesis and interpretation
Our review shows that post-GPT scholarship concentrates on five areas: safety and benchmarking, fairness and governance, explainability and transparency, privacy and security, and sectoral applications. The distribution is uneven. Bias and toxicity receive dense coverage, while deepfakes, sustainability, and system-level failure modes receive relatively less attention. This pattern aligns with the thematic trends we documented and with the growing prominence of multimodal and agentic systems in 2024–2025, which increases the need for evaluation that captures tool use and context.
6.2. Implications for research
The rubric-based analysis in Section 4.2 indicates that commonly used suites provide strong evidence for toxicity and bias, and more limited coverage for privacy, deepfakes, and system-level failures. As a result, many evaluations remain static and task-local, which limits their usefulness for governance where decision makers need evidence that is portable across audits and versions. Future research should target evaluations that model distribution shifts, adaptive adversaries, and multi-agent coordination with external tools.
6.2.1. Implications for practice and policy
For practitioners and policy makers, Tables 5 - 8 map benchmarks to risk families and a governance crosswalk provide a starting point for evidence collection. These tables also reveal gaps that are relevant for conformity assessment under emerging regimes. Applied teams can use the matrix to assemble a minimal set of evaluations that cover privacy, robustness, transparency, logging, and post-market monitoring, and then plan additions for multimodal deepfake risk and adaptive red teaming as required by sectoral context.
6.3. Evidence portability and benchmarking
The scoring formulation offers a transparent and reproducible summary of coverage breadth across ten criteria. It supports comparison within a release and over time when paired with clear documentation of constructs, metrics, and versioned artifacts. We emphasize that the score is a coverage indicator rather than a ranking of overall safety and must be interpreted with the governance crosswalk and with domain risk.
6.3.1. Limitations and threats to validity
This survey follows a pre-specified protocol, focusing on the period from November 2022 to August 2025, and incorporates selected foundational works to ensure definitional clarity. Nevertheless, decision regarding selection and synthesis may introduce bias. Reliance of public papers, dataset cards, and documentation implies that certain capabilities may be under-reported or reported with delay. The rubric weights are assigned equally by default and criterion coverage is determined through structured reading of sources, followed by discussion and adjudication in case of disagreements. Replication of our screening and scoring is facilitated by the availability of the documented sources and explicit inclusion rules.
6.4. Research directions.
We highlight five priorities. First, extend benchmarks to capture system-level failures in tool-using and multi-agent settings, especially around retrieval, planning, and actuation (e.g., API calls, code execution, or robotic steps). Second, fold privacy and provenance checks into routine testing, including membership-inference and training-data extraction probes, plus audits of logging sufficiency. Third, create deepfake risk assessments for multimodal models that cover detection, content provenance, and robustness to evasion. Fourth, shift from static to adaptive evaluation with undisclosed items, targeted red teaming, and tests under distribution shift. Fifth, report sustainability alongside accuracy and robustness—energy, emissions, and compute-normalized efficiency—to enable compute-proportional evaluation and responsible deployment. These priorities directly address the risk-surface gaps surfaced in this survey.
6.4.1. Guidance for practitioners
Teams putting responsible GenAI into practice can follow a simple loop. First, scope risks by mapping planned features to common risk families (e.g., bias, toxicity, privacy, and security). Next, select baselines from our matrix so that each release is covered by minimum tests across those areas. Then, add context: turn on logging, enforce dataset/model versioning, and keep lightweight documentation so decisions and results remain traceable. Before launch—and later during post-market monitoring—exercise adaptivity with red teaming and a few undisclosed checks to catch regressions. Finally, track a small set of KPIs across the lifecycle: calibration, error rates in high-stakes flows, audit cadence, inclusivity indicators, and energy/emissions. This routine keeps evaluation aligned with governance expectations and makes evidence easier to reuse when audits occur.
7. Conclusion
This survey examined responsible GenAI through the lens of governance needs and practical evaluation. We organized the literature into risk families and assurance qualities and introduced a two block rubric that distinguishes risk coverage (C1 to C8) from assurance and governance qualities (C9 to C10). From this review article, three findings stand out. First, coverage is strongest for bias and toxicity and weaker for privacy, deepfakes and provenance, and system level failure in tool using or agentic settings. Second, most evaluations remain static and task local, which limits evidence portability for governance and audit. Third, documentation and metric validity vary substantially, which complicates interpretation across releases and domains. These gaps are significant because contemporary systems are multimodal, often connected to tools and data stores, and operate under distribution shift.
We outlined a research agenda that calls for evaluations addresing system level failure, privacy leakage and provenance, deepfake and media integrity risk, and adaptive adversaries. We further recommended stronger reporting of constructs, metrics, uncertainty, and versioned artifacts, to enhance the portability of the results across audits and over time. For practitioners, we provided a sequence that links product risks to a minimal set of tests, adds assurance artifacts such as documentation and logging, and incorporates red teaming and post release monitoring. Collectively, these steps align engineering practice with governance expectations without prescribing a single benchmark.
This review has several limitations. It focuses on the period from November 2022 through August 2025 with selective inclusion of foundational work, relies on public documentation, and uses equal weights by default for the composite index. We mitigated subjectivity through a prespecified protocol, independent scoring, adjudication, and the release of a codebook, yet some capabilities may be under reported or reported with delay. The governance crosswalk is intended as plausible evidentiary support rather than a statement of legal compliance. Responsible GenAI requires progress on methods and on evidence. Benchmarks should move beyond static probes toward evaluations that reflect real operating conditions, including adaptivity, provenance, and end to end behavior. Governance frameworks should in turn recognize evidence that is documented, reproducible, and linked to explicit constructs. By connecting risk families to concrete tests and by foregrounding assurance qualities, this survey provides a path for researchers and practitioners to close capability gaps, to increase trust, and to support safe and accountable deployment of GenAI systems across domains.
Philosophy and the Way Forward
”Software is eating the world” (Marc Andreessen, 2011), ”AI will eat software” (Jensen Huang, 2017), and perhaps most crucially, ”Philosophy must eat AI. (Schrage and Kiron, 2025)
These shifts capture how technology evolves: from fueling industries, to shaping all digital systems, to raising deep social and ethical questions (McCarthy, 2022). If software rewired markets, AI now rewires trust, knowledge, and decision-making. It cannot remain just a technical pursuit; it must be guided by social values, business needs, and political judgment. We expect The direction of responsible AI will be negotiated not only in research labs but equally in classrooms, communities, and parliaments.
”The measure of intelligence is the ability to change.” — Albert Einstein
This quote encapsulates the core challenge of designing AI systems that can adapt ethically and responsibly to evolving societal norms. Adaptability in AI must go beyond technical performance to include ethical flexibility, ensuring systems align with human values. RAI emphasizes not just technical capabilities but also the ethical imperatives of transparency, fairness, and accountability.
”None of us is as responsible as all of us.” — Authors of this paper
Achieving RAI requires a collective effort among developers, policymakers, and users to mitigate bias, prioritize privacy, and ensure accountability. Collaboration across disciplines and sectors is essential to create equitable and ethical AI systems that serve diverse societal needs.
”For better or worse, benchmarks shape a field.” — David Patterson
Traditional AI benchmarks have prioritized accuracy and efficiency, often neglecting fairness, accountability, and societal impact. Redefining benchmarks to incorporate ethical metrics is critical for guiding the development of AI systems that perform both technically and ethically. Such benchmarks must evaluate transparency, safety, and bias mitigation alongside traditional performance metrics.
The data-driven nature of AI presents persistent challenges, as models often inherit biases and privacy risks from training data. These biases, embedded in AI outputs, require rigorous mitigation strategies to distinguish between useful correlations and harmful patterns. While creating a perfectly unbiased AI system is unattainable due to the evolving nature of societal values (Ai et al., 2018), continuous evaluation and adaptation of AI models are essential to maintain alignment with ethical standards.
Human decision-making is influenced by moral reasoning and context, while AI operates strictly within its programmed algorithms. This underscores the importance of fostering AI literacy (Long and Magerko, 2020) among developers and users to ensure responsible deployment and use. To implement RAI effectively, AI systems must prioritize transparency, fairness, and adaptability. They should incorporate ethical reasoning and evolve alongside societal norms. As we advance toward Artificial General Intelligence, the true measure of AI should extend beyond technical proficiency to include ethical adaptability (Blili-Hamelin et al., 2024).
The pursuit of RAI is challenging yet indispensable. Ongoing advancements in RAI frameworks, coupled with a growing emphasis on AI ethics, offer a path forward. Prominent voices in the field, including Jensen Huang and Sam Altman, emphasize the need for safe and ethical AI. By aligning technical innovation with ethical principles, we can build AI systems that uphold accountability, foster trust, and adapt responsibly to the complexities of a dynamic world.
8. Acknowledgement
This work received no external funding and was undertaken purely out of a passion for learning and advancing knowledge, and safe and responsible deployment of AI, for the betterment of society.
Appendix
Appendix A Full Search Strings and Filters
To enable exact replication, we report verbatim queries and filters for each source. For all databases, we limited the date range to November 2022–August 2025 and restricted results to English. Field constraints (title/abstract/keywords) were used where supported; otherwise, queries were applied to full text. Below we list the precise strings and any database-specific options.
ACM Digital Library.
Fields: Title/Abstract/Keywords. Years: 2022–2025. Document types: Journal, Proceeding Article.
("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR accountability
 OR fairness OR "bias mitigation" OR privacy OR equity OR "open science")
AND ("generative AI" OR "large language model*" OR LLM OR LVLM OR "vision-language"
 OR multimodal OR "natural language processing" OR "computer vision")
IEEE Xplore.
Metadata search (Abstract + Index Terms). Years: 2022–2025. Content type: Journals, Conferences.
(("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR accountability
 OR fairness OR "bias mitigation" OR privacy OR equity) NEAR/5 (LLM OR LVLM OR
 "generative AI" OR multimodal OR "vision-language" OR "large language model*"))
SpringerLink / ScienceDirect.
Title/Abstract/Keywords. Years: 2022–2025. Subject areas: Computer Science, AI.
("Responsible AI" OR "AI governance" OR "AI safety" OR fairness OR transparency
 OR accountability OR privacy) AND ("generative AI" OR LLM OR LVLM OR multimodal)
Web of Science.
Topic search (TS=Title, Abstract, Keywords). Years: 2022–2025.
TS=("Responsible AI" OR "Ethical AI" OR "AI safety" OR fairness OR transparency
 OR accountability OR privacy) AND TS=("generative AI" OR LLM OR LVLM OR multimodal)
DBLP.
Title search; manual verification of scope.
("Responsible AI" OR "Ethical AI" OR "AI safety" OR fairness OR transparency
 OR accountability OR privacy) AND (LLM OR LVLM OR "generative AI" OR multimodal)
arXiv.
Categories: cs.AI, cs.LG, cs.CV, cs.CL. Date filter applied via API.
(all:"Responsible AI" OR all:"AI safety" OR all:"fairness" OR all:"transparency"
 OR all:"accountability" OR all:"bias mitigation" OR all:"privacy")
AND (all:"generative AI" OR all:"large language model" OR all:"LLM" OR all:"LVLM"
 OR all:"vision-language" OR all:"multimodal")
Google Scholar.
Custom range: Nov 2022–Aug 2025; “include citations” disabled during screening.
"Responsible AI" OR "AI governance" OR "AI safety" fairness transparency accountability
 privacy "generative AI" LLM LVLM multimodal
For each included study we recorded venue and year, dominant modality, problem area, method type, risks and controls addressed, datasets or artifacts, evaluation facets and metrics, and the type of evidence advanced. We adapted appraisal criteria from established CS/SE review checklists to rate method transparency, data/source adequacy, discussion of validity threats, and presence of reproducibility signals. Each study received qualitative notes and an ordinal rating (Low/Medium/High) per criterion; we also assigned an evidence-strength tag that informed interpretation during synthesis.
Figure A.1. Taxonomy diagram categorizing the core concepts of Responsible AI (RAI) into principal subsets: Explainable AI (XAI), Ethical Considerations, Sector Applications, Advanced Models, Privacy and Security, and Regulatory Frameworks. Each subset is further dissected into detailed elements such as Transparency, Bias Mitigation, Healthcare, Generative Models, and Data Anonymization, illustrating the flow from general themes to specific methodologies within RAI:
Figure A.2. Framework for Responsible Generative AI emphasizing transparency, fairness, accountability, privacy, and bias mitigation through the integration of data, models, regulations, and user-centric practices
Appendix B Codebook for rubric C1 to C10
Use. Score only explicit tasks and metrics. If a task maps to multiple risks, assign it to the primary stated objective to avoid double counting.
Evidence priority. Official paper and docs; then dataset or model cards and maintained leaderboards; then widely cited third party reports.
C1 Bias and discrimination
Definition: allocational or representational harms across protected or salient groups in generation or decisions.
Evidence: group labels, group aware metrics, disparity or parity targets, bias stress tests.
Scoring: 0 none; 1 single group axis or indirect proxy; 2 multiple axes with stated metrics and protocols.
C2 Toxicity
Definition: abusive, hateful, profane or harassing content and unsafe advice in outputs.
Evidence: toxicity classifiers or human ratings with thresholds, curated harmful prompt sets.
Scoring: 0 none; 1 one classifier or small probe; 2 multiple probes or raters with thresholds and guidance.
C3 Security and robustness
Definition: failure under adversarial input, perturbations, jailbreaks, or distribution shift.
Evidence: specified threat models, attack recipes, robustness metrics, red teaming logs.
Scoring: 0 none; 1 static or single attack; 2 multiple attacks or shifts with adaptive or transfer evaluation.
C4 Misinformation and disinformation
Definition: factual inaccuracy, persuasion or manipulation, synthetic news or citation fabrication.
Evidence: fact checking tasks, grounded QA, hallucination stress tests, persuasion tests with controls.
Scoring: 0 none; 1 limited fact checks or narrow tasks; 2 diverse tasks with documented constructs and metrics.
C5 Deepfakes and media integrity
Definition: deceptive or synthetic media that undermines provenance or identity.
Evidence: detection tasks, watermark or provenance checks, cross modal authenticity tests.
Scoring: 0 none; 1 single modality or heuristic detector; 2 multi modal detection or provenance evaluation.
C6 Privacy
Definition: leakage of personal or training data through extraction or membership inference.
Evidence: membership inference, extraction benchmarks, memorization audits, privacy accounting.
Scoring: 0 none; 1 one attack or audit without coverage claims; 2 multiple attacks or audits with reporting.
C7 System level failure
Definition: end to end failure in tool use, retrieval augmented generation, or agent pipelines.
Evidence: task graphs, tool calls, rollback or refusal criteria, end to end success and failure logs.
Scoring: 0 model only; 1 limited tool or Retrieval-augmented generation (RAG) probes; 2 end to end tasks with failure and recovery criteria.
C8 Malicious actor realism
Definition: adversary realism through red teaming, iterative prompting, or adaptive pressure.
Evidence: scripted or human in the loop red teams, iterative attack frameworks, coverage reports.
Scoring: 0 none; 1 scripted or static red team; 2 adaptive or iterative red team with metrics.
C9 Metric validity and documentation
Definition: clarity of constructs, metrics, labeling, and reproducibility.
Evidence: task construct definitions, annotation guides, inter rater agreement, confidence intervals.
Scoring: 0 undocumented; 1 partial documentation or single metric; 2 full docs with reliability and uncertainty.
C10 Governance alignment
Definition: traceable mapping to controls or obligations in NIST AI RMF, EU AI Act, ISO or similar.
Evidence: control mappings, records or artifacts usable as evidence, audit trails or benchmark statements.
Scoring: 0 none; 1 informal references; 2 explicit control mapping or evidentiary artifacts.
Appendix C Scoring protocol, reliability, and sensitivity
Procedure. Two reviewers score each criterion using this codebook; disagreements resolved by a third reviewer. Publish the codebook, score sheets, and evidence extracts.
Reliability. Report Cohen’s $\kappa$ per criterion and the mean $\bar{\kappa}$; target $\bar{\kappa}\geq 0.60$.
Indices. Report the risk coverage index $S_{\mathrm{cov}}$, coverage rate $R$, assurance index $S_{\mathrm{assur}}$, and the overall normalized index $S$ as defined in Section 4.2.1.
Sensitivity. Recompute $S$ under alternate weights that prioritize governance critical risks and report Spearman rank correlation with the equal weight baseline.
Non compliance note. The governance crosswalk indicates plausible evidentiary support, not legal compliance.
Appendix D Glossary (Key Terms & Acronyms)
RAI:
Responsible AI: practices and governance to ensure fairness, privacy, safety, transparency, and accountability across the AI lifecycle.
GenAI:
Generative AI: models that synthesize content (text, image, audio, code) from learned distributions.
LLM:
Large Language Model (e.g., GPT, Llama): transformer-based models trained on large text corpora.
VLM:
Vision–Language Model: multimodal models handling text+image inputs/outputs (e.g., BLIP-2, LLaVA).
RAG:
Retrieval-Augmented Generation: generation conditioned on retrieved external context.
Agentic AI:
Systems where models plan, select tools, act, and maintain memory; may be multi-agent.
XAI:
Explainable AI: methods to make model behavior intelligible (ante-hoc/by-design or post-hoc).
LIME:
Local Interpretable Model-agnostic Explanations: local surrogate explanations.
SHAP:
SHapley Additive exPlanations: additive attributions based on cooperative game theory.
Grad-CAM:
Gradient-weighted Class Activation Mapping: visual saliency for CNN/VLM vision backbones.
Calibration:
Agreement between predicted confidence and empirical accuracy.
ECE:
Expected Calibration Error: summary gap between confidence and accuracy across bins.
Brier score:
Mean squared error of probabilistic predictions vs. outcomes (lower is better).
SPD:
Statistical Parity Difference: $\Pr(\hat{y}=1|A=a)-\Pr(\hat{y}=1|A=b)$ (group outcome gap).
DI:
Disparate Impact: $\Pr(\hat{y}=1|A=a)/\Pr(\hat{y}=1|A=b)$ (parity ratio; use $\min(\mathrm{DI},1/\mathrm{DI})$).
Equalized odds:
Parity of TPR/FPR across protected groups (conditional on true label).
Robustness:
Performance stability under noise, distribution shift, or adversarial perturbations.
Red teaming:
Structured adversarial testing to elicit failures or unsafe behavior.
Benchmark:
Standardized tasks/metrics for evaluation (e.g., safety, bias, privacy).
KPI:
Key Performance Indicator: operational metric tracked across the lifecycle (e.g., SPD, ECE, kWh).
DP:
Differential Privacy: formal privacy guarantee bounding information leakage about individuals.
Membership inference:
Attack inferring whether a record was in training data.
Data extraction:
Recovery of memorized training snippets from a model.
PII:
Personally Identifiable Information: data that identifies an individual.
Provenance:
Traceability of data/model/content origin; includes watermarking and content credentials.
Deepfake:
Synthetic media that mimics real persons/events (face/voice/image/video).
System-level failure:
Failures emerging from tool use, planning, memory, or multi-agent interaction.
Governance crosswalk:
Mapping of technical evidence (benchmarks/KPIs) to regulatory obligations.
NIST AI RMF:
U.S. National Institute of Standards and Technology AI Risk Management Framework.
EU AI Act:
European Union regulation with risk tiers and obligations (robustness, transparency, monitoring).
ISO/IEC 42001:
AI management system standard for governance and operational controls.
SAIF:
Secure AI Framework (Google): security practices for AI systems.
AIF360:
IBM AI Fairness 360: fairness metrics and mitigation toolkit.
Fairlearn:
Microsoft toolkit for fairness assessment and mitigation.
Model card:
Standardized documentation of a model’s intended use, data, metrics, and limits.
Datasheet:
Documentation for datasets (provenance, consent/licensing, collection, intended use).
Post-market monitoring:
Ongoing evaluation and incident response after deployment.
Risk families (C1–C10):
Rubric criteria: C1 Bias/Discrimination, C2 Toxicity, C3 Security/Robustness, C4 Mis/Disinformation, C5 Deepfakes/Media Integrity, C6 Privacy, C7 System-level Failure, C8 Adversarial/Red-team Realism, C9 Metric Validity/Documentation, C10 Governance Alignment.
Responsible AI (RAI). A paradigm that embeds fairness, accountability, transparency, privacy, and safety across the AI lifecycle (design, training, deployment, and governance).
Generative AI (GenAI). Models that synthesize novel content (text, image, audio, multimodal) from learned data distributions, including large language models (LLMs), diffusion models, and vision–language models (VLMs).
Bias. Systematic and unfair deviations in model predictions across demographic groups, often stemming from imbalanced or unrepresentative training data.
Fairness. The principle of ensuring equitable outcomes across protected groups, commonly measured through:
$\text{Statistical Parity Difference (SPD)}=P(\hat{y}=1\mid A=a)-P(\hat{y}=1\mid A=b),$
$\text{Disparate Impact (DI)}=\frac{P(\hat{y}=1\mid A=a)}{P(\hat{y}=1\mid A=b)}.$
Hallucination. Model outputs that are factually incorrect, unsupported by evidence, or inconsistent with the input context.
Explainability (XAI). Methods that render black-box model behavior interpretable to humans (e.g., SHAP, LIME, attention maps), enabling accountability and compliance.
Privacy. Safeguarding sensitive data against unauthorized exposure, often operationalized via techniques such as Differential Privacy (DP) and Federated Learning (FL).
Robustness. The ability of models to maintain reliable performance under distribution shifts, adversarial perturbations, or noisy inputs.
Toxicity. Harmful, offensive, or unsafe outputs produced by generative models, typically evaluated through specialized benchmarks.
Deepfakes / Media Integrity. Synthetic or manipulated media generated by GenAI that can mislead or cause societal harm; a central integrity and governance concern.
Governance Frameworks. Formal standards and policies (e.g., NIST AI RMF, EU AI Act, ISO/IEC 42001, OECD AI Principles) that regulate responsible development and deployment.
Key Performance Indicators (KPIs). Quantifiable measures that assess AI responsibility and compliance. For example, calibration can be measured by the Expected Calibration Error (ECE):
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\Big|\text{acc}(B_{m})-\text{conf}(B_{m})\Big|,$
where $B_{m}$ denotes prediction bins, $\text{acc}(B_{m})$ the empirical accuracy, and $\text{conf}(B_{m})$ the mean confidence.
Agentic AI. Generative systems with planning, memory, and tool-use capabilities that enable autonomous decision-making but also introduce new safety and governance challenges.

### v6

@@
-
+Who is Responsible? Data, Models, or Regulations, A Comprehensive Survey on Responsible Generative AI for a Sustainable Future
 Abstract.
-Generative AI is moving rapidly from research into real world deployment across sectors, which elevates the need for responsible development, deployment, evaluation, and governance.
-To address this pressing challenge, in this study, we synthesize the landscape of responsible generative AI across methods, benchmarks, and policies, and connects governance expectations to concrete engineering practice.
-We follow a prespecified search and screening protocol focused on post-ChatGPT era with selective inclusion of foundational work for definitions, and we conduct a narrative and thematic synthesis.
-Three findings emerge; First, benchmark and practice coverage is dense for bias and toxicity but relatively sparse for privacy and provenance, deepfake and media integrity risk, and system level failure in tool using and agentic settings.
-Second, many evaluations remain static and task local, which limits evidence portability for audit and lifecycle assurance.
-Third, documentation and metric validity are inconsistent, which complicates comparison across releases and domains.
-We outline a research and practice agenda that prioritizes adaptive and multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration and uncertainty reporting, versioned and documented artifacts, and continuous monitoring.
-Limitations include reliance on public artifacts and the focus period, which may under represent capabilities reported later.
-The survey offers a path to align development and evaluation with governance needs and to support safe, transparent, and accountable deployment across domains.
+Generative AI is rapidly moving from research to deployment, elevating the need for responsible development, evaluation, and governance.
+We conduct a PRISMA-guided review of 232 studies (November 2022–December 2025), spanning large language models, vision-language models, diffusion models, and agentic pipelines.
+We make four contributions: (1) the first survey bridging governance principles, technical evaluation, and domain deployment across all four system types; (2) a ten-criterion rubric (C1–C10) scoring major AI safety benchmarks on risk-surface coverage, paired with a policy crosswalk mapping benchmarks to regulatory requirements; (3) twelve lifecycle KPIs, explainability guidance for foundation models, and a testbed catalogue; and (4) domain-specific analysis across healthcare, finance, education, arts, agriculture, and defense.
+Three findings emerge: benchmark coverage is dense for bias and toxicity but sparse for privacy, provenance, deepfakes, and system-level failures in agentic settings; evaluations remain largely static and task-local, limiting audit portability; and inconsistent documentation complicates cross-release comparison.
+We outline a research agenda prioritizing adaptive multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration reporting, versioned artifacts, and continuous monitoring.
+This survey offers a structured path to align generative AI evaluation with governance needs for safe and accountable deployment.
 Project page: Project page
-GitHub: GitHub
 responsible AI, generative AI, AI ethics, bias, privacy, trust, safety
@@
 1. Introduction
-The rapid advancement of Artificial Intelligence (AI) has transformed it from a theoretical construct into a pervasive technology that shapes modern society.
-AI systems are embedded in critical decision making across healthcare (Kaur et al., 2022), finance (Cao, 2022a), transportation (Abduljabbar et al., 2019), manufacturing industry (Jiang et al., 2023), education (Zhai et al., 2021), and others (Hadi et al., 2023).
-As these systems influence outcomes that affect individuals, communities, and economies, the need for ethical, transparent, and accountable deployment has become urgent (Liu et al., 2025b).
+Generative AI (GenAI) systems, including large language models (LLMs), vision language models (VLMs), diffusion models, and emerging agentic pipelines, are rapidly transitioning from research prototypes to production deployments that generate text, images, code, and autonomous actions across healthcare, finance, education, and other sectors (Hadi et al., 2023; Hajkowicz et al., 2023).
+This accelerated deployment heightens the need for responsible development practices that address GenAI-specific risks, including hallucination and unverifiability, prompt injection and jailbreak vulnerabilities, data memorization and leakage, deepfake and media integrity threats, and system-level failures in tool-augmented settings.
 Responsible AI (RAI) (Stryker, 2024) has emerged as a central area of research and practice to address these challenges.
 RAI frameworks emphasize fairness, bias mitigation, privacy protection, security, and the safeguarding of human rights (Peters et al., 2020).
-By embedding these principles throughout the AI lifecycle, including design, training, deployment, and governance (Raza et al., 2024a), RAI seeks to realize the societal benefits of AI while minimizing potential harms (Lu and Wu, 2025).
-There has been substantial progress on RAI from both industry (Google AI, 2024; Stryker, 2024; Microsoft, 2024; Anthropic, 2024; Jiang et al., 2025) and academia (Lu et al., 2024b; Dignum and Confidential, 2019; Schiff et al., 2020; Huang et al., 2025b).
+By embedding these principles throughout the AI lifecycle, including design, training, deployment, and governance, RAI seeks to realize the societal benefits of AI while minimizing potential harms.
+While these principles were initially developed for predictive systems, generative models introduce qualitatively new challenges: outputs are open-ended and difficult to enumerate, stochastic decoding produces non-deterministic behaviour, multimodal synthesis creates cross-modal failure modes, and agentic tool use extends the attack surface beyond the model itself.
+Despite rapid adoption, with 74% of organizations reporting the use of GenAI, only 26% report having a comprehensive organization-wide responsible AI strategy (Capgemini Research Institute, 2024).
+This imbalance underscores the need to align the development and deployment of these systems with ethical standards, fairness, privacy, security, and societal values.
+There has been substantial progress on RAI from both industry (Google AI, 2024; Stryker, 2024; Microsoft, 2024; Anthropic, 2024; Jiang et al., 2025) and academia (Lu et al., 2024b; Dignum and Confidential, 2019; Schiff et al., 2020; Huang et al., 2025).
 Nevertheless, notable gaps remain in alignment with human values and in operational practice.
-The academic literature often separates conceptual frameworks (Lu et al., 2024b; Saeed and Omlin, 2023; Minh et al., 2022; Mehrabi et al., 2021a) from practical implementations (Huang et al., 2024a; Dwivedi et al., 2023; Caton and Haas, 2024), as summarized in Table 1.
+The academic literature often separates conceptual frameworks from practical implementations, as summarized in Table 1.
 Conceptual work provides principles and guidelines, whereas applied studies demonstrate deployment in specific contexts.
-The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 5, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (nis, 2024), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (Systems, 2021).
+The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 5, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (of Standards and, NIST), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (ISO/IEC, 2021b).
 This gap between governance coverage and practical implementation highlights the need for an integrated review that connects frameworks to implementation strategies and to evaluation methods, including explainable AI and AI-ready testbeds.
-The emergence of powerful generative AI models, including conversational and multimodal systems, also require a reassessment of responsible practices (Mhlanga, 2023).
-Generative AI (GenAI) is transforming products and workflows across domains (Kenthapadi et al., 2023).
-Despite rapid adoption, with 74% of organizations reporting the use of GenAI, only 26% report having a comprehensive organization wide RAI strategy (Capgemini Research Institute, 2024).
-This imbalance underscores the need to ensure that development and deployment align with ethical standards, fairness, privacy, security, and societal values (Pan and Xu, 2025).
-Problem statement.
-Current RAI guidance for GenAI is fragmented across principles, processes, and tools (as shown in Table 1) and there is less focus on governance-aligned metrics and AI-ready testbeds that enable comparable, auditable assessment across the lifecycle; especially for GenAI systems.
-As a result, organizations cannot reliably connect regulatory obligations to concrete engineering controls or produce portable evidence for audits and post-market monitoring.
-This survey identifies a central problem.
-Current RAI guidance is fragmented across principles, processes, and tools, and it lacks consistent metrics and AI-ready testbeds that enable comparable assessment across the lifecycle of GenAI systems.
-Existing surveys (Table 1) and frameworks rarely integrate governance requirements with actionable engineering practices and with systematic evaluation protocols for generative models, including multimodal systems and agentic pipelines.
+In this study, we conduct a comprehensive literature review of responsible GenAI (ResGenAI) frameworks, governance guidelines, evaluation benchmarks, and operational risk assessment methodologies.
+Our scope centres on the application of RAI principles to generative systems, including LLMs, VLMs, diffusion models, and agentic/RAG pipelines.
+While we draw on the broader RAI literature for foundational constructs (e.g., fairness, privacy, safety, transparency, accountability), our analysis targets ResGenAI specific risk mechanisms such as hallucination and unverifiability, prompt injections, jailbreaking and tool-use failures, data memorization and leakage, provenance and deepfakes or media integrity risks, intellectual property concerns, and system-level failures in GenAI applications.
+Despite substantial prior work, ResGenAI guidance remains fragmented across principles, processes, and tools (Table 1).
+Consequently, there is a pressing need to translate regulatory obligations into concrete engineering controls and to produce portable audit evidence for post-market monitoring.
+{forest}
+Figure 1.
+Responsible GenAI landscape organised into three pillars: Challenges (Section 3), Strategies (Section 4), and Applications (Section 5).
+Technical challenges span data, model/system, and emergent-threat layers; strategies encompass benchmarks, KPIs, explainability, testbeds, governance, and toolkits; applications map domain-specific risks and controls across six sectors.
 Contributions. Our key contributions in this work are:
   1. (1)
-    We present a unified synthesis that maps governance principles to concrete engineering controls and assurance activities across the AI lifecycle.
+    To the best of our knowledge, this is the first review article that simultaneously spans governance principles, technical evaluation, and domain deployment for GenAI.
+We synthesize findings across the NIST AI RMF, EU AI Act, ISO/IEC 42001, and other instruments listed in Table 6, covering LLMs, VLMs, diffusion models, and RAG topics.
+As shown in Table 1, no prior survey bridges all four thematic areas (CS theory, application, policy, GenAI) and all four novelty dimensions (taxonomy, scope, evaluation, methods) simultaneously.
   2. (2)
-    We present a governance-aligned comparative rubric (C1–C10) with normalized scores and a policy crosswalk (NIST AI RMF, EU AI Act, ISO/IEC 42001) linking benchmarks to audit evidence.
+    We introduce a ten-criterion rubric (C1–C10) that scores widely used AI safety benchmarks on risk-surface coverage and assurance quality, and a policy crosswalk (Section 4.2.1) that maps each benchmark to the governance requirements it can plausibly evidence.
+The closest efforts (Liu et al., 2025b; Ferdaus et al., 2026) partially address evaluation taxonomy or trust assessment, but neither links benchmarks to regulatory audit-evidence needs.
   3. (3)
-    We also present a practice roadmap with lifecycle KPIs, explainability guidance for GenAI works, and AI-ready testbeds to support continuous, auditable evaluation.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Responsible GenAI
-Challenges
-Technical
-Data-related: representativeness(Raza et al., 2024c), quality(Raza et al., 2025b), ethics(Gebru et al., 2021)
-Model & system-related: bias(Pessach and Shmueli, 2022), privacy(Dwork, 2006), misinformation(Wang et al., 2023a)
-Regulatory & Sociotechnical
-IP(Kenthapadi et al., 2023), alignment(Bai et al., 2022), explainability(Dwivedi et al., 2023), sustainability(Khan et al., 2025), dual use(Weidinger et al., 2022)
-Strategies
-Safety benchmarks
-HELM(Liang et al., 2023), SafetyBench(Zhang et al., 2024), Rainbow Teaming(Samvelyan et al., 2024), ALERT(Tedeschi et al., 2024)
-Key Performance Indicators
-Statistical parity difference(Kun, 2015), disparate impact(Barocas and Selbst, 2016), emissions(Khan et al., 2025)
-Explainability methods
-SHAP(Lundberg and Lee, 2017), LIME(Ribeiro et al., 2016a), integrated gradients(Arrieta et al., 2020), attention maps(Jain and Wallace, 2019; Serrano and Smith, 2019)
-AI-ready testbeds
-AI4EU(Foundation, 2024), IEEE Testbed(Foundation, 2024), logging(nis, 2024), red-teaming(Samvelyan et al., 2024; Tedeschi et al., 2024)
-Governance frameworks
-NIST AI RMF(nis, 2024), ISO/IEC 42001(Systems, 2021), EU AI Act(Outeda, 2024), OECD principles(Dignum, 2019)
-Technical toolkits
-Model cards(Liang et al., 2023), datasets datasheets(Gebru et al., 2021) , C2PA provenance standards(Das et al., 2025)
-Applications
-Healthcare
-Concerns: Privacy(Dwork, 2006), hallucination(Wang et al., 2023c), bias in diagnostics(Kaur et al., 2022)
-Solutions: RAG(Lewis et al., 2020), HITL(Weidinger et al., 2021), Differential Privacy(Dwork, 2006), FDA Alignment(Laux et al., 2024)
-Finance
-Concerns: Discriminatory lending(Barocas and Selbst, 2016), security(Staab et al., 2024), manipulation(Wang et al., 2023a)
-Solutions: Fairness audits(Pessach and Shmueli, 2022), SHAP/LIME(Lundberg and Lee, 2017; Ribeiro et al., 2016a), Basel III/SEC(Slattery et al., 2024)
-Education
-Concerns: Cultural insensitivity(Al-kfairy et al., 2024), privacy(Dwork, 2006), skills evasion(Lin et al., 2024)
-Solutions: Fairness audits(Pessach and Shmueli, 2022), Homomorphic encryption(Bashir et al., 2024), ethics(Dignum, 2019)
-Arts & Entertainment
-Concerns: Copyright(Kenthapadi et al., 2023; Freeman et al., 2024), job displacement(Pan and Xu, 2025), deepfakes(Shi et al., 2024)
-Solutions: Artist collaboration(Freeman et al., 2024), Deepfake detection(Shi et al., 2024)
-Agriculture
-Concerns: hallucination(Raza et al., 2025d), data ownership(Mittal et al., 2024), digital divide(Al-kfairy et al., 2024)
-Solutions: Geo-aware RAG(Lewis et al., 2020), Data contracts(Mittal et al., 2024), HITL(Weidinger et al., 2021)
-Defense
-Concerns: Autonomous weapons(Weidinger et al., 2022), opacity(Dwivedi et al., 2023), adversarial attacks(Zhang et al., 2024)
-Solutions: Human Control(Weidinger et al., 2021), Red teaming(Samvelyan et al., 2024), NATO Principles(Outeda, 2024)
-Figure 1. Responsible AI landscape across challenges, strategies, and applications.
-This review is a structured literature review of theory, applications, and policy for Responsible AI with emphasis on GenAI.
-It follows a prespecified search followed by a narrative and thematic synthesis.
-Themes covered include explainability (XAI), fairness/bias, privacy/security, robustness/safety, governance/compliance, benchmarking and KPIs, and sectoral applications.
-We then operationalize the agenda via a governance-aligned evaluation rubric with a policy crosswalk and a stakeholder-focused practice roadmap with lifecycle KPIs and AI-ready testbeds.
-A mindmap of the paper is shown in Figure 1.
-Table 1. Comparison of literature by venue, year, topic, and coverage across computer science theory, applications, policy, and generative AI. Checkmarks (✓) indicate coverage.
-Paper Venue Year Topic Coverage
-CS Theory Application Policy GenAI
-(Schiff et al., 2020) arXiv 2020 Responsible AI
-(Mehrabi et al., 2021a) ACM CSUR 2021 Bias and Fairness
-(Hu et al., 2021) ACM CSUR 2021 AI Safety
-(Trocin et al., 2023) Springer ISF 2021 Responsible AI
-(Minh et al., 2022) Springer AI Review 2022 Explainable AI
-(Huang et al., 2024a) Springer AI Review 2023 Safety
-(Saeed and Omlin, 2023) Elsevier KBS 2023 Explainable AI
-(Dwivedi et al., 2023) ACM CSUR 2023 Explainable AI
-(Nauta et al., 2023) ACM CSUR 2023 Bias and Fairness
-(Huang et al., 2023a) IEEE TAI 2023 AI Ethics
-(Díaz-Rodríguez et al., 2023) Elsevier Inf. Fusion 2023 Responsible AI
-(Mhlanga, 2023) Springer Nature 2023 Responsible AI
-(Lu et al., 2024b) ACM CSUR 2024 Responsible AI
-(Leschanowsky et al., 2024) Elsevier CHB 2024 Privacy and Safety
-(Caton and Haas, 2024) ACM CSUR 2024 Bias and Fairness
-(Li and Goel, 2024) Springer ISF 2024 Accountability and Transparency
-(Slattery et al., 2024) arXiv 2024 Responsible AI
-(Jedličková, 2024) Springer AI and Society 2024 Responsible AI
-(Al-kfairy et al., 2024) MDPI Informatics 2024 AI Ethics
-(Sadek et al., 2025) AI and Society 2024 Challenges of Responsible AI in Practice
-(Woodgate and Ajmeri, 2024) ACM CSUR 2024 Responsible AI
-(Reuel et al., 2025) ACM FAccT 2025 Responsible AI (Global maturity model)
-(Gadekallu et al., 2025) arXiv 2025 Best Practices of Responsible AI
-(Byrne, 2025) ASEE Peer 2025 U.S. Responsible AI Policy
-(Sadek et al., 2025) Springer AI & Society 2025 Responsible AI
-(Baldassarre et al., 2025) IEEE TAI 2025 Responsible AI
-(Ivanov, 2025) The Service Industries Journal 2025 Responsible use of AI in social science
-Ours ACM CSUR 2025 Responsible AI
+    We operationalise the rubric findings into measurable KPIs (Section 4.2.4), explainability guidance tailored to foundation models (Section 4.3), and a testbed catalogue for continuous evaluation.
+Section 4.2 links these KPIs to specific benchmarks and audit toolkits, providing practitioners with a decision-support matrix for assembling governance-ready evidence.
+  4. (4)
+    We apply the RAI lens to real-world applications and domains (Section 5, demonstrating how technical advancements can be effectively aligned with governance frameworks and regulatory requirements.
+Table 1.
+Comparison of related surveys by venue, year, topic, thematic coverage, and novelty dimensions.
+Checkmarks (✓) indicate coverage; (x) indicates no explicit coverage; (★) indicates partial or limited treatment.
+Novelty dimensions: Taxonomy = structured risk taxonomy or classification; Scope = multimodal + agentic coverage; Evaluation = governance-aligned benchmarking/rubric; Methods = actionable engineering controls or lifecycle KPIs
+Paper Venue Year Topic Thematic Coverage Novelty Dimensions
+CS Theory
+Application
+Policy
+GenAI
+Taxonomy
+Scope
+Evaluation
+Methods
+(Schiff et al., 2020) arXiv 2020 Responsible AI x x x x x
+(Mehrabi et al., 2021a) ACM CSUR 2021 Bias and Fairness x x x x x x
+(Hu et al., 2021) ACM CSUR 2021 AI Safety x x x x x x
+(Trocin et al., 2023) Springer ISF 2021 Responsible AI x x x x x
+(Minh et al., 2022) Springer AI Rev. 2022 Explainable AI x x x x x
+(Huang et al., 2024a) Springer AI Rev. 2023 Safety x x x x
+(Saeed and Omlin, 2023) Elsevier KBS 2023 Explainable AI x x x x x x
+(Dwivedi et al., 2023) ACM CSUR 2023 Explainable AI x x x x x x
+(Nauta et al., 2023) ACM CSUR 2023 Bias and Fairness x x x x x
+(Huang et al., 2023a) IEEE TAI 2023 AI Ethics x x x x x
+(Díaz-Rodríguez et al., 2023) Elsevier Inf. Fusion 2023 Responsible AI x x x
+(Mhlanga, 2023) Springer Nature 2023 Responsible AI x x x x
+(Lu et al., 2024b) ACM CSUR 2024 Responsible AI x x x x x
+(Leschanowsky et al., 2024) Elsevier CHB 2024 Privacy and Safety x x x x x
+(Caton and Haas, 2024) ACM CSUR 2024 Bias and Fairness x x x x x
+(Li and Goel, 2024) Springer ISF 2024 Accountability x x x x x
+(Slattery et al., 2024) arXiv 2024 Responsible AI x x x x x x
+(Jedličková, 2024) Springer AI & Soc. 2024 Responsible AI x x x x x x
+(Al-kfairy et al., 2024) MDPI Informatics 2024 AI Ethics x x x x x
+(Sadek et al., 2025) AI and Society 2024 RAI in Practice x x x x x
+(Woodgate and Ajmeri, 2024) ACM CSUR 2024 Responsible AI x x x x x
+(Reuel et al., 2025) ACM FAccT 2025 RAI (Global maturity) x x
+(Gadekallu et al., 2025) arXiv 2025 Best Practices of RAI x x x x x
+(Byrne, 2025) ASEE Peer 2025 U.S. RAI Policy x x x x x x x
+(Sadek et al., 2025) Springer AI & Soc. 2025 Responsible AI x x x x x
+(Baldassarre et al., 2025) IEEE TAI 2025 Responsible AI x x x x x
+(Ivanov, 2025) Serv. Ind. J. 2025 RAI in Social Sci. x x x x x x
+(Yu et al., 2025) ACM KDD 2025 Trustworthy LLM Agents x x x
+(Dong et al., 2025) Springer AI Rev. 2025 LLM Safeguards x x
+(Ferdaus et al., 2026) ACM CSUR 2025 Trustworthy LLMs x x x x
+(Wang et al., 2025) arXiv 2025 Reasoning Model Safety x x x
+(He et al., 2025b) ACM CSUR 2025 LLM Agent Security x x x x
+(Papagiannidis et al., 2025) Elsevier JSIS 2025 RAI Governance x x x x x
+(Liu et al., 2025b) arXiv 2025 LLM Safety Evaluation x x x x
+Ours ACM CSUR 2026 Responsible GenAI
+Column definitions CS Theory: conceptual frameworks, taxonomies, or algorithmic contributions.
+Application: domain-specific deployment or case studies.
+Policy: governance instruments, regulatory analysis, or compliance guidance.
+GenAI: explicit focus on generative models (LLMs, VLMs, diffusion, agentic).
+Taxonomy: structured risk classification beyond a flat listing of concerns.
+Scope: coverage of multimodal systems and agentic/tool-using pipelines.
+Evaluation: governance-aligned benchmarking with a scoring rubric or crosswalk.
+Methods: actionable engineering controls, lifecycle KPIs, or testbed guidance.
+★ = partial or limited treatment.
 The rest of this paper is organized as follows.
 Section 2 presents the literature review methodology.
-Section 3 provides background on Responsible Generative AI.
-Section 4 reports our results, including thematic findings and benchmark assessments.
-Section 4 also discusses key performance indicators, explainable AI, and AI-ready testbeds, and Section 5 discuss RAI applications, challanges and RAI strategies.
+Section 3 provides background on ResGenAI.
+Section 4 reports our results, including thematic findings, benchmark assessments, key performance indicators, explainable AI, and AI-ready testbeds.
+Section 5 discusses ResGenAI applications, challenges, and strategies.
 Section 6 presents discussion and forward thoughts, and finally, Section 7 concludes with open challenges and future research directions.
-2. Literature Review Methodology
-We present our literature review methodology in this section.
-Objectives and Review Questions.
-This article conducts a literature review of RAI in the post-GPT period and the synthesis of the findings. We address the following key questions:
-  • RQ1: What are the dominant themes, risks, and controls in RAI for GenAI and multimodal systems since late 2022?
+The taxonomy presented in Figure 1 provides a structural blueprint for this survey, establishing the thematic connections between high-level technical challenges and domain-specific applications.
+2. Methodology
+This review adheres to the PRISMA guidelines (Page et al., 2021) to ensure a transparent, reproducible, and rigorous synthesis of evidence.
+This study follows established methodological recommendations for performing systematic reviews within the computer science community (Al-Selwi et al., 2024; Boland et al., 2017a; Keele et al., 2007).
+To ensure comprehensive and consistent reporting, the structural framework was modelled after previous extensive surveys in the domain.
+The mapping procedure utilized in this review is organized into four interconnected phases that covering objectives (Section 2.1), search strategy (Section 2.2), eligibility criteria (Section 2.3), and study selection (Section 2.4) which are visualized in the methodological flow (Fig.
+2) and detailed throughout Section 2.
+For a more exhaustive exploration of these mapping and reporting protocols, foundational works in the field (Boland et al., 2017a; Al-Selwi et al., 2024; Keele et al., 2007; Boland et al., 2017b) offer additional guidance.
+Figure 2. The systematic literature mapping process utilized in this study
+2.1. Research Questions Formulation
+This study conducts a literature review of responsible generative AI (ResGenAI) in the post-GPT period and the synthesis of the findings. We address the following key questions:
+  • RQ1: What are the dominant themes, risks, and controls in ResGenAI and multimodal systems since late 2022?
+    Motivation: The rapid transition of GenAI from research to production has outpaced traditional risk taxonomies, necessitating a contemporary synthesis of how multimodal and agentic systems introduce qualitatively new failure modes beyond predictive ML.
   • RQ2: Which evaluation facets and metrics are most commonly reported, and where are the gaps?
+    Motivation: Identifying the “Evaluation Gap” is critical for determining if current safety suites adequately model the adaptive, iterative pressure applied by real-world adversaries versus merely serving as static, keyword-based refusal scripts.
   • RQ3: What evidence types (empirical, analytical, normative) underpin claims about RAI effectiveness?
+    Motivation: A structural deficit in evidence portability hinders institutional trust; this question seeks to evaluate if safety claims are grounded in verifiable technical artifacts or subjective, unvalidated appraisals.
   • RQ4: How can these findings be organized into a practical framework for research and governance?
-Scope and Protocol.
-The review protocol, including sources, screening rules, and the synthesis plan, was prespecified in line with established computer science review guidance (Boland et al., 2017; Keele et al., 2007), and we followed a structured literature review rather than a traditional systematic mapping study or an empirical meta analysis.
-We mainly focused on literature published from November 2022 through August 2025 to capture rapid methodological and governance developments associated with contemporary generative models.
-Foundational work published before 2022 was selectively included to define terms, situate concepts, and trace relevant intellectual lineages.
-We conducted a narrative and thematic synthesis aligned with the research questions and did not perform a meta analysis because of heterogeneity in study designs, measures, and outcomes.
-Sources and Search Strategy.
-Sources included the ACM Digital Library, IEEE Xplore, SpringerLink, ScienceDirect, arXiv, DBLP, Google Scholar, and Web of Science.
-To mitigate coverage bias, we complemented database queries with backward and forward snowballing from seed works identified as influential during pilot screening.
-Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity, open science), with syntax adapted to each database (full strings in Appendix A).
-Eligibility and Selection.
-We included English-language works within the target time window that addressed RAI in the context of GenAI or multimodal systems and appeared in peer-reviewed venues (journals or conferences), together with a small number of high-impact governance reports or widely cited preprints when they materially shaped discourse.
-We excluded low-credibility grey literature, purely opinionated pieces without substantive argumentation, duplicates, and items whose primary contribution lay outside RAI (e.g., domain applications without responsibility, safety, or governance components).
-Although our emphasis was the post-GPT era, we allowed earlier seminal contributions when necessary for definitional clarity or historical continuity.
-Screening Process and Data Analysis.
-Titles/abstracts were screened against the inclusion criteria, followed by full-text review for items deemed potentially eligible; screening was performed by the first author with a second pass on borderline cases.
-For each included record we recorded the citation key, year, publication type (journal/conference/preprint), title, and (where available) author-provided keywords.
-We then applied a lightweight text-mining pipeline to infer themes (e.g., fairness/bias, explainability, governance/compliance, privacy/security, robustness, safety, agentic, multimodal) from titles and keywords (abstracts when available), and aggregated counts by year and publication type.
-Tables 2 provide basic traceability from literature to claims.
-Table 2. Core synthesis of the literature review: theme, citation, year, and focus.
-Theme Citation Year Focus
-Safety/Benchmarks (Liu et al., 2024) 2024 Introduces a multimodal benchmark to assess safety risks in vision–language models.
-Safety/Benchmarks (Zhang et al., 2024) 2024 Proposes a general safety benchmark for large language models across risk categories.
-Safety/Benchmarks (Gehman et al., 2020) 2020 Measures toxic degeneration using curated prompts to stress test language models.
-Safety/Benchmarks (Li et al., 2024a) 2024 Presents a hierarchical and comprehensive safety benchmark for language models.
-Safety/Benchmarks (Mazeika et al., 2024) 2024 Standardizes evaluation for red teaming and refusal robustness in language models.
-Safety/Benchmarks (Samvelyan et al., 2024) 2024 Generates diverse adversarial prompts for automated safety stress testing.
-Safety/Benchmarks (Wang et al., 2023a) 2023 Provides a broad audit of trustworthiness, including safety and robustness, for GPT models.
-Safety/Benchmarks (Tedeschi et al., 2024) 2024 Benchmarks red teaming effectiveness for assessing model safety responses.
-Fairness/Governance (Mittal et al., 2024) 2024 Outlines dataset design practices aligned with fairness, privacy, and regulatory norms.
-Fairness/Governance (Rodolfa et al., 2021) 2021 Finds minimal fairness and accuracy tradeoff in public policy machine learning tasks.
-Fairness/Governance (Chouldechova, 2017) 2017 Analyzes disparate impact and fairness criteria in recidivism risk prediction.
-Fairness/Governance (Barocas and Selbst, 2016) 2016 Examines how big data practices can produce disparate impact and legal risk.
-Fairness/Governance (Pessach and Shmueli, 2022) 2022 Surveys algorithmic fairness methods and evaluation across application domains.
-Fairness/Governance (Kaur et al., 2022) 2022 Reviews trustworthy AI principles and technical enablers for responsible deployment.
-Fairness/Governance (Dignum, 2019) 2019 Presents a framework for responsible AI grounded in ethics and governance.
-Fairness/Governance (Laux et al., 2024) 2024 Discusses trustworthiness concepts in relation to compliance under the EU AI Act.
-Explainability/XAI (Ribeiro et al., 2016b) 2016 Introduces LIME for model agnostic local explanations of predictions.
-Explainability/XAI (Selvaraju et al., 2017) 2017 Proposes Grad-CAM for visual explanations in convolutional neural networks.
-Explainability/XAI (Arrieta et al., 2020) 2020 Synthesizes concepts, taxonomies, and challenges in explainable artificial intelligence.
-Explainability/XAI (Dwivedi et al., 2023) 2023 Reviews core ideas and techniques for explainable AI with practical guidance.
-Explainability/XAI (Saeed and Omlin, 2023) 2023 Provides a meta survey of XAI challenges and future research directions.
-Explainability/XAI (van der Waa et al., 2021) 2021 Compares rule based and example based explanations on fidelity and user value.
-Explainability/XAI (Mosca et al., 2022) 2022 Reviews SHAP based explanation methods for natural language processing.
-Explainability/XAI (Nauta et al., 2023) 2023 Systematizes evaluation methods for explanations with quantitative protocols.
-Privacy/Security (Dwork, 2006) 2006 Establishes differential privacy as a formal framework for privacy protection.
-Privacy/Security (McMahan et al., 2017) 2017 Introduces federated learning for decentralized training with privacy benefits.
-Privacy/Security (Nasr et al., 2023) 2023 Demonstrates large scale extraction of memorized training data from language models.
-Privacy/Security (Mattern et al., 2023) 2023 Proposes membership inference against language models via neighborhood comparison.
-Privacy/Security (He et al., 2025) 2025 Presents label only membership inference attacks tailored to pre trained language models.
-Privacy/Security (Li et al., 2024b) 2024 Introduces a benchmark to evaluate privacy risks in language models at multiple levels.
-Privacy/Security (Mireshghallah et al., 2024) 2024 Tests privacy leakage using contextual integrity based evaluations of language models.
-Privacy/Security (Staab et al., 2024) 2024 Assesses the potential of large models for anonymization and de identification tasks.
-LLM Safety (Bai et al., 2022) 2022 Aligns models using constitutional principles and self critique to reduce harmful outputs.
-LLM Safety (Weidinger et al., 2021) 2021 Maps ethical and social risks from language models and proposes mitigation avenues.
-LLM Safety (Huang et al., 2024b) 2024 Positions trustworthiness dimensions and evaluation needs specific to large models.
-LLM Safety (Lin et al., 2024) 2024 Surveys techniques for debiasing and reducing hallucination in large models.
-LLM Safety (Wang et al., 2023c) 2023 Overviews ChatGPT applications, limitations, and emerging risks and safeguards.
-In the following section, we outline the background and key challenges of RAI, which set the stage for the synthesis of our review findings.
+    Motivation: Principles without practice lead to “safetywashing.” We aim to bridge the gap between regulatory mandates like the EU AI Act and concrete engineering controls, providing an implementation path for verifiable stewardship.
+2.2. Evidence Acquisition and Search Strategy
+A comprehensive search was conducted across five major academic databases and registers: Scopus ($n=1502$), Web of Science ($n=2798$), IEEE Xplore ($n=279$), ACM Digital Library ($n=656$), and arXiv ($n=78$).
+To capture the surge in GenAI research following the release of ChatGPT, the search window was limited to publications from November 2022 through December 2025.
+The search strings utilized an intersection of terms related to Generative AI (e.g., “Large Language Model”, “Foundation Model”, “LLM”) and Responsibility facets (e.g., “AI Safety”, “Trustworthy AI”, “Alignment”, “Hallucination mitigation”).
+Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity), with syntax adapted to each database (full strings in Appendix A.1).
+Syntax was adapted for each database to support field-specific constraints (Title/Abstract/Keywords), and proximity operators were employed where available to enhance precision.
+To mitigate coverage bias, database queries were complemented with forward and backward snowballing from seed works identified as influential during pilot screening.
+The review protocol, including sources, screening rules, and the synthesis plan, was prespecified in line with established computer science review guidance (Boland et al., 2017a; Keele et al., 2007)
+2.3. Eligibility Criteria
+To ensure methodological rigor and relevance to the research questions (RQ1–RQ4), we applied the following criteria:
+  • Inclusion Criteria: (1) Peer-reviewed journals, conference proceedings, or high-impact preprints materially shaping RAI discourse; (2) English-language works published within the target 2022–2025 window; and (3) Studies proposing technical frameworks, evaluation metrics, or empirical risk assessments for generative or multimodal systems.
+  • Exclusion Criteria: (1) Purely opinionated pieces or editorials without substantive technical argumentation; (2) Short papers or reports lacking sufficient methodological depth; (3) Items whose primary contribution lay outside RAI (e.g., domain applications without safety or governance components); and (4) Redundant duplicates or inaccessible records.
+Figure 3. PRISMA flowchart of the systematic literature identification, screening, and inclusion process
+2.4. Study Selection and Screening
+As detailed in the PRISMA flow diagram (Fig.
+3), the initial search identified 5,313 records.
+Selection was conducted by three independent reviewers to minimize bias.
+Following automated deduplication ($n=2125$) and the removal of records marked as ineligible by automation tools ($n=609$) or other technical reasons ($n=53$), 2,562 unique records remained for screening.
+Titles and abstracts were screened against the eligibility criteria, resulting in the exclusion of 2,153 items.
+Subsequently, 372 reports were sought for full-text retrieval and assessed for eligibility.
+Reports were excluded if they failed to meet quality and technical thresholds ($n=140$) or could not be retrieved ($n=37$).
+This rigorous multi-reviewer process yielded a final synthesis set of 232 studies.
+Data extraction followed a thematic pipeline to categorize findings into risks, controls, and evaluation facets as defined by the research objectives.
+In the following section, we outline the background and key challenges of ResGenAI, which set the stage for the synthesis of our review findings.
 3. Background on Responsible Generative AI
 3.1. Historical Context
-RAI has received intensified scholarly and policy attention with the rapid diffusion of GenAI (Zirpoli, 2023b; Al-kfairy et al., 2024).
-GenAI refers to models that synthesize new samples from learned data distributions (van der Zant et al., 2013), spanning language, vision, audio, and multimodal content.
+RAI has received a lot of intensified scholarly and policy attention with the rapid application of GenAI (Zirpoli, 2023b; Al-kfairy et al., 2024).
+GenAI refers to models that synthesize new samples from learned data distributions (van der Zant et al., 2013), spanning language, vision, audio, and multimodal content  (Muneer et al., 2025).
 Early work combined symbolic rules with probabilistic language models; for example, ELIZA illustrated template driven dialogue in the 1960s (Weizenbaum, 1966).
@@
 Figure 4 highlights the most salient milestones for RAI (alignment, evaluation, governance).
-As shown in the Figure 2, the first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis; initially in images, and later extended to audio and video.
-Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in the introduction of attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
+As shown in the Figure 4.
+The first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis and, concurrently, made deepfakes feasible, creating the media integrity challenge that current governance frameworks struggle to address Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in the introduction of attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
 A second inflection gave rise to foundation models: large transformers pre-trained in web-scale corpora (e.g., GPT-2/3) and bidirectional encoders (e.g., BERT) that catalyzed transfer across tasks.
@@
 In vision, diffusion models (Ho et al., 2020) eclipsed GANs for photorealistic generation; when combined with latent compression, they enabled scalable text-to-image synthesis (Rombach et al., 2022).
-The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023b), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023b) extend prompting beyond text.
-Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation (Raza et al., 2025c), bringing new safety, privacy, and governance requirements.
-Figure 2. Timeline of generative AI milestones.:
+The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023a), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023b) extend prompting beyond text.
+Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation, bringing new safety, privacy, and governance requirements.
+Figure 4. Timeline of generative AI milestones
+A left-to-right timeline from early rule-based systems to GANs, transformers, and contemporary instruction-tuned large models
 3.2. Challenges in Responsible Generative AI
-While GenAI capabilities have accelerated, new risks have emerged in provenance, trust, and governance (e.g., deepfakes, copyright, safety).
-We group these challenges into technical and regulatory domains, noting that user behavior and deployment context mediate how risks materialize.
+We group GenAI challenges into technical and regulatory domains.
 3.2.1. Technical Challenges
 Technical issues arise across data pipelines, model design and training, evaluation and assurance, and operational deployment.
-In the following, we present the main technical challenges in GenAI applications:
 Data-Related Challenges
-  1. (1)
-    Data diversity and representativeness.
-AI models rely on diverse and representative training data to achieve equitable performance in demographic groups and application contexts.
-Underrepresentation or overrepresentation can induce biased predictions and exacerbate inequities (Raza et al., 2024c).
-Addressing this requires deliberate data collection across geographic, cultural, linguistic, and socio-economic axes.
-Targeted sampling, synthetic data, and augmentation can help fill gaps (Bellamy et al., 2019), yet even well-balanced datasets can embed systematic biases, necessitating ongoing monitoring and recalibration.
-Critics note that truly representative datasets may be infeasible in some domains due to scarcity, legal limits, or privacy concerns (Gebru et al., 2021; Paullada et al., 2021); indiscriminate oversampling can also introduce artifacts that degrade utility.
-  2. (2)
-    Quality, integrity, and timeliness.
-Noisy, incomplete, mislabeled, or stale data degrade reliability; adversarial poisoning can seed harmful patterns (Raza et al., 2025b).
-Robust validation, anomaly detection, and provenance tracking are critical (Das et al., 2025).
+AI models rely on diverse and representative training data to achieve equitable performance across demographic groups and application contexts.
+Underrepresentation or overrepresentation can induce biased predictions and exacerbate inequities.
+For GenAI models trained on web-scale corpora, these challenges are amplified by the opacity of crawl pipelines, the infeasibility of manual curation at trillions of tokens, and the risk that synthetic data from prior model generations contaminates future training sets (model collapse).
+Noisy, incomplete, mislabeled, or stale data also degrade reliability; therefore, adversarial poisoning can seed harmful patterns.
+Robust validation, anomaly detection, and provenance tracking are critical.
 Temporal validity is especially challenging in dynamic environments, where outdated data may encode obsolete norms or regulations, undermining relevance and fairness.
-  3. (3)
-    Ethical data governance.
-Responsible stewardship demands adherence to consent, privacy, and intellectual-property regimes (Mittal et al., 2024), including GDPR (Union, 2018) and emerging AI regulations (Díaz-Rodríguez et al., 2023), alongside sector-specific requirements (Bashir et al., 2024).
+ResGenAI stewardship, therefore, demands adherence to consent, privacy, and intellectual-property regimes, including emerging AI regulations (Table 5), alongside sector-specific requirements.
 Datasheets for datasets (Gebru et al., 2021) and data statements (Paullada et al., 2021) document distributions, licensing, collection methods, and intended uses, fostering transparency and accountability.
-Governance must also address power asymmetries, ensuring consent is informed, revocable, and non-coercive.
+These data challenges do not operate independently; they interact in ways that current literature undertheorizes.
+Representativeness goals can conflict with privacy regimes: expanding demographic coverage may require collecting sensitive attributes that GDPR restricts.
+Quality assurance assumes ground-truth labels, yet for generative training corpora at trillion-token scale, manual verification is infeasible, creating reliance on automated filters (e.g., toxicity classifiers) that themselves embed biases.
+Most critically, the literature treats these as pre-training concerns, but post-deployment data drift, where user interactions reshape fine-tuning distributions, remains largely unaddressed in the works we reviewed.
 Model- and System-Related Challenges
-  1. (1)
-    Explainability, transparency, and accountability.
-Large neural models are often treated as “black boxes,” limiting transparency and accountability in high-stakes settings (Meske et al., 2022).
-A recurring debate concerns a potential interpretability–performance trade-off (Rodolfa et al., 2021).
-Even where progress challenges the inevitability of this trade-off, scaling XAI across domains can add computational overhead (Khan et al., 2025), introduce subjectivity, and yield partial explanations that miss system-level logic.
-In practice, adoption requires balancing transparency, performance, and operational efficiency.
-  2. (2)
-    Bias, fairness, and privacy.
-Models inherit and can amplify data biases, producing disparate performance across groups.
-Fairness interventions interact with privacy constraints, and effects can vary by subgroup and domain (Pessach and Shmueli, 2022).
-Aggressive parity constraints may reduce accuracy in safety-critical contexts or shift harms (Raza and Ding, 2021).
-Differential privacy, federated learning, and secure multi-party computation support privacy (Bashir et al., 2024) but can obscure model logic and complicate audits.
-Achieving both fairness and privacy is context-dependent and often entails trade-offs between Responsible AI principles.
-  3. (3)
-    Robustness, generalization, and misuse.
-Performance often degrades under domain shift, noise, or adversarial perturbations (Chander et al., 2024).
-GenAI introduces misuse risks (deepfakes, misinformation, synthetic identity fraud).
-Mitigations include adversarial training, cross-domain evaluation, and stress testing in simulated deployment (Min et al., 2021), combined with safeguards and policy to deter malicious use.
-  4. (4)
-    Uncertainty calibration and selective abstention.
-Overconfidence amplifies harms from hallucination and biased reasoning.
-Calibration, abstention policies, and fallback to retrieval or human review reduce risk in high-stakes settings; systems should report calibration error and refusal behavior alongside accuracy (Pleiss et al., 2017).
-Security and supply-chain threats.
-As GenAI systems integrate tools (retrieval, code execution, browsing) and ingest third-party content, they inherit software supply-chain risk.
-Indirect prompt injection (Raza et al., 2025b) can subvert instructions via untrusted inputs; tool use can exfiltrate data or trigger harmful actions even when base models appear aligned.
-Threat modeling should include content provenance, capability gating, tool allow-lists, and input sanitization, in addition to adversarial robustness (Liu et al., 2023a).
-Privacy attacks beyond governance claims.
-Beyond policy statements, GenAI models face concrete leakage risks: membership inference and training-data extraction (Liu et al., 2025a).
-Mitigations (differential privacy; memorization audits; curation) trade off with utility and fairness; reporting should include empirical leakage checks, not governance alignment alone.
-Evaluation and assurance gaps.
-Static, one-shot benchmarks rarely capture adaptive adversaries, distribution shifts, or cross-modal failure modes.
-Without construct validity and task/metric documentation (Raza et al., 2024a), headline scores risk Goodhart effects.
-Goodhart’s law has been stated as: ”When a measure becomes a target, it ceases to be a good measure” (Strathern, 1997a).
-Safety claims should be framed as assurance cases with traceable evidence: coverage matrices, failure exemplars, and governance crosswalks, rather than scores alone.
-Post-deployment monitoring and incident response.
-Safety is non-stationary: ecosystems, user behavior, and model updates induce drift; adversarial tactics evolve.
-Continuous evaluation (canary tests, periodic red-teaming, shadow deployments), release gates tied to safety deltas, and incident response (rollback, hot-fixes, disclosure) are essential components of an RAI operating model (Breck et al., 2016).
-Misinformation and Media Integrity
-GenAI can produce plausible but false content (e.g., fabricated citations, numeric errors).
-Retrieval-augmented generation and automated fact-checking help, but adversaries adapt (Raza et al., 2025d).
-In vision and audio, high-fidelity synthesis enables identity theft and fraud (e.g., face swapping, voice cloning), amplifying social harm (Ganguli et al., 2022).
-Watermarking and content-provenance tools aim to bolster integrity (Jiang et al., 2023).
-Tables 3 and 4 summarize the biases observed in text-to-image generation, attributable to training data.
-For instance, the prompt “poor people” in Open AI DALLE-3  [^1] DALLE-3 frequently yielded depictions associated with poverty in African or South Asian contexts, reinforcing region-linked stereotypes; Gemini IMAGEN-3  [^2] Imagen-3also showed bias, with broader but still stereotypical portrayals.
-For “illiterate person”, IMAGEN-3 prominently depicted elderly women, reinforcing inaccurate age–gender associations; DALLE-3 outputs were more varied yet still included traditional stereotypes.
-These findings underscore the need for ethical data curation, bias auditing, and careful deployment.
-Table 3. Adversarial Examples of Image Generation with DALLE-3 (generated by the authors)
-Prompt Generated Images
-Poor person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-a b c
-Illiterate person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-a b c
-Rich person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-a b c
-Table 4. Adversarial Examples of Image Generation with IMAGEN-3 (generated by the authors)
-Prompt Generated Images
-Poor person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-a b c
-Illiterate person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-a b c
-Rich person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-a b c
-Literate person [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-a b c
-Google employee badge [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-a b c
+Model- and system-level risks span explainability, bias amplification, robustness under adversarial conditions, calibration, and post-deployment drift.
+LLMs are often treated as “black boxes,” limiting transparency and accountability in high-stakes settings.
+These models also inherit and can amplify data biases, producing disparate performance across groups.
+Aggressive parity constraints can reduce accuracy in safety-critical contexts.
+In generative settings, these biases manifest not only as allocational harms but as representational harms in generated text and images (e.g., stereotyped depictions, demographic skew, and culturally insensitive outputs), which are harder to enumerate and measure than disparate classification rates.
+Beyond governance claims, GenAI models also face concrete leakage risks including membership inference and verbatim training-data extraction (Liu et al., 2025a).
+Technical mitigations such as differential privacy and memorization audits often trade off with model utility (Bashir et al., 2024), necessitating that reporting includes empirical leakage checks rather than relying on governance alignment alone.
+Current static, one-shot benchmarks rarely capture these dynamics.
+Without construct validity and detailed task documentation, headline scores risk falling victim to Goodhart effects (Strathern, 1997a).
+Safety claims should instead be framed as assurance cases with traceable evidence (e.g., coverage matrices and governance crosswalks), which is a gap we address with our rubric in Section 4.2.1.
+3.2.2. Taxonomy of GenAI Hallucinations
+A pervasive challenge threatening the reliable deployment of Generative AI across high stakes domains is the phenomenon of hallucination.
+This refers to the generation of content that appears structurally fluent and plausible yet is factually incorrect, nonsensical, or unfaithful to source premises (Mamo et al., 2023).
+In the context of robust evaluation and governance, treating hallucinations as a monolithic error class is insufficient for effective risk management.
+They often stem from complex conflicts between a model’s parametric knowledge learned during pretraining and the provided contextual knowledge in prompts or RAG systems.
+To establish a systematic framework for evaluating misinformation specific to ResGenAI, we categorize hallucinations into a five part taxonomy based on technical and contextual failure modes (Banerjee et al., 2024).
+This taxonomy, visualized in Fig.
+5 clarifies the unverifiability risk that current benchmarks in Section 4.2.1 and sectoral applications in Section 5 must address.
+By mapping these types to underlying technical failure mechanisms and providing illustrative actual impacts, this framework enables practitioners to select appropriate detection metrics and mitigation techniques tailored to the specific nature of the generated error.
+Figure 5.
+A comprehensive taxonomy illustrating five distinct types of generative AI hallucinations, detailing their underlying technical failure mechanisms and providing illustrative real world impacts and cases.
 3.2.3. Regulatory and Broader Sociotechnical Challenges
 Regulatory challenges concern the design, implementation, and enforcement of frameworks that make GenAI ethical, accountable, privacy-preserving, and safe across domains and jurisdictions.
-We group key issues below, noting that user behavior and deployment context mediate how risks materialize.
-  1. (1)
-    Copyright and intellectual property.
-Using copyrighted material for training raises questions of ownership, licensing, and fair use (Kenthapadi et al., 2023).
-Reported reproduction and memorization incidents expose gaps in current regimes and motivate updated guidance (Freeman et al., 2024; Pasquale and Sun, 2024; Zirpoli, 2023b).
-  2. (2)
-    Safety and alignment obligations.
-GenAI can produce toxic or biased content, undermining trust (Weidinger et al., 2021).
-Regulatory strategies include bias-aware data curation, real-time detection/mitigation (Tango and Botta, 2013), independent audits (Laine et al., 2024), and adherence to fairness/accountability frameworks (Pagano et al., 2023).
-Incorporating user feedback and documenting limitations improves sociotechnical alignment.
-  3. (3)
-    Accountability, transparency, and record-keeping.
-Assigning responsibility is difficult when systems hallucinate or propagate training-set harms (Elliott et al., 2024).
-The EU AI Act, for example, proposes risk tiering and obligations for robustness, transparency, logging, and post-market monitoring (Outeda, 2024).
+User behavior and deployment context mediate how risks materialize.
+The EU AI Act proposes risk tiering and obligations for robustness, transparency, logging, and post-market monitoring (Outeda, 2024).
 Provenance transparency, especially for copyrighted or sensitive data, is critical.
-  4. (4)
-    Automated decision-making in high-stakes domains.
-Applications in hiring, credit, and healthcare raise concerns about discrimination and due process.
-Without recourse mechanisms and domain-specific guardrails, GenAI recommendations can entrench historical inequities.
-  5. (5)
-    Explainability mandates vs.
-model complexity.
-Legal and sectoral expectations for interpretability in high-risk settings meet practical limits as models scale (e.g., GPT-4), creating tension between compliance and performance (Slattery et al., 2024).
-Dynamic risks (e.g., election misinformation, deepfakes) further require evolving standards.
-  6. (6)
-    Codes of practice and operational maturity.
-Inconsistent processes for dataset governance, evaluation, deployment, and monitoring exacerbate bias and privacy risks (Slattery et al., 2024).
-Pragmatic, auditable codes of practice can harmonize controls and enable proportionate assurance across sectors.
-  7. (7)
-    Human oversight, recourse, and accessibility.
-Requirements for meaningful human oversight, user recourse, and disability access are unevenly specified (Raza et al., 2025a).
-Over-reliance on “assistive” GenAI can mask failure modes; documenting escalation paths to human review and accessible interaction patterns is essential in high-stakes contexts.
-  8. (8)
-    International alignment, conformity assessment, and enforcement.
-Divergent regimes (risk-management frameworks, sectoral law, emerging standards) create fragmentation across jurisdictions.
+Legal and sectoral expectations for interpretability in high-risk settings meet practical limits as models scale, creating tension between compliance and performance (Slattery et al., 2024); dynamic risks such as election misinformation and deepfakes further require evolving standards.
+Requirements for meaningful human oversight, user recourse, and disability access are unevenly specified.
+Over-reliance on assistive GenAI can mask failure modes; documenting escalation paths to human review and accessible interaction patterns is essential in high-stakes contexts.
+Divergent regimes, such as risk-management frameworks, sectoral law, and emerging standards, create fragmentation across jurisdictions.
 Conformity assessment, post-market monitoring, and enforcement capacity vary widely, and evidence portability remains an open problem for cross-border deployments.
-  9. (9)
-    Environmental and compute sustainability.
-Training and serving large large models have material energy and carbon costs, with equity implications for participation in research and deployment.
-Reporting energy/emissions, adopting efficiency techniques (distillation, sparsity), and compute-proportional evaluation can mitigate externalities (Khan et al., 2025).
-  10. (10)
-    Dual-use and malicious applications.
-The same GenAI capabilities that enable innovation can be weaponized for misinformation, cybercrime, biothreats, or automated exploitation.
-Regulatory regimes struggle to anticipate dual-use risks without stifling beneficial innovation, creating tension among openness, export controls, and responsible disclosure (Weidinger et al., 2022).
-Bridge to our evidence framework.
-Taken together, these challenges motivate proportionate, auditable controls and transparent evidence.
-In Section 4.2 we operationalize this via our comparative rubric, and provides a governance crosswalk mapping risks to evaluative evidence and compliance narratives.
+Inconsistent processes for dataset governance, evaluation, deployment, and monitoring further exacerbate bias and privacy risks (Slattery et al., 2024); pragmatic, auditable codes of practice can harmonize controls and enable proportionate assurance across sectors.
+3.3. Emergent Threats
+While Section 3.2 details established technical risks, we identify further ResGenAI emergent threats that represent a second generation risk landscape.
+These threats arise from the intersection of increasing model scale, autonomous capabilities, and decentralized deployment patterns.
+3.3.1. Adversarial Technical Evolution
+Recent evidence suggests that text embeddings can be used to reconstruct original, sensitive training data through representation inversion (Liu et al., 2025a).
+Unlike direct data extraction, these attacks exploit the vector space to recover high fidelity personal data, rendering traditional anonymization insufficient for models with high memorization rates.
+Concurrently, the landscape of jailbreaking is shifting from manual roleplay to automated, optimization driven techniques.
+Using gradient based optimization, adversaries can craft “universal” adversarial suffixes that destabilize safety layers at speed, often achieving success rates that manual engineering cannot reach (Samvelyan et al., 2024).
+3.3.2. Sociotechnical and Ecosystem Risks
+The rise of decentralized, open source deployments introduces shadow AI risks, where models are fine tuned to remove safety guardrails outside of institutional oversight.
+Furthermore, we identify the emergence of synthetic identity swarms that are coordinated networks of generative agents that create synthetic consensus cascades to shift human beliefs at an unprecedented scale (Lucas et al., 2023).
+These threats are compounded by broader sociotechnical concerns: training and serving large models incur material energy and carbon costs with equity implications for global research participation (Khan et al., 2025), while regulatory regimes struggle to anticipate dual use risks without stifling innovation (Weidinger et al., 2022).
+This complexity creates a “hydra effect”, where mitigating one failure mode, such as toxicity, may inadvertently expose others, including increased sycophancy or declining creative reasoning (Bai et al., 2022).
+While techniques like RLHF (Kaelbling et al., 1996) can suppress harmful surface level patterns, they often fail to remove underlying harmful representations.
+3.3.3. Agentic Specific Failure Vectors
+As GenAI evolves toward agentic workflows, where models autonomously plan, select tools, and execute actions and the attack surface expands beyond linguistic outputs to operational environments (Lu et al., 2024a).
+Unlike single turn interactions, agentic pipelines introduce multi step dependencies that can amplify failure modes.
+Table 2 systematizes these vectors to guide system level evaluation protocols.
+Table 2. Agentic Specific Failure Vectors: Risks in Autonomous Tool Use and Planning
+Failure Vector Technical Mechanism Governance and Security Impact
+Tool Misuse Unintended or unauthorized API calls resulting from prompt injection or planning errors (Lu et al., 2024a).
+Leads to financial loss, unauthorized data modification, or service disruption in production environments.
+Recursive Hallucination Compound errors where a hallucination in one step serves as the factual basis for subsequent planning (Alansari and Luqman, 2025).
+Destabilizes reasoning chains, making final autonomous actions entirely ungrounded.
+Authentication Leakage Inadvertent exposure of system keys, tokens, or PII during tool calling or log generation (Barth and De Jong, 2017).
+Creates critical security vulnerabilities that allow adversaries to hijack agentic permissions.
+Plan Hijacking Adversarial manipulation of intermediate Chain of Thought to redirect the goal of the autonomous system (Yu et al., 2025).
+Undermines human oversight by masking malicious intent within seemingly logical execution steps.
+This systemic fragility motivates the use of proportionate, auditable controls and transparent evidence.
+In Section 4.2, we operationalize this via our comparative rubric and provide a governance crosswalk mapping risks to evaluative evidence and compliance narratives.
 4. Results
-In this section, we present the results of our structured literature review on RAI.
+In this section, we present the results of our structured literature review on ResGenRAI.
 We first report the thematic findings that emerged from our synthesis.
 Then, we provide a comparative assessment of widely used AI safety benchmarks against our governance-aligned rubric, followed by an analysis of key performance indicators (KPIs), explainability methods, AI-ready testbeds, and applications across domains.
-4.1. Thematic Findings
-Figure 4 summarizes thematic distribution, temporal trends (including 2025), co-occurrence patterns, publication-type mix, and keyword salience.
-Figure 3 provides a compact view of annual volume and keyword evolution.
-Figure A.1 maps the broader concepts spanning explainable AI, ethics, sectoral applications, advanced models (including large vision–language models), privacy, and regulatory frameworks.
-Tables 2 list representative studies included in this review.
-Figure 3. Trends in Responsible AI Publications and Keyword Evolution Over Time. (A) Number of publications in RAI by Year. (B) Frequency of keywords by year.:
+The acronyms used in this work are reported in Appendix Table 1.
+Figure 6. Trends in Responsible AI Publications and Keyword Evolution Over Time. (A) Number of publications in RAI by Year. (B) Frequency of keywords by year
+Publications Trends in RAI by Year
 (a) Top themes
@@
 (f) Publications per year
-Figure 4. Thematic analysis overview on RAI literature.
-Our thematic synthesis highlights that RAI scholarship in the post-GPT period clusters around five dominant themes: (i) safety and benchmarking, with a surge of multimodal evaluation suites and red-teaming methods; (ii) fairness, bias, and governance, where debates center on aligning technical audits with emerging laws such as the EU AI Act and NIST AI RMF; (iii) explainability and transparency, underscoring the tension between interpretability and performance in large foundation models; (iv) privacy and security, particularly membership inference and data extraction risks; and (v) sectoral and sociotechnical applications, spanning healthcare, finance, education, and agriculture.
-Temporal analysis shows rapid growth from late 2022 onward, with multimodal and agentic systems gaining salience by 2024–2025.
+Figure 7. Thematic analysis overview on RAI literature
+4.1. Thematic Findings
+Figure 7 summarizes thematic distribution, temporal trends, co-occurrence patterns, publication-type mix, and keyword salience.
+Figure 6 provides a compact view of annual volume and keyword evolution.
+Appendix Table 2 list representative studies included in this review.
+Our thematic synthesis highlights that ResGenRAI scholarship in the post-GPT period clusters around five dominant themes: (i) safety and benchmarking, with a surge of multimodal evaluation suites and red-teaming methods; (ii) fairness, bias, and governance, where debates center on aligning technical audits with emerging laws such as the EU AI Act and NIST AI RMF and others (Table 5); (iii) explainability and transparency, underscoring the tension between interpretability and performance in large foundation models; (iv) privacy and security, particularly membership inference and data extraction risks; and (v) sectoral and sociotechnical applications, spanning healthcare, finance, education, and agriculture.
+Our temporal analysis shows rapid growth from late 2022 onward, with multimodal and agentic systems gaining salience by 2024–2026.
 These themes frequently co-occur, for example where explainability is invoked to mitigate fairness concerns, or where privacy and governance frameworks intersect in deployment contexts.
-Together, they map a research terrain that is expanding in scope but still uneven in coverage, with strong emphasis on bias/toxicity benchmarks but weaker treatment of deepfakes, sustainability, and system-level failures.
 4.2. Benchmarks: Comparative Assessment
-This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in the Tables 6 and 8.
+This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in Tables 4 and 6.
 4.2.1. Comparative Assessment Protocol
-We assess widely used AI safety benchmarks against a ten–item rubric aligned to risk families and governance needs: C1 Bias/Discrimination, C2 Toxicity, C3 Security/Robustness (including adversarial), C4 Misinformation/Disinformation, C5 Deepfakes/Media Integrity, C6 Privacy (membership/extraction), C7 System–level failure, C8 Malicious–actor realism (red–team/adaptive), C9 Metric validity/documentation, and C10 Governance alignment (e.g., NIST AI RMF, EU AI Act).
-The ten criteria (also given in Appendix B) reflect risk families that recur across governance instruments and risk repositories, including the NIST AI Risk Management Framework, the EU AI Act, and ISO/IEC 42001, as well as consolidated mappings such as the MIT AI Risk Repository (nis, 2024; Outeda, 2024; Systems, 2021; Slattery et al., 2024).
-We scored each benchmark after reviewing the official paper, public documentation, and widely cited usage reports; disagreements were resolved by discussion and, when needed, a third reviewer.
-Each benchmark is scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, 1 = partial, 2 = strong). Equal weights are used by default. The normalized score is
-$S(B)=\frac{\sum_{i=1}^{10}w_{i}\,s_{i}(B)}{2\sum_{i=1}^{10}w_{i}},\quad w_{i}>0.$
+Our benchmark assessment follows a structured qualitative expert appraisal with ordinal scoring, a well-established approach for rubric-based evaluation in computer science survey research (Harzing and Alakangas, 2016).
+It is neither a quantitative meta-analysis (no effect sizes or pooled statistics are computed) nor an unstructured qualitative review.
+Instead, we employ codebook-guided independent scoring that yields ordinal coverage indicators (0/1/2 per criterion), which are then normalized into a composite index.
+The resulting scores characterize the breadth of risk-surface coverage and documentation quality across benchmarks (they are not cardinal performance measures and should not be interpreted as rankings of overall safety).
+This approach is appropriate because the underlying evidence, e.g., benchmark papers, dataset cards, public documentation, and is heterogeneous in format and scope, precluding formal quantitative synthesis.
+Rubric design.
+We assess widely used AI safety benchmarks against a ten-item rubric aligned to risk families and governance needs: C1: Bias/Discrimination, C2 :Toxicity, C3 : Security/Robustness (including adversarial), C4: Misinformation/Disinformation, C5 : Deepfakes/Media Integrity, C6 : Privacy (membership/extraction), C7 : System-level failure, C8 : Malicious-actor realism (red-team/adaptive), C9 : Metric validity/documentation, and C10: Governance alignment (e.g., NIST AI RMF, EU AI Act).
+The ten criteria (also given in Appendix A.2) reflect risk families that recur across governance instruments and AI risk repositories.
+Scoring procedure and reliability.
+Each benchmark was scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, $1$ = partial, $2$ = strong) by two independent reviewers, following the codebook in Appendix A.2.
+The codebook specifies per-criterion construct definitions, qualifying evidence types, scoring anchors, and a disambiguation rule that assigns tasks spanning multiple risk families to the primary stated objective to prevent double-counting.
+Evidence was drawn according to a priority ordering: (i) official paper and documentation; (ii) dataset or model cards and maintained leaderboards; (iii) widely cited third-party reports.
+Disagreements were resolved through structured discussion and, where necessary, adjudication by a third reviewer.
+Equal weights are used by default.
+The normalised score is
+(1) $S(B)=\frac{\sum_{i=1}^{10}w_{i}\,s_{i}(B)}{2\,\sum_{i=1}^{10}w_{i}},\quad w_{i}>0.$
 This yields values in $[0,1]$, where 0 indicates no coverage and $1$ indicates full coverage across all criteria.
-We also provide a governance crosswalk indicating which requirements each benchmark can plausibly evidence (Tables 6 and 8).
-4.2.2. Comparative Assessment Findings
-Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs.
-Table 5 presents the risk surface matrix for criteria C1–C8.
-Criteria C9 (metric validity and documentation) and C10 (governance alignment) are summarized in Table 8.
-Table 5. Mapping benchmarks to risk surface areas. ✓ = supported; $\triangle$ = partial; – = no explicit coverage (from public papers and documentation).
-Benchmark Bias/Discr. Toxicity Security Misinfo/Disinfo Deepfakes Privacy System failure Malicious actors
-HarmBench (Mazeika et al., 2024) $\triangle$
-SALAD Bench (Li et al., 2024a) $\triangle$ $\triangle$ $\triangle$
-ALERT (Tedeschi et al., 2024) $\triangle$ $\triangle$
-Risk Taxonomy (Cui et al., 2024) $\triangle$ $\triangle$ $\triangle$
-TrustGPT (Huang et al., 2023b)
-RealToxicityPrompts (Gehman et al., 2020)
-DecodingTrust (Wang et al., 2023a) $\triangle$ $\triangle$ $\triangle$
-SafetyBench (Zhang et al., 2024) $\triangle$ $\triangle$ $\triangle$
-MM–SafetyBench (Liu et al., 2024) $\triangle$ $\triangle$ $\triangle$ $\triangle$
-XSTest (Röttger et al., 2024) $\triangle$
-Rainbow Teaming (Samvelyan et al., 2024) $\triangle$ $\triangle$
-MLCommons AI Safety v0.5 (Vidgen and et al., 2024) $\triangle$ $\triangle$ $\triangle$
-HELM (Liang et al., 2023) $\triangle$ $\triangle$
-SHIELD (Shi et al., 2024)
-BIG-bench (bench authors, 2023) $\triangle$ $\triangle$ $\triangle$ $\triangle$
-BEADs (Raza et al., 2024b) $\triangle$ $\triangle$
-PrivLM-Bench (Li et al., 2024c) $\triangle$ $\triangle$
-PrivacyLens (Shao and et al., 2024)
-HumaniBench (Raza et al., 2025a) $\triangle$ $\triangle$ $\triangle$
-Table 6. Comparative assessment rubric for AI safety benchmarks (0 = no coverage, 1 = partial, 2 = strong). Scores use equal weights and are normalized to $[0,1]$.
+To verify that our conclusions are not an artefact of the equal-weight assumption, we conducted a sensitivity analysis under alternate weight profiles that prioritise high-stakes criteria (e.g., C3 Security, C6 Privacy); the resulting Spearman rank correlation exceeded $\rho>0.85$, indicating that comparative breadth is a stable indicator of evaluative maturity.
+Table 3.
+Mapping benchmarks to risk surface areas.
+$\checkmark$ = supported; $\triangle$ = partial; – = no explicit coverage (from public papers and documentation).
+This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 4.
+Benchmark Bias/Discrimination Toxicity Security Misinfo Deepfakes Privacy System fail. Malicious
+HarmBench (Mazeika et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+SALAD-Bench (Li et al., 2024a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$
+ALERT (Tedeschi et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Risk Taxonomy (Cui et al., 2024) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
+TrustGPT (Huang et al., 2023b) $\checkmark$ $\checkmark$
+RealToxicityPrompts (Gehman et al., 2020) $\checkmark$
+DecodingTrust (Wang et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
+SafetyBench (Zhang et al., 2024) $\checkmark$ $\checkmark$ $\triangle$
+MM-SafetyBench (Liu et al., 2024) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\checkmark$
+XSTest (Röttger et al., 2024) $\triangle$ $\checkmark$ $\triangle$
+Rainbow Teaming (Samvelyan et al., 2024) $\triangle$ $\checkmark$ $\triangle$ $\checkmark$ $\checkmark$
+MLCommons AI Safety v0.5 (Vidgen and et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+HELM (Liang et al., 2023) $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$
+SHIELD(Shi et al., 2024) $\triangle$ $\triangle$ $\checkmark$
+BIG-bench (bench authors, 2023) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
+PrivLM-Bench (Li et al., 2024b) $\checkmark$
+PrivacyLens (Shao and et al., 2024) $\triangle$ $\checkmark$
+HumaniBench (Raza et al., 2025) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
+Table 4.
+Comparative assessment rubric for AI safety benchmarks (0 = no coverage, 1 = partial, 2 = strong).
+Scores use equal weights and are normalized to $[0,1]$.
+Per-criterion scoring anchors and evidence rules are provided in Appendix A.2.
+This table reports scores across all ten criteria (C1–C10), including C9 (metric validity/documentation) and C10 (governance alignment).
 Benchmark C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 Score
@@
 HELM (Liang et al., 2023) 1 1 1 1 0 0 0 0 1 1 0.30
-PrivLM Bench (Li et al., 2024c) 0 0 0 0 0 2 0 0 1 1 0.20
+PrivLM Bench (Li et al., 2024b) 0 0 0 0 0 2 0 0 1 1 0.20
 PrivacyLens (Shao and et al., 2024) 0 0 0 0 0 2 0 0 1 0 0.15
@@
 SHIELD (Shi et al., 2024) 0 0 1 0 2 0 0 0 1 0 0.20
-BEADs (Raza et al., 2024b) 1 1 0 2 0 0 0 0 1 0 0.25
 MFG for GenAI (Foundation, 2024) 1 0 1 1 0 0 0 1 0 0 0.20
@@
 TrustGPT (Huang et al., 2023b) 1 1 0 0 0 0 0 0 1 0 0.20
-HumaniBench (Raza et al., 2025a) 2 1 0 1 0 1 1 0 1 1 0.40
-Criteria (C1 to C10).
-C1 Bias/Discrimination; C2 Toxicity; C3 Security/Robustness; C4 Misinformation/Disinformation; C5 Deepfakes/Media Integrity; C6 Privacy (membership/extraction); C7 System level failure (tool/agentic); C8 Malicious actor realism (red team/adaptive); C9 Metric validity/documentation; C10 Governance alignment.
+HumaniBench (Raza et al., 2025) 2 1 0 1 0 1 1 0 1 1 0.40
+Criteria (C1 to C10) C1 Bias/Discrimination; C2 Toxicity; C3 Security/Robustness; C4 Misinformation/Disinformation; C5 Deepfakes/Media Integrity; C6 Privacy (membership/extraction); C7 System level failure (tool/agentic); C8 Malicious actor realism (red team/adaptive); C9 Metric validity/documentation; C10 Governance alignment.
 Normalization (equal weights): $S(B)=\frac{\sum_{i=1}^{10}s_{i}(B)}{20}$ with $s_{i}\in\{0,1,2\}$.
-Table 7. Governance-based Frameworks and Technical AI Tools used in RAI Practices.
+Table 5. Governance-based Frameworks and Technical AI Tools used in RAI Practices
 Category Governance-based Frameworks Technical AI Solutions
-Fairness & Bias Mitigation ISO/IEC TR 24027:2021 (3, 2021) : technical report on AI and data quality.
+Fairness & Bias Mitigation ISO/IEC TR 24027:2021 (ISO/IEC, 2021a): technical report on AI and data quality.
 NIST SP-1270 (Schwartz et al., 2022): framework for identifying and managing bias in AI.
-Framework for increasing diversity & inclusion in AI education, research, & policy: AI4ALL (AI4ALL, 2024) Evaluate for model fairness via visualizations: Google What-if Tool (Wexler et al., 2020) AI Fairness 360 Toolkit by IBM provides tools for evaluation and mitigation of bias: AIF360 (Bellamy et al., 2018b) Assess and enhance fairness of ML models: Microsoft Fairlearn (Weerts et al., 2023)
+AI4ALL (AI4ALL, [n.
+d.]): framework for increasing diversity & inclusion in AI education, research, & policy.
+African Union (AU) Data Policy Framework (African Union, 2022): promotes equitable, inclusive AI and data governance across member states.
+Singapore IMDA & PDPC’s AI Governance Framework (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): addresses fairness, human-centricity, and non-discrimination in AI decision-making.
+Google What-If Tool (Wexler et al., 2020): evaluate for model fairness via visualizations.
+IBM AI Fairness 360 (AIF360) (Bellamy et al., 2019): tools for evaluation and mitigation of bias.
+Microsoft Fairlearn (Weerts et al., 2023): assess and enhance fairness of ML models.
 Interpretability & Explainability IEEE P2976 (Polemi et al., 2024): standards for Explainable Artificial Intelligence.
 IEEE P2894 (Committee, 2024): guidelines for implementing explainable AI technologies.
-Explaining Decisions Made with AI: UK Info.
-Commissioner’s Office & The Alan Turning Institute (Leslie, 2022) Model interpretability and explainability tool: LIT (Tenney et al., 2020) Toolkit for measuring explainability using various models like LIME and SHAP: InterpretML (Sameki et al., 2020) Provide simple AI generated explanations: ELI5 (Deepgram, 2024) AI Explainability 360 toolkit by IBM provides tools for explainability of ML models: AIX360 (Arya et al., 2019)
-Transparency & Accountability U.S Government Accountability Office: AI Accountability Framework (U.S.
-Government Accountability Office, 2021) Organisation for Economic Co-operation and Development (OECD): AI Principle of Accountability ((, OECD) Evaluate, test and monitor ML models: Evidently (EvidentlyAI, 2024) Improve model transparency, reproducibility, and robustness: MLBench (MLBench, 2020)
-Privacy, Safety & Security NASA (Smith et al., 2020): hazard contribution modes of ML in space applications.
+UK ICO & The Alan Turing Institute: Explaining Decisions Made with AI (Information Commissioner’s Office and The Alan Turing Institute, 2022).
+China CAC Algorithm Recommendation Provisions (Cyberspace Administration of China, 2022b): mandates transparency and explainability requirements for algorithmic recommendation systems.
+Learning Interpretability Tool (LIT) (Google PAIR, [n.
+d.]): model interpretability and explainability.
+InterpretML (Microsoft Research): toolkit for measuring explainability using LIME and SHAP.
+IBM AI Explainability 360 (AIX360) (Arya et al., 2019): tools for explainability of ML models.
+Transparency & Accountability U.S.
+Government Accountability Office: AI Accountability Framework (U.S.
+Government Accountability Office, 2021).
+OECD AI Principle of Accountability (Organisation for Economic Co-operation and Development, 2019).
+China Measures for the Management of Generative AI Services (CAC, 2023) (Cyberspace Administration of China, 2023): requires service providers to disclose AI-generated content, maintain logs, and be accountable for model outputs.
+Brazil’s Proposed AI Act (PL 2338/2023) (Senado Federal do Brasil, 2023): mandates impact assessments, transparency obligations, and accountability measures for high-risk AI systems.
+SDAIA National AI Ethics Principles (Saudi Data and Artificial Intelligence Authority, 2023a): establishes accountability, auditability, and transparency requirements for AI deployments in Saudi Arabia.
+Evidently (Evidently AI, [n.
+d.]): evaluate, test and monitor ML models.
+MLBench (MLBench Team, [n.
+d.]): improve model transparency, reproducibility, and robustness.
+AI Verify (Infocomm Media Development Authority ([n.
+d.]), IMDA): Singapore’s IMDA-developed AI governance testing framework.
+Privacy, Safety & Security NASA (Smith and Kearns, 2020): hazard contribution modes of ML in space applications.
 IEEE P7009 (Farrell et al., 2021): standards for the fail-safe design of autonomous systems.
-ISO/IEC TS 27022:2021 (Systems, 2021): information security management systems guidelines.
-NIST AI 100-2 E2023 (Oprea and Vassilev, 2023) : specific cybersecurity practices for AI.
-UK Government AI Cybersecurity Code of Practice (Department of Science, Innovation & Technology, 2024).
-MIT Risk repo (Slattery et al., 2024): Database documenting AI risks and safety gaps.
-Google’s Secure AI Framework SAIF (Inc., 2024) Context-aware toxicity tracker: Unitary (Unitary, 2023) Adversarial Robustness Toolbox: ART Audit privacy of ML models: Privacy Meter Google Differential Privacy libraries: DP Privacy-preserving data analysis and ML toolkit Secret Flow Tensorflow privacy (TensorFlow, 2023) Privacy Enhancing Technology (PET) on Synthetic Data Generation: Betterdata.AI (inc., 2024)
-Ethical Guidelines & Compliance European General Data Protection Regulation: GDPR (Union, 2018) Canadian AI and Data Act: ISED (of Canada, 2023a) Canadian Voluntary Code of Conduct on the Responsible Development and Management of Advanced GenAI Systems: ISED (of Canada, 2023b) American Executive Order on Safe, Secure, and Trustworthy Artificial Intelligence: AEO (House, 2023) Google responsible AI practices (AI, 2023) Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, 2019) Microsoft’s Responsible AI Toolbox (Microsoft, 2024) Demonstrate trust, measure and manage risk, and compliance: OneTrust (OneTrust, 2023) World Health Organization (WHO): Ethics and governance of AI for health Guidance on large multi-modal models (Organization et al., 2024)
-Monitoring & Auditing Index of harms realized in the real world by AI systems: Incident Database (Database, 2023) Guidance provided by UK Information Commissioner’s Office AI Auditing Framework (UK Government Information Commissioner’s Office, 2020) Audit classification models for intersectional bias: FairVis (Cabrera et al., 2019) Bias and fairness audit toolkit: Aequitas (for Data Science and of Chicago, 2018) Auditing of bias in generalized ML applications: Audit AI (Inc., 2020)
-Table 8. Governance alignment crosswalk (illustrative). ✓= supported evidence; $\triangle$= partial; – = not covered.
+ISO/IEC TS 27022:2021 (ISO/IEC, 2021b): information security management systems guidelines.
+NIST AI 100-2 E2023 (Oprea and Vassilev, 2023b): specific cybersecurity practices for AI.
+UK Government AI Cybersecurity Code of Practice (Department of Science, 2024).
+MIT AI Risk Repository (Slattery et al., 2024): database documenting AI risks and safety gaps.
+Google’s Secure AI Framework (SAIF) (Google, 2023).
+Brazil LGPD (da Republica, Brasil): data protection law governing privacy, data processing, and individual rights in AI-driven systems.
+Nigeria Data Protection Act (NDPA, 2023) (Agency, 2023): establishes data privacy safeguards for AI systems processing personal data of Nigerian citizens.
+China Personal Information Protection Law (of China, 2021): comprehensive data privacy regulation for AI systems.
+UAE Federal Decree-Law No.
+45 on Personal Data Protection (Government, 2021): governs data privacy and individual rights in AI-driven data processing within the UAE.
+Unitary (Ltd., [n.
+d.]): context-aware toxicity tracker.
+Adversarial Robustness Toolbox (ART) (Research, [n.
+d.]).
+Privacy Meter (Privacy and Lab, [n.
+d.]): audit privacy of ML models.
+Google Differential Privacy Libraries (Google, [n.
+d.]).
+SecretFlow (Group, [n.
+d.]): privacy-preserving data analysis and ML toolkit.
+TensorFlow Privacy (Google, [n.
+d.]).
+Betterdata.AI (inc., 2024): PET on synthetic data generation.
+Ethical Guidelines & Compliance European General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016).
+Canadian AI and Data Act (Innovation, Science and Economic Development Canada, 2023a).
+Canadian Voluntary Code of Conduct on the Responsible Development and Management of Advanced GenAI Systems (Innovation, Science and Economic Development Canada, 2023b).
+American Executive Order on Safe, Secure, and Trustworthy AI (House, 2023).
+China’s New Generation AI Development Plan (AIDP, 2017) & AI Governance Principles (2019) (State Council of the People’s Republic of China, 2017; National New Generation AI Governance Expert Committee (2019), China): outline China’s national strategy for AI development with ethical principles for security, fairness, and human-centric AI.
+UAE National AI Strategy 2031 (United Arab Emirates Government, [n.
+d.]): comprehensive national strategy establishing ethical AI goals and responsible AI deployment.
+SDAIA Responsible AI Guidelines (Saudi Data and Artificial Intelligence Authority, 2023b): ethical guidelines for responsible AI use and regulatory compliance in Saudi Arabia.
+Singapore Model AI Governance Framework (2nd Ed., IMDA) (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): practical guidance on responsible AI deployment covering internal governance and human oversight.
+AU Continental AI Policy Framework (2024) (African Union, 2024): continent-wide framework for ethical, inclusive, and human rights-respecting AI governance.
+Colombia’s AI Ethics Framework (Ministerio de Tecnologías de la Información y las Comunicaciones (2019), MinTIC) & Chile National AI Policy (Ministerio de Ciencia, Tecnología, Conocimiento e Innovación, 2021): regional strategies for responsible AI adoption.
+Google Responsible AI Practices (Google, [n.
+d.]).
+Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, [n.
+d.]).
+Microsoft’s Responsible AI Toolbox (principles and approach, 2024).
+OneTrust (OneTrust LLC, [n.
+d.]): demonstrate trust, measure and manage risk, and compliance.
+WHO: Ethics and Governance of AI for Health (World Health Organization, 2024).
+Project Moonshot (AI Verify Foundation, 2024): open-source LLM evaluation and red-teaming toolkit to benchmark GenAI systems against safety and ethical compliance standards.
+Monitoring & Auditing AI Incident Database (Responsible AI Collaborative, [n.
+d.]): index of harms realized in the real world by AI systems.
+UK ICO AI Auditing Framework (UK Information Commissioner’s Office (2020), ICO).
+China CAC Measures for the Management of Deep Synthesis Internet Information Services (2022) (Cyberspace Administration of China, 2022a): requires labeling, monitoring, and auditing of AI-generated synthetic media.
+Smart Africa Alliance AI Audit Guidelines (Smart Africa, 2023): monitoring and audit guidance for AI systems in African public services.
+IDB AI Governance Toolkit (Inter-American Development Bank (2023), IDB): monitoring and assessment tools for responsible AI deployment across Latin American and Caribbean nations.
+FairVis (Poloclub, [n.
+d.]): audit classification models for intersectional bias.
+Aequitas (Center for Data Science and Public Policy, [n.
+d.]): bias and fairness audit toolkit.
+Audit-AI : auditing of bias in generalized ML applications.
+Table 6.
+Governance alignment crosswalk (illustrative).
+✓= supported evidence; $\triangle$= partial; –= not covered.
+A ✓ indicates plausible evidentiary support rather than a statement of legal compliance; see Appendix B for the non-compliance note.
+This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 4.
 Requirement
@@
 HumaniBench
-NIST: Privacy & Security (nis, 2020a, b) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
+NIST: Privacy & Security (of Standards and, NIST, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
 EU AI Act: Robustness/Resilience  [^1] https://artificialintelligenceact.eu/article/15 $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
 EU AI Act: Transparency/Record-keeping  [^2] https://artificialintelligenceact.eu/article/12/ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
-Multimodal Deepfake Risk (c2p, 2024) $\triangle$
-Adaptive/Red-team Evidence (nis, 2023) $\triangle$ $\triangle$ $\triangle$ $\triangle$
-Abbreviations. RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
+Multimodal Deepfake Risk (C2PA, 2024) $\triangle$
+Adaptive/Red-team Evidence (of Standards and, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$
+Abbreviations RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
+4.2.2. Comparative Assessment Findings.
+Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs.
+Table 3 presents the risk-surface matrix for criteria C1–C8.
+The full rubric scores, including C9 (metric validity and documentation) and C10 (governance alignment), are given in Table 4.
+Table 6 then provides a finer-grained governance crosswalk mapping each benchmark to specific regulatory requirements.
 Risk-surface coverage.
-Table 5 presents the representative benchmarks cover eight salient risk families (bias/discrimination, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure, and malicious-actor realism).
-Entries are based on public papers and documentation and reflect explicit task/metric coverage rather than incidental behavior.
-The results in Table 5 shows that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens).
+Table 3 presents the representative benchmarks across eight salient risk families (C1–C8).
+Entries are based on public papers and documentation and reflect explicit task/metric coverage rather than incidental behaviour.
+The results in Table 3 show that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens).
 Deepfake/media integrity is sparsely addressed (notably SHIELD), and only a few efforts incorporate adaptive adversaries or richer system context (Rainbow Teaming, ALERT).
@@
 We operationalize the ten-criterion rubric in Section 4.2.1 to produce a normalized $[0,1]$ score per benchmark.
-The aim is not to rank “best overall”, but to indicate breadth and governance-ready evidence across risk families and evaluation qualities (metric validity, documentation, policy alignment).
-The results in Table 6 shows that Rainbow Teaming attains the highest breadth score (0.50), reflecting strong coverage of security under adaptive attack and red-team realism.
+The aim is not to rank “best overall,” but to indicate breadth and governance-ready evidence across risk families and evaluation qualities (metric validity, documentation, policy alignment).
+The results in Table 4 show that Rainbow Teaming attains the highest breadth score (0.50), reflecting strong coverage of security under adaptive attack and red-team realism.
 Risk Taxonomy and HumaniBench (0.40 each) combine multi-risk awareness with governance-facing structure.
@@
 Table 5 situates the analysis within widely used frameworks and tools.
-Table 8 then maps each benchmark to selected governance requirements spanning NIST AI RMF privacy/security controls, EU AI Act robustness and transparency obligations, multimodal deepfake risk, and evidence for adaptive/red-team evaluation.
-A ✓indicates plausible evidentiary use; $\triangle$ indicates partial or indirect support; indicates no explicit coverage in public artifacts.
+Table 6 then maps each benchmark to selected governance requirements and risks such as multimodal deepfake risk, and evidence for adaptive/red-team evaluation.
 The results show that only a subset of suites provide tangible evidence for privacy/security controls (e.g., PrivLM-Bench).
@@
 Transparency/record-keeping signals are uneven outside of meta-benchmarks (HELM, MLCommons), which can hinder audit trails under the EU AI Act.
+Figure 8.
+Responsible GenAI governance landscape and linkage to our evaluation criteria.
+Technical factors correspond primarily to rubric items C1–C7 (fairness/bias, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure).
+Risk & assurance processes (anticipation, responsiveness, reflexivity, inclusion, knowledge management) support adaptive testing (C8), metric validity/record-keeping (C9) and Legal instruments (e.g., EU AI Act, NIST AI RMF) ground C10 (governance alignment).
+Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks, motivating the gaps observed in Tables 46
 4.2.3. Proposed Improvements to AI Safety Benchmarks
-The following suggestions are offered as constructive opportunities to complement existing benchmarks and to support broader governance use.
-To address the capability gaps identified above, we suggest:
-  1. (i)
-    Data quality and transparency.
-Encourage diverse and well documented corpora across regions, languages, and demographics, with dataset cards or data statements that capture provenance, consent and licensing, and curation criteria to enhance traceability and accountability (Eriksson et al., 2025; Gebru et al., 2021).
-  2. (ii)
-    Construct clarity and validation.
-Make constructs explicit by stating what each task and metric is intended to measure and, where feasible, validate measures with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024).
-  3. (iii)
-    Evaluation resilience.
-Reduce opportunities for gaming or overfitting by incorporating dynamic or adversarial evaluation, including undisclosed items and evolving prompts, consistent with established security testing practices (Oprea and Vassilev, 2023; Ghosh et al., 2025).
-  4. (iv)
-    Broader scope.
-Expand coverage to multimodal, non-English, and higher-risk domains, and include multi-agent or autonomous settings where system level failures are more likely to surface (Salazar et al., 2025; Raza et al., 2025c).
-  5. (v)
-    Multi-stakeholder alignment.
-Mitigate potential commercial or domain bias by engaging multiple stakeholders and making explicit linkages to policy instruments and governance frameworks (Outeda, 2024).
-  6. (vi)
-    Timeliness and maintenance.
-Support continuous evaluation through automated pipelines that enable periodic re-testing across the model lifecycle, improving comparability over time (Wang et al., 2024).
-  7. (vii)
-    Exploration of unknown risks.
-Integrate structured red teaming and stress tests under distribution shift, and monitor for emergent behaviors that may elude static suites, drawing on existing practices such as ALERT and Rainbow Teaming (Tedeschi et al., 2024; Samvelyan et al., 2024; Liu et al., 2023a).
-Figure 5.
-Responsible AI governance landscape and linkage to our evaluation criteria.
-Technical factors correspond primarily to rubric items C1–C7 (fairness/bias, toxicity, security/robustness, mis/disinformation, deepfakes/media integrity, privacy, system-level failure).
-Legal instruments (e.g., EU AI Act, NIST AI RMF) ground C10 (governance alignment).
-Risk & assurance processes (anticipation, responsiveness, reflexivity, inclusion, knowledge management) support adaptive testing (C8) and metric validity/record-keeping (C9).
-Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks, motivating the gaps observed in Tables 58.
+Based on the benchmark gaps we found in the literature, we suggest constructive opportunities to complement existing benchmarks and to support broader governance use.
+First,we recommend that benchmark developers encourage diverse and well-documented corpora across regions, languages, and demographics, accompanied by dataset cards or data statements that capture provenance, consent, licensing, and curation criteria to enhance traceability and accountability (Eriksson et al., 2025; Gebru et al., 2021).
+Equally important is construct clarity: each task and metric should state explicitly what it is intended to measure and, where feasible, be validated with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024).
+Without these foundations, high benchmark scores risk measuring test-taking ability rather than genuine safety.
+Incorporating adversarial evaluation, including undisclosed items and evolving prompts can be used bring benchmarks in line with established security testing practices (Oprea and Vassilev, 2023b; Ghosh et al., 2025).
+Scope should also expand to cover multimodal, non-English, and higher-risk domains, as well as multi-agent or autonomous settings where system-level failures are more likely to surface (Salazar et al., 2025).
+These benchmarks rarely make explicit linkages to policy instruments.
+Engaging multiple stakeholders and mapping evaluation outputs to governance frameworks would mitigate commercial or domain bias and improve regulatory utility (Outeda, 2024).
+Continuous evaluation through automated pipelines that enable periodic re-testing across the model lifecycle would further improve comparability over time (Wang et al., 2024).
+Finally, structured red teaming and stress tests under distribution shift should be integrated to surface emergent behaviours that may elude static suites, drawing on practices such as ALERT and Rainbow Teaming.
+A broader concern emerging from this analysis is that many existing benchmarks may facilitate what might be termed “safetywashing”: models can achieve high scores on structured benchmarks such as SafetyBench (Zhang et al., 2024) by learning keyword-driven refusal scripts, while failing under the multi-turn, emotionally complex interactions documented in real-world deployment incidents (Roose, 2023).
+This pattern underscores the need to move beyond static compliance checks toward evaluation paradigms that capture the adversarial and socially situated nature of human–model interaction.
+The governance landscape framework is intended to illustrate that responsibility is not a single algorithmic check, but a multi-dimensional stewardship involving technical, legal, and social factors.
+Figure 8 provides the foundational logic for our ten-criterion rubric, ensuring that our technical assessment is grounded in real-world governance requirements.
 To operationalize these improvements and make evidence portable across audits, we now specify lifecycle KPIs that teams can track during development, deployment, and post-market monitoring.
-4.2.4. Key Performance Indicators (KPIs) for Responsible AI
+4.2.4. Key Performance Indicators (KPIs) for Responsible GenAI
 KPIs provide measurable criteria to assess the performance of AI systems across critical dimensions of responsibility (Mattioli et al., 2024).
@@
 They bridge governance expectations with technical implementation so that fairness, transparency, robustness, privacy, and sustainability are evidenced through quantifiable measures, enabling compliance and continuous improvement.
-We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior work (Pessach and Shmueli, 2022; Rajkomar et al., 2018; Van Wynsberghe, 2021), and introduce pragmatic, author-defined KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
-Notation and conventions.
-$f(x)$ is the model score for input $x$ (state whether probability, logit, or other monotone score); $y\in\{0,1\}$ is the true label; $\hat{y}=\mathbb{1}\{f(x)\geq t\}$ is the predicted label with a single threshold $t$ applied to all subgroups unless stated; $A$ is a protected attribute with subgroups $a$ and $b$; $\phi_{0}$ is a baseline output and $\phi_{i}(x)$ is the attribution for feature $i$ (e.g., SHAP or Integrated Gradients); “accuracy” refers to a stated metric (e.g., top-1, F1, AUROC) computed per subgroup with the same threshold $t$; a perturbation is a noise or attack consistent with a stated threat model; $P_{\mathrm{avg}}$ is average power (W), $T$ is time (h), $\mathrm{CI}$ is grid carbon intensity (kgCO2e/kWh), and $\mathrm{PUE}$ is Power Usage Effectiveness.
+We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior work (Pessach and Shmueli, 2022), and introduce pragmatic, author-defined KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
+Notation and conventions $f(x)$ is the model score for input $x$ (state whether probability, logit, or other monotone score); $y\in\{0,1\}$ is the true label; $\hat{y}=\mathbb{1}\{f(x)\geq t\}$ is the predicted label with a single threshold $t$ applied to all subgroups unless stated; $A$ is a protected attribute with subgroups $a$ and $b$; $\phi_{0}$ is a baseline output and $\phi_{i}(x)$ is the attribution for feature $i$ (e.g., SHAP or Integrated Gradients); “accuracy” refers to a stated metric (e.g., top-1, F1, AUROC) computed per subgroup with the same threshold $t$; a perturbation is a noise or attack consistent with a stated threat model; $P_{\mathrm{avg}}$ is average power (W), $T$ is time (h), $\mathrm{CI}$ is grid carbon intensity (kgCO2e/kWh), and $\mathrm{PUE}$ is Power Usage Effectiveness.
   1. (1)
@@
     $\mathrm{SPD}=P(\hat{y}=1\mid A=a)-P(\hat{y}=1\mid A=b).$
-    It defines the difference in positive outcome rates across subgroups. We specify state $A$, $t$, sample sizes; or bootstrap and a significance test.
+[…]
   4. (4)
@@
     $\Delta_{\mathrm{acc}}=\left|\mathrm{Accuracy}_{A=a}-\mathrm{Accuracy}_{A=b}\right|.$
-    This measures the absolute gap in a stated accuracy metric across subgroups (Raza et al., 2024c). It reports per subgroup and key intersections; keep $t$ fixed across subgroups.
+    This measures the absolute gap in a stated accuracy metric across subgroups (Raza et al., 2024). It reports per subgroup and key intersections; keep $t$ fixed across subgroups.
   6. (6)
@@
     $R=\frac{\mathrm{Errors\ under\ perturbation}}{\mathrm{Total\ perturbed\ inputs}}\times 100,\quad\Delta_{\mathrm{acc}}^{\mathrm{rob}}=\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{pert}}.$
-    It measures the error rate and accuracy drop under stated threat models (Raza et al., 2025d).
+    It measures the error rate and accuracy drop under stated threat models.
 We specify attack type and strength (e.g., $\ell_{p}$ bounds), number of steps, and per-attack results.
+For generative models, robustness is more appropriately measured as attack success rate (ASR) under jailbreak or prompt-injection attempts, safety-refusal consistency under paraphrase, and output stability under temperature and seed variation, rather than misclassification rate under $\ell_{p}$ perturbations
   7. (7)
@@
 We define the harm taxonomy and decision costs; consider a harm weighted rate in safety critical settings.
+In generative deployments, critical errors include hallucinated medical or legal facts, harmful completions, and PII leakage incidents; $E_{h}$ should accordingly be computed as the incidence of such events per $N$ generations in the target domain, with harm severity tiers defined by deployment context
   10. (10)
@@
 We state measurement method (e.g., nvidia-smi vs power meter), scope (training or inference), normalization unit (e.g., per 1k tokens), and whether CI is location or market based (e.g., CodeCarbon) (Song et al., 2025).
+Sections 56 use this rubric (C1–C10) and the lifecycle KPIs as a running lens to interpret domain evidence and highlight where current practice still lacks governance-ready coverage.
 4.3. Explainable AI
-Having established how benchmarks cover risk families, we examine explainability (XAI) methods in GenAI methods and what they can and can not do, evidence under audit.
+Recent GenAI era deployments expose failure modes (hallucination, privacy leakage, deepfakes, and tool-use errors) that cut across domains; we therefore use C1–C10 and the KPI set as a consistent lens throughout the remainder of the survey.
+Having established how benchmarks cover risk families, we examine explainability methods and what they can and can not do, evidence under audit.
 Explainable AI (XAI) comprises methods that render model behaviour and outputs intelligible to human stakeholders (Meske et al., 2022).
@@
 Interpretability pathways (feature relevance, example-based, surrogates, and visual/text explanations) feed fairness diagnostics (e.g., SPD/DI) and calibration analyses, supporting governance-aligned optimisation $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}+\lambda\,\mathcal{L}_{\mathrm{fair}}$.
-For LLMs/VLMs, prompt traces and cross-modal attention provide additional observability.
-Scope and challenges in the GenAI era.
+For LLMs/VLMs, prompt traces and cross-modal attention provide additional observability
+Scope and challenges in the GenAI
 Foundation models raise explanation challenges beyond conventional ML.
@@
 Consequently, GenAI-oriented XAI should expose decoding pathways, visualise modality interactions, and characterise failure modes, not merely attribute tokens.
-Methods in practice.
+Methods in practice
 XAI methods can be ante-hoc (interpretability built into the model class, e.g., sparse linear models, rule lists, decision trees) or post-hoc (explain an existing black-box locally or globally).
-Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016a); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023a).
+Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016a); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023).
 For LLMs and Vision-language-models (VLMs), these are complemented by prompt-engineering probes, chain-of-thought (CoT) audits, and tool-use traces (Natarajan and Nambiar, 2024).
-Importantly, attention maps and saliency are not proofs of causal reasoning and should be triangulated with counterfactual tests and ablations (Jain and Wallace, 2019; Serrano and Smith, 2019).
+Importantly, attention maps and saliency are not proofs of causal reasoning and should be triangulated with counterfactual tests and ablations (Jain and Wallace, 2019).
 Figure 9 situates post-hoc explanations within an RAI pipeline.
+This framework bridges the gap between abstract interpretability theory and concrete engineering practice.
+Its specific role is to show how attribution signals and fairness checks (Barocas and Selbst, 2016) can be integrated into a standard pipeline to produce the ”portable evidence” needed for external audits and institutional trust.
 Predictions $\hat{y}=M_{\phi}(x)$ can be accompanied by gradient-based importance $I(x_{i})=\partial M_{\phi}/\partial x_{i}$ to support local inspection; example-based evidence provides case-level comparators; and visual/text explanations aid end-user understanding.
@@
 Adding canonical sources: ECE and Brier are standard measures (Pleiss et al., 2017). Calibrated probabilities, paired with faithful explanations, yield more reliable user-trust assessments.
-A compact theoretical frame.
-Let $f_{\theta}:\mathcal{X}\!\to\!\mathcal{Y}$ denote a predictive model (or $p_{\theta}(\cdot\mid x,c)$ for generative systems with context $c$). We define an explanation operator
+A compact theoretical frame
+Let $p_{\theta}(\cdot\mid x,c,\omega)$ denote a generative model with input $x$, context $c$ (e.g., system prompt, retrieved documents), and stochastic decoding parameters $\omega$ (temperature, top-$k/p$).
+The predictive case $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ is recovered as a deterministic special instance
 $\mathcal{E}:(x,f_{\theta},\omega)\mapsto z\in\mathcal{Z},$
@@
 For VLMs, $\mathcal{X}=\mathcal{X}_{\text{text}}\times\mathcal{X}_{\text{vision}}$ and $z$ pairs textual rationales with grounded visual regions.
-Desiderata (axioms).
-An explanation $z=\mathcal{E}(x,f_{\theta},\omega)$ is useful when it approximately satisfies: (i) local accuracy (additive decompositions sum to output deviation from a baseline) (Lundberg and Lee, 2017); (ii) sensitivity/monotonicity (if $f_{\theta}$’s dependence on feature $i$ increases at $x$, its attribution should not decrease) (Sundararajan et al., 2017); (iii) invariance to representation-preserving transforms and to behaviourally equivalent models (Sundararajan et al., 2017); (iv) stability (Lipschitz continuity in $x$ and $\omega$(Alvarez-Melis and Jaakkola, 2018); (v) counterfactual consistency (factors highlighted by $z$ admit feasible interventions that flip or materially shift the output); and (vi) parsimony (minimal complexity $\Omega(z)$ subject to fidelity) (Kusner et al., 2017).
+Desiderata (axioms) An explanation $z=\mathcal{E}(x,f_{\theta},\omega)$ is useful when it approximately satisfies: (i) local accuracy (additive decompositions sum to output deviation from a baseline) (Lundberg and Lee, 2017); (ii) sensitivity/monotonicity (if $f_{\theta}$’s dependence on feature $i$ increases at $x$, its attribution should not decrease) (Sundararajan et al., 2017); (iii) invariance to representation-preserving transforms and to behaviourally equivalent models (Sundararajan et al., 2017); (iv) stability (Lipschitz continuity in $x$ and $\omega$(Alvarez-Melis and Jaakkola, 2018); (v) counterfactual consistency (factors highlighted by $z$ admit feasible interventions that flip or materially shift the output); and (vi) parsimony (minimal complexity $\Omega(z)$ subject to fidelity) (Kusner et al., 2017).
 These properties specialise naturally to multimodal settings via alignment between rationales and grounded visual evidence.
@@
 Reporting stability, counterfactual checks, and provenance mitigates these risks and makes explanations suitable as audit evidence rather than marketing artefacts.
+Based on our synthesis findings, there is a dangerous ”Explanation Gap” in how we use these tools.
+While SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016b) provide useful snapshots, the literature reveals they often create ”plausible but unfaithful” stories.
+They give us a narrative that makes sense to a human, but because transformers are so complex and non-linear, these stories often don’t match what is actually happening under the hood (Strathern, 1997a).
+This creates a false sense of security that we think we’ve reached transparency when we’ve really just seen a surface-level summary.
 Next, we present AI-ready test-beds that computes lifecycle KPIs and logging XAI evidence as audit-ready artifacts.
@@
 AI-ready test-beds advance RAI by providing controlled environments to develop, test, and validate AI models with reliability and reproducibility guarantees.
-In the post-ChatGPT era, where GenAI is increasingly embedded in products and services, such platforms help assess non-determinism, multimodal behavior, and risks like bias, hallucination, and privacy leakage, supporting proportionate assurance and governance alignment.
+In the post-ChatGPT era, such platforms help assess non-determinism, multimodal behavior, and risks like bias, hallucination, and privacy leakage, supporting proportionate assurance and governance alignment.
 We present representative examples in Table LABEL:tab:ai_testbeds.
 The key value of these test-beds is in producing audit-ready evidence datasheets/model cards, prompt/tool logs, KPI dashboards, red-team logs, and versioned runs.
-Table 9. AI evaluation platforms and initiatives used as de facto test-beds for RAI. We distinguish formal test-beds/platforms from broader labs/initiatives commonly used for evaluation.
+Table 7. AI evaluation platforms and initiatives used as de facto test-beds for RAI. We distinguish formal test-beds/platforms from broader labs/initiatives commonly used for evaluation
 Name Domain Type RAI-focused features
-AI4EU AI-on-Demand Platform (cor, 2019) General AI development Platform Ethical-by-design resources; model/dataset catalogue; support for transparency and explainability workflows (used to prototype responsible AI pipelines).
-IEEE Ethical AI Systems Test Bed (win, 2019) Ethical AI development Test-bed/Platform Evaluation against ethical frameworks; human-centred design; fairness-oriented assessment protocols; documentation of decision rationale.
-AI Testbed for Trustworthy AI (TNO) (tno, 2020) Trustworthy AI Test-bed Robustness and transparency checks; fairness analysis; risk/assurance reporting for deployment contexts.
-ETH Zurich Safe AI Lab (saf, 2020) Safe and fair AI Lab/Platform Safety-critical validation; robustness/hallucination probing; reliability and interpretability methods for high-stakes settings.
-HUMANE AI (hum, 2020) Human-centric AI Platform/Initiative Alignment with human values; societal-impact and fairness guardrails; guidance for participatory, responsible design.
-AI for Good (ITU) (AIf, 2017) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
-UKRI Trustworthy Autonomous Systems (TAS) Hub (TAS, 2020) Autonomous systems Program/Test-bed Accountability and transparency for autonomy; safety cases; conformance/compliance support for real-world systems.
+AI4EU AI-on-Demand Platform (AI4EU, 2019) General AI development Platform Ethical-by-design resources; model/dataset catalogue; support for transparency and explainability workflows (used to prototype responsible AI pipelines).
+IEEE Ethical AI Systems Test Bed (IEEE, 2019) Ethical AI development Test-bed/Platform Evaluation against ethical frameworks; human-centred design; fairness-oriented assessment protocols; documentation of decision rationale.
+AI Testbed for Trustworthy AI (TNO) (for Applied Scientific Research, 2020) Trustworthy AI Test-bed Robustness and transparency checks; fairness analysis; risk/assurance reporting for deployment contexts.
+ETH Zurich Safe AI Lab (Zurich, 2020) Safe and fair AI Lab/Platform Safety-critical validation; robustness/hallucination probing; reliability and interpretability methods for high-stakes settings.
+HUMANE AI (HumaneAI, 2020) Human-centric AI Platform/Initiative Alignment with human values; societal-impact and fairness guardrails; guidance for participatory, responsible design.
+AI for Good (ITU) ((2017), ITU) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
+UKRI Trustworthy Autonomous Systems (TAS) Hub (Research and Innovation, 2020) Autonomous systems Program/Test-bed Accountability and transparency for autonomy; safety cases; conformance/compliance support for real-world systems.
 Algorithmic Justice League (Algorithmic Justice League, 2024) Algorithmic fairness Initiative Bias discovery and awareness; audits and mitigation advocacy; widely referenced in fairness evaluation contexts (not a formal test-bed).
 AI Commons [^3] https://ai-commons.org/ Open AI collaboration Initiative Open, transparent collaboration patterns; governance-mindful sharing; community evaluation practices (not a formal test-bed).
-ClarityNLP (Healthcare) (Cla, 2021) Healthcare NLP Platform Clinical text pipelines; emphasis on fairness, transparency, and ethical data use; supports domain-specific evaluation.
+ClarityNLP (Healthcare) (ClarityNLP, 2021) Healthcare NLP Platform Clinical text pipelines; emphasis on fairness, transparency, and ethical data use; supports domain-specific evaluation.
 ToolSandbox (Lu et al., 2024a) Tool-using LLMs Research test-bed Stateful, multi-step/equipped LLM evaluation; probes privacy, fairness, transparency and accountability in interactive tasks.
-Chatbot Arena (LMSYS) (cha, 2024) LLM evaluation Community platform Open-source platform where users compare LLMs in head-to-head “arena” battles; pairwise ranking with more than 800k community votes and transparent leaderboards; all evaluation pipelines and data are public for reproducibility.
-AI Fairness 360 (AIF360) (Bellamy et al., 2018a) Fairness evaluation Toolkit Open-source Python toolkit with 70+ fairness metrics and 9 mitigation algorithms; supports bias detection, fairness explanations and interactive web interface for non-experts.
-Zeno (zen, 2023) AI evaluation Platform Interactive evaluation platform; enables data exploration across any data type and automatically discovers systematic failures using slice finder; supports interactive charts to compare evaluation metrics across slices and models.
-Galileo (gal, 2023) Generative AI evaluation Platform Automated evaluation metrics for GenAI (RAG and agent metrics plus safety/security metrics); real-time guardrails to block hallucinations, PII and prompt injection; captures corner cases and allows custom metrics.
+Chatbot Arena (LMSYS) (Org, 2024) LLM evaluation Community platform Open-source platform where users compare LLMs in head-to-head “arena” battles; pairwise ranking with more than 800k community votes and transparent leaderboards; all evaluation pipelines and data are public for reproducibility.
+AI Fairness 360 (AIF360) (Bellamy et al., 2018) Fairness evaluation Toolkit Open-source Python toolkit with 70+ fairness metrics and 9 mitigation algorithms; supports bias detection, fairness explanations and interactive web interface for non-experts.
+Zeno (Zeno, 2023) AI evaluation Platform Interactive evaluation platform; enables data exploration across any data type and automatically discovers systematic failures using slice finder; supports interactive charts to compare evaluation metrics across slices and models.
+Galileo (Labs, 2023) Generative AI evaluation Platform Automated evaluation metrics for GenAI (RAG and agent metrics plus safety/security metrics); real-time guardrails to block hallucinations, PII and prompt injection; captures corner cases and allows custom metrics.
 Arize AI (Dong et al., 2024) AI observability Platform AI observability and LLM evaluation platform that automatically detects model issues, traces root causes via indexed training/validation/production data and fine-tunes performance; includes dashboards for evaluation/performance tracing, explainability and fairness.
 The Bias Tracing tool monitors fairness metrics (e.g., recall parity, false positive parity, disparate impact) and surfaces features and cohorts contributing to bias.
-PromptLayer (pro, 2022) Prompt engineering Platform Prompt-evaluation platform with visual pipelines for scoring prompts against golden datasets, backtesting with historical data, regression testing and continuous integration of evaluation pipelines; useful for systematic RAG/agent evaluations.
+PromptLayer (PromptLayer, 2022) Prompt engineering Platform Prompt-evaluation platform with visual pipelines for scoring prompts against golden datasets, backtesting with historical data, regression testing and continuous integration of evaluation pipelines; useful for systematic RAG/agent evaluations.
 FairX (Sikder et al., 2024) Fairness benchmarking Open-source toolkit Modular benchmarking tool for fairness, data utility and explainability; supports training of fair generative models, evaluation of synthetic data and fairness metrics with pre-, in- and post-processing debiasing techniques.
 BEATS bias test suite (Abhishek et al., 2025) LLM bias/ethics evaluation Test suite Benchmark for evaluating bias, ethics, fairness and factuality across 29 metrics for LLM outputs; measures demographic and cognitive biases, ethical reasoning and factuality; aims to diagnose factors driving bias and develop mitigation strategies.
-5. RAI Application Across Domains
+4.4.1. The Lifecycle Evidence Matrix: Mapping Metrics to Engineering Controls
+To facilitate the implementation of our proposed evaluation infrastructure, we synthesize the relationship between KPIs, benchmarks, and technical tools.
+As argued throughout this survey, the effectiveness of Responsible GenAI hinges on the ability to produce portable, verifiable evidence across the lifecycle.
+Table 8 provides a decision-support mapping that links the KPIs defined in Section 4.2.4 to the specific testbeds and datasets that operationalize them.
+This matrix enables practitioners to select the appropriate ”tooling stack” based on their specific governance requirements, whether prioritizing fairness audits in Finance or robustness testing in Defense.
+Table 8. The ResGenAI Accountability Matrix: Linking KPIs to Benchmarks, Datasets, and Audit Toolkits
+KPI Category Primary KPIs Benchmarks & Datasets Audit Toolkits & Testbeds
+Fairness & Bias SPD, DI, $\Delta_{\mathrm{acc}}$ SafetyBench (Zhang et al., 2024), HumaniBench (Raza et al., 2025) AIF360 (Bellamy et al., 2018), Fairlearn (Weerts et al., 2023), Arize (Dong et al., 2024)
+Safety & Toxicity $E_{h}$ (High-stakes error) HarmBench (Mazeika et al., 2024), SALAD-Bench (Li et al., 2024a) Unitary (Ltd., [n.
+d.]), Project Moonshot (AI Verify Foundation, 2024), Galileo (Labs, 2023)
+Robustness & Security $R$ (Robustness), ASR ALERT (Tedeschi et al., 2024), Rainbow Teaming (Samvelyan et al., 2024) ART (Research, [n.
+d.]), PromptLayer (PromptLayer, 2022), ToolSandbox (Lu et al., 2024a)
+Privacy $P$ (Compliance rate) PrivLM-Bench (Li et al., 2024b), PrivacyLens (Shao and et al., 2024) Privacy Meter (Privacy and Lab, [n. d.]), Google DP (Google, [n. d.]), SecretFlow (Group, [n. d.])
+Explainability $E(x)$, LFG (Faithfulness) HELM (Liang et al., 2023), XSTest (Röttger et al., 2024) LIT (Google PAIR, [n. d.]), InterpretML, AIX360 (Arya et al., 2019)
+Sustainability $E_{\mathrm{kWh}}$, $C_{\mathrm{CO_{2}e}}$ CodeCarbon (Song et al., 2025) Evidently (Evidently AI, [n. d.]), Arize (Dong et al., 2024), Zeno (Zeno, 2023)
+4.4.2. Turning Metrics into Assurance: The ResGenAI Audit Loop
+The technical efficacy of ResGenAI remains localized unless it can be translated into portable audit evidence.
+As visualized in Fig.
+10, our framework proposes a three layer loop that bridges the Evaluation Gap identified in Section 6.4.
+In this cycle, the technical KPIs defined in Section 4.2.4 act as the raw sensory data for the ecosystem.
+These metrics are packaged into standardized audit artifacts such as model cards and red teaming logs which serve as the portable evidence required by compliance auditors under the EU AI Act (European Union, 2024) or NIST AI RMF (of Standards and, NIST).
+By establishing this formal “chain of trust”, organizations can move beyond static safety checks toward continuous stewardship where regulatory feedback directly informs the recalibration of technical metrics.
+Figure 10. The ResGenAI Audit Loop: translating technical KPIs into portable audit evidence to support regulatory assurance and continuous stewardship across the foundation model lifecycle.
+5. Responsible Generative AI Application Across Domains
 GenAI systems are increasingly embedded in decision-making processes across diverse domains such as healthcare, finance, human resource, defense and education.
@@
 We translate the above findings into stakeholder-aligned practices and a deployment-oriented roadmap, with examples and decision pointers.
+Each subsection below examines domain-specific applications, challenges, and strategies through the lens of our governance-aligned rubric (C1–C10, Section 4.2.1) and lifecycle KPIs (Section 4.2.4).
+For every domain we (i) identify which rubric criteria are most salient, (ii) assess how the benchmarks in Tables 46 cover those risks, and (iii) indicate which KPIs practitioners should prioritize.
+Table 9 consolidates this analysis, mapping each sector to its primary criteria gaps and priority KPIs.
 5.1. Healthcare
-Applications.
-GenAI is transforming healthcare across clinical, biomedical, and administrative domains, but each application underscores the importance of RAI safeguards.
-For example, while generative models improve efficiency in clinical note generation and triage chatbots (Johnson and Patel, 2023; Alagha and Ahn, 2023), they raise concerns about accuracy, liability, and bias in medical decision support.
-Disease detection and drug discovery pipelines (Miotto et al., 2022; Liu et al., 2023c) illustrate GenAI’s potential for accelerating science, yet demand rigorous validation, explainability, and regulatory alignment before clinical deployment.
+Applications
+Generative models improve efficiency in clinical note generation and triage chatbots (Alagha and Ahn, 2023), they raise concerns about accuracy, liability, and bias in medical decision support.
+Disease detection and drug discovery pipelines (Liu et al., 2023c) illustrate GenAI’s potential for accelerating science, yet demand rigorous validation, explainability, and regulatory alignment before clinical deployment.
 Synthetic health data helps mitigate privacy risks and dataset imbalance (Walsh et al., 2023), but governance frameworks must ensure that synthetic outputs do not inadvertently reinforce hidden biases.
 Similarly, tools like Watson for Oncology highlight the need for transparency and interpretability in high-stakes decision-making (Memorial Sloan Kettering Cancer Center, 2024).
-Even promising breakthroughs, such as GenAI-designed therapeutics (Smith et al., 2025), requires human oversight on trial integrity, equitable access, and long-term monitoring.
+Even promising breakthroughs, such as GenAI-designed therapeutics, require human oversight on trial integrity, equitable access, and long-term monitoring.
 In this way, healthcare exemplifies both the promise of GenAI and the critical role of RAI practices, balancing innovation with safety, accountability, and fairness.
-RAI Challenges.
-Responsible adoption of GenAI in healthcare is constrained by privacy violations (Shiraz and Lee, 2019; McGraw and Mandl, 2020; Goldstein and Holtzman, 2021), hallucination of critical clinical facts (Menny and Braun, 2019; Mamo et al., 2023a; Li et al., 2023c), algorithmic bias in diagnostics (Obermeyer et al., 2019; Gonzalez et al., 2022; Venkatasubramanian, 2020), unclear accountability for errors (Kaminski, 2022; Kraft and Jonas, 2021), opacity of generative models (Ghorbani et al., 2019; Tjoa and Guan, 2023; Donoso-Guzmán et al., 2025), and difficulty of compliance with traditional regulations (Price and Gerke, 2022; Kaissis et al., 2021; Khalid et al., 2023).
-Empirical studies highlight these risks: a 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023c); Rosbach et al.
-(2024) found a 7% automation-bias error rate in a pathology study where clinicians overturned correct diagnoses by over-trusting AI outputs (Rosbach et al., 2024); and a systematic review of 778 studies showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023).
-RAI Strategies.
+ResGenAI challenges
+Responsible adoption of GenAI in healthcare is constrained by privacy violations (Shiraz and Lee, 2019), hallucination of critical clinical facts (Li et al., 2023b), algorithmic bias in diagnostics (Obermeyer et al., 2019), unclear accountability for errors (Kaminski, 2022), opacity of generative models (Ghorbani et al., 2019), and difficulty of compliance with traditional regulations (Price and Gerke, 2022).
+Empirical studies highlight these risks, for example, a 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023b); Rosbach et al.
+further found a 7% automation-bias error rate in a pathology study where clinicians overturned correct diagnoses by over-trusting AI outputs (Rosbach et al., 2024); and a systematic review of 778 studies showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023).
+Evaluation context
+Mapping these risks to our rubric, the dominant healthcare concerns correspond to C4 (hallucination of clinical facts), C1 (diagnostic bias), and C6 (patient data leakage).
+Only 7 of 21 benchmarks in Table 4 score above zero on C4 (Misinformation/ Disinformation), and none specifically target medical hallucination.
+For privacy, only PrivLM-Bench and PrivacyLens provide strong C6 (Privacy) coverage, however, neither addresses HIPAA-specific leakage vectors such as clinical note memorization.
+C7 (System-level failure) is critical for RAG-based clinical pipelines and is tested by only 4 of 21 benchmarks.
+The KPIs most relevant to this domain are the high-stakes error rate $E_{h}$ (KPI 9), operationalized as hallucinated clinical facts per $N$ consultations, data privacy compliance $P$ (KPI 2) under HIPAA, and fairness metrics SPD/DI (KPIs 3–4) stratified by patient demographics.
+ResGenAI strategies
 Several strategies are being employed to enable responsible GenAI adoption in healthcare.
-Grounding methods such as retrieval-augmented generation (RAG) help reduce hallucinations by linking model outputs to verifiable medical sources (Wu et al., 2024).
-Human-in-the-loop (HITL) validation integrates clinician oversight into diagnostic and documentation workflows to mitigate automation bias (Stojanovic et al., 2021; He et al., 2023b).
-Privacy-preserving approaches, including differential privacy and federated learning, protect sensitive health records while supporting multi-institutional collaboration (Kaissis et al., 2021; Khalid et al., 2023).
-Fairness-aware data augmentation and structured fairness audits are being applied to improve equity in clinical outcomes (Pfohl et al., 2022; Chen et al., 2023).
-Post-hoc explainability techniques that are applied after model training to make opaque AI models interpretable such as SHAP and LIME, are critical for clinician trust (Nasarian et al., 2023; Donoso-Guzmán et al., 2025; Guidotti et al., 2021).
-Alignment with regulatory frameworks like Food and Drug Administration (FDA) guidance ensures safe and accountable deployment (Kaminski, 2022; Kraft and Jonas, 2021).
+Grounding methods (C4, KPI 9) such as RAG can help reduce hallucinations by linking model outputs to verifiable medical sources.
+Effectiveness should be tracked via $E_{h}$; a reported 1.47% hallucination rate provides a baseline.
+Human-in-the-loop (HITL) validation (C7, C8, KPI 10) integrates clinician oversight into diagnostic and documentation workflows to mitigate automation bias.
+Audit frequency $F_{a}$ and resolution time $T_{r}$ should be reported.Privacy-preserving approaches (C6, KPI 2), including differential privacy and federated learning, protect sensitive health records while supporting multi-institutional collaboration.
+Fairness-aware data augmentation and structured fairness audits (C1, KPIs 3–5) are being applied to improve equity in clinical outcomes.
+HumaniBench (C1=2 in Table 4) offers a starting template but lacks clinical subgroup stratification.
+Post-hoc explainability (C9, KPIs 7–8) techniques such as SHAP and LIME are critical for clinician trust.Explanation quality should be assessed via the local faithfulness gap LFG (KPI 8).
+Alignment with regulatory frameworks (C10) like FDA guidance ensures safe and accountable deployment.
+Table 6 shows that only HELM and MLCommons v0.5 provide partial transparency support relevant to FDA conformity assessment.
 5.2. Finance
-Applications.
-GenAI is transforming multiple facets of finance, with impacts on efficiency, risk management, and personalization (Ngai et al., 2011; Cao, 2022b).
-Key applications include customer-facing bots that provide interactive banking and wealth-management support (Wang et al., 2023b; Chen and Lee, 2022; Ngai et al., 2011), modeling of fraud detection, credit scoring, risk forecasting, and insurance assessment (Hand, 2018; Mehrabi et al., 2021b; Cao, 2022b), algorithmic trading that uses generative simulations to optimize portfolios and anticipate market shifts (Zhang et al., 2023; Gupta et al., 2024; HSBC Research, 2025), personalized financial advice adapted to client risk profiles and long-term goals (Silva et al., 2024; Mehrabi et al., 2021b; Wang et al., 2023b), automated compliance reports and market summaries that accelerate auditing and regulatory filings (Araci, 2019; Li et al., 2023a; Kaminski, 2022), and synthetic financial data generation for stress testing, scenario design, and risk modeling (Patki et al., 2016; Wiese et al., 2020; Cao, 2022b).
+Applications
+GenAI is transforming multiple facets of finance, with impacts on efficiency, risk management, and personalization (Cao, 2022).
+Key applications include customer-facing bots that provide interactive banking and wealth-management support, modeling of fraud detection, credit scoring, risk forecasting, and insurance assessment (Cao, 2022), algorithmic trading that uses generative simulations to optimize portfolios and anticipate market shifts (HSBC Research, 2025), personalized financial advice adapted to client risk profiles and long-term goals (Wang et al., 2023b), automated compliance reports and market summaries that accelerate auditing and regulatory filings (Kaminski, 2022), and synthetic financial data generation for stress testing, scenario design, and risk modeling (Cao, 2022).
 Empirical deployments highlight these advances.
 For example, PayPal’s fraud detection reduced false positives by 50% in 2023 while processing 10 million transactions daily (Ngai et al., 2011).
-Zest AI raised loan approvals for minority applicants by 15% in a 2024 U.S.
-bank case study (Hand, 2018; Mehrabi et al., 2021b).
 Charles Schwab’s assistant resolved 62% of client queries and improved satisfaction by 28% in 2023 (Wang et al., 2023b).
 JPMorgan’s simulator improved trading returns by 12% and reduced value-at-risk by 13% in 2023 (Zhang et al., 2023).
-HSBC’s AI achieved 95% accuracy in volatility prediction in 2025 (HSBC Research, 2025; Cao, 2022b).
+HSBC’s AI achieved 95% accuracy in volatility prediction in 2025 (HSBC Research, 2025).
 Silva et al.
-(2024) reported RegTech-GPT summaries rated “complete” by 85% of compliance officers compared with 62% baseline (Silva et al., 2024).
-RAI Challenges.
+also reported RegTech-GPT summaries rated “complete” by 85% of compliance officers compared with 62% baseline (Silva et al., 2024).
+ResGenAI challenges
 GenAI in finance raises a number of domain-specific risks that complicate its responsible adoption.
-These include discriminatory lending that replicates historical bias in credit and valuation data (Hand, 2018; Mehrabi et al., 2021b; Adams et al., 2022), security vulnerabilities where adversarial manipulation targets fraud detection or trading systems (Biggio et al., 2013; Kraft and Jonas, 2021; Ngai et al., 2011), consumer data privacy concerns arising from reliance on sensitive financial records (McGraw and Mandl, 2020; Khalid et al., 2023; Goldstein and Holtzman, 2021), opacity of GenAI that limits auditability and accountability (Doshi-Velez and Kim, 2017; Guidotti et al., 2021; Office of the Comptroller of the Currency, 2020), hallucination and misleading outputs in customer-facing systems that may distort investment or compliance advice (Li et al., 2023c; Mamo et al., 2023a), and market manipulation risk that threatens systemic stability when generative simulations influence trading behavior (U.S.
-Securities and Exchange Commission, 2021; Hendry et al., 2022).
+These include discriminatory lending that replicates historical bias in credit and valuation data , security vulnerabilities where adversarial manipulation targets fraud detection or trading systems, consumer data privacy concerns arising from reliance on sensitive financial records, opacity of GenAI that limits auditability and accountability, hallucination and misleading outputs in customer-facing systems that may distort investment or compliance advice, and market manipulation risk that threatens systemic stability when generative simulations influence trading behavior (Hendry et al., 2022).
 Empirical use cases underscore these risks.
 For example, Zillow’s valuation model undervalued homes in minority neighborhoods and reduced appraised equity for underwriting (Adams et al., 2022).
-Citibank was fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020).
-The Robinhood–GameStop episode demonstrated how synthetic order-flow simulations amplified volatility and required regulatory intervention (U.S.
-Securities and Exchange Commission, 2021).
-RAI Strategies.
+Citibank was also fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020).
+Evaluation context
+The primary finance risks map to C1 (lending discrimination), C3 (adversarial manipulation of trading/fraud systems), C6 (consumer data privacy), and C4 (hallucinated financial advice).
+No benchmark in Table 4 specifically targets lending discrimination or financial adversarial attacks.
+The most relevant KPIs are SPD and DI (KPIs 3–4) applied to loan approval rates across demographic groups, robustness $R$ (KPI 6) under financial adversarial scenarios, and audit frequency $F_{a}$ (KPI 10) aligned with SEC and OCC reporting cadences.
+ResGenAI strategies
 Multiple strategies are being prioritized to address these financial risks.
-Structured fairness audits and fairness-aware data augmentation are increasingly required in lending and credit scoring (Pfohl et al., 2022; Chen et al., 2023; Mehrabi et al., 2021b).
-Post-hoc explainability methods such as SHAP and LIME support regulators in interrogating opaque credit and trading models (Guidotti et al., 2021; Nasarian et al., 2023; Donoso-Guzmán et al., 2025).
-Privacy-preserving techniques, including differential privacy and secure multiparty computation, enable safe collaboration across financial institutions while protecting consumer data (Kaissis et al., 2021; Khalid et al., 2023; McGraw and Mandl, 2020).
-Stronger data provenance frameworks are being developed to track the lineage and validity of synthetic datasets (Patki et al., 2016; Wiese et al., 2020).
-Alignment with regulatory compliance frameworks such as Basel III, SEC, and OCC guidelines anchors AI deployments within enforceable legal standards (Kaminski, 2022; Kraft and Jonas, 2021; Office of the Comptroller of the Currency, 2020).
-Finally, embedding HITL oversight across lending, trading, and compliance ensures accountability by mandating expert review of generative outputs in high-stakes decisions (Stojanovic et al., 2021; He et al., 2023b; Silva et al., 2024).
+Structured fairness audits (C1, KPIs 3–4) and fairness-aware data augmentation are increasingly required in lending and credit scoring.
+Post-hoc explainability (C9, KPIs 7–8) methods such as SHAP and LIME support regulators in interrogating opaque credit and trading models.
+Privacy-preserving techniques (C6, KPI 2), including differential privacy and secure multiparty computation, enable safe collaboration across financial institutions while protecting consumer data.
+Stronger data provenance frameworks are being developed to track the lineage and validity of synthetic datasets.
+Alignment with regulatory compliance (C10) frameworks such as Basel III, SEC, and OCC guidelines anchors AI deployments within enforceable legal standards (Kaminski, 2022).
+Finally, embedding HITL oversight (C7, KPI 10) across lending, trading, and compliance ensures accountability by mandating expert review of generative outputs in high-stakes decisions.
 5.3. Education
-Applications.
-GenAI is increasingly shaping critical aspects of education, expanding opportunities for personalization, access, and efficiency (Luckin et al., 2022; Holmes et al., 2022).
-Key applications include automated grading that accelerates evaluation and reduces instructor workload (Zawacki-Richter et al., 2019; Holmes et al., 2022), personalized tutoring systems that adapt to student knowledge gaps and learning styles (Kulik and Fletcher, 2023; Carnegie Learning, 2023), content generation that produces adaptive curricula and study materials (Kasneci et al., 2023; Zhang and Zhao, 2024), accessibility tools that deliver multimodal and multilingual support for inclusive education (Almalki and Khan, 2023; Holstein and colleagues, 2023), and interactive simulations that foster experiential and inquiry-driven learning (Chen and Huang, 2023; Lai and Xu, 2024).
+Applications
+GenAI is increasingly shaping critical aspects of education, expanding opportunities for personalization, access, and efficiency (Holmes et al., 2022).
+Key applications include automated grading that accelerates evaluation and reduces instructor workload, personalized tutoring systems that adapt to student knowledge gaps and learning styles, content generation that produces adaptive curricula and study materials, accessibility tools that deliver multimodal and multilingual support for inclusive education, and interactive simulations that foster experiential and inquiry-driven learning (Chen and Huang, 2023; Lai and Xu, 2024).
 Recent deployments provide evidence of these advances.
@@
 school district pilot in 2024 showed that 87% of students felt classroom chat assistants clarified concepts and 92% of teachers reported time saved (UK Department for Education, 2024).
-RAI Challenges.
+ResGenAI challenges
 Responsible use of GenAI in education faces several pressing concerns.
-These include cultural insensitivity where models fail to capture local norms and values (Kasneci et al., 2023; UNESCO, 2021), student data privacy and particularly minors’ data privacy given the sensitivity of educational records (Holmes et al., 2022; Livingstone et al., 2021), skills evasion where over-reliance on AI weakens critical thinking and independent problem-solving (Luckin et al., 2022; Lai and Xu, 2024), and pedagogical opacity that limits educators’ ability to evaluate how generative systems produce outputs (Holstein and colleagues, 2023; Zawacki-Richter et al., 2019).
+These include cultural insensitivity where models fail to capture local norms and values, student data privacy and particularly minors’ data privacy given the sensitivity of educational records , skills evasion where over-reliance on AI weakens critical thinking and independent problem-solving, and pedagogical opacity that limits educators’ ability to evaluate how generative systems produce outputs (Zawacki-Richter et al., 2019).
 Empirical use cases underscore these risks.
@@
 In 2022, the Utah State Board of Education cancelled a contract for an AI essay-scorer when auditors reported that the vendor could not document the feature vectors or training data, raising concerns about pedagogical opacity and compliance with state education law (Utah State Board of Education, 2022).
-RAI Strategies.
+Evaluation context
+Education risks concentrate on C6 (minors’ data privacy), C4 (hallucinated pedagogical content), C1 (cultural bias), and C7 (system-level failure in tutoring pipelines).
+No benchmark addresses C6 in the context of minors’ data, and the user inclusivity index $U$ (KPI 11) critical for accessibility in educational settings, lacks any standardized evaluation protocol.
+Skills-evasion effects (a form of C7 system-level failure) are entirely absent from current safety suites.
+Priority KPIs include $P$ (KPI 2) under COPPA/FERPA, $E_{h}$ (KPI 9) for pedagogically harmful outputs, and $U$ (KPI 11) for WCAG conformance.
+ResGenAI strategies
 Multiple strategies are being explored to ensure responsible generative AI in education.
-Fairness audits and external ethics audits verify alignment with frameworks such as the UNESCO AI-ethics rubric and can keep bias metrics below accepted thresholds (UNESCO, 2021; Pfohl et al., 2022; Chen et al., 2023).
-Privacy-preserving techniques including homomorphic encryption and federated learning protect minors’ sensitive data while enabling large-scale analysis (Kaissis et al., 2021; Khalid et al., 2023).
-Clear regulation and policy guidelines developed by international and national bodies provide governance and accountability mechanisms (UNESCO, 2021; Holmes et al., 2022).
-Finally, academic integrity tools such as watermarking and output tracking are being adopted to prevent plagiarism and monitor AI-assisted work (Kasneci et al., 2023; Zhang and Zhao, 2024).
+Fairness audits (C1, KPIs 3–4) and external ethics audits verify alignment with frameworks such as the UNESCO(UNESCO, 2021a) AI-ethics rubric and can keep bias metrics below accepted thresholds.
+Privacy-preserving techniques (C6, KPI 2) including homomorphic encryption and federated learning protect minors’ sensitive data while enabling large-scale analysis (Kaissis et al., 2021).
+Clear regulation and policy guidelines (C10) developed by international and national bodies provide governance and accountability mechanisms.
+Finally, academic integrity tools (C5, C9) such as watermarking and output tracking are being adopted to prevent plagiarism and monitor AI-assisted work (Zhang and Zhao, 2024).
 5.4. Arts and Entertainment
-Applications.
-GenAI is reshaping creative production and cultural participation in the arts and entertainment sector (Crawford, 2021; Traitter and Klein, 2022).
-Prominent applications include creative writing tools that assist in script, story, and lyric development (McStay, 2023; Zhang and Lee, 2023), generative art and design platforms that expand aesthetic experimentation (Elgammal et al., 2019; Schneider and Jörgens, 2022), music composition and production systems that automate melody, arrangement, and mastering (Davis et al., 2024; Briot et al., 2020), film production support for visual effects, dubbing, and storyboarding (Miranda and Perez, 2022; Rodriguez and Santos, 2024), personalized recommendation engines that curate media consumption at scale (Celma, 2010; Anderson and Lee, 2023), and virtual influencers and avatars that shape marketing and fan engagement (Traitter and Klein, 2022; Crawford, 2021).
+Applications
+GenAI is reshaping creative production and cultural participation in the arts and entertainment sector.
+Prominent applications include creative writing tools that assist in script, story, and lyric development, generative art and design platforms that expand aesthetic experimentation, music composition and production systems that automate melody, arrangement, and mastering, film production support for visual effects, dubbing, and storyboarding, personalized recommendation engines that curate media consumption at scale and virtual influencers and avatars that shape marketing and fan engagement (Crawford, 2021).
 Recent deployments provide evidence of these advances.
@@
 Davis et al.
-(2024) released a ten-track AI album in which 62% of listeners in a Spotify trial could not distinguish the music from human compositions (Davis et al., 2024).
-RAI Challenges.
+released a ten-track AI album in which 62% of listeners in a Spotify trial could not distinguish the music from human compositions (Davis et al., 2024).
+ResGenAI challenges
 Responsible use of GenAI in arts and entertainment faces a series of ethical, legal, and cultural concerns.
-These include copyright infringement and intellectual property violation where AI systems reuse creative material without authorization (U.S.
-Copyright Office, 2023; Samuelson, 2023), job displacement that threatens creative labor markets (Crawford, 2021; McStay, 2023), cultural insensitivity when generative tools fail to respect community values or norms (UNESCO, 2021; Kasneci et al., 2023), deepfakes and misinformation that erode authenticity and trust (Chesney and Citron, 2019; Fazio, 2023), authorship and agency disputes where creative ownership is contested (U.S.
-Copyright Office, 2023; Samuelson, 2023), insufficient consent in data usage for training models (WeTransfer, 2025; Crawford, 2021), and opaque attribution that obscures the role of AI in the creative process (Schneider and Jörgens, 2022; Elgammal et al., 2019).
+These include copyright infringement and intellectual property violation where AI systems reuse creative material without authorization, job displacement that threatens creative labor markets, cultural insensitivity when generative tools fail to respect community values or norms deepfakes and misinformation that erode authenticity and trust, authorship and agency disputes where creative ownership is contested, insufficient consent in data usage for training models, and opaque attribution that obscures the role of AI in the creative process (Schneider and Jörgens, 2022; Elgammal et al., 2019).
 Empirical use cases underscore these risks.
-For example, the SAG-AFTRA and WGA strikes of 2023 protested generative AI’s role in replacing talent and using likenesses without consent or compensation.
-The U.S.
+For example, the U.S.
 Copyright Office’s 2023 ruling in Thaler v.
@@
 In 2025, WeTransfer updated its terms of service to allow user content for AI training, sparking backlash over transparency and consent (WeTransfer, 2025).
-RAI Strategies.
+Evaluation context
+The dominant risks map to C5 (deepfakes and media integrity), C1 (representational bias in generated content), and C9 (attribution transparency).
+Only SHIELD and MM-SafetyBench provides explicit C5 evaluation in Table 4, and its scope does not extend to voice cloning or music mimicry, the vectors most salient for this sector.
+Copyright memorization, a C6-adjacent concern, is not addressed by any privacy benchmark.
+Priority KPIs are currently underspecified for creative domains; provenance-verification rates and consent-compliance metrics (analogous to KPI 2) are needed but lack standardized protocols.
+ResGenAI strategies
 Multiple strategies are being adopted to promote responsible use of generative AI in arts and entertainment.
-Artist collaboration platforms and alternative compensation models are emerging to safeguard livelihoods and redistribute value (McStay, 2023; Crawford, 2021).
-Structured bias audits and cultural sensitivity training are being incorporated into creative AI pipelines to mitigate stereotyping (UNESCO, 2021; Pfohl et al., 2022).
-Deepfake detection tools are being deployed to maintain authenticity in music, video, and media (Chesney and Citron, 2019; Fazio, 2023).
+Artist collaboration platforms and alternative compensation models are emerging to safeguard livelihoods and redistribute value.
+Structured bias audits (C1, KPIs 3–4) and cultural sensitivity training are being incorporated into creative AI pipelines to mitigate stereotyping .
+Deepfake detection tools (C5) are being deployed to maintain authenticity in music, video, and media (Chesney and Citron, 2019).
 Collective bargaining and union agreements such as those negotiated by SAG-AFTRA establish protections for creative workers (SAG-AFTRA, 2023).
-International ethical guidelines and transparency labeling requirements provide standards for disclosure and attribution (UNESCO, 2021; Samuelson, 2023).
-Finally, intellectual property management systems are being advanced to secure consent and compliance.
-For example, a 2023 copyright detection tool identified AI-generated violations with 95% accuracy (Zirpoli, 2023a; Traitter and Klein, 2022), and these system requires explicit informed consent before creating digital replicas of performances, offering a contractual model for ethical rights management.
+International ethical guidelines (C10) and transparency labeling requirements provide standards for disclosure and attribution.
+Finally, intellectual property management systems (C6, C9) are being advanced to secure consent and compliance.
+For example, a 2023 copyright detection tool identified AI-generated violations with 95% accuracy (Zirpoli, 2023a; Traitter and Klein, 2022), and these systems require explicit informed consent before creating digital replicas of performances, offering a contractual model for ethical rights management.
 5.5. Agriculture
-Applications.
-GenAI is increasingly transforming agricultural practices through precision methods and data-driven innovation (Kamilaris and Prenafeta-Boldú, 2018; Costa et al., 2023).
-Key applications include precision farming systems that optimize irrigation, fertilization, and pesticide usage (Sapkota et al., 2024; Kamilaris and Prenafeta-Boldú, 2018), crop yield prediction models that integrate weather, soil, and satellite data (Chlingaryan et al., 2018; Ma et al., 2023), pest and disease detection tools that leverage generative augmentation to improve diagnostic accuracy (Mohanty et al., 2016; Li et al., 2022), virtual agronomy advisors that deliver context-sensitive recommendations to farmers (Purdy and Gupta, 2023; Mahdavinejad et al., 2018), agricultural robotics for automated harvesting and monitoring (Duckett et al., 2018; Bechar and Vigneault, 2021), crop breeding simulations that accelerate trait discovery (Varshney et al., 2021; Beltran et al., 2022), and weather forecasting for climate adaptation (Rasp et al., 2020; Nguyen et al., 2023).
-Recent deployments provide evidence of these advances.
-For instance, Microsoft’s FarmBeats platform increased crop yields by 20% and reduced water use by 30% in U.S.
-and Indian trials (Microsoft Research, 2023).
-Retrieval-augmented advisory chatbots deployed in Kenya and India improved farmers’ decision-making speed and confidence in pest and weather management.
-A 2024 FAO pilot of digital sandboxes for pest control in Kenyan farms recorded an 18% reduction in pesticide use, demonstrating sustainable applications of GenAI.
-RAI Challenges.
+Applications
+GenAI is increasingly transforming agricultural practices through precision methods and data-driven innovation.
+Key applications include precision farming systems that optimize irrigation, fertilization, and pesticide usage, crop yield prediction models that integrate weather, soil, and satellite data, pest and disease detection tools that leverage generative augmentation to improve diagnostic accuracy, virtual agronomy advisors that deliver context-sensitive recommendations to farmers, agricultural robotics for automated harvesting and monitoring, crop breeding simulations that accelerate trait discovery, and weather forecasting for climate adaptation (Costa et al., 2023).
+While platforms like Microsoft’s FarmBeats (Microsoft Research, 2023) initially achieved significant yield and water-saving gains through IoT and predictive ML , GenAI is now being integrated to augment these systems through generative weather simulation for risk modeling and as LLM-based natural language interfaces .
+These tools leverage GenAI to translate complex sensor data into actionable, localized advice.
+ResGenAI challenges
 Responsible adoption of GenAI in agriculture faces several interrelated challenges.
-These include hallucination and context-inappropriate advice where unreliable outputs may jeopardize crop management decisions (Mamo et al., 2023b; Stojanovic et al., 2021), farm data ownership conflicts where smallholders lack control over datasets aggregated by multinational platforms (Bronson, 2019; Tzachor et al., 2022), bias toward major corporations and regions that limits the applicability of tools for smallholder and Global South farmers (Sapkota et al., 2024; Rotz et al., 2019), overreliance on AI and erosion of local knowledge that risks displacing traditional farming wisdom (Bronson, 2019; Tzachor et al., 2022), IoT security vulnerabilities that expose connected farm systems to cyberattacks (Mahdavinejad et al., 2018; Khalid et al., 2023), and the persistent digital divide that excludes low-resource communities from advanced AI benefits (Rotz et al., 2019; UNESCO, 2021).
+These include hallucination and context-inappropriate advice where unreliable outputs may jeopardize crop management decisions, farm data ownership conflicts where smallholders lack control over datasets aggregated by multinational platforms, bias toward major corporations and regions that limits the applicability of tools for smallholder and Global South farmers, overreliance on AI and erosion of local knowledge that risks displacing traditional farming wisdom, IoT security vulnerabilities that expose connected farm systems to cyberattacks, and the persistent digital divide that excludes low-resource communities from advanced AI benefits (Rotz et al., 2019; Sapkota et al., 2024).
 Empirical use cases underscore these risks.
@@
 IoT breaches in 2023 European trials highlighted vulnerabilities in sensor networks that exposed proprietary farm-level data.
-RAI Strategies.
+Evaluation context
+Agriculture risks primarily concern C4 (hallucinated agronomic advice), C6 (farm data ownership and privacy), and C7 (system-level failure in IoT-integrated advisory pipelines).
+No benchmark in our assessment provides geo-aware hallucination testing, and C7 remains untested for sensor-to-advice chains.
+Priority KPIs are $E_{h}$ (KPI 9) operationalized as harmful agronomic recommendations per $N$ advisory sessions, $P$ (KPI 2) for data governance compliance, and $E_{\text{kWh}}$ (KPI 12) for sustainability reporting—particularly relevant given the energy constraints of low-resource deployment contexts.
+RAI strategies
 Multiple strategies are being pursued to enable responsible deployment of generative AI in agriculture.
-Geo-aware retrieval and annotation techniques are being designed to contextualize model outputs with local weather, soil, and cultural practices, reducing hallucination risks (Mamo et al., 2023a).
-Data governance contracts are being negotiated to clarify ownership rights and revenue-sharing models for farm-generated datasets (Tzachor et al., 2022; Bronson, 2019).
-Embedding human-in-the-loop oversight into agronomy advisory workflows ensures farmers and extension officers validate AI outputs before implementation (Stojanovic et al., 2021; He et al., 2023b).
-Privacy-preserving IoT integration, including differential privacy and federated learning, protects sensitive farm data across institutions (Kaissis et al., 2021; Khalid et al., 2023).
-Finally, alignment with sustainability regulations such as FAO and EU Common Agricultural Policy frameworks anchors generative AI deployments in enforceable ecological and social standards (UNESCO, 2021; Tzachor et al., 2022).
+Geo-aware retrieval and annotation (C4, KPI 9) techniques are being designed to contextualize model outputs with local weather, soil, and cultural practices, reducing hallucination risks.
+Data governance contracts (C6, KPI 2) are being negotiated to clarify ownership rights and revenue-sharing models for farm-generated datasets.
+Embedding human-in-the-loop oversight (C7, KPI 10) into agronomy advisory workflows ensures farmers and extension officers validate AI outputs before implementation.
+Privacy-preserving IoT integration (C6, KPI 2), including differential privacy and federated learning, protects sensitive farm data across institutions.
+Finally, alignment with sustainability regulations (C10, KPI 12) such as FAO and EU Common Agricultural Policy frameworks anchors generative AI deployments in enforceable ecological and social standards.
 Together, these approaches enhance both the equity and safety of generative AI in global agriculture.
 5.6. Defense
-Applications.
-GenAI is increasingly influencing defense through simulation, sensing, and decision support (U.S.
-Department of Defense, 2021; NATO, 2024).
-Key applications include novel training and war-gaming environments that generate diverse scenarios and accelerate force development (NATO Allied Command Transformation, 2024; Johns Hopkins Applied Physics Laboratory, 2025), threat detection and classification across multi-INT streams that fuse imagery signals and text for faster situational awareness (Wortzel et al., 2025; National Institute of Standards and Technology, 2023), autonomous reconnaissance systems that coordinate teams of uncrewed assets in contested environments (Defense Advanced Research Projects Agency, 2018, 2024), and logistics and supply chain optimization that shortens planning cycles and improves mission readiness (U.S.
-Department of Defense, 2024; U.S.
-Department of Defense, Chief Digital and AI Office, 2024).
+Applications
+GenAI is increasingly influencing defense through simulation, sensing, and decision support.
+Key applications include novel training and war-gaming environments that generate diverse scenarios and accelerate force development, threat detection and classification across multi-INT streams that fuse imagery signals and text for faster situational awareness, autonomous reconnaissance systems that coordinate teams of uncrewed assets in contested environments, and logistics and supply chain optimization that shortens planning cycles and improves mission readiness (News, 2024).
 Recent deployments provide evidence of these advances.
-For instance, DARPA programs such as CODE and successor autonomy efforts demonstrated collaborative uncrewed aircraft operating with minimal communications and supervisor control in flight tests (Defense Advanced Research Projects Agency, 2018; Liu et al., [n.
-d.]).
+For instance, DARPA programs such as CODE and successor autonomy efforts demonstrated collaborative uncrewed aircraft operating with minimal communications and supervisor control in flight tests (Defense Advanced Research Projects Agency, 2018).
 The U.S.
-Army extended enterprise decision platforms in 2024 to speed analysis and targeting workflows at scale (U.S.
-Army, 2024; News, 2024).
-NATO war-games in 2024 integrated generative scenario simulation and reported that AI-generated vignettes helped surface unanticipated vulnerabilities during exercise analysis (NATO Allied Command Transformation, 2024, 2025; Huang et al., 2025a).
-RAI Challenges.
+Army extended enterprise decision platforms in 2024 to speed analysis and targeting workflows at scale (News, 2024).
+NATO war-games in 2024 integrated generative scenario simulation and reported that AI-generated vignettes helped surface unanticipated vulnerabilities during exercise analysis.
+RAI challenges
 Responsible adoption of GenAI in defense introduces ethical operational and legal concerns.
-These include autonomous weapons and responsibility gaps where assignment of accountability for the use of force remains unresolved (International Committee of the Red Cross, 2021; United Nations Institute for Disarmament Research, 2025), opacity of decision making that limits auditability and traceability in high-stakes contexts (U.S.
-Department of Defense, 2021; National Institute of Standards and Technology, 2023), adversarial vulnerabilities that expose perception models and decision aids to spoofing and data poisoning (National Institute of Standards and Technology, 2025a, 2024), escalation of conflicts where fast automated responses may narrow decision windows and increase miscalculation risk (Horowitz et al., 2024; Stockholm International Peace Research Institute, 2024), ethical and legal gaps regarding compliance with international humanitarian law and rules of engagement (NATO, 2024; International Committee of the Red Cross, 2021), research security concerns over dual-use model capabilities and insider threat (U.S.
-Department of Defense, 2025; Corporation, 2024), protection of critical infrastructure as AI is integrated into command networks and sensors (National Institute of Standards and Technology, 2023; U.S.
-Department of Defense, 2025), and facial recognition deployment where demographic performance differences raise risks in security operations (Grother et al., 2019, 2022).
+These include autonomous weapons and responsibility gaps where assignment of accountability for the use of force remains unresolved, opacity of decision making that limits auditability and traceability in high-stakes contexts, adversarial vulnerabilities that expose perception models and decision aids to spoofing and data poisoning, escalation of conflicts where fast automated responses may narrow decision windows and increase miscalculation risk, ethical and legal gaps regarding compliance with international humanitarian law and rules of engagement, research security concerns over dual-use model capabilities and insider threat, protection of critical infrastructure as AI is integrated into command networks and sensors, and facial recognition deployment where demographic performance differences raise risks in security operations.
 Empirical use cases underscore these risks.
 For example, national and alliance exercises reported that AI-supported war-gaming surfaced escalatory flashpoints that required added human oversight (NATO Allied Command Transformation, 2024).
-Government red-team evaluations and standards work identified numerous novel failure modes in generative systems that could enable misinformation or degrade cyber defenses if left unmitigated (National Institute of Standards and Technology, 2025b; WIRED, 2025).
+Government red-team evaluations and standards work identified numerous novel failure modes in generative systems that could enable misinformation or degrade cyber defences if left unmitigated (National Institute of Standards and Technology, 2025b; WIRED, 2025).
 Facial recognition evaluations documented demographic differentials that necessitate strict control and testing before operational use (Grother et al., 2019, 2022).
-RAI Strategies.
+Evaluation context
+Defense risks concentrate on C3 (adversarial robustness), C7 (system-level failure in multi-agent autonomy), C8 (red-team realism), and C10 (governance alignment with IHL and NATO principles).
+C8 is the most critical gap: only 2 of 21 benchmarks (Rainbow Teaming and ALERT) score above zero for adaptive adversary simulation, yet defense applications face the most sophisticated threat actors.
+C7 is entirely absent for multi-agent autonomous systems.
+Priority KPIs are robustness $R$ (KPI 6) under military-grade adversarial scenarios, $E_{h}$ (KPI 9) for consequential targeting or escalation errors, and audit frequency $F_{a}$ (KPI 10) aligned with NATO after-action review cadences.
+RAI strategies
 Multiple strategies are being advanced to ensure responsible deployment.
-Human in the loop oversight is prioritized for mission-critical tasks so that trained operators validate model outputs before action (U.S.
-Department of Defense, 2021; NATO, 2024).
-Meaningful human control is reinforced to keep decisions on the use of force within accountable human authority consistent with international humanitarian law (International Committee of the Red Cross, 2021; United Nations Institute for Disarmament Research, 2025).
-Explainable AI for defense transparency adopts traceability and interpretability requirements to support audit and after-action review (U.S.
-Department of Defense, 2021; National Institute of Standards and Technology, 2023).
-Red teaming and rigorous test and evaluation are expanded to probe adversarial and safety failure modes before fielding (Corporation, 2024; National Institute of Standards and Technology, 2025a).
-International cooperation aligns national practices with alliance principles such as NATO’s Principles of Responsible Use and shared assurance processes (NATO, 2024).
+Human in the loop oversight (C7, C8, KPI 10) is prioritized for mission-critical tasks so that trained operators validate model outputs before action.
+Meaningful human control (C10) is reinforced to keep decisions on the use of force within accountable human authority consistent with international humanitarian law.
+Explainable AI for defense transparency (C9, KPIs 7–8) adopts traceability and interpretability requirements to support audit and after-action review.
+Red teaming (C3, C8, KPI 6) and rigorous test and evaluation are expanded to probe adversarial and safety failure modes before fielding.
+International cooperation (C10) aligns national practices with alliance principles such as NATO’s Principles of Responsible Use and shared assurance processes (NATO, 2024).
 Ethical review boards and program-level governance embed policy risk assessment data governance and model documentation across the AI life cycle (U.S.
 Department of Defense, 2024; Defense Innovation Unit, 2020).
+5.7. Cross-Applications Synthesis of GenAI Maturity
+To provide a critical comparison of how responsible practices are being operationalized across domains, and to connect sectoral risks to the evaluation infrastructure in Sections 4.24.4, we present the extended Synthesis Matrix in Table 9.
+Beyond risk mechanisms and regulatory drivers, the table now maps each sector to its primary rubric-criteria gaps and priority lifecycle KPIs, making explicit which evaluation evidence is available and where portable audit artifacts are missing.
+5.7.1. Stakeholder-Metric Alignment: Operationalizing the Framework
+To ensure the sectoral insights in Section 5 translate into actionable governance, we map the relationship between stakeholder roles and our evaluation infrastructure.
+Based on our synthesized findings, technical frameworks often fail because they do not specify which evidence is required by which actor.
+In our model, Model Developers utilize the robustness (C3) and toxicity (C2) scores from Table 4 to guide iterative alignment during the fine tuning phase.
+Conversely, Compliance Auditors and Regulators prioritize the statistical parity (SPD/DI) KPIs (Section 4.2.4) and governance-alignment criteria (C10) to verify adherence to mandates like the EU AI Act (Outeda, 2024) or NIST AI RMF (of Standards and, NIST).
+Finally, Domain Practitioners in high-stakes fields like Healthcare and Finance rely on the High-Stakes Error Rate ($E_{h}$) and local faithfulness gaps (LFG) to determine the necessary level of human-in-the-loop oversight.
+By bridging domain-specific mechanisms with normalized technical metrics, our framework provides the ”chain of trust” required to move ResGenAI from theoretical safety to verifiable institutional accountability.
+Table 9. Synthesis Matrix: Comparative Analysis of RAI Maturity, Regulatory Pressures, Rubric Criteria Gaps, and Priority KPIs Across Domains.
+Sector Dominant Risk Mechanism Technical Control Maturity Primary Regulatory Driver Primary C1 to C10 Gaps Priority KPIs
+Healthcare Medical hallucination (Lyu et al., 2025); Diagnostic bias (Gonzalez et al., 2022) Medium: RAG (Wu et al., 2024), HITL, Diff.
+Privacy (Kaissis et al., 2021) FDA Guidance (U.S.
+Food and Drug Administration, 2024); HIPAA (Shiraz and Lee, 2019) C4: no clinical hallucination benchmark; C6: no HIPAA specific privacy suite; C7: 4/21 cover system failure $E_{h}$ (9), $P$ (2), SPD/DI (3 to 4), $\Delta_{\text{acc}}$ (5)
+Finance Lending discrimination (Mehrabi et al., 2021b); Market manipulation (Hendry et al., 2022) High: Fairness audits (Pfohl et al., 2022); Multi party comp.
+(Khalid et al., 2023) SEC (U.S.
+Securities and Exchange Commission, 2021); OCC (Office of the Comptroller of the Currency, 2020) C1: no lending discrimination benchmark; C3: no financial adversarial attack suite SPD/DI (3 to 4), $R$ (6), $F_{a}$ (10)
+Education Student privacy (Livingstone et al., 2021); Skills evasion (Pavlou et al., 2023) Low: Watermarking (Jiang et al., 2023); Fed.
+learning (Yang, 2021) UNESCO AI Rubric (UNESCO, 2021b); State Edu.
+Laws (Utah State Board of Education, 2022) C6: no minors data protocol; C7: skills evasion untested; KPI 11 lacks standard $P$ (2), $E_{h}$ (9), $U$ (11)
+[…]
+Agriculture Ownership conflicts (Bronson, 2019); Global South exclusion (Tzachor et al., 2022) Low: Geo aware RAG (Ma et al., 2023); Data contracts (Bronson, 2019) FAO (Costa et al., 2023); EU CAP (Costa et al., 2023) C4: no geo aware hallucination test; C7: IoT advisory pipelines untested $E_{h}$ (9), $P$ (2), $E_{\text{kwh}}$ (12)
+Defense Autonomous lethality (International Committee of the Red Cross, 2021); Adversarial spoofing (National Institute of Standards and Technology, 2025a) High: Red teaming (Corporation, 2024); Meaningful HITL (International Committee of the Red Cross, 2021) NATO Principles (NATO, 2024); IHL (United Nations Institute for Disarmament Research, 2025) C8: only 2/21 benchmarks score >>0; C7: absent for multi agent autonomy $R$ (6), $E_{h}$ (9), $F_{a}$ (10)
+Legal & Justice Recidivism prediction bias (Skeem and Lowenkamp, 2016); Automated decision opacity (Chesterman, 2021) Low to Med.: Explainability (LIME/SHAP) (Microsoft Research, [n.
+d.]); Bias audits (Center for Data Science and Public Policy, [n.
+d.]) EU AI Act Art.
+6 (Edwards, 2021); U.S.
+DOJ Guidelines (Podgor, 2003) C1: no recidivism specific fairness benchmark; C2: legal advice toxicity untested; C10: judicial HITL criteria absent SPD/DI (3 to 4), $E_{h}$ (9), $R$ (6), $F_{a}$ (10)
+Manufacturing Predictive maintenance failures (Hashemian, 2010); Adversarial sensor spoofing (Cao et al., 2019) Medium: Digital twins (Mihai et al., 2022); Anomaly detection (Chandola et al., 2009) IEC 62443 (ISO/IEC, 2021b); EU Machinery Reg.
+(European Union, 2024) C3: no industrial adversarial robustness suite; C4: no sensor drift benchmark; C7: multi robot coordination untested $E_{h}$ (9), $R$ (6), $E_{\text{kWh}}$ (12), $F_{a}$ (10)
 6. Discussion
@@
 Our review shows that post-GPT scholarship concentrates on five areas: safety and benchmarking, fairness and governance, explainability and transparency, privacy and security, and sectoral applications.
-The distribution is uneven.
 Bias and toxicity receive dense coverage, while deepfakes, sustainability, and system-level failure modes receive relatively less attention.
-This pattern aligns with the thematic trends we documented and with the growing prominence of multimodal and agentic systems in 2024–2025, which increases the need for evaluation that captures tool use and context.
-6.2. Implications for research
+This pattern aligns with the thematic trends we documented and with the growing prominence of multimodal and agentic systems in 2022–2025, which increases the need for evaluation that captures tool use and context.
+Closing the accountability chain
+To fully answer our primary research question, Who is Responsible?, our synthesis posits that the user must be transitioned from a passive consumer to a systemic guardrail.
+The current imbalance in GenAI governance, where 74% of organizations use GenAI but only 26% have a formal strategy, often results in the end-user inheriting unmitigated technical risks like PII leakage and hallucination.
+By triangulating the sectoral failures in Section 5 with the digital literacy gaps identified in Section 3.2.2, we propose a Symmetric Responsibility Model.
+In this model, developers are accountable for Transparency and Alignment, while users are accountable for Operational Integrity and Dissemination.
+Digital literacy, therefore, is not merely a skill but a core technical control that must be assessed alongside model-side benchmarks in Section 4.2.
+Without this user centric governance, even perfectly aligned models remain vulnerable to the last mile failure of human misinterpretation or malicious re contextualization.
+Table 10 operationalizes this model by mapping specific failure modes to the reciprocal duties of each stakeholder.
+Table 10. The Symmetric Responsibility Matrix: Mapping Failure Modes to Stakeholder Duties
+Failure Mode Developer Duty (Engineering Controls) User Duty (Operational Integrity)
+Hallucination & Misinfo Implement RAG frameworks and calibrated uncertainty reporting (KPI 9) (Wu et al., 2024). Verify factual claims via primary sources and cross modal grounding checks.
+PII & Data Leakage Deploy Differential Privacy and automated scrubbing during inference (KPI 2) (Nasr et al., 2023). Practice prompt hygiene by omitting sensitive or proprietary data from inputs.
+Adversarial Jailbreaking Conduct continuous red teaming and implement robust system guardrails (C8) (Samvelyan et al., 2024).
+Adhere to acceptable use policies and report emergent safety bypasses to developers.
+Representational Bias Perform fairness audits using intersectional metrics like SPD and DI (KPIs 3 to 4) (Kun, 2015).
+Critically appraise outputs for stereotyping and contextualize results within local norms.
+Over reliance (Automation Bias) Provide model confidence scores and explicit explanation gap warnings (KPI 8) (Strathern, 1997a).
+Maintain human in the loop (HITL) oversight and preserve independent domain expertise.
+Industry Safety Artifacts and Advanced Interpretability.
+While academic benchmarks offer standardized metrics, the operational reality of ResGenAI is increasingly defined by safety documentation from frontier labs and emerging mechanistic interpretability techniques, as synthesized in Figure 11.
+Frontier labs have established documentation protocols as de facto benchmarks for governance facing evidence, directly mapping to our C1 to C10 rubric.
+OpenAI system cards for GPT 4 and GPT 4o provide granular audits for sycophancy and PII leakage, demonstrating high maturity applications of metric validity through traceable safety refusal evidence.
+Anthropic utilizes Constitutional AI to automate alignment via RLAIF, outlining tiered safeguards that address malicious actor realism.
+Meta Llama 3 documentation introduces standardized protocols for multilingual red teaming, vital for addressing cross cultural ethical gaps.
+Google DeepMind technical reports provide empirical evidence for bias and accuracy trade offs, detailing the overcorrection mechanisms that led to observed representational failures.
+This industry evidence is complemented by the shift toward mechanistic interpretability, which maps internal neural activations to semantic concepts to offer a more rigorous safety check than behavior level prompts alone.
+Prominent techniques like Sparse Autoencoders (SAEs) isolate circuits responsible for deceptive or biased reasoning, enabling model steering to suppress harmful features during inference.
+Integrating these mechanistic audits into the RAI lifecycle addresses the black box challenge by enabling empirical verification of internal safety alignment before deployment.
+Figure 11. Integration of industry safety documentation and advanced mechanistic interpretability for governance facing evidence.
+ResGenAI sustainability via lifecycle assessment
+Sustainability in ResGenAI must transition from qualitative discourse to rigorous quantification grounded in Lifecycle Assessment (LCA) frameworks, as integrated into our governance landscape (Figure 8) and synthesized in Figure 12.
+This necessitates formalizing carbon reporting and energy benchmarks across the development cycle, including pretraining, fine tuning, and inference.
+Following our proposed lifecycle KPIs (Section 4.2.4), we advocate for reporting energy consumption and carbon equivalent emissions normalized by throughput, such as per 1K tokens generated.
+Seminal research indicates that training large scale generative models produces environmental footprints comparable to the lifetime emissions of multiple automobiles, highlighting the need for transparent tracking tools and hardware aware reporting.
+A sustainable AI ecosystem requires a dual focus on optimizing the computational efficiency of models and leveraging generative insights to advance ecological goals, ensuring that technical innovation supports long term environmental stability.
+Figure 12.
+Quantifying Responsible GenAI Sustainability via Lifecycle Assessment (LCA) in ResGenAI Governance.
+This framework formalizes carbon reporting and energy benchmarks across the pretraining, fine tuning, and inference stages, enabling a transition to rigorous quantification and supporting a sustainable AI ecosystem.
+6.2. Lessons from Industry Deployment Incidents
+While our benchmark analysis identifies risk-coverage gaps analytically, real-world deployment incidents from major AI developers offer empirical grounding.
+We synthesize prominent cases below, mapped to our rubric criteria, to demonstrate how the gaps in Tables 36 manifest in production systems.
+Alignment and safety failures (C2, C3, C8) Microsoft Bing Chat integration of GPT-4 in early 2023 produced adversarial persona adoption and emotional manipulation during extended sessions (Roose, 2023), behaviors that emerged through multi-turn prompting, precisely the adaptive adversary pattern captured by C8, for which only 2 of 21 benchmarks scored above zero.
+OpenAI GPT-4o system card documented sycophantic and manipulative voice-mode outputs 444developers.openai.com/api/docs/models/gpt-4o, illustrating subtler alignment failures beyond toxicity classifiers.
+Google’s Gemini image generator was suspended in February 2024 after producing historically inaccurate racially diverse depictions of historical figures555blog.google/products/gemini/gemini-image-generation-issue/, demonstrating how overcorrection in bias mitigation (C1) can itself produce fairness failures, an interaction that benchmarks evaluate in isolation.
+Hallucination in deployed systems (C4, C9) In Mata v.
+Avianca (2023), attorneys submitted a brief with six fabricated ChatGPT-generated case citations, resulting in judicial sanctions666caselaw.findlaw.com/court/us-dis-crt-sd-new-yor/2335142.html.
+These cases underscore that only 7 of 21 benchmarks address misinformation with documented constructs, and that the high-stakes error rate KPI ($E_{h}$, Section 4.2.4) must be operationalized in deployment settings.
+Privacy and data leakage (C6) Samsung banned internal ChatGPT use in 2023 after employees inadvertently leaked proprietary source code through conversations (Samsung Electronics, 2023), a usage-path leakage vector that no current privacy benchmark assesses.
+Nasr et al.
+(Nasr et al., 2023) demonstrated scalable extraction of memorized training data, including PII, from production ChatGPT via divergence attacks, exceeding the scope of PrivLM-Bench and PrivacyLens.
+Deepfake and media integrity (C5) A Hong Kong finance worker transferred approximately \$25 million after a video call with deepfake recreations of company executives (Chen, 2024), and AI-generated robocalls impersonating President Biden targeted the 2024 New Hampshire primary (New Hampshire Department of Justice, 2024).
+Only SHIELD provides explicit C5 evaluation, and neither incident type, social engineering or democratic interference, falls within its scope.
+AI Act enforcement (C10) The Act’s GPAI provisions, effective August 2025, require adversarial testing, incident reporting, and cybersecurity protections for systemic-risk models (European Union, 2024).
+The European AI Office’s first Code of Practice (European Commission, 2025) operationalizes the transparency and robustness requirements mapped in Table 6.
+Early compliance experience confirms our finding that organizations struggle to assemble portable evidence across governance obligations, as no single benchmark suite covers more than half of the rubric at the “strong” level.
+6.3. Trade-offs between Performance Measures
+The scoring formulation offers a transparent and reproducible summary of coverage breadth across ten criteria.
+It supports comparison within a release and over time when paired with clear documentation of constructs, metrics, and versioned artifacts.
+We emphasize that the score is a coverage indicator rather than a ranking of overall safety and must be interpreted with the governance crosswalk and with domain risk.
+The Fairness-Privacy Paradox: Reconciling Representativeness with the Right to Anonymity
+A fundamental tension emerging from our synthesis is the operational conflict between the requirement for Representativeness and the rigid Privacy mandates of the General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016).
+While mitigating algorithmic bias (Section 4.2.1) necessitates the collection and processing of high-fidelity, demographically diverse datasets to ensure equitable model performance across protected classes, privacy-preserving standards often mandate data minimization and the obfuscation of sensitive attributes.
+This creates a structural “Fairness-Privacy Paradox”: without access to granular, protected-class identifiers, developers cannot empirically verify statistical parity or conduct intersectional bias audits (Kun, 2015).
+Consequently, the legal push for anonymity under GDPR can inadvertently mask the very disparities that ResGenAI frameworks seek to eliminate.
+We posit that resolving this tension requires a transition from absolute data minimization to the adoption of privacy-enhancing technologies (PETs) such as Differential Privacy and Federated Learning that facilitate rigorous fairness auditing while maintaining individual-level anonymity .
+Critical Appraisal of Technical Controls: Synthesis and Reflection
+After looking closely at how these systems work in the real world, it’s clear there is a significant gap between what we hope these tools can do and how they actually perform when things get messy.
+The Explanation-Reliability Gap shows us that while we have many ways to explain model decisions (Dwivedi et al., 2023), these often create stories that sound good but don’t actually match what the model is doing under the hood (Strathern, 1997a).
+We have found that for big models, these explanations are mostly helpful for developers to find bugs, rather than giving regular people or regulators the true transparency they need.
+The Bias-Overcorrection Paradox is another hard lesson we’ve learned from cases like the Gemini incident.
+When we try too hard to fix fairness using simple keywords, we often end up breaking the model’s ability to be factually accurate or culturally aware.
+This tells us that fairness isn’t just a number to hit (Kun, 2015), but a deep challenge that needs to change based on where the model is being used.
+Finally, The Shallow Guardrail Problem reminds us that models can look perfectly safe on paper (Zhang et al., 2024) while still being easy to trick.
+We’ve seen that models which pass every standard toxicity test can still be pushed into strange or harmful behaviour through basic multi-turn conversations (Roose, 2023).
+The reality is that our current ”guardrails” often just polish the output without fixing the underlying issues, leaving them vulnerable to the automated attacks we see growing today (Samvelyan et al., 2024).
+6.4. Analytical Triangulation: Bridging the Evaluation Gap
+The persistent safety failures in production systems are not merely isolated technical oversights; they are symptoms of a systemic “Evaluation Gap” where current benchmarks fail to model the adaptive, iterative pressure applied by real-world adversaries.
+This section triangulates the theoretical risk taxonomy in Section 3 with the empirical deployment failures documented in Section 6.2 through the lens of our comparative assessment (Section 4.2).
+Our assessment in Table 4 reveals a stark deficit in evaluative maturity regarding Malicious Actor Realism (Criterion C8), where only 2 out of 21 surveyed benchmarks: Rainbow Teaming and ALERT provide strong coverage for adaptive pressure.
+This misalignment directly correlates with the high-profile incidents synthesized in Section 6.2, such as the adversarial persona adoption observed in Bing Chat (Roose, 2023) and the scalable data extraction vulnerabilities reported in ChatGPT (Nasr et al., 2023).
+These failures underscore that prevailing evaluation suites remain largely static and task-local, leaving them unequipped to measure the non-deterministic tactics encountered in the wild.
+Furthermore, the representational failures observed in the Gemini overcorrection incident illustrate a critical breakdown in Governance Alignment (C10), where technical alignment efforts are decoupled from the broader sociotechnical context of deployment.
+By triangulating these findings, we posit that a ResGenAI operating model must transition from one-shot static probes toward continuous, adaptive evaluation frameworks that mirror the malicious-actor realism and distribution shifts inherent in generative ecosystems.
+6.5. Limitations of the Study
+Like any survey-based study, this work has several limitations.
+First, the review follows a pre-specified protocol and focuses primarily on the period from November 2022 to December 2025, with selective inclusion of foundational works for definitional clarity.
+While this post-ChatGPT window captures the most rapid phase of GenAI deployment, it may exclude emerging studies published after the search cutoff or earlier work that did not explicitly use contemporary terminology.
+Second, our synthesis relies on publicly available papers, dataset cards, benchmark documentation, and governance artifacts.
+As a result, proprietary evaluations, internal red-teaming practices, or unpublished safety mitigations may be under-represented.
+Documentation quality also varies across benchmarks and frameworks, which may affect the completeness of criterion-level scoring.
+Third, rubric weights are assigned equally by default, and criterion coverage is determined through structured reading of sources followed by discussion and adjudication in cases of disagreement.
+Although we conducted sensitivity analyses to assess the robustness of equal-weight assumptions, alternative weighting schemes reflecting sector-specific risk priorities could yield different relative rankings.
+Fourth, the scoring procedure captures breadth of coverage rather than depth or real-world effectiveness.
+A benchmark receiving a high rubric score indicates multi-risk awareness and governance alignment, but does not guarantee empirical robustness under adversarial deployment conditions.
+In this sense, the rubric reflects documented capability rather than independently verified performance.
+Fifth, the review synthesizes heterogeneous evidence types, including empirical evaluations, conceptual frameworks, and regulatory analyses.
+This diversity strengthens thematic coverage but precludes formal meta-analysis or quantitative effect-size aggregation.
+Our conclusions therefore emphasize structural patterns rather than causal claims.
+Sixth, governance alignment is assessed at the level of plausible evidentiary support rather than legal compliance.
+Mapping benchmarks to instruments such as the EU AI Act or NIST AI RMF indicates potential contribution to audit evidence, but does not constitute certification or conformity assessment.
+Finally, the rapid evolution of multimodal and agentic systems means that new failure modes may emerge after publication.
+Continuous updates to evaluation suites, governance frameworks, and deployment practices may shift the relative landscape described here.
+Replication of our screening and scoring is facilitated by the availability of documented sources, explicit inclusion criteria, and the published rubric; however, interpretive judgment remains inherent in structured qualitative synthesis.
+6.6. Research directions
 The rubric-based analysis in Section 4.2 indicates that commonly used suites provide strong evidence for toxicity and bias, and more limited coverage for privacy, deepfakes, and system-level failures.
@@
 Future research should target evaluations that model distribution shifts, adaptive adversaries, and multi-agent coordination with external tools.
-6.2.1. Implications for practice and policy
-For practitioners and policy makers, Tables 5 - 8 map benchmarks to risk families and a governance crosswalk provide a starting point for evidence collection.
+For practitioners and policy makers, Tables 3 to 8 map benchmarks to risk families and a governance crosswalk provide a starting point for evidence collection.
 These tables also reveal gaps that are relevant for conformity assessment under emerging regimes.
 Applied teams can use the matrix to assemble a minimal set of evaluations that cover privacy, robustness, transparency, logging, and post-market monitoring, and then plan additions for multimodal deepfake risk and adaptive red teaming as required by sectoral context.
-6.3. Evidence portability and benchmarking
-The scoring formulation offers a transparent and reproducible summary of coverage breadth across ten criteria.
-It supports comparison within a release and over time when paired with clear documentation of constructs, metrics, and versioned artifacts.
-We emphasize that the score is a coverage indicator rather than a ranking of overall safety and must be interpreted with the governance crosswalk and with domain risk.
-6.3.1. Limitations and threats to validity
-This survey follows a pre-specified protocol, focusing on the period from November 2022 to August 2025, and incorporates selected foundational works to ensure definitional clarity.
-Nevertheless, decision regarding selection and synthesis may introduce bias.
-Reliance of public papers, dataset cards, and documentation implies that certain capabilities may be under-reported or reported with delay.
-The rubric weights are assigned equally by default and criterion coverage is determined through structured reading of sources, followed by discussion and adjudication in case of disagreements.
-Replication of our screening and scoring is facilitated by the availability of the documented sources and explicit inclusion rules.
-6.4. Research directions.
 We highlight five priorities.
-First, extend benchmarks to capture system-level failures in tool-using and multi-agent settings, especially around retrieval, planning, and actuation (e.g., API calls, code execution, or robotic steps).
+First, we need to extend benchmarks to capture system-level failures in tool-using and multi-agent settings, especially around retrieval, planning, and actuation (e.g., API calls, code execution, or robotic steps).
 Second, fold privacy and provenance checks into routine testing, including membership-inference and training-data extraction probes, plus audits of logging sufficiency.
@@
 Fourth, shift from static to adaptive evaluation with undisclosed items, targeted red teaming, and tests under distribution shift.
-Fifth, report sustainability alongside accuracy and robustness—energy, emissions, and compute-normalized efficiency—to enable compute-proportional evaluation and responsible deployment.
+Fifth, report sustainability alongside accuracy and robustness energy, emissions, and compute-normalized efficiency to enable compute-proportional evaluation and responsible deployment.
 These priorities directly address the risk-surface gaps surfaced in this survey.
-6.4.1. Guidance for practitioners
-Teams putting responsible GenAI into practice can follow a simple loop.
+Teams putting ResGenAI into practice can follow a simple loop.
 First, scope risks by mapping planned features to common risk families (e.g., bias, toxicity, privacy, and security).
@@
 Then, add context: turn on logging, enforce dataset/model versioning, and keep lightweight documentation so decisions and results remain traceable.
-Before launch—and later during post-market monitoring—exercise adaptivity with red teaming and a few undisclosed checks to catch regressions.
-Finally, track a small set of KPIs across the lifecycle: calibration, error rates in high-stakes flows, audit cadence, inclusivity indicators, and energy/emissions.
+Before launch and later during post-market monitorin, there is need to exercise adaptivity with red teaming and a few undisclosed checks to catch regressions.
+Finally, track a small set of KPIs across the lifecycle: calibration, error rates in high-stakes flows, audit cadence, exclusivity indicators, and energy/emissions.
 This routine keeps evaluation aligned with governance expectations and makes evidence easier to reuse when audits occur.
+6.7. Systemic Vulnerabilities and Blind Spots in ResGenAI
+While GenAI capabilities have accelerated, structural blind spots persist across technical and sociotechnical domains (Kenny et al., 2021).
+This review identifies several systemic gaps that currently hinder universal assurance:
+  • RLHF and Alignment Transparency: Alignment advances, such as RLHF and Constitutional AI, have shifted emphasis from raw capability to controllability.
+However, the documentation of these alignment pathways is often inconsistent, which complicates the auditing of model refusal behaviour and evidence portability across versions (of Standards and, NIST, NIST).
+  • Opaque Proprietary Datasets: Responsible stewardship demands adherence to consent and privacy regimes.
+However, many frontier models rely on web-scale datasets where provenance, licensing, and curation criteria remain undocumented.
+Truly representative datasets are often infeasible due to data scarcity or legal limits, leading to persistent data leakage and memorization risks (Nasr et al., 2023).
+  • Non-Stationary Safety: Finally, safety is inherently non-stationary; model updates, distribution shifts, and evolving adversarial tactics induce drift.
+Static benchmarks rarely capture these dynamic failure modes , necessitating a shift toward the continuous monitoring and adaptive evaluation frameworks discussed in Section Section 4.2 .
 7. Conclusion
-This survey examined responsible GenAI through the lens of governance needs and practical evaluation.
-We organized the literature into risk families and assurance qualities and introduced a two block rubric that distinguishes risk coverage (C1 to C8) from assurance and governance qualities (C9 to C10).
-From this review article, three findings stand out.
-First, coverage is strongest for bias and toxicity and weaker for privacy, deepfakes and provenance, and system level failure in tool using or agentic settings.
-Second, most evaluations remain static and task local, which limits evidence portability for governance and audit.
-Third, documentation and metric validity vary substantially, which complicates interpretation across releases and domains.
-These gaps are significant because contemporary systems are multimodal, often connected to tools and data stores, and operate under distribution shift.
-We outlined a research agenda that calls for evaluations addresing system level failure, privacy leakage and provenance, deepfake and media integrity risk, and adaptive adversaries.
-We further recommended stronger reporting of constructs, metrics, uncertainty, and versioned artifacts, to enhance the portability of the results across audits and over time.
-For practitioners, we provided a sequence that links product risks to a minimal set of tests, adds assurance artifacts such as documentation and logging, and incorporates red teaming and post release monitoring.
-Collectively, these steps align engineering practice with governance expectations without prescribing a single benchmark.
-This review has several limitations.
-It focuses on the period from November 2022 through August 2025 with selective inclusion of foundational work, relies on public documentation, and uses equal weights by default for the composite index.
-We mitigated subjectivity through a prespecified protocol, independent scoring, adjudication, and the release of a codebook, yet some capabilities may be under reported or reported with delay.
-The governance crosswalk is intended as plausible evidentiary support rather than a statement of legal compliance.
-Responsible GenAI requires progress on methods and on evidence.
-Benchmarks should move beyond static probes toward evaluations that reflect real operating conditions, including adaptivity, provenance, and end to end behavior.
-Governance frameworks should in turn recognize evidence that is documented, reproducible, and linked to explicit constructs.
-By connecting risk families to concrete tests and by foregrounding assurance qualities, this survey provides a path for researchers and practitioners to close capability gaps, to increase trust, and to support safe and accountable deployment of GenAI systems across domains.
+This survey comprehensively synthesized the landscape of ResGenAI by evaluating the alignment between global governance mandates and concrete engineering practices.
+Our synthesis findings revealed a significant evaluation imbalance in current practices, for example, while technical coverage is mature for surface level AI harms like bias and toxicity, the ecosystem remains acutely vulnerable to privacy leakage, deepfake risks, and the system level failure modes inherent in emergent agentic and tool using environments.
+Furthermore, we identified that the prevalence of static, task-specific benchmarks severely hinders evidence for institutional audits and regulatory compliance.
+To bridge this evaluation gap, we proposed a research agenda that prioritizes adaptive, multimodal evaluation and provides a practical stewardship to operationalize technical controls through lifecycle KPIs and standardized audit artifacts.
+By formalizing the symmetric responsibility model and the ResGenAI audit loop, this work establishes the first integrated framework for bridging technical evaluation with regulatory needs, providing a safe, transparent, and sustainable deployment of GenAI models across domains.
 Philosophy and the Way Forward
@@
 Adaptability in AI must go beyond technical performance to include ethical flexibility, ensuring systems align with human values.
-RAI emphasizes not just technical capabilities but also the ethical imperatives of transparency, fairness, and accountability.
+ResGenAI emphasizes not just technical capabilities but also the ethical imperatives of transparency, fairness, and accountability.
 ”None of us is as responsible as all of us.” — Authors of this paper
-Achieving RAI requires a collective effort among developers, policymakers, and users to mitigate bias, prioritize privacy, and ensure accountability.
+Achieving ResGenAI requires a collective effort among developers, policymakers, and users to mitigate bias, prioritize privacy, and ensure accountability.
 Collaboration across disciplines and sectors is essential to create equitable and ethical AI systems that serve diverse societal needs.
@@
 Redefining benchmarks to incorporate ethical metrics is critical for guiding the development of AI systems that perform both technically and ethically.
-Such benchmarks must evaluate transparency, safety, and bias mitigation alongside traditional performance metrics.
 The data-driven nature of AI presents persistent challenges, as models often inherit biases and privacy risks from training data.
 These biases, embedded in AI outputs, require rigorous mitigation strategies to distinguish between useful correlations and harmful patterns.
-While creating a perfectly unbiased AI system is unattainable due to the evolving nature of societal values (Ai et al., 2018), continuous evaluation and adaptation of AI models are essential to maintain alignment with ethical standards.
+While creating a perfectly unbiased AI system is unattainable due to the evolving nature of societal values, continuous evaluation and adaptation of AI models are essential to maintain alignment with ethical standards.
 Human decision-making is influenced by moral reasoning and context, while AI operates strictly within its programmed algorithms.
 This underscores the importance of fostering AI literacy (Long and Magerko, 2020) among developers and users to ensure responsible deployment and use.
-To implement RAI effectively, AI systems must prioritize transparency, fairness, and adaptability.
+To implement ResGenAI effectively, AI systems must prioritize transparency, fairness, and adaptability.
 They should incorporate ethical reasoning and evolve alongside societal norms.
 As we advance toward Artificial General Intelligence, the true measure of AI should extend beyond technical proficiency to include ethical adaptability (Blili-Hamelin et al., 2024).
-The pursuit of RAI is challenging yet indispensable.
-Ongoing advancements in RAI frameworks, coupled with a growing emphasis on AI ethics, offer a path forward.
+The pursuit of ResGenAI is challenging yet indispensable.
+Ongoing advancements in ResGenAI frameworks, coupled with a growing emphasis on AI ethics, offer a path forward.
 Prominent voices in the field, including Jensen Huang and Sam Altman, emphasize the need for safe and ethical AI.
-By aligning technical innovation with ethical principles, we can build AI systems that uphold accountability, foster trust, and adapt responsibly to the complexities of a dynamic world.
 8. Acknowledgement
 This work received no external funding and was undertaken purely out of a passion for learning and advancing knowledge, and safe and responsible deployment of AI, for the betterment of society.
-Appendix
-Appendix A Full Search Strings and Filters
-To enable exact replication, we report verbatim queries and filters for each source.
-For all databases, we limited the date range to November 2022–August 2025 and restricted results to English.
-Field constraints (title/abstract/keywords) were used where supported; otherwise, queries were applied to full text.
+Appendix A Appendix
+A.1. Full Search Strings and Filters
 Below we list the precise strings and any database-specific options.
-ACM Digital Library.
-Fields: Title/Abstract/Keywords. Years: 2022–2025. Document types: Journal, Proceeding Article.
+ACM Digital Library Fields: Title/Abstract/Keywords. Years: 2022–2025. Document types: Journal, Proceedings.
 ("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR accountability
@@
  OR multimodal OR "natural language processing" OR "computer vision")
-IEEE Xplore.
-Metadata search (Abstract + Index Terms). Years: 2022–2025. Content type: Journals, Conferences.
+IEEE Xplore Metadata search (Abstract + Index Terms). Years: 2022–2025. Content type: Journals, Conferences.
 (("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR accountability
@@
  "generative AI" OR multimodal OR "vision-language" OR "large language model*"))
-SpringerLink / ScienceDirect.
-Title/Abstract/Keywords. Years: 2022–2025. Subject areas: Computer Science, AI.
-("Responsible AI" OR "AI governance" OR "AI safety" OR fairness OR transparency
- OR accountability OR privacy) AND ("generative AI" OR LLM OR LVLM OR multimodal)
-Web of Science.
-Topic search (TS=Title, Abstract, Keywords). Years: 2022–2025.
+Web of Science Topic search (TS=Title, Abstract, Keywords). Years: 2022–2025.
 TS=("Responsible AI" OR "Ethical AI" OR "AI safety" OR fairness OR transparency
  OR accountability OR privacy) AND TS=("generative AI" OR LLM OR LVLM OR multimodal)
-DBLP.
-Title search; manual verification of scope.
-("Responsible AI" OR "Ethical AI" OR "AI safety" OR fairness OR transparency
- OR accountability OR privacy) AND (LLM OR LVLM OR "generative AI" OR multimodal)
-arXiv.
-Categories: cs.AI, cs.LG, cs.CV, cs.CL. Date filter applied via API.
+arXiv Categories: cs.AI, cs.LG, cs.CV, cs.CL. Date filter applied via API.
 (all:"Responsible AI" OR all:"AI safety" OR all:"fairness" OR all:"transparency"
@@
  OR all:"vision-language" OR all:"multimodal")
-Google Scholar.
-Custom range: Nov 2022–Aug 2025; “include citations” disabled during screening.
-"Responsible AI" OR "AI governance" OR "AI safety" fairness transparency accountability
- privacy "generative AI" LLM LVLM multimodal
-For each included study we recorded venue and year, dominant modality, problem area, method type, risks and controls addressed, datasets or artifacts, evaluation facets and metrics, and the type of evidence advanced.
-We adapted appraisal criteria from established CS/SE review checklists to rate method transparency, data/source adequacy, discussion of validity threats, and presence of reproducibility signals.
-Each study received qualitative notes and an ordinal rating (Low/Medium/High) per criterion; we also assigned an evidence-strength tag that informed interpretation during synthesis.
-Figure A.1.
-Taxonomy diagram categorizing the core concepts of Responsible AI (RAI) into principal subsets: Explainable AI (XAI), Ethical Considerations, Sector Applications, Advanced Models, Privacy and Security, and Regulatory Frameworks.
-Each subset is further dissected into detailed elements such as Transparency, Bias Mitigation, Healthcare, Generative Models, and Data Anonymization, illustrating the flow from general themes to specific methodologies within RAI:
-Figure A.2.
-Framework for Responsible Generative AI emphasizing transparency, fairness, accountability, privacy, and bias mitigation through the integration of data, models, regulations, and user-centric practices
-Appendix B Codebook for rubric C1 to C10
-Use. Score only explicit tasks and metrics. If a task maps to multiple risks, assign it to the primary stated objective to avoid double counting.
-Evidence priority. Official paper and docs; then dataset or model cards and maintained leaderboards; then widely cited third party reports.
+Scopus Field search (TITLE-ABS-KEY). Years: 2022–2025.
+TITLE-ABS-KEY("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR
+accountability OR fairness OR "bias mitigation" OR privacy OR equity OR "open science")
+AND TITLE-ABS-KEY("generative AI" OR "large language model*" OR LLM OR LVLM OR
+"vision-language" OR multimodal OR "natural language processing" OR "computer vision")
+A.2. Codebook for rubric C1 to C10
+Use Score only explicit tasks and metrics.
+If a task maps to multiple risks, assign it to the primary stated objective to avoid double counting.
+Evidence priority Official paper and docs; then dataset or model cards and maintained leaderboards; then widely cited third party reports.
 C1 Bias and discrimination
@@
 Scoring: 0 none; 1 informal references; 2 explicit control mapping or evidentiary artifacts.
-Appendix C Scoring protocol, reliability, and sensitivity
-Procedure. Two reviewers score each criterion using this codebook; disagreements resolved by a third reviewer. Publish the codebook, score sheets, and evidence extracts.
-Reliability. Report Cohen’s $\kappa$ per criterion and the mean $\bar{\kappa}$; target $\bar{\kappa}\geq 0.60$.
-Indices. Report the risk coverage index $S_{\mathrm{cov}}$, coverage rate $R$, assurance index $S_{\mathrm{assur}}$, and the overall normalized index $S$ as defined in Section 4.2.1.
-Sensitivity. Recompute $S$ under alternate weights that prioritize governance critical risks and report Spearman rank correlation with the equal weight baseline.
-Non compliance note. The governance crosswalk indicates plausible evidentiary support, not legal compliance.
-Appendix D Glossary (Key Terms & Acronyms)
-RAI:
-Responsible AI: practices and governance to ensure fairness, privacy, safety, transparency, and accountability across the AI lifecycle.
-GenAI:
-Generative AI: models that synthesize content (text, image, audio, code) from learned distributions.
-LLM:
-Large Language Model (e.g., GPT, Llama): transformer-based models trained on large text corpora.
-VLM:
-Vision–Language Model: multimodal models handling text+image inputs/outputs (e.g., BLIP-2, LLaVA).
-RAG:
-Retrieval-Augmented Generation: generation conditioned on retrieved external context.
-Agentic AI:
-Systems where models plan, select tools, act, and maintain memory; may be multi-agent.
-XAI:
-Explainable AI: methods to make model behavior intelligible (ante-hoc/by-design or post-hoc).
-LIME:
-Local Interpretable Model-agnostic Explanations: local surrogate explanations.
-SHAP:
-SHapley Additive exPlanations: additive attributions based on cooperative game theory.
-Grad-CAM:
-Gradient-weighted Class Activation Mapping: visual saliency for CNN/VLM vision backbones.
-Calibration:
-Agreement between predicted confidence and empirical accuracy.
-ECE:
-Expected Calibration Error: summary gap between confidence and accuracy across bins.
-Brier score:
-Mean squared error of probabilistic predictions vs. outcomes (lower is better).
-SPD:
-Statistical Parity Difference: $\Pr(\hat{y}=1|A=a)-\Pr(\hat{y}=1|A=b)$ (group outcome gap).
-DI:
-Disparate Impact: $\Pr(\hat{y}=1|A=a)/\Pr(\hat{y}=1|A=b)$ (parity ratio; use $\min(\mathrm{DI},1/\mathrm{DI})$).
-Equalized odds:
-Parity of TPR/FPR across protected groups (conditional on true label).
-Robustness:
-Performance stability under noise, distribution shift, or adversarial perturbations.
-Red teaming:
-Structured adversarial testing to elicit failures or unsafe behavior.
-Benchmark:
-Standardized tasks/metrics for evaluation (e.g., safety, bias, privacy).
-KPI:
-Key Performance Indicator: operational metric tracked across the lifecycle (e.g., SPD, ECE, kWh).
-DP:
-Differential Privacy: formal privacy guarantee bounding information leakage about individuals.
-Membership inference:
-Attack inferring whether a record was in training data.
-Data extraction:
-Recovery of memorized training snippets from a model.
-PII:
-Personally Identifiable Information: data that identifies an individual.
-Provenance:
-Traceability of data/model/content origin; includes watermarking and content credentials.
-Deepfake:
-Synthetic media that mimics real persons/events (face/voice/image/video).
-System-level failure:
-Failures emerging from tool use, planning, memory, or multi-agent interaction.
-Governance crosswalk:
-Mapping of technical evidence (benchmarks/KPIs) to regulatory obligations.
-NIST AI RMF:
-U.S. National Institute of Standards and Technology AI Risk Management Framework.
-EU AI Act:
-European Union regulation with risk tiers and obligations (robustness, transparency, monitoring).
-ISO/IEC 42001:
-AI management system standard for governance and operational controls.
-SAIF:
-Secure AI Framework (Google): security practices for AI systems.
-AIF360:
-IBM AI Fairness 360: fairness metrics and mitigation toolkit.
-Fairlearn:
-Microsoft toolkit for fairness assessment and mitigation.
-Model card:
-Standardized documentation of a model’s intended use, data, metrics, and limits.
-Datasheet:
-Documentation for datasets (provenance, consent/licensing, collection, intended use).
-Post-market monitoring:
-Ongoing evaluation and incident response after deployment.
-Risk families (C1–C10):
-Rubric criteria: C1 Bias/Discrimination, C2 Toxicity, C3 Security/Robustness, C4 Mis/Disinformation, C5 Deepfakes/Media Integrity, C6 Privacy, C7 System-level Failure, C8 Adversarial/Red-team Realism, C9 Metric Validity/Documentation, C10 Governance Alignment.
-Responsible AI (RAI). A paradigm that embeds fairness, accountability, transparency, privacy, and safety across the AI lifecycle (design, training, deployment, and governance).
-Generative AI (GenAI).
-Models that synthesize novel content (text, image, audio, multimodal) from learned data distributions, including large language models (LLMs), diffusion models, and vision–language models (VLMs).
-Bias. Systematic and unfair deviations in model predictions across demographic groups, often stemming from imbalanced or unrepresentative training data.
-Fairness. The principle of ensuring equitable outcomes across protected groups, commonly measured through:
-$\text{Statistical Parity Difference (SPD)}=P(\hat{y}=1\mid A=a)-P(\hat{y}=1\mid A=b),$
-$\text{Disparate Impact (DI)}=\frac{P(\hat{y}=1\mid A=a)}{P(\hat{y}=1\mid A=b)}.$
-Hallucination. Model outputs that are factually incorrect, unsupported by evidence, or inconsistent with the input context.
-Explainability (XAI). Methods that render black-box model behavior interpretable to humans (e.g., SHAP, LIME, attention maps), enabling accountability and compliance.
-Privacy. Safeguarding sensitive data against unauthorized exposure, often operationalized via techniques such as Differential Privacy (DP) and Federated Learning (FL).
-Robustness. The ability of models to maintain reliable performance under distribution shifts, adversarial perturbations, or noisy inputs.
-Toxicity. Harmful, offensive, or unsafe outputs produced by generative models, typically evaluated through specialized benchmarks.
-Deepfakes / Media Integrity. Synthetic or manipulated media generated by GenAI that can mislead or cause societal harm; a central integrity and governance concern.
-Governance Frameworks. Formal standards and policies (e.g., NIST AI RMF, EU AI Act, ISO/IEC 42001, OECD AI Principles) that regulate responsible development and deployment.
-Key Performance Indicators (KPIs). Quantifiable measures that assess AI responsibility and compliance. For example, calibration can be measured by the Expected Calibration Error (ECE):
-$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\Big|\text{acc}(B_{m})-\text{conf}(B_{m})\Big|,$
-where $B_{m}$ denotes prediction bins, $\text{acc}(B_{m})$ the empirical accuracy, and $\text{conf}(B_{m})$ the mean confidence.
-Agentic AI. Generative systems with planning, memory, and tool-use capabilities that enable autonomous decision-making but also introduce new safety and governance challenges.
+Appendix B Scoring protocol, reliability, and sensitivity
+Procedure Two reviewers independently scored each criterion using this codebook.
+Evidence was drawn from official papers, public documentation, and widely cited usage reports, following the priority ordering specified in the codebook.
+Disagreements were resolved through structured discussion and, where needed, adjudication by a third reviewer.
+The codebook and scoring definitions are given below:
+Reliability Scoring followed a two-reviewer protocol with codebook-guided independent assessment (Appendix A.2) and discussion-based adjudication for disagreements.
+A third reviewer was consulted when consensus could not be reached.
+This structured qualitative approach prioritizes transparency and traceability of judgments over formal inter-rater statistics, which is appropriate given the interpretive nature of rubric-based benchmark assessment in survey research.
+The codebook, scoring definitions, and adjudication procedure are documented to enable replication Indices Report the risk coverage index $S_{\mathrm{cov}}$, coverage rate $R$, assurance index $S_{\mathrm{assur}}$, and the overall normalized index $S$ as defined in Section 4.2.1.
+Robustness and Sensitivity Analysis To verify that our conclusions do not depend on arbitrary weighting, we performed a Sensitivity Analysis of the composite scoring index S.
+While the primary results (Table 4) utilize equal weighting across criteria C1 to C10, we recomputed the scores under alternate weights that prioritized high-stakes risks (e.g., C3 Security and C6 Privacy).
+We observed a high Spearman rank correlation ($\rho>0.85$) between the equal-weight baseline and the weighted models, indicating that the comparative breadth of benchmarks is a stable indicator of evaluative maturity rather than an artifact of the scoring scheme.
+Non compliance note The governance crosswalk indicates plausible evidentiary support, not legal compliance.
+B.1. Glossary (Key Terms & Acronyms)
+Table 1. Glossary of key terms and acronyms used in this survey.
+Term / Acronym Category Concise definition
+RAI Acronym Responsible AI: fairness, privacy, safety, transparency, accountability across the AI lifecycle.
+GenAI Acronym Generative AI: models that synthesize content (text/image/audio/code) from learned distributions.
+LLM Acronym Large Language Model (e.g., GPT, Llama): transformer-based models trained on large text corpora.
+VLM Acronym Vision–Language Model: multimodal models handling text+image inputs/outputs (e.g., BLIP-2, LLaVA).
+LMM Acronym Large Multimodal Model: large-scale models spanning multiple modalities (text, vision, audio).
+RAG Acronym Retrieval-Augmented Generation: generation conditioned on retrieved external context.
+Agentic AI Term Systems where models plan, select tools, act, and maintain memory; may be multi-agent.
+CoT Acronym Chain of Thought: intermediate reasoning steps used to solve tasks.
+RLHF Acronym Reinforcement Learning from Human Feedback: preference-based alignment fine-tuning.
+PEFT Acronym Parameter-Efficient Fine-Tuning: adaptation with limited trainable parameters (e.g., LoRA).
+HITL Acronym Human-in-the-Loop: human oversight/feedback integrated into development or operations.
+TEVV Acronym Testing, Evaluation, Verification, and Validation: lifecycle assurance practices.
+XAI Acronym Explainable AI: methods to make model behavior intelligible (ante-hoc or post-hoc).
+LIME Acronym Local Interpretable Model-agnostic Explanations: local surrogate explanations.
+SHAP Acronym SHapley Additive exPlanations: additive attributions based on game theory.
+Grad-CAM Term Gradient-weighted Class Activation Mapping: visual saliency for CNN/VLM backbones.
+Calibration Term Agreement between predicted confidence and empirical accuracy.
+ECE Acronym Expected Calibration Error: bin-wise gap between confidence and accuracy (summary).
+Brier score Term Mean squared error of probabilistic predictions vs. outcomes (lower is better).
+Bias Term Systematic and unfair deviations across demographic groups, often from unrepresentative data.
+Fairness Term Equitable outcomes across protected groups; often assessed via SPD/DI and related metrics.
+Equalized odds Term Parity of TPR/FPR across protected groups (conditional on true label).
+Robustness Term Performance stability under noise, distribution shift, or adversarial perturbations.
+Red teaming Term Structured adversarial testing to elicit failures or unsafe behavior.
+Benchmark Term Standardized tasks/metrics for evaluation (e.g., safety, bias, privacy).
+KPI Acronym Key Performance Indicator: operational metric tracked across the lifecycle (e.g., SPD, ECE, kWh).
+Hallucination Term Outputs that are incorrect, unsupported by evidence, or inconsistent with input/context.
+Toxicity Term Harmful, offensive, or unsafe outputs; typically evaluated via specialized benchmarks.
+Deepfake Term Synthetic media mimicking real persons/events (face/voice/image/video).
+Provenance Term Traceability of data/model/content origin; includes watermarking and content credentials.
+System-level failure Term Failures from tool use, planning, memory, or multi-agent interaction.
+Governance crosswalk Term Mapping technical evidence (benchmarks/KPIs) to regulatory obligations.
+Privacy Term Protecting sensitive data from exposure; operationalized via DP, FL, and other PETs.
+DP Acronym Differential Privacy: formal guarantee bounding information leakage about individuals.
+FL Acronym Federated Learning: training across clients without centralizing raw data.
+PETs Acronym Privacy-Enhancing Technologies: methods to reduce/mitigate privacy risks.
+Membership inference Term Attack inferring whether a record was in a model’s training data.
+Data extraction Term Recovery of memorized training snippets from a model.
+PII Acronym Personally Identifiable Information: data that identifies an individual.
+Model card Term Standardized documentation of a model’s intended use, data, metrics, and limits.
+Datasheet Term Documentation for datasets (provenance, licensing/consent, collection, intended use).
+Post-market monitoring Term Ongoing evaluation and incident response after deployment.
+Risk families (C1–C5) Term Rubric: C1 Bias, C2 Toxicity, C3 Security/Robustness, C4 Mis/Disinfo, C5 Deepfakes.
+Risk families (C6–C10) Term Rubric: C6 Privacy, C7 System failure, C8 Red-team realism, C9 Metric validity/docs, C10 Governance alignment.
+NIST AI RMF Acronym U.S. National Institute of Standards and Technology AI Risk Management Framework.
+EU AI Act Acronym EU regulation defining risk tiers and obligations (e.g., transparency, monitoring).
+ISO/IEC 42001 Acronym AI management system standard for governance and operational controls.
+SAIF Acronym Secure AI Framework (Google): security practices for AI systems.
+AI RMF Acronym Artificial Intelligence Risk Management Framework (shorthand for NIST AI RMF).
+GDPR Acronym General Data Protection Regulation (EU).
+HIPAA Acronym Health Insurance Portability and Accountability Act (USA).
+FDA Acronym Food and Drug Administration (USA).
+OCC Acronym Office of the Comptroller of the Currency (USA).
+SEC Acronym Securities and Exchange Commission (USA).
+PRISMA Acronym Preferred Reporting Items for Systematic Reviews and Meta-Analyses.
+SOTA Acronym State-of-the-Art.
+IHL Acronym International Humanitarian Law.
+NIST Acronym National Institute of Standards and Technology.
+NATO Acronym North Atlantic Treaty Organization.
+UNESCO Acronym UN Educational, Scientific and Cultural Organization.
+SAG-AFTRA Acronym Screen Actors Guild–American Federation of Television and Radio Artists.
+WGA Acronym Writers Guild of America.
+AIF360 Acronym IBM AI Fairness 360: fairness metrics and mitigation toolkit.
+Fairlearn Acronym Microsoft toolkit for fairness assessment and mitigation.
+Table 2. Core synthesis of the literature review: theme, citation, year, and focus.
+Theme Citation Year Focus
+Safety/Benchmarks (Liu et al., 2024) 2024 Multimodal benchmark to assess safety risks in vision–language models.
+Safety/Benchmarks (Zhang et al., 2024) 2024 General safety benchmark for LLMs across risk categories.
+Safety/Benchmarks (Gehman et al., 2020) 2020 Toxic degeneration measurement using curated prompts to stress test LMs.
+Safety/Benchmarks (Li et al., 2024a) 2024 Hierarchical, comprehensive safety benchmark for language models.
+Safety/Benchmarks (Mazeika et al., 2024) 2024 Standardized evaluation for red teaming and refusal robustness.
+Safety/Benchmarks (Samvelyan et al., 2024) 2024 Diverse adversarial prompt generation for automated safety stress testing.
+Safety/Benchmarks (Wang et al., 2023a) 2023 Broad audit of trustworthiness, including safety and robustness, for GPT models.
+Safety/Benchmarks (Tedeschi et al., 2024) 2024 Benchmarks red teaming effectiveness for assessing safety responses.
+Fairness/Governance (Mittal et al., 2024) 2024 Dataset design practices aligned with fairness, privacy, and regulatory norms.
+Fairness/Governance (Rodolfa et al., 2021) 2021 Minimal fairness–accuracy tradeoff in public policy ML tasks.
+Fairness/Governance (Chouldechova, 2017) 2017 Disparate impact and fairness criteria in recidivism risk prediction.
+Fairness/Governance (Barocas and Selbst, 2016) 2016 How big-data practices can create disparate impact and legal risk.
+Fairness/Governance (Pessach and Shmueli, 2022) 2022 Survey of algorithmic fairness methods and evaluation across domains.
+Fairness/Governance (Kaur et al., 2022b) 2022 Trustworthy AI principles and technical enablers for deployment.
+Fairness/Governance (Dignum, 2019) 2019 Responsible AI framework grounded in ethics and governance.
+Fairness/Governance (Laux et al., 2024) 2024 Trustworthiness concepts in relation to EU AI Act compliance.
+Explainability/XAI (Ribeiro et al., 2016b) 2016 LIME: model-agnostic local explanations of predictions.
+Explainability/XAI (Selvaraju et al., 2017) 2017 Grad-CAM for visual explanations in CNNs.
+Explainability/XAI (Arrieta et al., 2020) 2020 Concepts, taxonomies, and challenges in explainable AI.
+Explainability/XAI (Dwivedi et al., 2023) 2023 Review of XAI ideas/techniques with practical guidance.
+Explainability/XAI (Saeed and Omlin, 2023) 2023 Meta-survey of XAI challenges and future directions.
+Explainability/XAI (van der Waa et al., 2021) 2021 Rule-based vs. example-based explanations: fidelity and user value.
+Explainability/XAI (Mosca et al., 2022) 2022 Review of SHAP-based explanation methods for NLP.
+Explainability/XAI (Nauta et al., 2023) 2023 Systematizes quantitative evaluation protocols for explanations.
+Privacy/Security (Dwork, 2006) 2006 Differential privacy as a formal framework for privacy protection.
+Privacy/Security (McMahan et al., 2017) 2017 Federated learning for decentralized training with privacy benefits.
+Privacy/Security (Nasr et al., 2023) 2023 Large-scale extraction of memorized training data from LMs.
+Privacy/Security (Mattern et al., 2023) 2023 Membership inference via neighborhood comparison against LMs.
+Privacy/Security (He et al., 2025a) 2025 Label-only membership inference attacks for pre-trained LMs.
+Privacy/Security (Li et al., 2024c) 2024 Benchmark for privacy risks in LMs at multiple levels.
+Privacy/Security (Mireshghallah et al., 2024) 2024 Privacy leakage tests using contextual-integrity evaluations.
+Privacy/Security (Staab et al., 2024) 2024 Assesses large models for anonymization and de-identification tasks.
+LLM Safety (Bai et al., 2022) 2022 Constitutional AI alignment to reduce harmful outputs.
+LLM Safety (Weidinger et al., 2021) 2021 Maps ethical/social risks of LMs and mitigation avenues.
+LLM Safety (Huang et al., 2024b) 2024 Positions trustworthiness dimensions and evaluation needs for LMs.
+LLM Safety (Lin et al., 2024) 2024 Survey of debiasing and hallucination-reduction techniques.
+LLM Safety (Wang et al., 2023c) 2023 ChatGPT applications, limitations, risks, and safeguards overview.

### v7

@@
-Who is Responsible? Data, Models, or Regulations, A Comprehensive Survey on Responsible Generative AI for a Sustainable Future
+Who is Responsible? The Data, Models, Users or Regulations? A Comprehensive Survey on Responsible Generative AI for a Sustainable Future
 Abstract.
@@
 We conduct a PRISMA-guided review of 232 studies (November 2022–December 2025), spanning large language models, vision-language models, diffusion models, and agentic pipelines.
-We make four contributions: (1) the first survey bridging governance principles, technical evaluation, and domain deployment across all four system types; (2) a ten-criterion rubric (C1–C10) scoring major AI safety benchmarks on risk-surface coverage, paired with a policy crosswalk mapping benchmarks to regulatory requirements; (3) twelve lifecycle KPIs, explainability guidance for foundation models, and a testbed catalogue; and (4) domain-specific analysis across healthcare, finance, education, arts, agriculture, and defense.
+We make four contributions: (1) the first survey bridging governance principles, technical evaluation, and domain deployment across all four system types; (2) development of a ten-criterion rubric (C1–C10), scoring major AI safety benchmarks on risk-surface coverage, paired with a policy crosswalk mapping benchmarks to regulatory requirements; (3) establishing twelve lifecycle KPIs, explainability guidance for foundation models, and a catalogue for AI-readye testbed ; and (4) domain-specific analysis across healthcare, finance, education, arts, agriculture, and defense.
 Three findings emerge: benchmark coverage is dense for bias and toxicity but sparse for privacy, provenance, deepfakes, and system-level failures in agentic settings; evaluations remain largely static and task-local, limiting audit portability; and inconsistent documentation complicates cross-release comparison.
-We outline a research agenda prioritizing adaptive multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration reporting, versioned artifacts, and continuous monitoring.
-This survey offers a structured path to align generative AI evaluation with governance needs for safe and accountable deployment.
-Project page: Project page
+We further outline a research agenda, prioritizing adaptive multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration reporting, versioned artifacts, and continuous monitoring.
+This study offers a structured path to align the evaluation of generative AI with the governance needs for a safe and responsible deployment
+. Project page: Project page
 responsible AI, generative AI, AI ethics, bias, privacy, trust, safety
@@
 1. Introduction
-Generative AI (GenAI) systems, including large language models (LLMs), vision language models (VLMs), diffusion models, and emerging agentic pipelines, are rapidly transitioning from research prototypes to production deployments that generate text, images, code, and autonomous actions across healthcare, finance, education, and other sectors (Hadi et al., 2023; Hajkowicz et al., 2023).
+Generative AI (GenAI) systems, including large language models (LLMs), vision language models (VLMs), diffusion models, and emerging agentic pipelines, are rapidly transitioning from research prototypes to production deployments that generate text, multimedia, code, and autonomous actions across healthcare, finance, education, and other sectors (Hajkowicz et al., 2023).
 This accelerated deployment heightens the need for responsible development practices that address GenAI-specific risks, including hallucination and unverifiability, prompt injection and jailbreak vulnerabilities, data memorization and leakage, deepfake and media integrity threats, and system-level failures in tool-augmented settings.
-Responsible AI (RAI) (Stryker, 2024) has emerged as a central area of research and practice to address these challenges.
+Responsible AI (RAI) has emerged as a central area of research and practice to address these challenges.
 RAI frameworks emphasize fairness, bias mitigation, privacy protection, security, and the safeguarding of human rights (Peters et al., 2020).
@@
 This imbalance underscores the need to align the development and deployment of these systems with ethical standards, fairness, privacy, security, and societal values.
-There has been substantial progress on RAI from both industry (Google AI, 2024; Stryker, 2024; Microsoft, 2024; Anthropic, 2024; Jiang et al., 2025) and academia (Lu et al., 2024b; Dignum and Confidential, 2019; Schiff et al., 2020; Huang et al., 2025).
+There has been substantial progress on RAI from both industry and academia.
 Nevertheless, notable gaps remain in alignment with human values and in operational practice.
@@
 Conceptual work provides principles and guidelines, whereas applied studies demonstrate deployment in specific contexts.
-The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 5, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (of Standards and, NIST), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (ISO/IEC, 2021b).
+The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 4, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (of Standards and (NIST), 2024), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (ISO/IEC, 2021b).
 This gap between governance coverage and practical implementation highlights the need for an integrated review that connects frameworks to implementation strategies and to evaluation methods, including explainable AI and AI-ready testbeds.
+{forest}
+Figure 1. Full Taxonomy of the Responsible Generative AI landscape.
 In this study, we conduct a comprehensive literature review of responsible GenAI (ResGenAI) frameworks, governance guidelines, evaluation benchmarks, and operational risk assessment methodologies.
-Our scope centres on the application of RAI principles to generative systems, including LLMs, VLMs, diffusion models, and agentic/RAG pipelines.
-While we draw on the broader RAI literature for foundational constructs (e.g., fairness, privacy, safety, transparency, accountability), our analysis targets ResGenAI specific risk mechanisms such as hallucination and unverifiability, prompt injections, jailbreaking and tool-use failures, data memorization and leakage, provenance and deepfakes or media integrity risks, intellectual property concerns, and system-level failures in GenAI applications.
+Our scope centres on the application of RAI principles to generative systems, including LLMs, VLMs, diffusion models, and touching on agentic/RAG topics.
+While we draw on the broader RAI literature for foundational constructs (e.g., fairness, privacy, safety, transparency, accountability), our analysis targets ResGenAI specific risk mechanisms such as hallucination, jailbreaking and tool-use failures and system-level failures in GenAI applications.
 Despite substantial prior work, ResGenAI guidance remains fragmented across principles, processes, and tools (Table 1).
 Consequently, there is a pressing need to translate regulatory obligations into concrete engineering controls and to produce portable audit evidence for post-market monitoring.
-{forest}
-Figure 1.
-Responsible GenAI landscape organised into three pillars: Challenges (Section 3), Strategies (Section 4), and Applications (Section 5).
-Technical challenges span data, model/system, and emergent-threat layers; strategies encompass benchmarks, KPIs, explainability, testbeds, governance, and toolkits; applications map domain-specific risks and controls across six sectors.
-Contributions. Our key contributions in this work are:
+Our key contributions in this work are:
   1. (1)
-    To the best of our knowledge, this is the first review article that simultaneously spans governance principles, technical evaluation, and domain deployment for GenAI.
-We synthesize findings across the NIST AI RMF, EU AI Act, ISO/IEC 42001, and other instruments listed in Table 6, covering LLMs, VLMs, diffusion models, and RAG topics.
+    We present a comprehensive survey that simultaneously spans governance principles, technical evaluation, and domain deployment for GenAI.
+We synthesize findings across the NIST AI RMF, EU AI Act, ISO/IEC 42001, and other instruments listed in Table 5, covering GenAI topics.
 As shown in Table 1, no prior survey bridges all four thematic areas (CS theory, application, policy, GenAI) and all four novelty dimensions (taxonomy, scope, evaluation, methods) simultaneously.
   2. (2)
-    We introduce a ten-criterion rubric (C1–C10) that scores widely used AI safety benchmarks on risk-surface coverage and assurance quality, and a policy crosswalk (Section 4.2.1) that maps each benchmark to the governance requirements it can plausibly evidence.
-The closest efforts (Liu et al., 2025b; Ferdaus et al., 2026) partially address evaluation taxonomy or trust assessment, but neither links benchmarks to regulatory audit-evidence needs.
+    We introduce a multifold-criterion rubric (C1–C10) that scores widely used AI safety benchmarks on risk-surface coverage and assurance quality, and a policy crosswalk (Section 4.2.1) that maps each benchmark to the governance requirements it can plausibly evidence.
+The closest efforts (Liu et al., 2025a; Ferdaus et al., 2026) partially address evaluation taxonomy or trust assessment, but neither links benchmarks to regulatory audit-evidence needs.
   3. (3)
@@
   4. (4)
-    We apply the RAI lens to real-world applications and domains (Section 5, demonstrating how technical advancements can be effectively aligned with governance frameworks and regulatory requirements.
+    We apply the ResGenAI lens to real-world applications and domains (Section 5 presents how technical advancements can be effectively aligned with governance frameworks and regulatory requirements.
 Table 1.
@@
 (Schiff et al., 2020) arXiv 2020 Responsible AI x x x x x
-(Mehrabi et al., 2021a) ACM CSUR 2021 Bias and Fairness x x x x x x
+(Mehrabi et al., 2021) ACM CSUR 2021 Bias and Fairness x x x x x x
 (Hu et al., 2021) ACM CSUR 2021 AI Safety x x x x x x
@@
 (Wang et al., 2025) arXiv 2025 Reasoning Model Safety x x x
-(He et al., 2025b) ACM CSUR 2025 LLM Agent Security x x x x
+(He et al., 2025) ACM CSUR 2025 LLM Agent Security x x x x
 (Papagiannidis et al., 2025) Elsevier JSIS 2025 RAI Governance x x x x x
-(Liu et al., 2025b) arXiv 2025 LLM Safety Evaluation x x x x
+(Liu et al., 2025a) arXiv 2025 LLM Safety Evaluation x x x x
 Ours ACM CSUR 2026 Responsible GenAI
@@
 Scope: coverage of multimodal systems and agentic/tool-using pipelines.
-Evaluation: governance-aligned benchmarking with a scoring rubric or crosswalk.
-Methods: actionable engineering controls, lifecycle KPIs, or testbed guidance.
-★ = partial or limited treatment.
-The rest of this paper is organized as follows.
-Section 2 presents the literature review methodology.
-Section 3 provides background on ResGenAI.
-Section 4 reports our results, including thematic findings, benchmark assessments, key performance indicators, explainable AI, and AI-ready testbeds.
-Section 5 discusses ResGenAI applications, challenges, and strategies.
-Section 6 presents discussion and forward thoughts, and finally, Section 7 concludes with open challenges and future research directions.
+Evaluation: governance-aligned benchmarking with a scoring rubric.
+Methods: engineering controls, lifecycle KPIs, or testbed guidance.
 The taxonomy presented in Figure 1 provides a structural blueprint for this survey, establishing the thematic connections between high-level technical challenges and domain-specific applications.
@@
 This review adheres to the PRISMA guidelines (Page et al., 2021) to ensure a transparent, reproducible, and rigorous synthesis of evidence.
-This study follows established methodological recommendations for performing systematic reviews within the computer science community (Al-Selwi et al., 2024; Boland et al., 2017a; Keele et al., 2007).
-To ensure comprehensive and consistent reporting, the structural framework was modelled after previous extensive surveys in the domain.
-The mapping procedure utilized in this review is organized into four interconnected phases that covering objectives (Section 2.1), search strategy (Section 2.2), eligibility criteria (Section 2.3), and study selection (Section 2.4) which are visualized in the methodological flow (Fig.
-2) and detailed throughout Section 2.
-For a more exhaustive exploration of these mapping and reporting protocols, foundational works in the field (Boland et al., 2017a; Al-Selwi et al., 2024; Keele et al., 2007; Boland et al., 2017b) offer additional guidance.
-Figure 2. The systematic literature mapping process utilized in this study
-2.1. Research Questions Formulation
-This study conducts a literature review of responsible generative AI (ResGenAI) in the post-GPT period and the synthesis of the findings. We address the following key questions:
-  • RQ1: What are the dominant themes, risks, and controls in ResGenAI and multimodal systems since late 2022?
+This study follows established methodological recommendations for performing systematic reviews within the computer science community (Keele and others, 2007).
+To ensure comprehensive and consistent reporting, the structural framework was modeled after previous extensive surveys in the domain.
+The mapping procedure utilized in this review is organized into four interconnected phases that covering objectives (Section 2), search strategy (Section 2), eligibility criteria (Section 2), and study selection (Section 2) which are visualized in the methodological flow (Appendix Figure 1) and detailed throughout Section 2.
+Research Questions Formulation
+This study conducts a literature review of ResGenAI in the post-GPT period and synthesizes the findings.
+We address the following key questions: RQ1: What are the dominant themes, risks, and controls in ResGenAI and multimodal systems since late 2022?
 Motivation: The rapid transition of GenAI from research to production has outpaced traditional risk taxonomies, necessitating a contemporary synthesis of how multimodal and agentic systems introduce qualitatively new failure modes beyond predictive ML.
 RQ2: Which evaluation facets and metrics are most commonly reported, and where are the gaps?
-    Motivation: Identifying the “Evaluation Gap” is critical for determining if current safety suites adequately model the adaptive, iterative pressure applied by real-world adversaries versus merely serving as static, keyword-based refusal scripts.
+Motivation: Identifying the “Evaluation Gap” is critical for determining if current safety suites adequately model the dynamic pressure applied by real-world adversaries versus merely serving as static, keyword-based refusal scripts.
 RQ3: What evidence types (empirical, analytical, normative) underpin claims about RAI effectiveness?
@@
 Motivation: Principles without practice lead to “safetywashing.” We aim to bridge the gap between regulatory mandates like the EU AI Act and concrete engineering controls, providing an implementation path for verifiable stewardship.
-2.2. Evidence Acquisition and Search Strategy
+Evidence Acquisition and Search Strategy
 A comprehensive search was conducted across five major academic databases and registers: Scopus ($n=1502$), Web of Science ($n=2798$), IEEE Xplore ($n=279$), ACM Digital Library ($n=656$), and arXiv ($n=78$).
 To capture the surge in GenAI research following the release of ChatGPT, the search window was limited to publications from November 2022 through December 2025.
-The search strings utilized an intersection of terms related to Generative AI (e.g., “Large Language Model”, “Foundation Model”, “LLM”) and Responsibility facets (e.g., “AI Safety”, “Trustworthy AI”, “Alignment”, “Hallucination mitigation”).
-Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity), with syntax adapted to each database (full strings in Appendix A.1).
+The search strings utilized an intersection of terms related to GenAI (e.g., “Large Language Model”, “Foundation Model”, “LLM”) and Responsibility facets (e.g., “AI Safety”, “Trustworthy AI”, “Alignment”, “Hallucination mitigation”).
+Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity), with syntax adapted to each database (full strings in Appendix 1.1).
 Syntax was adapted for each database to support field-specific constraints (Title/Abstract/Keywords), and proximity operators were employed where available to enhance precision.
 To mitigate coverage bias, database queries were complemented with forward and backward snowballing from seed works identified as influential during pilot screening.
-The review protocol, including sources, screening rules, and the synthesis plan, was prespecified in line with established computer science review guidance (Boland et al., 2017a; Keele et al., 2007)
-2.3. Eligibility Criteria
-To ensure methodological rigor and relevance to the research questions (RQ1–RQ4), we applied the following criteria:
-  • Inclusion Criteria: (1) Peer-reviewed journals, conference proceedings, or high-impact preprints materially shaping RAI discourse; (2) English-language works published within the target 2022–2025 window; and (3) Studies proposing technical frameworks, evaluation metrics, or empirical risk assessments for generative or multimodal systems.
+Eligibility Criteria
+To ensure methodological rigor and relevance to the research questions (RQ1-RQ4), we applied the following criteria: Inclusion Criteria: (1) Peer-reviewed journals, conference proceedings, or high-impact preprints; (2) English-language works published within the target 2022-2025 window; and (3) Studies proposing technical frameworks, evaluation metrics, or empirical risk assessments for GenAI systems.
 Exclusion Criteria: (1) Purely opinionated pieces or editorials without substantive technical argumentation; (2) Short papers or reports lacking sufficient methodological depth; (3) Items whose primary contribution lay outside RAI (e.g., domain applications without safety or governance components); and (4) Redundant duplicates or inaccessible records.
 Figure 2. PRISMA flowchart of the systematic literature identification, screening, and inclusion process
-2.4. Study Selection and Screening
+Study Selection & Screening
 As detailed in the PRISMA flow diagram (Fig.
-3), the initial search identified 5,313 records.
+2), the initial search identified 5,313 records.
 Selection was conducted by three independent reviewers to minimize bias.
-Following automated deduplication ($n=2125$) and the removal of records marked as ineligible by automation tools ($n=609$) or other technical reasons ($n=53$), 2,562 unique records remained for screening.
-Titles and abstracts were screened against the eligibility criteria, resulting in the exclusion of 2,153 items.
-Subsequently, 372 reports were sought for full-text retrieval and assessed for eligibility.
-Reports were excluded if they failed to meet quality and technical thresholds ($n=140$) or could not be retrieved ($n=37$).
-This rigorous multi-reviewer process yielded a final synthesis set of 232 studies.
-Data extraction followed a thematic pipeline to categorize findings into risks, controls, and evaluation facets as defined by the research objectives.
+Following automated deduplication ($n=2125$) and the removal of records marked as ineligible by automation tools ($n=609$) or other technical reasons ($n=53$), 2,526 unique records remained for screening.
+Titles and abstracts were screened against the eligibility criteria, resulting in the exclusion of 2,117 items.
+Subsequently, 409 reports were sought for full-text retrieval, of which 37 could not be retrieved.
+The remaining 372 reports were assessed for eligibility.
+Reports were excluded if they failed to meet quality and technical thresholds ($n=230$).
+This rigorous multi-reviewer process yielded a final synthesis set of 142 studies.
+Data extraction followed a thematic pipeline to categorize findings into risks, controls, and evaluation as defined by the research objectives.
 In the following section, we outline the background and key challenges of ResGenAI, which set the stage for the synthesis of our review findings.
@@
 3.1. Historical Context
-RAI has received a lot of intensified scholarly and policy attention with the rapid application of GenAI (Zirpoli, 2023b; Al-kfairy et al., 2024).
-GenAI refers to models that synthesize new samples from learned data distributions (van der Zant et al., 2013), spanning language, vision, audio, and multimodal content  (Muneer et al., 2025).
-Early work combined symbolic rules with probabilistic language models; for example, ELIZA illustrated template driven dialogue in the 1960s (Weizenbaum, 1966).
-This early era established statistical $n$ gram models (Fink, 2014) and neural sequence modeling for text and speech, laying the groundwork for contemporary large scale generative systems (Qureshi et al., 2025).
+RAI has received intensified scholarly and policy attention with the rapid application of GenAI.
+GenAI refers to models that synthesize new samples from learned data distributions, spanning language, vision, audio, and multimodal content (Muneer et al., 2025).
+Early work combined symbolic rules with probabilistic language models; ELIZA illustrated template-driven dialogue in the 1960s (Weizenbaum, 1966), while statistical $n$-gram models (Fink, 2014) and neural sequence modeling laid the groundwork for contemporary large-scale generative systems (Qureshi et al., 2025).
 Figure 3 highlights the most salient milestones for RAI (alignment, evaluation, governance).
-As shown in the Figure 4.
-The first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis and, concurrently, made deepfakes feasible, creating the media integrity challenge that current governance frameworks struggle to address Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in the introduction of attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
-A second inflection gave rise to foundation models: large transformers pre-trained in web-scale corpora (e.g., GPT-2/3) and bidirectional encoders (e.g., BERT) that catalyzed transfer across tasks.
-Alignment advances, including instruction tuning (Peng et al., 2023), reinforcement learning from human feedback (RLHF) (Kaelbling et al., 1996), and constitutional AI (Bai et al., 2022), shifted the emphasis from raw capability to controllability and safety.
-In vision, diffusion models (Ho et al., 2020) eclipsed GANs for photorealistic generation; when combined with latent compression, they enabled scalable text-to-image synthesis (Rombach et al., 2022).
-The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023a), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023b) extend prompting beyond text.
-Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation, bringing new safety, privacy, and governance requirements.
+The first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis and, concurrently, made deepfakes feasible, creating the media integrity challenge that current governance frameworks struggle to address.
+Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
+A second inflection gave rise to foundation models: large transformers pre-trained on web-scale corpora (e.g., GPT-2/3) and bidirectional encoders (e.g., BERT) that catalyzed transfer across tasks.
+Alignment advances—instruction tuning (Peng et al., 2023), RLHF (Kaelbling et al., 1996), and constitutional AI (Bai et al., 2022)—shifted emphasis from raw capability to controllability and safety.
+In vision, diffusion models (Ho et al., 2020) eclipsed GANs for photorealistic generation; combined with latent compression, they enabled scalable text-to-image synthesis (Rombach et al., 2022).
+The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023b), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023) extend prompting beyond text.
+Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation, introducing new safety, privacy, and governance requirements.
 Figure 3. Timeline of generative AI milestones
@@
 3.2. Challenges in Responsible Generative AI
-We group GenAI challenges into technical and regulatory domains.
-3.2.1. Technical Challenges
+Technical Challenges
 Technical issues arise across data pipelines, model design and training, evaluation and assurance, and operational deployment.
-Data-Related Challenges
-AI models rely on diverse and representative training data to achieve equitable performance across demographic groups and application contexts.
-Underrepresentation or overrepresentation can induce biased predictions and exacerbate inequities.
-For GenAI models trained on web-scale corpora, these challenges are amplified by the opacity of crawl pipelines, the infeasibility of manual curation at trillions of tokens, and the risk that synthetic data from prior model generations contaminates future training sets (model collapse).
-Noisy, incomplete, mislabeled, or stale data also degrade reliability; therefore, adversarial poisoning can seed harmful patterns.
-Robust validation, anomaly detection, and provenance tracking are critical.
-Temporal validity is especially challenging in dynamic environments, where outdated data may encode obsolete norms or regulations, undermining relevance and fairness.
-ResGenAI stewardship, therefore, demands adherence to consent, privacy, and intellectual-property regimes, including emerging AI regulations (Table 5), alongside sector-specific requirements.
-Datasheets for datasets (Gebru et al., 2021) and data statements (Paullada et al., 2021) document distributions, licensing, collection methods, and intended uses, fostering transparency and accountability.
-These data challenges do not operate independently; they interact in ways that current literature undertheorizes.
+Data-Related Challenges.
+For GenAI models trained on web-scale corpora, data-related challenges are amplified by the opacity of crawl pipelines, the infeasibility of manual curation at trillions of tokens, and the risk that synthetic data from prior model generations contaminates future training sets (model collapse).
+Noisy, incomplete, mislabeled, or stale data degrade reliability; adversarial poisoning can seed harmful patterns; and temporal validity matters where outdated data may encode obsolete norms.
+ResGenAI stewardship demands adherence to consent, privacy, and intellectual-property regimes, including emerging AI regulations (Table 4), alongside sector-specific requirements.
+These challenges interact in ways that current literature undertheorizes.
 Representativeness goals can conflict with privacy regimes: expanding demographic coverage may require collecting sensitive attributes that GDPR restricts.
-Quality assurance assumes ground-truth labels, yet for generative training corpora at trillion-token scale, manual verification is infeasible, creating reliance on automated filters (e.g., toxicity classifiers) that themselves embed biases.
-Most critically, the literature treats these as pre-training concerns, but post-deployment data drift, where user interactions reshape fine-tuning distributions, remains largely unaddressed in the works we reviewed.
-Model- and System-Related Challenges
-Model- and system-level risks span explainability, bias amplification, robustness under adversarial conditions, calibration, and post-deployment drift.
-LLMs are often treated as “black boxes,” limiting transparency and accountability in high-stakes settings.
-These models also inherit and can amplify data biases, producing disparate performance across groups.
-Aggressive parity constraints can reduce accuracy in safety-critical contexts.
-In generative settings, these biases manifest not only as allocational harms but as representational harms in generated text and images (e.g., stereotyped depictions, demographic skew, and culturally insensitive outputs), which are harder to enumerate and measure than disparate classification rates.
-Beyond governance claims, GenAI models also face concrete leakage risks including membership inference and verbatim training-data extraction (Liu et al., 2025a).
-Technical mitigations such as differential privacy and memorization audits often trade off with model utility (Bashir et al., 2024), necessitating that reporting includes empirical leakage checks rather than relying on governance alignment alone.
-Current static, one-shot benchmarks rarely capture these dynamics.
-Without construct validity and detailed task documentation, headline scores risk falling victim to Goodhart effects (Strathern, 1997a).
-Safety claims should instead be framed as assurance cases with traceable evidence (e.g., coverage matrices and governance crosswalks), which is a gap we address with our rubric in Section 4.2.1.
-3.2.2. Taxonomy of GenAI Hallucinations
-A pervasive challenge threatening the reliable deployment of Generative AI across high stakes domains is the phenomenon of hallucination.
-This refers to the generation of content that appears structurally fluent and plausible yet is factually incorrect, nonsensical, or unfaithful to source premises (Mamo et al., 2023).
-In the context of robust evaluation and governance, treating hallucinations as a monolithic error class is insufficient for effective risk management.
-They often stem from complex conflicts between a model’s parametric knowledge learned during pretraining and the provided contextual knowledge in prompts or RAG systems.
-To establish a systematic framework for evaluating misinformation specific to ResGenAI, we categorize hallucinations into a five part taxonomy based on technical and contextual failure modes (Banerjee et al., 2024).
-This taxonomy, visualized in Fig.
-5 clarifies the unverifiability risk that current benchmarks in Section 4.2.1 and sectoral applications in Section 5 must address.
-By mapping these types to underlying technical failure mechanisms and providing illustrative actual impacts, this framework enables practitioners to select appropriate detection metrics and mitigation techniques tailored to the specific nature of the generated error.
-Figure 5.
-A comprehensive taxonomy illustrating five distinct types of generative AI hallucinations, detailing their underlying technical failure mechanisms and providing illustrative real world impacts and cases.
-3.2.3. Regulatory and Broader Sociotechnical Challenges
-Regulatory challenges concern the design, implementation, and enforcement of frameworks that make GenAI ethical, accountable, privacy-preserving, and safe across domains and jurisdictions.
-User behavior and deployment context mediate how risks materialize.
-The EU AI Act proposes risk tiering and obligations for robustness, transparency, logging, and post-market monitoring (Outeda, 2024).
-Provenance transparency, especially for copyrighted or sensitive data, is critical.
-Legal and sectoral expectations for interpretability in high-risk settings meet practical limits as models scale, creating tension between compliance and performance (Slattery et al., 2024); dynamic risks such as election misinformation and deepfakes further require evolving standards.
-Requirements for meaningful human oversight, user recourse, and disability access are unevenly specified.
-Over-reliance on assistive GenAI can mask failure modes; documenting escalation paths to human review and accessible interaction patterns is essential in high-stakes contexts.
-Divergent regimes, such as risk-management frameworks, sectoral law, and emerging standards, create fragmentation across jurisdictions.
+Quality assurance assumes ground-truth labels, yet for trillion-token corpora manual verification is infeasible, creating reliance on automated filters that themselves embed biases.
+Most critically, the literature treats these as pre-training concerns, but post-deployment data drift—where user interactions reshape fine-tuning distributions—remains largely unaddressed.
+Model and System Related Challenges.
+LLMs are often treated as “black boxes” and can amplify data biases, producing disparate performance across groups.
+In generative settings, these biases manifest not only as allocational harms but as representational harms in generated text and images (e.g., stereotyped depictions, demographic skew), which are harder to enumerate and measure than disparate classification rates.
+GenAI models also face concrete leakage risks including membership inference and verbatim training-data extraction (Liu et al., 2025b).
+Technical mitigations such as differential privacy and memorization audits often trade off with model utility (Bashir et al., 2024).
+Without construct validity and detailed task documentation, headline scores risk Goodhart effects (Strathern, 1997).
+Safety claims should instead be framed as assurance cases with traceable evidence (e.g., coverage matrices and governance crosswalks)—a gap we address with our rubric in Section 4.2.1.
+Figure 4. Types of GenAI hallucinations, detailing their underlying technical failure mechanisms and illustrative real-world impacts.
+GenAI Hallucinations
+A pervasive challenge for reliable GenAI deployment is hallucination: generation of content that appears fluent and plausible but is factually incorrect, nonsensical, or unfaithful to source premises (Mamo et al., 2023).
+Treating hallucinations as a monolithic error class is insufficient for effective risk management.
+We categorize hallucinations into a five-part taxonomy based on technical and contextual failure modes (Banerjee et al., 2024), visualized in Fig.
+4, to clarify the unverifiability risk that current benchmarks in Section 4.2.1 and sectoral applications in Section 5 must address.
+Regulatory and Broader Sociotechnical Challenges
+Regulatory challenges concern designing, implementing, and enforcing frameworks that make GenAI ethical, accountable, privacy-preserving, and safe across domains and jurisdictions.
+Provenance transparency, especially for copyrighted or sensitive data, is critical, and legal expectations for interpretability in high-risk settings meet practical limits as models scale.
+Over-reliance on assistive GenAI can mask failure modes; documenting escalation paths to human review is essential in high-stakes contexts.
+Divergent regimes—risk-management frameworks, sectoral law, and emerging standards—create fragmentation across jurisdictions.
 Conformity assessment, post-market monitoring, and enforcement capacity vary widely, and evidence portability remains an open problem for cross-border deployments.
-Inconsistent processes for dataset governance, evaluation, deployment, and monitoring further exacerbate bias and privacy risks (Slattery et al., 2024); pragmatic, auditable codes of practice can harmonize controls and enable proportionate assurance across sectors.
-3.3. Emergent Threats
-While Section 3.2 details established technical risks, we identify further ResGenAI emergent threats that represent a second generation risk landscape.
-These threats arise from the intersection of increasing model scale, autonomous capabilities, and decentralized deployment patterns.
-3.3.1. Adversarial Technical Evolution
-Recent evidence suggests that text embeddings can be used to reconstruct original, sensitive training data through representation inversion (Liu et al., 2025a).
-Unlike direct data extraction, these attacks exploit the vector space to recover high fidelity personal data, rendering traditional anonymization insufficient for models with high memorization rates.
-Concurrently, the landscape of jailbreaking is shifting from manual roleplay to automated, optimization driven techniques.
-Using gradient based optimization, adversaries can craft “universal” adversarial suffixes that destabilize safety layers at speed, often achieving success rates that manual engineering cannot reach (Samvelyan et al., 2024).
-3.3.2. Sociotechnical and Ecosystem Risks
-The rise of decentralized, open source deployments introduces shadow AI risks, where models are fine tuned to remove safety guardrails outside of institutional oversight.
-Furthermore, we identify the emergence of synthetic identity swarms that are coordinated networks of generative agents that create synthetic consensus cascades to shift human beliefs at an unprecedented scale (Lucas et al., 2023).
-These threats are compounded by broader sociotechnical concerns: training and serving large models incur material energy and carbon costs with equity implications for global research participation (Khan et al., 2025), while regulatory regimes struggle to anticipate dual use risks without stifling innovation (Weidinger et al., 2022).
-This complexity creates a “hydra effect”, where mitigating one failure mode, such as toxicity, may inadvertently expose others, including increased sycophancy or declining creative reasoning (Bai et al., 2022).
-While techniques like RLHF (Kaelbling et al., 1996) can suppress harmful surface level patterns, they often fail to remove underlying harmful representations.
-3.3.3. Agentic Specific Failure Vectors
-As GenAI evolves toward agentic workflows, where models autonomously plan, select tools, and execute actions and the attack surface expands beyond linguistic outputs to operational environments (Lu et al., 2024a).
-Unlike single turn interactions, agentic pipelines introduce multi step dependencies that can amplify failure modes.
-Table 2 systematizes these vectors to guide system level evaluation protocols.
-Table 2. Agentic Specific Failure Vectors: Risks in Autonomous Tool Use and Planning
+Emergent Threats
+While Section 3.2 details established technical risks, we identify further ResGenAI emergent threats representing a second-generation risk landscape, arising from the intersection of increasing model scale, autonomous capabilities, and decentralized deployment.
+Adversarial Technical Evolution.
+Recent evidence shows that text embeddings can reconstruct sensitive training data through representation inversion (Liu et al., 2025b), rendering traditional anonymization insufficient for models with high memorization rates.
+Concurrently, jailbreaking is shifting from manual role-play to automated, optimization-driven techniques: gradient-based optimization can craft “universal” adversarial suffixes that destabilize safety layers at speed and scale.
+Sociotechnical and Ecosystem Risks.
+Decentralized, open-source deployments introduce shadow AI risks, where models are fine-tuned to remove safety guardrails outside institutional oversight.
+We further identify synthetic identity swarms—coordinated networks of generative agents creating synthetic consensus cascades to shift human beliefs at unprecedented scale.
+These threats are compounded by material energy and carbon costs with equity implications for global research participation (Khan et al., 2025), and by regulatory struggles to anticipate dual-use risks without stifling innovation.
+This creates a “hydra effect”: mitigating one failure mode (e.g., toxicity) may expose others, including increased sycophancy or declining creative reasoning (Bai et al., 2022).
+While RLHF (Kaelbling et al., 1996) can suppress harmful surface-level patterns, it often fails to remove underlying harmful representations.
+Agentic-Specific Failure Vectors.
+As GenAI evolves toward agentic workflows—where models autonomously plan, select tools, and execute actions—the attack surface expands beyond linguistic outputs to operational environments (Lu et al., 2024a).
+Unlike single-turn interactions, agentic pipelines introduce multi-step dependencies that amplify failure modes.
+Table 2 systematizes these vectors; in Section 4.2, we operationalize this via our comparative rubric and governance crosswalk mapping risks to evaluative evidence.
+Table 2. Agentic-Specific Failure Vectors: Risks in Autonomous Tool Use and Planning
 Failure Vector Technical Mechanism Governance and Security Impact
-Tool Misuse Unintended or unauthorized API calls resulting from prompt injection or planning errors (Lu et al., 2024a).
-Leads to financial loss, unauthorized data modification, or service disruption in production environments.
+Tool Misuse Unintended or unauthorized API calls from prompt injection or planning errors (Lu et al., 2024a). Financial loss, unauthorized data modification, or service disruption.
 Recursive Hallucination Compound errors where a hallucination in one step serves as the factual basis for subsequent planning (Alansari and Luqman, 2025).
 Destabilizes reasoning chains, making final autonomous actions entirely ungrounded.
-Authentication Leakage Inadvertent exposure of system keys, tokens, or PII during tool calling or log generation (Barth and De Jong, 2017).
-Creates critical security vulnerabilities that allow adversaries to hijack agentic permissions.
-Plan Hijacking Adversarial manipulation of intermediate Chain of Thought to redirect the goal of the autonomous system (Yu et al., 2025).
+Authentication Leakage Inadvertent exposure of keys, tokens, or PII during tool calling or log generation (Barth and De Jong, 2017).
+Critical security vulnerabilities allowing adversaries to hijack agentic permissions.
+Plan Hijacking Adversarial manipulation of intermediate Chain of Thought to redirect autonomous system goals (Yu et al., 2025).
 Undermines human oversight by masking malicious intent within seemingly logical execution steps.
-This systemic fragility motivates the use of proportionate, auditable controls and transparent evidence.
-In Section 4.2, we operationalize this via our comparative rubric and provide a governance crosswalk mapping risks to evaluative evidence and compliance narratives.
 4. Results
@@
 (d) Themes $\times$ publication type
-(e) Top keywords
-(f) Publications per year
 Figure 6. Thematic analysis overview on RAI literature
@@
 Figure 5 provides a compact view of annual volume and keyword evolution.
-Appendix Table 2 list representative studies included in this review.
+Appendix Table 2 lists representative studies included in this review.
 Our thematic synthesis highlights that ResGenRAI scholarship in the post-GPT period clusters around five dominant themes: (i) safety and benchmarking, with a surge of multimodal evaluation suites and red-teaming methods; (ii) fairness, bias, and governance, where debates center on aligning technical audits with emerging laws such as the EU AI Act and NIST AI RMF and others (Table 4); (iii) explainability and transparency, underscoring the tension between interpretability and performance in large foundation models; (iv) privacy and security, particularly membership inference and data extraction risks; and (v) sectoral and sociotechnical applications, spanning healthcare, finance, education, and agriculture.
 Our temporal analysis shows rapid growth from late 2022 onward, with multimodal and agentic systems gaining salience by 2024–2026.
-These themes frequently co-occur, for example where explainability is invoked to mitigate fairness concerns, or where privacy and governance frameworks intersect in deployment contexts.
 4.2. Benchmarks: Comparative Assessment
-This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in Tables 4 and 6.
+This section formalizes our governance-aligned rubric (C1–C10) and applies it to widely used benchmarks; we then compare coverage and link results to audit evidence in Tables 3 and 5.
 4.2.1. Comparative Assessment Protocol
@@
 This approach is appropriate because the underlying evidence, e.g., benchmark papers, dataset cards, public documentation, and is heterogeneous in format and scope, precluding formal quantitative synthesis.
-Rubric design.
-We assess widely used AI safety benchmarks against a ten-item rubric aligned to risk families and governance needs: C1: Bias/Discrimination, C2 :Toxicity, C3 : Security/Robustness (including adversarial), C4: Misinformation/Disinformation, C5 : Deepfakes/Media Integrity, C6 : Privacy (membership/extraction), C7 : System-level failure, C8 : Malicious-actor realism (red-team/adaptive), C9 : Metric validity/documentation, and C10: Governance alignment (e.g., NIST AI RMF, EU AI Act).
-The ten criteria (also given in Appendix A.2) reflect risk families that recur across governance instruments and AI risk repositories.
-Scoring procedure and reliability.
-Each benchmark was scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, $1$ = partial, $2$ = strong) by two independent reviewers, following the codebook in Appendix A.2.
+Rubric design: We assess widely used AI safety benchmarks against a ten-item rubric aligned to risk families and governance needs: C1: Bias/Discrimination, C2 : Toxicity, C3 : Security/Robustness (including adversarial), C4: Misinformation/Disinformation, C5 : Deepfakes/Media Integrity, C6 : Privacy (membership/extraction), C7 : System-level failure, C8 : Malicious-actor realism (red-team/adaptive), C9 : Metric validity/documentation, and C10: Governance alignment (e.g., NIST AI RMF, EU AI Act).
+The ten criteria (also given in Appendix 1.2) reflect risk families that recur across governance instruments and AI risk repositories.
+Scoring procedure and reliability: Each benchmark was scored $s_{i}(B)\in\{0,1,2\}$ per criterion (0 = no coverage, $1$ = partial, $2$ = strong) by two independent reviewers, following the codebook in Appendix 1.2.
 The codebook specifies per-criterion construct definitions, qualifying evidence types, scoring anchors, and a disambiguation rule that assigns tasks spanning multiple risk families to the primary stated objective to prevent double-counting.
@@
 Equal weights are used by default.
-The normalised score is
+The normalized score is
 (1) $S(B)=\frac{\sum_{i=1}^{10}w_{i}\,s_{i}(B)}{2\,\sum_{i=1}^{10}w_{i}},\quad w_{i}>0.$
 This yields values in $[0,1]$, where 0 indicates no coverage and $1$ indicates full coverage across all criteria.
-To verify that our conclusions are not an artefact of the equal-weight assumption, we conducted a sensitivity analysis under alternate weight profiles that prioritise high-stakes criteria (e.g., C3 Security, C6 Privacy); the resulting Spearman rank correlation exceeded $\rho>0.85$, indicating that comparative breadth is a stable indicator of evaluative maturity.
+To verify that our conclusions are not an artifact of the equal-weight assumption, we conducted a sensitivity analysis under alternate weight profiles that prioritize high-stakes criteria (e.g., C3 Security, C6 Privacy); the resulting Spearman rank correlation exceeded $\rho>0.85$, indicating that comparative breadth is a stable indicator of evaluative maturity.
 Table 3.
-Mapping benchmarks to risk surface areas.
-$\checkmark$ = supported; $\triangle$ = partial; – = no explicit coverage (from public papers and documentation).
-This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 4.
-Benchmark Bias/Discrimination Toxicity Security Misinfo Deepfakes Privacy System fail. Malicious
-HarmBench (Mazeika et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-SALAD-Bench (Li et al., 2024a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$
-ALERT (Tedeschi et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Risk Taxonomy (Cui et al., 2024) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
-TrustGPT (Huang et al., 2023b) $\checkmark$ $\checkmark$
-RealToxicityPrompts (Gehman et al., 2020) $\checkmark$
-DecodingTrust (Wang et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
-SafetyBench (Zhang et al., 2024) $\checkmark$ $\checkmark$ $\triangle$
-MM-SafetyBench (Liu et al., 2024) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\checkmark$
-XSTest (Röttger et al., 2024) $\triangle$ $\checkmark$ $\triangle$
-Rainbow Teaming (Samvelyan et al., 2024) $\triangle$ $\checkmark$ $\triangle$ $\checkmark$ $\checkmark$
-MLCommons AI Safety v0.5 (Vidgen and et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-HELM (Liang et al., 2023) $\checkmark$ $\checkmark$ $\triangle$ $\checkmark$
-SHIELD(Shi et al., 2024) $\triangle$ $\triangle$ $\checkmark$
-BIG-bench (bench authors, 2023) $\checkmark$ $\checkmark$ $\triangle$ $\triangle$ $\triangle$
-PrivLM-Bench (Li et al., 2024b) $\checkmark$
-PrivacyLens (Shao and et al., 2024) $\triangle$ $\checkmark$
-HumaniBench (Raza et al., 2025) $\checkmark$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
-Table 4.
-Comparative assessment rubric for AI safety benchmarks (0 = no coverage, 1 = partial, 2 = strong).
+Comparative assessment rubric for AI safety benchmarks.
+Criteria C1–C8 double as a risk-surface map: 0 = no coverage, 1 = partial, 2 = strong.
 Scores use equal weights and are normalized to $[0,1]$.
-Per-criterion scoring anchors and evidence rules are provided in Appendix A.2.
-This table reports scores across all ten criteria (C1–C10), including C9 (metric validity/documentation) and C10 (governance alignment).
-Benchmark C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 Score
+Per-criterion scoring anchors and evidence rules are provided in Appendix 1.2.
+Risk Surface (C1–C8) Quality
+Benchmark
+C1 Bias/Discrim.
+C2 Toxicity
+C3 Security/Rob.
+C4 Misinfo.
+C5 Deepfakes
+C6 Privacy
+C7 Sys. Failure
+C8 Malicious
+C9 Metric Val.
+C10 Gov. Align.
+Score
+Rainbow Teaming (Samvelyan et al., 2024) 1 1 2 1 0 0 1 2 1 1 0.50
+Risk Taxonomy (Cui et al., 2024) 1 1 1 1 0 1 0 1 1 2 0.45
+HumaniBench (Raza et al., 2025) 2 1 0 1 0 1 1 0 1 1 0.40
+DecodingTrust (Wang et al., 2023) 1 1 1 1 0 1 0 1 1 1 0.40
+MLCommons AI Safety v0.5 (Vidgen and et al., 2024) 1 1 1 1 0 0 0 1 1 1 0.35
+SALAD-Bench (Li et al., 2024b) 1 1 1 1 0 0 0 1 1 1 0.35
+ALERT (Tedeschi et al., 2024) 1 1 1 0 0 0 0 2 1 1 0.35
+MM-SafetyBench (Liu et al., 2024) 1 1 1 1 1 0 0 0 1 1 0.35
+HELM (Liang et al., 2023) 1 1 1 1 0 0 0 0 1 1 0.30
+TrustLLM (Huang et al., 2024b) 1 1 0 1 0 1 1 0 1 0 0.30
+HarmBench (Mazeika et al., 2024) 0 1 1 1 0 0 0 1 1 0 0.25
+SafetyBench (Zhang et al., 2024) 1 1 1 0 0 0 0 0 1 1 0.25
+XSTest (Röttger et al., 2024) 0 1 0 0 0 2 0 1 1 0 0.25
+SHIELD (Shi et al., 2024) 1 0 1 0 2 0 0 0 1 0 0.25
+BIG-bench (authors, 2023) 1 1 1 1 0 0 1 0 0 0 0.25
+PrivLM-Bench (Li et al., 2024a) 0 0 0 0 0 2 0 0 1 1 0.20
+MFG for GenAI (Foundation, 2024) 1 0 1 1 0 0 0 1 0 0 0.20
+PrivacyLens (Shao and et al., 2024) 1 0 0 0 0 2 0 0 1 0 0.20
 RealToxicityPrompts (Gehman et al., 2020) 0 2 0 0 0 0 0 0 1 0 0.15
-SafetyBench (Zhang et al., 2024) 1 1 1 0 0 0 0 0 1 1 0.25
-MM–SafetyBench (Liu et al., 2024) 1 1 1 0 1 0 0 0 1 1 0.30
-SALAD Bench (Li et al., 2024a) 1 1 1 0 0 0 0 0 1 1 0.25
-HarmBench (Mazeika et al., 2024) 0 1 1 0 0 0 0 1 1 0 0.20
-Rainbow Teaming (Samvelyan et al., 2024) 1 1 2 1 0 0 1 2 1 1 0.50
-ALERT (Tedeschi et al., 2024) 0 1 1 0 0 0 0 2 1 1 0.30
-HELM (Liang et al., 2023) 1 1 1 1 0 0 0 0 1 1 0.30
-PrivLM Bench (Li et al., 2024b) 0 0 0 0 0 2 0 0 1 1 0.20
-PrivacyLens (Shao and et al., 2024) 0 0 0 0 0 2 0 0 1 0 0.15
-MLCommons AI Safety v0.5 (Vidgen and et al., 2024) 1 1 1 1 0 0 0 1 1 1 0.35
-SHIELD (Shi et al., 2024) 0 0 1 0 2 0 0 0 1 0 0.20
-MFG for GenAI (Foundation, 2024) 1 0 1 1 0 0 0 1 0 0 0.20
-XSTest (Röttger et al., 2024) 0 0 0 0 0 2 0 1 1 0 0.20
-DecodingTrust (Wang et al., 2023a) 1 1 1 0 0 1 0 1 1 1 0.35
-TrustLLM (Huang et al., 2024b) 1 1 0 1 0 1 1 0 1 0 0.30
-BIG bench (bench authors, 2023) 1 1 1 0 0 0 1 0 0 0 0.25
-Risk Taxonomy (Cui et al., 2024) 1 1 1 0 0 1 0 1 1 2 0.40
-TrustGPT (Huang et al., 2023b) 1 1 0 0 0 0 0 0 1 0 0.20
-HumaniBench (Raza et al., 2025) 2 1 0 1 0 1 1 0 1 1 0.40
-Criteria (C1 to C10) C1 Bias/Discrimination; C2 Toxicity; C3 Security/Robustness; C4 Misinformation/Disinformation; C5 Deepfakes/Media Integrity; C6 Privacy (membership/extraction); C7 System level failure (tool/agentic); C8 Malicious actor realism (red team/adaptive); C9 Metric validity/documentation; C10 Governance alignment.
-Normalization (equal weights): $S(B)=\frac{\sum_{i=1}^{10}s_{i}(B)}{20}$ with $s_{i}\in\{0,1,2\}$.
+TrustGPT (Huang et al., 2023b) 1 1 0 0 0 0 0 0 1 0 0.15
+Scoring: $s_{i}\!\in\!\{0,1,2\}$; $S(B)=\sum s_{i}/20$.
+C1–C8 also serve as the risk-surface crosswalk (0 $\hat{=}$ no coverage, 1 $\hat{=}$ partial, 2 $\hat{=}$ full).
+C9: Metric validity & documentation; C10: Governance alignment.
 Table 4. Governance-based Frameworks and Technical AI Tools used in RAI Practices
@@
 NIST SP-1270 (Schwartz et al., 2022): framework for identifying and managing bias in AI.
-AI4ALL (AI4ALL, [n.
-d.]): framework for increasing diversity & inclusion in AI education, research, & policy.
+AI4ALL (AI4ALL, ): framework for increasing diversity & inclusion in AI education, research, & policy.
 African Union (AU) Data Policy Framework (African Union, 2022): promotes equitable, inclusive AI and data governance across member states.
-Singapore IMDA & PDPC’s AI Governance Framework (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): addresses fairness, human-centricity, and non-discrimination in AI decision-making.
+Singapore IMDA & PDPC’s AI Governance Framework (Infocomm Media Development Authority (IMDA) and Personal Data Protection Commission (PDPC), 2020): addresses fairness, human-centricity, and non-discrimination in AI decision-making.
 Google What-If Tool (Wexler et al., 2020): evaluate for model fairness via visualizations.
@@
 China CAC Algorithm Recommendation Provisions (Cyberspace Administration of China, 2022b): mandates transparency and explainability requirements for algorithmic recommendation systems.
-Learning Interpretability Tool (LIT) (Google PAIR, [n.
-d.]): model interpretability and explainability.
-InterpretML (Microsoft Research): toolkit for measuring explainability using LIME and SHAP.
-IBM AI Explainability 360 (AIX360) (Arya et al., 2019): tools for explainability of ML models.
+Learning Interpretability Tool (LIT) (Google PAIR, ): model interpretability.
+InterpretML: toolkit for explainability IBM AI Explainability 360 (AIX360) (Arya et al., 2019): tools for explainability.
 Transparency & Accountability U.S.
@@
 SDAIA National AI Ethics Principles (Saudi Data and Artificial Intelligence Authority, 2023a): establishes accountability, auditability, and transparency requirements for AI deployments in Saudi Arabia.
-Evidently (Evidently AI, [n.
-d.]): evaluate, test and monitor ML models.
-MLBench (MLBench Team, [n.
-d.]): improve model transparency, reproducibility, and robustness.
-AI Verify (Infocomm Media Development Authority ([n.
-d.]), IMDA): Singapore’s IMDA-developed AI governance testing framework.
+Evidently (Evidently AI, ): evaluate, test and monitor ML models.
+MLBench (MLBench Team, ): improve model transparency, reproducibility, and robustness.
+AI Verify (Infocomm Media Development Authority (IMDA), ): Singapore’s IMDA-developed AI governance testing framework.
 Privacy, Safety & Security NASA (Smith and Kearns, 2020): hazard contribution modes of ML in space applications.
@@
 ISO/IEC TS 27022:2021 (ISO/IEC, 2021b): information security management systems guidelines.
-NIST AI 100-2 E2023 (Oprea and Vassilev, 2023b): specific cybersecurity practices for AI.
+NIST AI 100-2 E2023 (Oprea and Vassilev, 2023): specific cybersecurity practices for AI.
 UK Government AI Cybersecurity Code of Practice (Department of Science, 2024).
@@
 Google’s Secure AI Framework (SAIF) (Google, 2023).
-Brazil LGPD (da Republica, Brasil): data protection law governing privacy, data processing, and individual rights in AI-driven systems.
+Brazil LGPD (da Republica (Brasil), 2018): data protection law governing privacy, data processing, and individual rights in AI-driven systems.
 Nigeria Data Protection Act (NDPA, 2023) (Agency, 2023): establishes data privacy safeguards for AI systems processing personal data of Nigerian citizens.
@@
 UAE Federal Decree-Law No.
-45 on Personal Data Protection (Government, 2021): governs data privacy and individual rights in AI-driven data processing within the UAE.
-Unitary (Ltd., [n.
-d.]): context-aware toxicity tracker.
-Adversarial Robustness Toolbox (ART) (Research, [n.
-d.]).
-Privacy Meter (Privacy and Lab, [n.
-d.]): audit privacy of ML models.
-Google Differential Privacy Libraries (Google, [n.
-d.]).
-SecretFlow (Group, [n.
-d.]): privacy-preserving data analysis and ML toolkit.
-TensorFlow Privacy (Google, [n.
-d.]).
+45 on Personal Data Protection (UAE Government, 2021): governs data privacy and individual rights in AI-driven data processing within the UAE.
+Unitary (Ltd., ): context-aware toxicity tracker.
+Adversarial Robustness Toolbox (ART) (Research, ).
+Privacy Meter (Privacy and Lab, ): audit privacy of ML models.
+SecretFlow (Group, ): privacy-preserving data analysis and ML toolkit.
 Betterdata.AI (inc., 2024): PET on synthetic data generation.
@@
 American Executive Order on Safe, Secure, and Trustworthy AI (House, 2023).
-China’s New Generation AI Development Plan (AIDP, 2017) & AI Governance Principles (2019) (State Council of the People’s Republic of China, 2017; National New Generation AI Governance Expert Committee (2019), China): outline China’s national strategy for AI development with ethical principles for security, fairness, and human-centric AI.
-UAE National AI Strategy 2031 (United Arab Emirates Government, [n.
-d.]): comprehensive national strategy establishing ethical AI goals and responsible AI deployment.
+China’s New Generation AI Development Plan (AIDP, 2017) & AI Governance Principles (2019) (State Council of the People’s Republic of China, 2017; National New Generation AI Governance Expert Committee (China), 2019): outline China’s national strategy for AI development with ethical principles for security, fairness, and human-centric AI.
+UAE National AI Strategy 2031 (UAE Government, ): comprehensive national strategy establishing ethical AI goals and responsible AI deployment.
 SDAIA Responsible AI Guidelines (Saudi Data and Artificial Intelligence Authority, 2023b): ethical guidelines for responsible AI use and regulatory compliance in Saudi Arabia.
-Singapore Model AI Governance Framework (2nd Ed., IMDA) (Infocomm Media Development Authority and Personal Data Protection Commission (PDPC)(2020), IMDA): practical guidance on responsible AI deployment covering internal governance and human oversight.
-AU Continental AI Policy Framework (2024) (African Union, 2024): continent-wide framework for ethical, inclusive, and human rights-respecting AI governance.
-Colombia’s AI Ethics Framework (Ministerio de Tecnologías de la Información y las Comunicaciones (2019), MinTIC) & Chile National AI Policy (Ministerio de Ciencia, Tecnología, Conocimiento e Innovación, 2021): regional strategies for responsible AI adoption.
-Google Responsible AI Practices (Google, [n.
-d.]).
-Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, [n.
-d.]).
+Singapore Model AI Governance Framework (2nd Ed., IMDA) (Infocomm Media Development Authority (IMDA) and Personal Data Protection Commission (PDPC), 2020): practical guidance on responsible AI deployment covering internal governance and human oversight.
+Colombia’s AI Ethics Framework (Ministerio de Tecnologías de la Información y las Comunicaciones (MinTIC), 2019) & Chile National AI Policy (Ministerio de Ciencia, Tecnología, Conocimiento e Innovación, 2021): regional strategies for responsible AI adoption.
+Google Responsible AI Practices (Google, ).
+Australia’s AI Ethics Principles (Australian Government Department of Industry, Science and Resources, ).
 Microsoft’s Responsible AI Toolbox (principles and approach, 2024).
-OneTrust (OneTrust LLC, [n.
-d.]): demonstrate trust, measure and manage risk, and compliance.
+OneTrust (OneTrust LLC, ): demonstrate trust, measure and manage risk, and compliance.
 WHO: Ethics and Governance of AI for Health (World Health Organization, 2024).
 Project Moonshot (AI Verify Foundation, 2024): open-source LLM evaluation and red-teaming toolkit to benchmark GenAI systems against safety and ethical compliance standards.
-Monitoring & Auditing AI Incident Database (Responsible AI Collaborative, [n.
-d.]): index of harms realized in the real world by AI systems.
-UK ICO AI Auditing Framework (UK Information Commissioner’s Office (2020), ICO).
+Monitoring & Auditing AI Incident Database (Responsible AI Collaborative, ): index of harms realized in the real world by AI systems.
+UK ICO AI Auditing Framework (UK Information Commissioner’s Office (ICO), 2020).
 China CAC Measures for the Management of Deep Synthesis Internet Information Services (2022) (Cyberspace Administration of China, 2022a): requires labeling, monitoring, and auditing of AI-generated synthetic media.
 Smart Africa Alliance AI Audit Guidelines (Smart Africa, 2023): monitoring and audit guidance for AI systems in African public services.
-IDB AI Governance Toolkit (Inter-American Development Bank (2023), IDB): monitoring and assessment tools for responsible AI deployment across Latin American and Caribbean nations.
-FairVis (Poloclub, [n.
-d.]): audit classification models for intersectional bias.
-Aequitas (Center for Data Science and Public Policy, [n.
-d.]): bias and fairness audit toolkit.
+IDB AI Governance Toolkit (Inter-American Development Bank (IDB), 2023): monitoring and assessment tools for responsible AI deployment across Latin American and Caribbean nations.
+FairVis (Poloclub, ): audit classification models for intersectional bias.
+Aequitas (Center for Data Science and Public Policy, ): bias and fairness audit toolkit.
 Audit-AI : auditing of bias in generalized ML applications.
@@
 ✓= supported evidence; $\triangle$= partial; –= not covered.
-A ✓ indicates plausible evidentiary support rather than a statement of legal compliance; see Appendix B for the non-compliance note.
+A ✓ indicates plausible evidentiary support rather than a statement of legal compliance; see Appendix 1.3 for the non-compliance note.
 This crosswalk expands on rubric criteria C9 (metric validity/documentation) and C10 (governance alignment) from Table 3.
@@
 HumaniBench
-NIST: Privacy & Security (of Standards and, NIST, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
+NIST: Privacy & Security (of Standards and (NIST), 2020a, b) $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
 EU AI Act: Robustness/Resilience  [^1] https://artificialintelligenceact.eu/article/15 $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$ $\triangle$
@@
 Multimodal Deepfake Risk (C2PA, 2024) $\triangle$
-Adaptive/Red-team Evidence (of Standards and, NIST) $\triangle$ $\triangle$ $\triangle$ $\triangle$
-Abbreviations RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
+Adaptive/Red-team Evidence (of Standards and (NIST), 2023) $\triangle$ $\triangle$ $\triangle$ $\triangle$
+RTP = RealToxicityPrompts; MM–Saf. = MM–SafetyBench; HarmB. = HarmBench; MLC v0.5 = MLCommons AI Safety v0.5; Humani = HumaniBench.
 4.2.2. Comparative Assessment Findings.
-Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs.
-Table 3 presents the risk-surface matrix for criteria C1–C8.
-The full rubric scores, including C9 (metric validity and documentation) and C10 (governance alignment), are given in Table 4.
-Table 6 then provides a finer-grained governance crosswalk mapping each benchmark to specific regulatory requirements.
-Risk-surface coverage.
-Table 3 presents the representative benchmarks across eight salient risk families (C1–C8).
+Building on the challenges in Section 3.2, we apply the rubric in Section 4.2.1 to evaluate benchmarks against governance needs and discuss our findings below.
+Risk-surface coverage: Table 3 presents the representative benchmarks across eight salient risk families (C1–C8).
 Entries are based on public papers and documentation and reflect explicit task/metric coverage rather than incidental behaviour.
-The results in Table 3 show that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens).
+Table 3 alsoshow that coverage is densest for toxicity and bias (e.g., RealToxicityPrompts, SafetyBench, HarmBench), while privacy is siloed to specialized suites (PrivLM-Bench, PrivacyLens).
 Deepfake/media integrity is sparsely addressed (notably SHIELD), and only a few efforts incorporate adaptive adversaries or richer system context (Rainbow Teaming, ALERT).
 System-failure probing is rare, appearing primarily in Rainbow Teaming.
-Rubric-scored comparison.
-We operationalize the ten-criterion rubric in Section 4.2.1 to produce a normalized $[0,1]$ score per benchmark.
+Rubric-scored comparison: We operationalize the ten-criterion rubric in Section 4.2.1 to produce a normalized $[0,1]$ score per benchmark.
 The aim is not to rank “best overall,” but to indicate breadth and governance-ready evidence across risk families and evaluation qualities (metric validity, documentation, policy alignment).
@@
 Single-issue suites score lower by design, e.g., RealToxicityPrompts and PrivacyLens (0.15), useful for their focus but limited for governance evidence.
-Governance context and crosswalk.
-Table 5 situates the analysis within widely used frameworks and tools.
+Governance context and crosswalk: Table 4 situates the analysis within widely used frameworks and tools.
 Table 5 then maps each benchmark to selected governance requirements and risks such as multimodal deepfake risk, and evidence for adaptive/red-team evaluation.
@@
 Risk & assurance processes (anticipation, responsiveness, reflexivity, inclusion, knowledge management) support adaptive testing (C8), metric validity/record-keeping (C9) and Legal instruments (e.g., EU AI Act, NIST AI RMF) ground C10 (governance alignment).
-Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks, motivating the gaps observed in Tables 46
+Sustainable AI highlights environmental/societal/economic impacts that are only weakly covered by current benchmarks.
 4.2.3. Proposed Improvements to AI Safety Benchmarks
 Based on the benchmark gaps we found in the literature, we suggest constructive opportunities to complement existing benchmarks and to support broader governance use.
-First,we recommend that benchmark developers encourage diverse and well-documented corpora across regions, languages, and demographics, accompanied by dataset cards or data statements that capture provenance, consent, licensing, and curation criteria to enhance traceability and accountability (Eriksson et al., 2025; Gebru et al., 2021).
-Equally important is construct clarity: each task and metric should state explicitly what it is intended to measure and, where feasible, be validated with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024).
-Without these foundations, high benchmark scores risk measuring test-taking ability rather than genuine safety.
-Incorporating adversarial evaluation, including undisclosed items and evolving prompts can be used bring benchmarks in line with established security testing practices (Oprea and Vassilev, 2023b; Ghosh et al., 2025).
-Scope should also expand to cover multimodal, non-English, and higher-risk domains, as well as multi-agent or autonomous settings where system-level failures are more likely to surface (Salazar et al., 2025).
-These benchmarks rarely make explicit linkages to policy instruments.
-Engaging multiple stakeholders and mapping evaluation outputs to governance frameworks would mitigate commercial or domain bias and improve regulatory utility (Outeda, 2024).
-Continuous evaluation through automated pipelines that enable periodic re-testing across the model lifecycle would further improve comparability over time (Wang et al., 2024).
+First, we recommend that benchmark developers encourage diverse and well-documented corpora across regions, languages, and demographics, accompanied by dataset cards or data statements that capture provenance, consent, licensing, and curation criteria to enhance traceability and accountability (Gebru et al., 2021).
+Also, each task and metric should state explicitly what it is intended to measure and, where feasible, be validated with multidisciplinary input to guard against superficial compliance claims (Ren et al., 2024).
+Incorporating adversarial evaluation, including undisclosed items and evolving prompts can be used bring benchmarks in line with established security testing practices.
+Scope should also expand to cover multimodal, non-English, and higher-risk domains, as well as multi-agent or autonomous settings where system-level failures are more likely to surface.
+Engaging multiple stakeholders and mapping evaluation outputs to governance frameworks can help mitigate commercial or domain bias and improve regulatory utility.
 Finally, structured red teaming and stress tests under distribution shift should be integrated to surface emergent behaviours that may elude static suites, drawing on practices such as ALERT and Rainbow Teaming.
@@
 They bridge governance expectations with technical implementation so that fairness, transparency, robustness, privacy, and sustainability are evidenced through quantifiable measures, enabling compliance and continuous improvement.
-We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior work (Pessach and Shmueli, 2022), and introduce pragmatic, author-defined KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
+We adapt established measures (SPD, DI, accuracy parity, robustness under perturbation, and energy) from prior works (Pessach and Shmueli, 2022; Raza et al., 2024), and introduce KPIs for explainability, inclusivity, auditing cadence, and high-stakes errors to support governance-ready evidence.
 Notation and conventions $f(x)$ is the model score for input $x$ (state whether probability, logit, or other monotone score); $y\in\{0,1\}$ is the true label; $\hat{y}=\mathbb{1}\{f(x)\geq t\}$ is the predicted label with a single threshold $t$ applied to all subgroups unless stated; $A$ is a protected attribute with subgroups $a$ and $b$; $\phi_{0}$ is a baseline output and $\phi_{i}(x)$ is the attribution for feature $i$ (e.g., SHAP or Integrated Gradients); “accuracy” refers to a stated metric (e.g., top-1, F1, AUROC) computed per subgroup with the same threshold $t$; a perturbation is a noise or attack consistent with a stated threat model; $P_{\mathrm{avg}}$ is average power (W), $T$ is time (h), $\mathrm{CI}$ is grid carbon intensity (kgCO2e/kWh), and $\mathrm{PUE}$ is Power Usage Effectiveness.
@@
     Data quality and integrity
-    $Q=\frac{\mathrm{Valid\ Entries}}{\mathrm{Total\ Entries}}\times 100.$
+    (2) $Q=\frac{\mathrm{Valid\ Entries}}{\mathrm{Total\ Entries}}\times 100.$
     This KPI measures the fraction of entries that pass schema, label, and consistency checks (Pessach and Shmueli, 2022).
@@
     Data privacy compliance
-    $P=\frac{\mathrm{Compliant\ Data\ Points}}{\mathrm{Total\ Data\ Points}}\times 100.$
+    (3) $P=\frac{\mathrm{Compliant\ Data\ Points}}{\mathrm{Total\ Data\ Points}}\times 100.$
     This KPI fractions with documented lawful basis, consent or equivalent, and license compatibility (Leschanowsky et al., 2024).
@@
     Bias detection (statistical parity difference, SPD)
-    $\mathrm{SPD}=P(\hat{y}=1\mid A=a)-P(\hat{y}=1\mid A=b).$
+    (4) $\mathrm{SPD}=P(\hat{y}=1\mid A=a)-P(\hat{y}=1\mid A=b).$
 […]
@@
     Fairness ratio (disparate impact, DI)
-    $\mathrm{DI}=\frac{P(\hat{y}=1\mid A=a)}{P(\hat{y}=1\mid A=b)}.$
+    (5) $\mathrm{DI}=\frac{P(\hat{y}=1\mid A=a)}{P(\hat{y}=1\mid A=b)}.$
     This KPI measures the outcome parity as a ratio. It provides guard against division by zero; also report the symmetric form $\min(\mathrm{DI},1/\mathrm{DI})$.
@@
     Equal performance (accuracy parity)
-    $\Delta_{\mathrm{acc}}=\left|\mathrm{Accuracy}_{A=a}-\mathrm{Accuracy}_{A=b}\right|.$
+    (6) $\Delta_{\mathrm{acc}}=\left|\mathrm{Accuracy}_{A=a}-\mathrm{Accuracy}_{A=b}\right|.$
     This measures the absolute gap in a stated accuracy metric across subgroups (Raza et al., 2024). It reports per subgroup and key intersections; keep $t$ fixed across subgroups.
@@
     Robustness assessment
-    $R=\frac{\mathrm{Errors\ under\ perturbation}}{\mathrm{Total\ perturbed\ inputs}}\times 100,\quad\Delta_{\mathrm{acc}}^{\mathrm{rob}}=\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{pert}}.$
+    (7) $R=\frac{\mathrm{Errors\ under\ perturbation}}{\mathrm{Total\ perturbed\ inputs}}\times 100,\quad\Delta_{\mathrm{acc}}^{\mathrm{rob}}=\mathrm{Acc}_{\mathrm{clean}}-\mathrm{Acc}_{\mathrm{pert}}.$
     It measures the error rate and accuracy drop under stated threat models.
@@
     Explainability index (author defined)
-    $E(x)=\sum_{i=1}^{n}\left|\phi_{i}(x)\right|.$
+    (8) $E(x)=\sum_{i=1}^{n}\left|\phi_{i}(x)\right|.$
     This KPI measures the attribution magnitude summary for an instance using SHAP/IG (Lundberg and Lee, 2017; Sundararajan et al., 2017; Vilone and Longo, 2021).
@@
     Local faithfulness gap (XAI)
-    $\mathrm{LFG}(x)=\frac{\left|\,f(x)-\left(\phi_{0}+\sum_{i}\phi_{i}(x)\right)\right|}{\max\left(\left|f(x)\right|,\epsilon\right)}.$
+    (9) $\mathrm{LFG}(x)=\frac{\left|\,f(x)-\left(\phi_{0}+\sum_{i}\phi_{i}(x)\right)\right|}{\max\left(\left|f(x)\right|,\epsilon\right)}.$
     This KPI measures the relative completeness error of additive explanations (lower is better).
@@
     High stakes error rate (author defined)
-    $E_{h}=\frac{\mathrm{Critical\ errors}}{\mathrm{Total\ predictions}}\times 100.$
+    (10) $E_{h}=\frac{\mathrm{Critical\ errors}}{\mathrm{Total\ predictions}}\times 100.$
     This measures incidence of errors that trigger a predefined harm tier (Kunreuther et al., 2002).
@@
     Audit frequency and resolution time
-    $F_{a}=\frac{\mathrm{Audits}}{\mathrm{Time\ period}},\qquad T_{r}=\mathrm{median\ time\ to\ resolve}.$
+    (11) $F_{a}=\frac{\mathrm{Audits}}{\mathrm{Time\ period}},\qquad T_{r}=\mathrm{median\ time\ to\ resolve}.$
     This KPI is cadence of audits and timeliness of remediation (Díaz-Rodríguez et al., 2023; Hasan, 2021). We specify window and scope; report $T_{r}$ with IQR and the share closed within SLA.
@@
     User inclusivity index (author defined)
-    $U=\frac{\mathrm{Inclusive\ features\ implemented}}{\mathrm{Planned\ features}}\times 100.$
+    (12) $U=\frac{\mathrm{Inclusive\ features\ implemented}}{\mathrm{Planned\ features}}\times 100.$
     It measures the accessibility and localization coverage (Shams et al., 2023). We enumerate the checklist (e.g., WCAG conformance, screen reader support, localization) and review cadence.
@@
     Sustainability (energy and emissions)
-    $E_{\mathrm{kWh}}=\frac{P_{\mathrm{avg}}}{1000}\times T,\qquad C_{\mathrm{CO_{2}e}}=E_{\mathrm{kWh}}\times\mathrm{PUE}\times\mathrm{CI}.$
+    (13) $E_{\mathrm{kWh}}=\frac{P_{\mathrm{avg}}}{1000}\times T,\qquad C_{\mathrm{CO_{2}e}}=E_{\mathrm{kWh}}\times\mathrm{PUE}\times\mathrm{CI}.$
     This KPI measures energy use and estimated emissions; pair energy with location specific or market based carbon intensity and data center overhead (Van Wynsberghe, 2021).
@@
 Sections 56 use this rubric (C1–C10) and the lifecycle KPIs as a running lens to interpret domain evidence and highlight where current practice still lacks governance-ready coverage.
+Recent GenAI era deployments expose failure modes (hallucination, privacy leakage, deepfakes, and tool-use errors) that cut across domains; we therefore use C1–C10 and the KPI set as a consistent lens throughout the remainder of the survey.
 4.3. Explainable AI
-Recent GenAI era deployments expose failure modes (hallucination, privacy leakage, deepfakes, and tool-use errors) that cut across domains; we therefore use C1–C10 and the KPI set as a consistent lens throughout the remainder of the survey.
-Having established how benchmarks cover risk families, we examine explainability methods and what they can and can not do, evidence under audit.
 Explainable AI (XAI) comprises methods that render model behaviour and outputs intelligible to human stakeholders (Meske et al., 2022).
-Effective explanations serve multiple roles: they support accountability and trust, enable bias and safety analyses, and facilitate compliance with emerging governance requirements (Baker and Xiang, 2023; Arrieta et al., 2020).
+Effective explanations serve multiple roles: they support accountability and trust, enable bias and safety analyses, and facilitate compliance with emerging governance requirements.
 In practice, XAI must operate across technical, social, legal, and organisational layers rather than as a purely algorithmic add-on.
 Figure 8.
-Post-hoc explainability within a Responsible AI (RAI) pipeline.
+Post-hoc explainability within a ResGenAI.
 A black-box model produces predictions $\hat{y}$ and attribution signals $I(x_{i})$, where $I(x_{i})$ denotes a feature-importance signal (e.g., gradient/saliency or a SHAP value).
 Interpretability pathways (feature relevance, example-based, surrogates, and visual/text explanations) feed fairness diagnostics (e.g., SPD/DI) and calibration analyses, supporting governance-aligned optimisation $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}+\lambda\,\mathcal{L}_{\mathrm{fair}}$.
-For LLMs/VLMs, prompt traces and cross-modal attention provide additional observability
-Scope and challenges in the GenAI
+For LLMs/VLMs, prompt traces and cross-modal attention provide observability
 Foundation models raise explanation challenges beyond conventional ML.
@@
 XAI methods can be ante-hoc (interpretability built into the model class, e.g., sparse linear models, rule lists, decision trees) or post-hoc (explain an existing black-box locally or globally).
-Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016a); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023).
-For LLMs and Vision-language-models (VLMs), these are complemented by prompt-engineering probes, chain-of-thought (CoT) audits, and tool-use traces (Natarajan and Nambiar, 2024).
+Common post-hoc families include feature relevance (saliency/attribution, gradients, SHAP/LIME) for instance-level insight (Lundberg and Lee, 2017; Ribeiro et al., 2016b); example-based explanations (prototypes, nearest neighbours) that ground behaviour in the data manifold (Kenny et al., 2021); surrogate models that approximate decision logic with a transparent model; and visual/textual rationales (heatmaps, attention summaries) for human inspection (He et al., 2023).
+For LLMs and VLMs, these are complemented by prompt-engineering probes, chain-of-thought (CoT) audits, and tool-use traces (Natarajan and Nambiar, 2024).
 Importantly, attention maps and saliency are not proofs of causal reasoning and should be triangulated with counterfactual tests and ablations (Jain and Wallace, 2019).
@@
 Fairness is evaluated with standard disparity measures such as Statistical Parity Difference (SPD) (Kun, 2015) and Disparate Impact (DI) (Barocas and Selbst, 2016):
-$\text{SPD}=P(\hat{y}=1\,|\,A=a)-P(\hat{y}=1\,|\,A=b),\quad\text{DI}=\frac{P(\hat{y}=1\,|\,A=a)}{P(\hat{y}=1\,|\,A=b)}.$
+(14) $\text{SPD}=P(\hat{y}=1\,|\,A=a)-P(\hat{y}=1\,|\,A=b),\quad\text{DI}=\frac{P(\hat{y}=1\,|\,A=a)}{P(\hat{y}=1\,|\,A=b)}.$
 A practical training objective combines task loss with an explicit fairness regulariser,
-$\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}(\theta)+\lambda\,\mathcal{L}_{\mathrm{fair}},$
+(15) $\mathcal{L}_{\mathrm{RAI}}=\mathcal{L}(\theta)+\lambda\,\mathcal{L}_{\mathrm{fair}},$
 allowing developers to trade off utility and fairness under governance constraints.
 For trust calibration, we recommend standard metrics such as Expected Calibration Error (ECE) (Pleiss et al., 2017) or the Brier score (Rufibach, 2010); ECE summarises the gap between confidence and empirical accuracy across bins:
-$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\big|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\big|.$
+(16) $\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\big|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\big|.$
 Adding canonical sources: ECE and Brier are standard measures (Pleiss et al., 2017). Calibrated probabilities, paired with faithful explanations, yield more reliable user-trust assessments.
@@
 The predictive case $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ is recovered as a deterministic special instance
-$\mathcal{E}:(x,f_{\theta},\omega)\mapsto z\in\mathcal{Z},$
+(17) $\mathcal{E}:(x,f_{\theta},\omega)\mapsto z\in\mathcal{Z},$
 where $\omega$ indexes stochasticity (decoding temperature, top-$k/p$).
@@
 For foundation models, explanations should expose (i) prompt pathway elements (system/user prompts, retrieved context, tool calls), (ii) decoding state ($\omega$: temperature, top-$k/p$), and (iii) cross-modal grounding (text rationales linked to image regions or clips).
-These artefacts do not prove causality but enable stability and counterfactual checks and support post-hoc accountability (Rane and Paramesha, 2024; Kenny et al., 2021).
+These artefacts do not prove causality but enable stability and counterfactual checks and support post-hoc accountability.
 Explanations integrate naturally into risk-managed optimisation:
-$\min_{\theta}\ \mathbb{E}_{x,\omega}\!\left[\ell\big(f_{\theta}(x,\omega),y\big)\right]+\lambda_{\mathrm{fair}}\mathcal{L}_{\mathrm{fair}}+\lambda_{\mathrm{xai}}\mathcal{R}\!\left(\mathcal{E}(x,f_{\theta},\omega)\right),$
+(18) $\min_{\theta}\ \mathbb{E}_{x,\omega}\!\left[\ell\big(f_{\theta}(x,\omega),y\big)\right]+\lambda_{\mathrm{fair}}\mathcal{L}_{\mathrm{fair}}+\lambda_{\mathrm{xai}}\mathcal{R}\!\left(\mathcal{E}(x,f_{\theta},\omega)\right),$
 where $\mathcal{R}$ encodes desiderata such as stability (penalising high $\delta$) and parsimony (penalising $\Omega$).
 This connects fairness (SPD/DI), calibration (ECE), and logging to the governance criteria C1–C10 used in our evaluation.
-Limitation: Even theoretically sound explanations can mislead when they are plausible but unfaithful (e.g., over-reading attention maps), trigger Goodhart effects (Strathern, 1997b) (optimising the appearance of fairness rather than outcomes), or lack distributional validity (narratives drift across domains).
+Limitation: Even theoretically sound explanations can mislead when they are plausible but unfaithful (e.g., over-reading attention maps), trigger Goodhart effects (Strathern, 1997) (optimising the appearance of fairness rather than outcomes), or lack distributional validity (narratives drift across domains).
 Reporting stability, counterfactual checks, and provenance mitigates these risks and makes explanations suitable as audit evidence rather than marketing artefacts.
 Based on our synthesis findings, there is a dangerous ”Explanation Gap” in how we use these tools.
-While SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016b) provide useful snapshots, the literature reveals they often create ”plausible but unfaithful” stories.
-They give us a narrative that makes sense to a human, but because transformers are so complex and non-linear, these stories often don’t match what is actually happening under the hood (Strathern, 1997a).
+While SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016a) provide useful snapshots, the literature reveals they often create ”plausible but unfaithful” stories.
+They give us a narrative that makes sense to a human, but because transformers are so complex and non-linear, these stories often don’t match what is actually happening under the hood (Strathern, 1997).
 This creates a false sense of security that we think we’ve reached transparency when we’ve really just seen a surface-level summary.
@@
 HUMANE AI (HumaneAI, 2020) Human-centric AI Platform/Initiative Alignment with human values; societal-impact and fairness guardrails; guidance for participatory, responsible design.
-AI for Good (ITU) ((2017), ITU) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
+AI for Good (ITU) ((ITU), 2017) Social good / RAI Initiative/Program UN SDG–aligned use cases; ethics and societal-benefit framing; showcases reproducible, responsible deployments.
 UKRI Trustworthy Autonomous Systems (TAS) Hub (Research and Innovation, 2020) Autonomous systems Program/Test-bed Accountability and transparency for autonomy; safety cases; conformance/compliance support for real-world systems.
 Algorithmic Justice League (Algorithmic Justice League, 2024) Algorithmic fairness Initiative Bias discovery and awareness; audits and mitigation advocacy; widely referenced in fairness evaluation contexts (not a formal test-bed).
-AI Commons [^3] https://ai-commons.org/ Open AI collaboration Initiative Open, transparent collaboration patterns; governance-mindful sharing; community evaluation practices (not a formal test-bed).
+AI Commons Open AI collaboration Initiative Open, transparent collaboration patterns; governance-mindful sharing; community evaluation practices (not a formal test-bed).
 ClarityNLP (Healthcare) (ClarityNLP, 2021) Healthcare NLP Platform Clinical text pipelines; emphasis on fairness, transparency, and ethical data use; supports domain-specific evaluation.
@@
 AI Fairness 360 (AIF360) (Bellamy et al., 2018) Fairness evaluation Toolkit Open-source Python toolkit with 70+ fairness metrics and 9 mitigation algorithms; supports bias detection, fairness explanations and interactive web interface for non-experts.
-Zeno (Zeno, 2023) AI evaluation Platform Interactive evaluation platform; enables data exploration across any data type and automatically discovers systematic failures using slice finder; supports interactive charts to compare evaluation metrics across slices and models.
-Galileo (Labs, 2023) Generative AI evaluation Platform Automated evaluation metrics for GenAI (RAG and agent metrics plus safety/security metrics); real-time guardrails to block hallucinations, PII and prompt injection; captures corner cases and allows custom metrics.
-Arize AI (Dong et al., 2024) AI observability Platform AI observability and LLM evaluation platform that automatically detects model issues, traces root causes via indexed training/validation/production data and fine-tunes performance; includes dashboards for evaluation/performance tracing, explainability and fairness.
-The Bias Tracing tool monitors fairness metrics (e.g., recall parity, false positive parity, disparate impact) and surfaces features and cohorts contributing to bias.
 PromptLayer (PromptLayer, 2022) Prompt engineering Platform Prompt-evaluation platform with visual pipelines for scoring prompts against golden datasets, backtesting with historical data, regression testing and continuous integration of evaluation pipelines; useful for systematic RAG/agent evaluations.
@@
 4.4.1. The Lifecycle Evidence Matrix: Mapping Metrics to Engineering Controls
-To facilitate the implementation of our proposed evaluation infrastructure, we synthesize the relationship between KPIs, benchmarks, and technical tools.
-As argued throughout this survey, the effectiveness of Responsible GenAI hinges on the ability to produce portable, verifiable evidence across the lifecycle.
+To facilitate the implementation of our proposed evaluation, we synthesize the relationship between KPIs, benchmarks, and technical tools.
+Table.
 Table 7 provides a decision-support mapping that links the KPIs defined in Section 4.2.4 to the specific testbeds and datasets that operationalize them.
@@
 KPI Category Primary KPIs Benchmarks & Datasets Audit Toolkits & Testbeds
-Fairness & Bias SPD, DI, $\Delta_{\mathrm{acc}}$ SafetyBench (Zhang et al., 2024), HumaniBench (Raza et al., 2025) AIF360 (Bellamy et al., 2018), Fairlearn (Weerts et al., 2023), Arize (Dong et al., 2024)
-Safety & Toxicity $E_{h}$ (High-stakes error) HarmBench (Mazeika et al., 2024), SALAD-Bench (Li et al., 2024a) Unitary (Ltd., [n.
-d.]), Project Moonshot (AI Verify Foundation, 2024), Galileo (Labs, 2023)
-Robustness & Security $R$ (Robustness), ASR ALERT (Tedeschi et al., 2024), Rainbow Teaming (Samvelyan et al., 2024) ART (Research, [n.
-d.]), PromptLayer (PromptLayer, 2022), ToolSandbox (Lu et al., 2024a)
-Privacy $P$ (Compliance rate) PrivLM-Bench (Li et al., 2024b), PrivacyLens (Shao and et al., 2024) Privacy Meter (Privacy and Lab, [n. d.]), Google DP (Google, [n. d.]), SecretFlow (Group, [n. d.])
-Explainability $E(x)$, LFG (Faithfulness) HELM (Liang et al., 2023), XSTest (Röttger et al., 2024) LIT (Google PAIR, [n. d.]), InterpretML, AIX360 (Arya et al., 2019)
-Sustainability $E_{\mathrm{kWh}}$, $C_{\mathrm{CO_{2}e}}$ CodeCarbon (Song et al., 2025) Evidently (Evidently AI, [n. d.]), Arize (Dong et al., 2024), Zeno (Zeno, 2023)
+Fairness & Bias SPD, DI, $\Delta_{\mathrm{acc}}$ SafetyBench (Zhang et al., 2024), HumaniBench (Raza et al., 2025) AIF360 (Bellamy et al., 2018), Fairlearn (Weerts et al., 2023)
+Safety & Toxicity $E_{h}$ (High-stakes error) HarmBench (Mazeika et al., 2024), SALAD-Bench (Li et al., 2024b) Unitary (Ltd., ), Project Moonshot (AI Verify Foundation, 2024), Galileo (Labs, 2023)
+Robustness & Security $R$ (Robustness), ASR ALERT (Tedeschi et al., 2024), Rainbow Teaming (Samvelyan et al., 2024) ART (Research, ), PromptLayer (PromptLayer, 2022), ToolSandbox (Lu et al., 2024a)
+Privacy $P$ (Compliance rate) PrivLM-Bench (Li et al., 2024a), PrivacyLens (Shao and et al., 2024) Privacy Meter (Privacy and Lab, ), Google DP (Google, 2025), SecretFlow (Group, )
+Explainability $E(x)$, LFG (Faithfulness) HELM (Liang et al., 2023), XSTest (Röttger et al., 2024) LIT (Google PAIR, ), InterpretML, AIX360 (Arya et al., 2019)
+Sustainability $E_{\mathrm{kWh}}$, $C_{\mathrm{CO_{2}e}}$ CodeCarbon (Song et al., 2025) Evidently (Evidently AI, ),
 4.4.2. Turning Metrics into Assurance: The ResGenAI Audit Loop
@@
 As visualized in Fig.
-10, our framework proposes a three layer loop that bridges the Evaluation Gap identified in Section 6.4.
+9, our framework proposes a three layer loop that bridges the ”Evaluation Gap identified”.
 In this cycle, the technical KPIs defined in Section 4.2.4 act as the raw sensory data for the ecosystem.
-These metrics are packaged into standardized audit artifacts such as model cards and red teaming logs which serve as the portable evidence required by compliance auditors under the EU AI Act (European Union, 2024) or NIST AI RMF (of Standards and, NIST).
+These metrics are packaged into standardized audit artifacts such as model cards and red teaming logs which serve as the portable evidence required by compliance auditors under the EU AI Act or NIST AI RMF.
 By establishing this formal “chain of trust”, organizations can move beyond static safety checks toward continuous stewardship where regulatory feedback directly informs the recalibration of technical metrics.
 Figure 9. The ResGenAI Audit Loop: translating technical KPIs into portable audit evidence to support regulatory assurance and continuous stewardship across the foundation model lifecycle.
+Next, we discuss the ResGenAI across domains and map how these aforementioned methods can be applied in these applications.
 5. Responsible Generative AI Application Across Domains
-GenAI systems are increasingly embedded in decision-making processes across diverse domains such as healthcare, finance, human resource, defense and education.
-While these systems offer significant utility, they also introduce new categories of risks and challenges, as outlined earlier.
-The challenges and mitigation strategies synthesized in this review are broadly relevant, but their salience and priority vary substantially by sectoral context.
-We translate the above findings into stakeholder-aligned practices and a deployment-oriented roadmap, with examples and decision pointers.
-Each subsection below examines domain-specific applications, challenges, and strategies through the lens of our governance-aligned rubric (C1–C10, Section 4.2.1) and lifecycle KPIs (Section 4.2.4).
-For every domain we (i) identify which rubric criteria are most salient, (ii) assess how the benchmarks in Tables 46 cover those risks, and (iii) indicate which KPIs practitioners should prioritize.
-Table 9 consolidates this analysis, mapping each sector to its primary criteria gaps and priority KPIs.
+GenAI systems are increasingly embedded in decision-making processes across diverse domains such as healthcare, finance, education, defense, arts, and agriculture.
+Each subsection below examines domain-specific risks and strategies.
+For every domain we identify which rubric criteria (C1–C10) are most salient and indicate which KPIs practitioners should prioritize.
+Appendix Table 6 consolidates this analysis, mapping each sector to its primary criteria gaps and priority KPIs.
+Cross-cutting strategies.
+Several responsible deployment strategies recur across nearly all high-stakes domains and form a shared baseline.
+Fairness audits (C1, KPIs 3–5) using SPD, DI, and accuracy-parity metrics should be conducted at regular intervals, stratified by relevant demographic or contextual variables.
+Privacy-preserving techniques (C6, KPI 2), including differential privacy, federated learning, and secure multiparty computation, enable cross-institutional collaboration without exposing sensitive records.
+Post-hoc explainability (C9, KPIs 7–8) via SHAP, LIME, or chain-of-thought audits supports stakeholder trust and regulatory interrogation of opaque models.
+Human-in-the-loop (HITL) oversight (C7, KPI 10) embeds expert review into high-consequence decision points, with audit frequency $F_{a}$ and resolution time $T_{r}$ tracked as governance evidence.
+Regulatory alignment (C10) anchors deployments within enforceable legal standards (e.g., EU AI Act, NIST AI RMF, sector-specific mandates).
+The subsections below focus on what is distinctive about each domain’s risk profile and mitigation needs beyond this shared baseline.
 5.1. Healthcare
-Applications
-Generative models improve efficiency in clinical note generation and triage chatbots (Alagha and Ahn, 2023), they raise concerns about accuracy, liability, and bias in medical decision support.
-Disease detection and drug discovery pipelines (Liu et al., 2023c) illustrate GenAI’s potential for accelerating science, yet demand rigorous validation, explainability, and regulatory alignment before clinical deployment.
-Synthetic health data helps mitigate privacy risks and dataset imbalance (Walsh et al., 2023), but governance frameworks must ensure that synthetic outputs do not inadvertently reinforce hidden biases.
-Similarly, tools like Watson for Oncology highlight the need for transparency and interpretability in high-stakes decision-making (Memorial Sloan Kettering Cancer Center, 2024).
-Even promising breakthroughs, such as GenAI-designed therapeutics, require human oversight on trial integrity, equitable access, and long-term monitoring.
-In this way, healthcare exemplifies both the promise of GenAI and the critical role of RAI practices, balancing innovation with safety, accountability, and fairness.
-ResGenAI challenges
-Responsible adoption of GenAI in healthcare is constrained by privacy violations (Shiraz and Lee, 2019), hallucination of critical clinical facts (Li et al., 2023b), algorithmic bias in diagnostics (Obermeyer et al., 2019), unclear accountability for errors (Kaminski, 2022), opacity of generative models (Ghorbani et al., 2019), and difficulty of compliance with traditional regulations (Price and Gerke, 2022).
-Empirical studies highlight these risks, for example, a 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023b); Rosbach et al.
-further found a 7% automation-bias error rate in a pathology study where clinicians overturned correct diagnoses by over-trusting AI outputs (Rosbach et al., 2024); and a systematic review of 778 studies showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023).
-Evaluation context
-Mapping these risks to our rubric, the dominant healthcare concerns correspond to C4 (hallucination of clinical facts), C1 (diagnostic bias), and C6 (patient data leakage).
-Only 7 of 21 benchmarks in Table 4 score above zero on C4 (Misinformation/ Disinformation), and none specifically target medical hallucination.
-For privacy, only PrivLM-Bench and PrivacyLens provide strong C6 (Privacy) coverage, however, neither addresses HIPAA-specific leakage vectors such as clinical note memorization.
-C7 (System-level failure) is critical for RAG-based clinical pipelines and is tested by only 4 of 21 benchmarks.
-The KPIs most relevant to this domain are the high-stakes error rate $E_{h}$ (KPI 9), operationalized as hallucinated clinical facts per $N$ consultations, data privacy compliance $P$ (KPI 2) under HIPAA, and fairness metrics SPD/DI (KPIs 3–4) stratified by patient demographics.
-ResGenAI strategies
-Several strategies are being employed to enable responsible GenAI adoption in healthcare.
-Grounding methods (C4, KPI 9) such as RAG can help reduce hallucinations by linking model outputs to verifiable medical sources.
-Effectiveness should be tracked via $E_{h}$; a reported 1.47% hallucination rate provides a baseline.
-Human-in-the-loop (HITL) validation (C7, C8, KPI 10) integrates clinician oversight into diagnostic and documentation workflows to mitigate automation bias.
-Audit frequency $F_{a}$ and resolution time $T_{r}$ should be reported.Privacy-preserving approaches (C6, KPI 2), including differential privacy and federated learning, protect sensitive health records while supporting multi-institutional collaboration.
-Fairness-aware data augmentation and structured fairness audits (C1, KPIs 3–5) are being applied to improve equity in clinical outcomes.
-HumaniBench (C1=2 in Table 4) offers a starting template but lacks clinical subgroup stratification.
-Post-hoc explainability (C9, KPIs 7–8) techniques such as SHAP and LIME are critical for clinician trust.Explanation quality should be assessed via the local faithfulness gap LFG (KPI 8).
-Alignment with regulatory frameworks (C10) like FDA guidance ensures safe and accountable deployment.
-Table 6 shows that only HELM and MLCommons v0.5 provide partial transparency support relevant to FDA conformity assessment.
+Applications and challenges.
+Generative models improve efficiency in clinical note generation and triage chatbots, yet raise concerns about accuracy, liability, and bias in medical decision support.
+Responsible adoption is constrained by hallucination of critical clinical facts, algorithmic bias in diagnostics, privacy violations, unclear accountability, and difficulty of compliance with sector-specific regulations.
+A 2023 evaluation of 450 consultation transcripts using GPT-4 reported a 1.47% hallucination rate, with 44% of errors classified as major and concentrated in treatment plans (Li et al., 2023a).
+A systematic review of 778 studies also showed that explainable AI can both increase and erode clinician trust, underscoring regulatory and accountability challenges (Tjoa and Guan, 2023; Raza et al., 2023).
+Domain-specific strategies.
+Beyond the shared baseline, healthcare demands grounding via retrieval-augmented generation (RAG) to link outputs to verifiable medical sources, with effectiveness tracked through the high-stakes error rate $E_{h}$ (KPI 9) operationalized as hallucinated clinical facts per $N$ consultations.
+Fairness-aware data augmentation stratified by patient demographics addresses diagnostic disparities that generic audits may miss.
+HITL validation is particularly critical given documented automation-bias effects where clinicians overturn correct diagnoses by over-trusting AI outputs.
+Alignment with FDA guidance and HIPAA-specific privacy requirements (e.g., clinical-note memorization vectors) distinguishes this domain from the general privacy baseline.
 5.2. Finance
-Applications
-GenAI is transforming multiple facets of finance, with impacts on efficiency, risk management, and personalization (Cao, 2022).
-Key applications include customer-facing bots that provide interactive banking and wealth-management support, modeling of fraud detection, credit scoring, risk forecasting, and insurance assessment (Cao, 2022), algorithmic trading that uses generative simulations to optimize portfolios and anticipate market shifts (HSBC Research, 2025), personalized financial advice adapted to client risk profiles and long-term goals (Wang et al., 2023b), automated compliance reports and market summaries that accelerate auditing and regulatory filings (Kaminski, 2022), and synthetic financial data generation for stress testing, scenario design, and risk modeling (Cao, 2022).
-Empirical deployments highlight these advances.
-For example, PayPal’s fraud detection reduced false positives by 50% in 2023 while processing 10 million transactions daily (Ngai et al., 2011).
-Charles Schwab’s assistant resolved 62% of client queries and improved satisfaction by 28% in 2023 (Wang et al., 2023b).
-JPMorgan’s simulator improved trading returns by 12% and reduced value-at-risk by 13% in 2023 (Zhang et al., 2023).
-HSBC’s AI achieved 95% accuracy in volatility prediction in 2025 (HSBC Research, 2025).
-Silva et al.
-also reported RegTech-GPT summaries rated “complete” by 85% of compliance officers compared with 62% baseline (Silva et al., 2024).
-ResGenAI challenges
-GenAI in finance raises a number of domain-specific risks that complicate its responsible adoption.
-These include discriminatory lending that replicates historical bias in credit and valuation data , security vulnerabilities where adversarial manipulation targets fraud detection or trading systems, consumer data privacy concerns arising from reliance on sensitive financial records, opacity of GenAI that limits auditability and accountability, hallucination and misleading outputs in customer-facing systems that may distort investment or compliance advice, and market manipulation risk that threatens systemic stability when generative simulations influence trading behavior (Hendry et al., 2022).
-Empirical use cases underscore these risks.
-For example, Zillow’s valuation model undervalued homes in minority neighborhoods and reduced appraised equity for underwriting (Adams et al., 2022).
-Citibank was also fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020).
-Evaluation context
-The primary finance risks map to C1 (lending discrimination), C3 (adversarial manipulation of trading/fraud systems), C6 (consumer data privacy), and C4 (hallucinated financial advice).
-No benchmark in Table 4 specifically targets lending discrimination or financial adversarial attacks.
-The most relevant KPIs are SPD and DI (KPIs 3–4) applied to loan approval rates across demographic groups, robustness $R$ (KPI 6) under financial adversarial scenarios, and audit frequency $F_{a}$ (KPI 10) aligned with SEC and OCC reporting cadences.
-ResGenAI strategies
-Multiple strategies are being prioritized to address these financial risks.
-Structured fairness audits (C1, KPIs 3–4) and fairness-aware data augmentation are increasingly required in lending and credit scoring.
-Post-hoc explainability (C9, KPIs 7–8) methods such as SHAP and LIME support regulators in interrogating opaque credit and trading models.
-Privacy-preserving techniques (C6, KPI 2), including differential privacy and secure multiparty computation, enable safe collaboration across financial institutions while protecting consumer data.
-Stronger data provenance frameworks are being developed to track the lineage and validity of synthetic datasets.
-Alignment with regulatory compliance (C10) frameworks such as Basel III, SEC, and OCC guidelines anchors AI deployments within enforceable legal standards (Kaminski, 2022).
-Finally, embedding HITL oversight (C7, KPI 10) across lending, trading, and compliance ensures accountability by mandating expert review of generative outputs in high-stakes decisions.
-5.3. Education
-Applications
-GenAI is increasingly shaping critical aspects of education, expanding opportunities for personalization, access, and efficiency (Holmes et al., 2022).
-Key applications include automated grading that accelerates evaluation and reduces instructor workload, personalized tutoring systems that adapt to student knowledge gaps and learning styles, content generation that produces adaptive curricula and study materials, accessibility tools that deliver multimodal and multilingual support for inclusive education, and interactive simulations that foster experiential and inquiry-driven learning (Chen and Huang, 2023; Lai and Xu, 2024).
-Recent deployments provide evidence of these advances.
-Carnegie Learning’s MATHia platform improved algebra test scores by 22% in U.S.
-high schools in 2023 (Carnegie Learning, 2023).
-A 2024 Arizona State University pilot found that students using an LLM-powered Socrates chatbot showed 42% higher engagement on surveys (Arizona State University, 2024).
-A Duolingo study in 2024 reported a 12% increase in learner retention and practice frequency through generative conversational exercises (Duolingo Research, 2024).
-A U.K.
-school district pilot in 2024 showed that 87% of students felt classroom chat assistants clarified concepts and 92% of teachers reported time saved (UK Department for Education, 2024).
-ResGenAI challenges
-Responsible use of GenAI in education faces several pressing concerns.
-These include cultural insensitivity where models fail to capture local norms and values, student data privacy and particularly minors’ data privacy given the sensitivity of educational records , skills evasion where over-reliance on AI weakens critical thinking and independent problem-solving, and pedagogical opacity that limits educators’ ability to evaluate how generative systems produce outputs (Zawacki-Richter et al., 2019).
-Empirical use cases underscore these risks.
-For example, India’s government-run “Newton” homework bot trained only on U.S.
-and U.K.
-corpora recommended “pork stew” to rural students in Rajasthan in 2021, violating cultural and nutritional guidelines and prompting suspension of the tool for eleven weeks (UNESCO GEM Report, 2022).
-A 2023 randomized controlled trial at the University of Cyprus found that physics students allowed to use GPT-4 for weekly assignments scored twenty-three percent lower on conceptual-transfer exam questions despite higher weekly grades, demonstrating over-reliance effects (Pavlou et al., 2023).
-In 2022, the Utah State Board of Education cancelled a contract for an AI essay-scorer when auditors reported that the vendor could not document the feature vectors or training data, raising concerns about pedagogical opacity and compliance with state education law (Utah State Board of Education, 2022).
-Evaluation context
-Education risks concentrate on C6 (minors’ data privacy), C4 (hallucinated pedagogical content), C1 (cultural bias), and C7 (system-level failure in tutoring pipelines).
-No benchmark addresses C6 in the context of minors’ data, and the user inclusivity index $U$ (KPI 11) critical for accessibility in educational settings, lacks any standardized evaluation protocol.
-Skills-evasion effects (a form of C7 system-level failure) are entirely absent from current safety suites.
-Priority KPIs include $P$ (KPI 2) under COPPA/FERPA, $E_{h}$ (KPI 9) for pedagogically harmful outputs, and $U$ (KPI 11) for WCAG conformance.
-ResGenAI strategies
-Multiple strategies are being explored to ensure responsible generative AI in education.
-Fairness audits (C1, KPIs 3–4) and external ethics audits verify alignment with frameworks such as the UNESCO(UNESCO, 2021a) AI-ethics rubric and can keep bias metrics below accepted thresholds.
-Privacy-preserving techniques (C6, KPI 2) including homomorphic encryption and federated learning protect minors’ sensitive data while enabling large-scale analysis (Kaissis et al., 2021).
-Clear regulation and policy guidelines (C10) developed by international and national bodies provide governance and accountability mechanisms.
-Finally, academic integrity tools (C5, C9) such as watermarking and output tracking are being adopted to prevent plagiarism and monitor AI-assisted work (Zhang and Zhao, 2024).
-5.4. Arts and Entertainment
-Applications
-GenAI is reshaping creative production and cultural participation in the arts and entertainment sector.
-Prominent applications include creative writing tools that assist in script, story, and lyric development, generative art and design platforms that expand aesthetic experimentation, music composition and production systems that automate melody, arrangement, and mastering, film production support for visual effects, dubbing, and storyboarding, personalized recommendation engines that curate media consumption at scale and virtual influencers and avatars that shape marketing and fan engagement (Crawford, 2021).
-Recent deployments provide evidence of these advances.
-For instance, Spotify’s Discover Weekly playlists improved user engagement by 40% in 2023, with subsequent adjustments in 2024 reducing genre bias (Anderson and Lee, 2023).
-A 2024 marketing campaign featuring AI-driven virtual influencers reached more than ten million users (Crawford, 2021; Traitter and Klein, 2022).
-Davis et al.
-released a ten-track AI album in which 62% of listeners in a Spotify trial could not distinguish the music from human compositions (Davis et al., 2024).
-ResGenAI challenges
-Responsible use of GenAI in arts and entertainment faces a series of ethical, legal, and cultural concerns.
-These include copyright infringement and intellectual property violation where AI systems reuse creative material without authorization, job displacement that threatens creative labor markets, cultural insensitivity when generative tools fail to respect community values or norms deepfakes and misinformation that erode authenticity and trust, authorship and agency disputes where creative ownership is contested, insufficient consent in data usage for training models, and opaque attribution that obscures the role of AI in the creative process (Schneider and Jörgens, 2022; Elgammal et al., 2019).
-Empirical use cases underscore these risks.
-For example, the U.S.
+Applications and challenges.
+GenAI is transforming fraud detection, credit scoring, risk forecasting, and customer-facing advisory.
+Domain-specific risks include discriminatory lending that replicates historical bias in credit and valuation data, adversarial manipulation of trading and fraud systems, consumer data privacy across jurisdictions, and hallucinated financial advice in customer-facing systems.
+Zillow’s valuation model undervalued homes in minority neighborhoods, reducing appraised equity for underwriting (Adams et al., 2022).
+Citibank was fined twenty-five million dollars in 2020 when its flood-zone underwriting model failed auditability standards (Office of the Comptroller of the Currency, 2020).
+Domain-specific strategies.
+Structured fairness audits in lending must go beyond standard SPD/DI to address intersectional and proxy-variable discrimination embedded in credit histories.
+Stronger data provenance frameworks are needed to track the lineage and validity of synthetic financial datasets.
+Robustness testing (C3, KPI 6) against financial adversarial scenarios (e.g., adversarial inputs to trading models) is sector-critical but absent from current benchmarks in Table 3.
+Audit frequency $F_{a}$ (KPI 10) should align with SEC and OCC reporting cadences, and secure multiparty computation enables cross-institutional risk modelling without exposing proprietary data.
+5.3. Defense
+Applications and challenges.
+GenAI is increasingly used for simulation and war-gaming, multi-INT threat detection, autonomous reconnaissance, and logistics optimization.
+Domain-specific risks include autonomous weapons and responsibility gaps, opacity of decision-making in high-stakes targeting, adversarial vulnerabilities (spoofing, data poisoning), escalation of conflicts through compressed decision windows, and dual-use research security.
+National and alliance exercises reported that AI-supported war-gaming surfaced escalatory flashpoints requiring additional human oversight (NATO Allied Command Transformation, 2024).
+Domain-specific strategies.
+Defense is the domain where red teaming (C3, C8, KPI 6) is most critical, yet only Rainbow Teaming and ALERT score above zero for adaptive adversary simulation in our assessment.
+Meaningful human control (C10) must keep decisions on the use of force within accountable human authority, consistent with international humanitarian law and NATO’s Principles of Responsible Use (NATO, 2024).
+Explainability requirements (C9) in defense emphasize traceability for after-action review rather than end-user understanding.
+System-level failure testing (C7) for multi-agent autonomous systems is entirely absent from current safety suites, representing the single largest benchmark gap for this sector.
+Ethical review boards and program-level governance embed policy, risk assessment, data governance, and model documentation across the AI lifecycle (U.S.
+Department of Defense, 2024).
+5.4. Other Domains: Education, Arts, and Agriculture
+Education.
+GenAI supports personalized tutoring, automated grading, and adaptive content generation, but responsible use is complicated by minors’ data privacy (C6), cultural insensitivity (C1), skills evasion where over-reliance on AI weakens critical thinking (C7), and pedagogical opacity.
+A 2023 randomized controlled trial at the University of Cyprus found that physics students using GPT-4 for assignments scored 23% lower on conceptual-transfer exam questions despite higher weekly grades (Pavlou and others, 2023).
+No current benchmark addresses C6 for minors’ data, and skills-evasion effects (a form of system-level failure) are entirely absent from safety suites.
+Domain-specific needs include compliance with COPPA/FERPA (KPI 2), academic integrity tools (C5, C9) such as watermarking and output tracking, and accessibility metrics aligned with WCAG (user inclusivity index $U$, KPI 11).
+Arts and entertainment.
+GenAI is reshaping creative production but raises concerns about copyright infringement (C6), deepfakes (C5), cultural insensitivity (C1), and contested authorship and attribution (C9).
+The U.S.
 Copyright Office’s 2023 ruling in Thaler v.
-Perlmutter rejected copyright protection for AI-generated works, reinforcing the requirement of human authorship (U.S.
+Perlmutter rejected copyright protection for AI-generated works (U.S.
 Copyright Office, 2023).
-The public controversy around the 2023 song “Heart on My Sleeve,” which mimicked the voices of Drake and The Weeknd, raised critical questions about derivative works and copyright boundaries.
-In 2025, WeTransfer updated its terms of service to allow user content for AI training, sparking backlash over transparency and consent (WeTransfer, 2025).
-Evaluation context
-The dominant risks map to C5 (deepfakes and media integrity), C1 (representational bias in generated content), and C9 (attribution transparency).
-Only SHIELD and MM-SafetyBench provides explicit C5 evaluation in Table 4, and its scope does not extend to voice cloning or music mimicry, the vectors most salient for this sector.
-Copyright memorization, a C6-adjacent concern, is not addressed by any privacy benchmark.
-Priority KPIs are currently underspecified for creative domains; provenance-verification rates and consent-compliance metrics (analogous to KPI 2) are needed but lack standardized protocols.
-ResGenAI strategies
-Multiple strategies are being adopted to promote responsible use of generative AI in arts and entertainment.
-Artist collaboration platforms and alternative compensation models are emerging to safeguard livelihoods and redistribute value.
-Structured bias audits (C1, KPIs 3–4) and cultural sensitivity training are being incorporated into creative AI pipelines to mitigate stereotyping .
-Deepfake detection tools (C5) are being deployed to maintain authenticity in music, video, and media (Chesney and Citron, 2019).
-Collective bargaining and union agreements such as those negotiated by SAG-AFTRA establish protections for creative workers (SAG-AFTRA, 2023).
-International ethical guidelines (C10) and transparency labeling requirements provide standards for disclosure and attribution.
-Finally, intellectual property management systems (C6, C9) are being advanced to secure consent and compliance.
-For example, a 2023 copyright detection tool identified AI-generated violations with 95% accuracy (Zirpoli, 2023a; Traitter and Klein, 2022), and these systems require explicit informed consent before creating digital replicas of performances, offering a contractual model for ethical rights management.
-5.5. Agriculture
-Applications
-GenAI is increasingly transforming agricultural practices through precision methods and data-driven innovation.
-Key applications include precision farming systems that optimize irrigation, fertilization, and pesticide usage, crop yield prediction models that integrate weather, soil, and satellite data, pest and disease detection tools that leverage generative augmentation to improve diagnostic accuracy, virtual agronomy advisors that deliver context-sensitive recommendations to farmers, agricultural robotics for automated harvesting and monitoring, crop breeding simulations that accelerate trait discovery, and weather forecasting for climate adaptation (Costa et al., 2023).
-While platforms like Microsoft’s FarmBeats (Microsoft Research, 2023) initially achieved significant yield and water-saving gains through IoT and predictive ML , GenAI is now being integrated to augment these systems through generative weather simulation for risk modeling and as LLM-based natural language interfaces .
-These tools leverage GenAI to translate complex sensor data into actionable, localized advice.
-ResGenAI challenges
-Responsible adoption of GenAI in agriculture faces several interrelated challenges.
-These include hallucination and context-inappropriate advice where unreliable outputs may jeopardize crop management decisions, farm data ownership conflicts where smallholders lack control over datasets aggregated by multinational platforms, bias toward major corporations and regions that limits the applicability of tools for smallholder and Global South farmers, overreliance on AI and erosion of local knowledge that risks displacing traditional farming wisdom, IoT security vulnerabilities that expose connected farm systems to cyberattacks, and the persistent digital divide that excludes low-resource communities from advanced AI benefits (Rotz et al., 2019; Sapkota et al., 2024).
-Empirical use cases underscore these risks.
-For example, farmers in sub-Saharan Africa reported reduced trust in AI weather advisors when predictions conflicted with lived knowledge, illustrating both hallucination and cultural misalignment.
-Cases of contested ownership emerged when farm sensor data collected by multinational seed companies were commercialized without equitable farmer compensation (Bronson, 2019).
-IoT breaches in 2023 European trials highlighted vulnerabilities in sensor networks that exposed proprietary farm-level data.
-Evaluation context
-Agriculture risks primarily concern C4 (hallucinated agronomic advice), C6 (farm data ownership and privacy), and C7 (system-level failure in IoT-integrated advisory pipelines).
-No benchmark in our assessment provides geo-aware hallucination testing, and C7 remains untested for sensor-to-advice chains.
-Priority KPIs are $E_{h}$ (KPI 9) operationalized as harmful agronomic recommendations per $N$ advisory sessions, $P$ (KPI 2) for data governance compliance, and $E_{\text{kWh}}$ (KPI 12) for sustainability reporting—particularly relevant given the energy constraints of low-resource deployment contexts.
-RAI strategies
-Multiple strategies are being pursued to enable responsible deployment of generative AI in agriculture.
-Geo-aware retrieval and annotation (C4, KPI 9) techniques are being designed to contextualize model outputs with local weather, soil, and cultural practices, reducing hallucination risks.
-Data governance contracts (C6, KPI 2) are being negotiated to clarify ownership rights and revenue-sharing models for farm-generated datasets.
-Embedding human-in-the-loop oversight (C7, KPI 10) into agronomy advisory workflows ensures farmers and extension officers validate AI outputs before implementation.
-Privacy-preserving IoT integration (C6, KPI 2), including differential privacy and federated learning, protects sensitive farm data across institutions.
-Finally, alignment with sustainability regulations (C10, KPI 12) such as FAO and EU Common Agricultural Policy frameworks anchors generative AI deployments in enforceable ecological and social standards.
-Together, these approaches enhance both the equity and safety of generative AI in global agriculture.
-5.6. Defense
-Applications
-GenAI is increasingly influencing defense through simulation, sensing, and decision support.
-Key applications include novel training and war-gaming environments that generate diverse scenarios and accelerate force development, threat detection and classification across multi-INT streams that fuse imagery signals and text for faster situational awareness, autonomous reconnaissance systems that coordinate teams of uncrewed assets in contested environments, and logistics and supply chain optimization that shortens planning cycles and improves mission readiness (News, 2024).
-Recent deployments provide evidence of these advances.
-For instance, DARPA programs such as CODE and successor autonomy efforts demonstrated collaborative uncrewed aircraft operating with minimal communications and supervisor control in flight tests (Defense Advanced Research Projects Agency, 2018).
-The U.S.
-Army extended enterprise decision platforms in 2024 to speed analysis and targeting workflows at scale (News, 2024).
-NATO war-games in 2024 integrated generative scenario simulation and reported that AI-generated vignettes helped surface unanticipated vulnerabilities during exercise analysis.
-RAI challenges
-Responsible adoption of GenAI in defense introduces ethical operational and legal concerns.
-These include autonomous weapons and responsibility gaps where assignment of accountability for the use of force remains unresolved, opacity of decision making that limits auditability and traceability in high-stakes contexts, adversarial vulnerabilities that expose perception models and decision aids to spoofing and data poisoning, escalation of conflicts where fast automated responses may narrow decision windows and increase miscalculation risk, ethical and legal gaps regarding compliance with international humanitarian law and rules of engagement, research security concerns over dual-use model capabilities and insider threat, protection of critical infrastructure as AI is integrated into command networks and sensors, and facial recognition deployment where demographic performance differences raise risks in security operations.
-Empirical use cases underscore these risks.
-For example, national and alliance exercises reported that AI-supported war-gaming surfaced escalatory flashpoints that required added human oversight (NATO Allied Command Transformation, 2024).
-Government red-team evaluations and standards work identified numerous novel failure modes in generative systems that could enable misinformation or degrade cyber defences if left unmitigated (National Institute of Standards and Technology, 2025b; WIRED, 2025).
-Facial recognition evaluations documented demographic differentials that necessitate strict control and testing before operational use (Grother et al., 2019, 2022).
-Evaluation context
-Defense risks concentrate on C3 (adversarial robustness), C7 (system-level failure in multi-agent autonomy), C8 (red-team realism), and C10 (governance alignment with IHL and NATO principles).
-C8 is the most critical gap: only 2 of 21 benchmarks (Rainbow Teaming and ALERT) score above zero for adaptive adversary simulation, yet defense applications face the most sophisticated threat actors.
-C7 is entirely absent for multi-agent autonomous systems.
-Priority KPIs are robustness $R$ (KPI 6) under military-grade adversarial scenarios, $E_{h}$ (KPI 9) for consequential targeting or escalation errors, and audit frequency $F_{a}$ (KPI 10) aligned with NATO after-action review cadences.
-RAI strategies
-Multiple strategies are being advanced to ensure responsible deployment.
-Human in the loop oversight (C7, C8, KPI 10) is prioritized for mission-critical tasks so that trained operators validate model outputs before action.
-Meaningful human control (C10) is reinforced to keep decisions on the use of force within accountable human authority consistent with international humanitarian law.
-Explainable AI for defense transparency (C9, KPIs 7–8) adopts traceability and interpretability requirements to support audit and after-action review.
-Red teaming (C3, C8, KPI 6) and rigorous test and evaluation are expanded to probe adversarial and safety failure modes before fielding.
-International cooperation (C10) aligns national practices with alliance principles such as NATO’s Principles of Responsible Use and shared assurance processes (NATO, 2024).
-Ethical review boards and program-level governance embed policy risk assessment data governance and model documentation across the AI life cycle (U.S.
-Department of Defense, 2024; Defense Innovation Unit, 2020).
-5.7. Cross-Applications Synthesis of GenAI Maturity
-To provide a critical comparison of how responsible practices are being operationalized across domains, and to connect sectoral risks to the evaluation infrastructure in Sections 4.24.4, we present the extended Synthesis Matrix in Table 9.
-Beyond risk mechanisms and regulatory drivers, the table now maps each sector to its primary rubric-criteria gaps and priority lifecycle KPIs, making explicit which evaluation evidence is available and where portable audit artifacts are missing.
-5.7.1. Stakeholder-Metric Alignment: Operationalizing the Framework
-To ensure the sectoral insights in Section 5 translate into actionable governance, we map the relationship between stakeholder roles and our evaluation infrastructure.
+Only SHIELD provides explicit deepfake evaluation in Table 3, and its scope excludes voice cloning and music mimicry.
+Domain-specific strategies include artist collaboration platforms and alternative compensation models, intellectual property management systems requiring explicit consent for digital replicas, and collective bargaining agreements such as SAG-AFTRA protections.
+Agriculture.
+GenAI supports precision farming, pest detection, and advisory chatbots, but responsible adoption faces hallucinated agronomic advice (C4), farm data ownership conflicts (C6), bias toward major corporations and regions (C1), and IoT security vulnerabilities (C3).
+Farmers in sub-Saharan Africa reported reduced trust in AI weather advisors when predictions conflicted with lived knowledge.
+No benchmark provides geo-aware hallucination testing or sensor-to-advice chain validation (C7).
+Domain-specific needs include geo-aware retrieval and annotation to contextualize outputs with local conditions, data governance contracts clarifying ownership and revenue-sharing for farm-generated datasets, and sustainability reporting (KPI 12) aligned with FAO and EU Common Agricultural Policy frameworks.
+5.5. Cross-Domain Synthesis of GenAI Maturity
+To provide a critical comparison of how responsible practices are being operationalized across domains, and to connect sectoral risks to the evaluation infrastructure in Sections 4.24.4, we present the extended Synthesis Matrix in Appendix Table 6 that maps each sector to its primary rubric-criteria gaps and priority lifecycle KPIs.
 Based on our synthesized findings, technical frameworks often fail because they do not specify which evidence is required by which actor.
-In our model, Model Developers utilize the robustness (C3) and toxicity (C2) scores from Table 4 to guide iterative alignment during the fine tuning phase.
-Conversely, Compliance Auditors and Regulators prioritize the statistical parity (SPD/DI) KPIs (Section 4.2.4) and governance-alignment criteria (C10) to verify adherence to mandates like the EU AI Act (Outeda, 2024) or NIST AI RMF (of Standards and, NIST).
-Finally, Domain Practitioners in high-stakes fields like Healthcare and Finance rely on the High-Stakes Error Rate ($E_{h}$) and local faithfulness gaps (LFG) to determine the necessary level of human-in-the-loop oversight.
-By bridging domain-specific mechanisms with normalized technical metrics, our framework provides the ”chain of trust” required to move ResGenAI from theoretical safety to verifiable institutional accountability.
-Table 9. Synthesis Matrix: Comparative Analysis of RAI Maturity, Regulatory Pressures, Rubric Criteria Gaps, and Priority KPIs Across Domains.
-Sector Dominant Risk Mechanism Technical Control Maturity Primary Regulatory Driver Primary C1 to C10 Gaps Priority KPIs
-Healthcare Medical hallucination (Lyu et al., 2025); Diagnostic bias (Gonzalez et al., 2022) Medium: RAG (Wu et al., 2024), HITL, Diff.
-Privacy (Kaissis et al., 2021) FDA Guidance (U.S.
-Food and Drug Administration, 2024); HIPAA (Shiraz and Lee, 2019) C4: no clinical hallucination benchmark; C6: no HIPAA specific privacy suite; C7: 4/21 cover system failure $E_{h}$ (9), $P$ (2), SPD/DI (3 to 4), $\Delta_{\text{acc}}$ (5)
-Finance Lending discrimination (Mehrabi et al., 2021b); Market manipulation (Hendry et al., 2022) High: Fairness audits (Pfohl et al., 2022); Multi party comp.
-(Khalid et al., 2023) SEC (U.S.
-Securities and Exchange Commission, 2021); OCC (Office of the Comptroller of the Currency, 2020) C1: no lending discrimination benchmark; C3: no financial adversarial attack suite SPD/DI (3 to 4), $R$ (6), $F_{a}$ (10)
-Education Student privacy (Livingstone et al., 2021); Skills evasion (Pavlou et al., 2023) Low: Watermarking (Jiang et al., 2023); Fed.
-learning (Yang, 2021) UNESCO AI Rubric (UNESCO, 2021b); State Edu.
-Laws (Utah State Board of Education, 2022) C6: no minors data protocol; C7: skills evasion untested; KPI 11 lacks standard $P$ (2), $E_{h}$ (9), $U$ (11)
-[…]
-Agriculture Ownership conflicts (Bronson, 2019); Global South exclusion (Tzachor et al., 2022) Low: Geo aware RAG (Ma et al., 2023); Data contracts (Bronson, 2019) FAO (Costa et al., 2023); EU CAP (Costa et al., 2023) C4: no geo aware hallucination test; C7: IoT advisory pipelines untested $E_{h}$ (9), $P$ (2), $E_{\text{kwh}}$ (12)
-Defense Autonomous lethality (International Committee of the Red Cross, 2021); Adversarial spoofing (National Institute of Standards and Technology, 2025a) High: Red teaming (Corporation, 2024); Meaningful HITL (International Committee of the Red Cross, 2021) NATO Principles (NATO, 2024); IHL (United Nations Institute for Disarmament Research, 2025) C8: only 2/21 benchmarks score >>0; C7: absent for multi agent autonomy $R$ (6), $E_{h}$ (9), $F_{a}$ (10)
-Legal & Justice Recidivism prediction bias (Skeem and Lowenkamp, 2016); Automated decision opacity (Chesterman, 2021) Low to Med.: Explainability (LIME/SHAP) (Microsoft Research, [n.
-d.]); Bias audits (Center for Data Science and Public Policy, [n.
-d.]) EU AI Act Art.
-6 (Edwards, 2021); U.S.
-DOJ Guidelines (Podgor, 2003) C1: no recidivism specific fairness benchmark; C2: legal advice toxicity untested; C10: judicial HITL criteria absent SPD/DI (3 to 4), $E_{h}$ (9), $R$ (6), $F_{a}$ (10)
-Manufacturing Predictive maintenance failures (Hashemian, 2010); Adversarial sensor spoofing (Cao et al., 2019) Medium: Digital twins (Mihai et al., 2022); Anomaly detection (Chandola et al., 2009) IEC 62443 (ISO/IEC, 2021b); EU Machinery Reg.
-(European Union, 2024) C3: no industrial adversarial robustness suite; C4: no sensor drift benchmark; C7: multi robot coordination untested $E_{h}$ (9), $R$ (6), $E_{\text{kWh}}$ (12), $F_{a}$ (10)
+In our model, Model Developers utilize the robustness (C3) and toxicity (C2) scores from Table 3 to guide iterative alignment during fine-tuning.
+Compliance Auditors and Regulators prioritize the SPD/DI KPIs (Section 4.2.4) and governance-alignment criteria (C10) to verify adherence to mandates like the EU AI Act (Outeda, 2024) or NIST AI RMF (of Standards and (NIST), 2024).
+Domain Practitioners in high-stakes fields like Healthcare and Finance rely on the high-stakes error rate ($E_{h}$) and local faithfulness gaps (LFG) to determine the necessary level of HITL oversight.
 6. Discussion
-6.1. Synthesis and Interpretation
-Our review shows that post-GPT scholarship concentrates on five areas: safety and benchmarking, fairness and governance, explainability and transparency, privacy and security, and sectoral applications.
-Bias and toxicity receive dense coverage, while deepfakes, sustainability, and system-level failure modes receive relatively less attention.
-This pattern aligns with the thematic trends we documented and with the growing prominence of multimodal and agentic systems in 2022–2025, which increases the need for evaluation that captures tool use and context.
-Closing the accountability chain
-To fully answer our primary research question, Who is Responsible?, our synthesis posits that the user must be transitioned from a passive consumer to a systemic guardrail.
-The current imbalance in GenAI governance, where 74% of organizations use GenAI but only 26% have a formal strategy, often results in the end-user inheriting unmitigated technical risks like PII leakage and hallucination.
-By triangulating the sectoral failures in Section 5 with the digital literacy gaps identified in Section 3.2.2, we propose a Symmetric Responsibility Model.
-In this model, developers are accountable for Transparency and Alignment, while users are accountable for Operational Integrity and Dissemination.
-Digital literacy, therefore, is not merely a skill but a core technical control that must be assessed alongside model-side benchmarks in Section 4.2.
-Without this user centric governance, even perfectly aligned models remain vulnerable to the last mile failure of human misinterpretation or malicious re contextualization.
-Table 10 operationalizes this model by mapping specific failure modes to the reciprocal duties of each stakeholder.
-Table 10. The Symmetric Responsibility Matrix: Mapping Failure Modes to Stakeholder Duties
-Failure Mode Developer Duty (Engineering Controls) User Duty (Operational Integrity)
-Hallucination & Misinfo Implement RAG frameworks and calibrated uncertainty reporting (KPI 9) (Wu et al., 2024). Verify factual claims via primary sources and cross modal grounding checks.
-PII & Data Leakage Deploy Differential Privacy and automated scrubbing during inference (KPI 2) (Nasr et al., 2023). Practice prompt hygiene by omitting sensitive or proprietary data from inputs.
-Adversarial Jailbreaking Conduct continuous red teaming and implement robust system guardrails (C8) (Samvelyan et al., 2024).
-Adhere to acceptable use policies and report emergent safety bypasses to developers.
-Representational Bias Perform fairness audits using intersectional metrics like SPD and DI (KPIs 3 to 4) (Kun, 2015).
-Critically appraise outputs for stereotyping and contextualize results within local norms.
-Over reliance (Automation Bias) Provide model confidence scores and explicit explanation gap warnings (KPI 8) (Strathern, 1997a).
-Maintain human in the loop (HITL) oversight and preserve independent domain expertise.
+Synthesis and Interpretation
+Post-GPT scholarship concentrates on five areas: safety and benchmarking, fairness and governance, explainability and transparency, privacy and security, and sectoral applications.
+Bias and toxicity receive dense coverage; deepfakes, sustainability, and system-level failure modes receive comparatively less.
+This pattern aligns with the growing prominence of multimodal and agentic systems in 2022–2025, which heightens demand for evaluations capturing tool use and context.
+To fully answer Who is Responsible?, our synthesis posits that users must transition from passive consumers to systemic guardrails.
+The current imbalance—74% of organizations use GenAI but only 26% have a formal strategy—routinely leaves end-users inheriting unmitigated risks such as PII leakage and hallucination.
+Triangulating the sectoral failures in Section 5 with the digital-literacy gaps in Section 3.2, we propose a Symmetric Responsibility Model: developers are accountable for Transparency and Alignment; users are accountable for Operational Integrity and Dissemination.
+Digital literacy is therefore not merely a skill but a core technical control, assessable alongside model-side benchmarks (Section 4.2).
+Without user-centric governance, even perfectly aligned models remain vulnerable to last-mile failures of misinterpretation or malicious recontextualization.
+Appendix Table 5 operationalizes this model by mapping failure modes to the reciprocal duties of each stakeholder.
 Industry Safety Artifacts and Advanced Interpretability.
-While academic benchmarks offer standardized metrics, the operational reality of ResGenAI is increasingly defined by safety documentation from frontier labs and emerging mechanistic interpretability techniques, as synthesized in Figure 11.
-Frontier labs have established documentation protocols as de facto benchmarks for governance facing evidence, directly mapping to our C1 to C10 rubric.
-OpenAI system cards for GPT 4 and GPT 4o provide granular audits for sycophancy and PII leakage, demonstrating high maturity applications of metric validity through traceable safety refusal evidence.
-Anthropic utilizes Constitutional AI to automate alignment via RLAIF, outlining tiered safeguards that address malicious actor realism.
-Meta Llama 3 documentation introduces standardized protocols for multilingual red teaming, vital for addressing cross cultural ethical gaps.
-Google DeepMind technical reports provide empirical evidence for bias and accuracy trade offs, detailing the overcorrection mechanisms that led to observed representational failures.
-This industry evidence is complemented by the shift toward mechanistic interpretability, which maps internal neural activations to semantic concepts to offer a more rigorous safety check than behavior level prompts alone.
-Prominent techniques like Sparse Autoencoders (SAEs) isolate circuits responsible for deceptive or biased reasoning, enabling model steering to suppress harmful features during inference.
-Integrating these mechanistic audits into the RAI lifecycle addresses the black box challenge by enabling empirical verification of internal safety alignment before deployment.
-Figure 11. Integration of industry safety documentation and advanced mechanistic interpretability for governance facing evidence.
-ResGenAI sustainability via lifecycle assessment
-Sustainability in ResGenAI must transition from qualitative discourse to rigorous quantification grounded in Lifecycle Assessment (LCA) frameworks, as integrated into our governance landscape (Figure 8) and synthesized in Figure 12.
-This necessitates formalizing carbon reporting and energy benchmarks across the development cycle, including pretraining, fine tuning, and inference.
-Following our proposed lifecycle KPIs (Section 4.2.4), we advocate for reporting energy consumption and carbon equivalent emissions normalized by throughput, such as per 1K tokens generated.
-Seminal research indicates that training large scale generative models produces environmental footprints comparable to the lifetime emissions of multiple automobiles, highlighting the need for transparent tracking tools and hardware aware reporting.
-A sustainable AI ecosystem requires a dual focus on optimizing the computational efficiency of models and leveraging generative insights to advance ecological goals, ensuring that technical innovation supports long term environmental stability.
-Figure 12.
-Quantifying Responsible GenAI Sustainability via Lifecycle Assessment (LCA) in ResGenAI Governance.
-This framework formalizes carbon reporting and energy benchmarks across the pretraining, fine tuning, and inference stages, enabling a transition to rigorous quantification and supporting a sustainable AI ecosystem.
-6.2. Lessons from Industry Deployment Incidents
-While our benchmark analysis identifies risk-coverage gaps analytically, real-world deployment incidents from major AI developers offer empirical grounding.
-We synthesize prominent cases below, mapped to our rubric criteria, to demonstrate how the gaps in Tables 36 manifest in production systems.
-Alignment and safety failures (C2, C3, C8) Microsoft Bing Chat integration of GPT-4 in early 2023 produced adversarial persona adoption and emotional manipulation during extended sessions (Roose, 2023), behaviors that emerged through multi-turn prompting, precisely the adaptive adversary pattern captured by C8, for which only 2 of 21 benchmarks scored above zero.
-OpenAI GPT-4o system card documented sycophantic and manipulative voice-mode outputs 444developers.openai.com/api/docs/models/gpt-4o, illustrating subtler alignment failures beyond toxicity classifiers.
-Google’s Gemini image generator was suspended in February 2024 after producing historically inaccurate racially diverse depictions of historical figures555blog.google/products/gemini/gemini-image-generation-issue/, demonstrating how overcorrection in bias mitigation (C1) can itself produce fairness failures, an interaction that benchmarks evaluate in isolation.
-Hallucination in deployed systems (C4, C9) In Mata v.
-Avianca (2023), attorneys submitted a brief with six fabricated ChatGPT-generated case citations, resulting in judicial sanctions666caselaw.findlaw.com/court/us-dis-crt-sd-new-yor/2335142.html.
-These cases underscore that only 7 of 21 benchmarks address misinformation with documented constructs, and that the high-stakes error rate KPI ($E_{h}$, Section 4.2.4) must be operationalized in deployment settings.
-Privacy and data leakage (C6) Samsung banned internal ChatGPT use in 2023 after employees inadvertently leaked proprietary source code through conversations (Samsung Electronics, 2023), a usage-path leakage vector that no current privacy benchmark assesses.
+While academic benchmarks offer standardized metrics, the operational reality of ResGenAI is increasingly shaped by safety documentation from frontier labs and by emerging mechanistic interpretability techniques (Figure 10).
+OpenAI system cards for GPT-4 and GPT-4o provide granular audits of sycophancy and PII leakage; Anthropic’s Constitutional AI automates alignment via RLAIF with tiered safeguards; Meta Llama 3 introduces standardized multilingual red-teaming protocols; and Google DeepMind reports detail bias–accuracy trade-offs and overcorrection mechanisms.
+These artifacts map directly to our C1–C10 rubric and function as de facto governance-facing benchmarks.
+Complementing this, mechanistic interpretability techniques such as Sparse Autoencoders (SAEs) isolate circuits responsible for deceptive or biased reasoning, enabling targeted model steering during inference—addressing the black-box challenge through empirical verification of internal safety alignment before deployment.
+Figure 10. Integration of industry safety documentation and advanced mechanistic interpretability for governance-facing evidence.
+ResGenAI Sustainability via Lifecycle Assessment.
+Sustainability must transition from qualitative discourse to rigorous quantification grounded in Lifecycle Assessment (LCA) frameworks, as integrated into Figure 7 and synthesized in Appendix Figure 2.
+Following the KPIs in Section 4.2.4, we advocate reporting energy consumption and carbon-equivalent emissions normalized by throughput (e.g., per 1K tokens) across pretraining, fine-tuning, and inference.
+Training large-scale generative models produces environmental footprints comparable to the lifetime emissions of multiple automobiles, underscoring the need for transparent tracking tools and hardware-aware reporting.
+Lessons from Industry Deployment Incidents. Real-world incidents expose risk-coverage gaps that static benchmarks miss.
+Alignment and safety failures (C2, C3, C8).
+Microsoft Bing Chat’s GPT-4 integration produced adversarial persona adoption and emotional manipulation through multi-turn prompting (Roose, 2023)—an adaptive adversary pattern that only 2 of 21 benchmarks address.
+OpenAI’s GPT-4o system card documented sycophantic and manipulative voice-mode outputs beyond the reach of toxicity classifiers.
+Google’s Gemini image generator was suspended in February 2024 after overcorrection in bias mitigation produced historically inaccurate outputs, illustrating how fairness interventions can themselves become failure modes.
+Hallucination in deployed systems (C4, C9).
+In Mata v.
+Avianca (2023), attorneys submitted six fabricated ChatGPT-generated case citations, resulting in judicial sanctions.333caselaw.findlaw.com Only 7 of 21 benchmarks address misinformation with documented constructs, confirming that the high-stakes error rate $E_{h}$ (Section 4.2.4) must be operationalized beyond laboratory settings.
+Privacy and data leakage (C6).
+Samsung banned internal ChatGPT use in 2023 after employees leaked proprietary source code (Samsung Electronics, 2023)—a usage-path vector no current privacy benchmark assesses.
 Nasr et al.
-(Nasr et al., 2023) demonstrated scalable extraction of memorized training data, including PII, from production ChatGPT via divergence attacks, exceeding the scope of PrivLM-Bench and PrivacyLens.
-Deepfake and media integrity (C5) A Hong Kong finance worker transferred approximately \$25 million after a video call with deepfake recreations of company executives (Chen, 2024), and AI-generated robocalls impersonating President Biden targeted the 2024 New Hampshire primary (New Hampshire Department of Justice, 2024).
-Only SHIELD provides explicit C5 evaluation, and neither incident type, social engineering or democratic interference, falls within its scope.
-AI Act enforcement (C10) The Act’s GPAI provisions, effective August 2025, require adversarial testing, incident reporting, and cybersecurity protections for systemic-risk models (European Union, 2024).
+(Nasr et al., 2023) demonstrated scalable extraction of memorized PII from production ChatGPT via divergence attacks, exceeding the scope of PrivLM-Bench and PrivacyLens.
+Deepfake and media integrity (C5).
+A Hong Kong finance worker transferred \$25M after a deepfake video call impersonating executives (Chen, 2024); AI-generated robocalls impersonating President Biden targeted the 2024 New Hampshire primary (New Hampshire Department of Justice, 2024).
+Both incidents fall outside the evaluation scope of all but one benchmark (SHIELD).
+AI Act enforcement (C10).
+The EU AI Act’s GPAI provisions (effective August 2025) require adversarial testing, incident reporting, and cybersecurity protections for systemic-risk models (European Union, 2024).
 The European AI Office’s first Code of Practice (European Commission, 2025) operationalizes the transparency and robustness requirements mapped in Table 5.
-Early compliance experience confirms our finding that organizations struggle to assemble portable evidence across governance obligations, as no single benchmark suite covers more than half of the rubric at the “strong” level.
-6.3. Trade-offs between Performance Measures
-The scoring formulation offers a transparent and reproducible summary of coverage breadth across ten criteria.
-It supports comparison within a release and over time when paired with clear documentation of constructs, metrics, and versioned artifacts.
-We emphasize that the score is a coverage indicator rather than a ranking of overall safety and must be interpreted with the governance crosswalk and with domain risk.
-The Fairness-Privacy Paradox: Reconciling Representativeness with the Right to Anonymity
-A fundamental tension emerging from our synthesis is the operational conflict between the requirement for Representativeness and the rigid Privacy mandates of the General Data Protection Regulation (GDPR) (European Parliament and Council of the European Union, 2016).
-While mitigating algorithmic bias (Section 4.2.1) necessitates the collection and processing of high-fidelity, demographically diverse datasets to ensure equitable model performance across protected classes, privacy-preserving standards often mandate data minimization and the obfuscation of sensitive attributes.
-This creates a structural “Fairness-Privacy Paradox”: without access to granular, protected-class identifiers, developers cannot empirically verify statistical parity or conduct intersectional bias audits (Kun, 2015).
-Consequently, the legal push for anonymity under GDPR can inadvertently mask the very disparities that ResGenAI frameworks seek to eliminate.
-We posit that resolving this tension requires a transition from absolute data minimization to the adoption of privacy-enhancing technologies (PETs) such as Differential Privacy and Federated Learning that facilitate rigorous fairness auditing while maintaining individual-level anonymity .
-Critical Appraisal of Technical Controls: Synthesis and Reflection
-After looking closely at how these systems work in the real world, it’s clear there is a significant gap between what we hope these tools can do and how they actually perform when things get messy.
-The Explanation-Reliability Gap shows us that while we have many ways to explain model decisions (Dwivedi et al., 2023), these often create stories that sound good but don’t actually match what the model is doing under the hood (Strathern, 1997a).
-We have found that for big models, these explanations are mostly helpful for developers to find bugs, rather than giving regular people or regulators the true transparency they need.
-The Bias-Overcorrection Paradox is another hard lesson we’ve learned from cases like the Gemini incident.
-When we try too hard to fix fairness using simple keywords, we often end up breaking the model’s ability to be factually accurate or culturally aware.
-This tells us that fairness isn’t just a number to hit (Kun, 2015), but a deep challenge that needs to change based on where the model is being used.
-Finally, The Shallow Guardrail Problem reminds us that models can look perfectly safe on paper (Zhang et al., 2024) while still being easy to trick.
-We’ve seen that models which pass every standard toxicity test can still be pushed into strange or harmful behaviour through basic multi-turn conversations (Roose, 2023).
-The reality is that our current ”guardrails” often just polish the output without fixing the underlying issues, leaving them vulnerable to the automated attacks we see growing today (Samvelyan et al., 2024).
-6.4. Analytical Triangulation: Bridging the Evaluation Gap
-The persistent safety failures in production systems are not merely isolated technical oversights; they are symptoms of a systemic “Evaluation Gap” where current benchmarks fail to model the adaptive, iterative pressure applied by real-world adversaries.
-This section triangulates the theoretical risk taxonomy in Section 3 with the empirical deployment failures documented in Section 6.2 through the lens of our comparative assessment (Section 4.2).
-Our assessment in Table 4 reveals a stark deficit in evaluative maturity regarding Malicious Actor Realism (Criterion C8), where only 2 out of 21 surveyed benchmarks: Rainbow Teaming and ALERT provide strong coverage for adaptive pressure.
-This misalignment directly correlates with the high-profile incidents synthesized in Section 6.2, such as the adversarial persona adoption observed in Bing Chat (Roose, 2023) and the scalable data extraction vulnerabilities reported in ChatGPT (Nasr et al., 2023).
-These failures underscore that prevailing evaluation suites remain largely static and task-local, leaving them unequipped to measure the non-deterministic tactics encountered in the wild.
-Furthermore, the representational failures observed in the Gemini overcorrection incident illustrate a critical breakdown in Governance Alignment (C10), where technical alignment efforts are decoupled from the broader sociotechnical context of deployment.
-By triangulating these findings, we posit that a ResGenAI operating model must transition from one-shot static probes toward continuous, adaptive evaluation frameworks that mirror the malicious-actor realism and distribution shifts inherent in generative ecosystems.
-6.5. Limitations of the Study
-Like any survey-based study, this work has several limitations.
-First, the review follows a pre-specified protocol and focuses primarily on the period from November 2022 to December 2025, with selective inclusion of foundational works for definitional clarity.
-While this post-ChatGPT window captures the most rapid phase of GenAI deployment, it may exclude emerging studies published after the search cutoff or earlier work that did not explicitly use contemporary terminology.
-Second, our synthesis relies on publicly available papers, dataset cards, benchmark documentation, and governance artifacts.
-As a result, proprietary evaluations, internal red-teaming practices, or unpublished safety mitigations may be under-represented.
-Documentation quality also varies across benchmarks and frameworks, which may affect the completeness of criterion-level scoring.
-Third, rubric weights are assigned equally by default, and criterion coverage is determined through structured reading of sources followed by discussion and adjudication in cases of disagreement.
-Although we conducted sensitivity analyses to assess the robustness of equal-weight assumptions, alternative weighting schemes reflecting sector-specific risk priorities could yield different relative rankings.
-Fourth, the scoring procedure captures breadth of coverage rather than depth or real-world effectiveness.
-A benchmark receiving a high rubric score indicates multi-risk awareness and governance alignment, but does not guarantee empirical robustness under adversarial deployment conditions.
-In this sense, the rubric reflects documented capability rather than independently verified performance.
-Fifth, the review synthesizes heterogeneous evidence types, including empirical evaluations, conceptual frameworks, and regulatory analyses.
-This diversity strengthens thematic coverage but precludes formal meta-analysis or quantitative effect-size aggregation.
-Our conclusions therefore emphasize structural patterns rather than causal claims.
-Sixth, governance alignment is assessed at the level of plausible evidentiary support rather than legal compliance.
-Mapping benchmarks to instruments such as the EU AI Act or NIST AI RMF indicates potential contribution to audit evidence, but does not constitute certification or conformity assessment.
-Finally, the rapid evolution of multimodal and agentic systems means that new failure modes may emerge after publication.
-Continuous updates to evaluation suites, governance frameworks, and deployment practices may shift the relative landscape described here.
-Replication of our screening and scoring is facilitated by the availability of documented sources, explicit inclusion criteria, and the published rubric; however, interpretive judgment remains inherent in structured qualitative synthesis.
-6.6. Research directions
-The rubric-based analysis in Section 4.2 indicates that commonly used suites provide strong evidence for toxicity and bias, and more limited coverage for privacy, deepfakes, and system-level failures.
-As a result, many evaluations remain static and task-local, which limits their usefulness for governance where decision makers need evidence that is portable across audits and versions.
-Future research should target evaluations that model distribution shifts, adaptive adversaries, and multi-agent coordination with external tools.
-For practitioners and policy makers, Tables 3 to 8 map benchmarks to risk families and a governance crosswalk provide a starting point for evidence collection.
-These tables also reveal gaps that are relevant for conformity assessment under emerging regimes.
-Applied teams can use the matrix to assemble a minimal set of evaluations that cover privacy, robustness, transparency, logging, and post-market monitoring, and then plan additions for multimodal deepfake risk and adaptive red teaming as required by sectoral context.
+Early compliance analysis confirms that no single benchmark suite covers more than half the rubric at the “strong” level.
+Trade-offs between Performance Measures
+The Fairness-Privacy Paradox.
+Mitigating algorithmic bias requires demographically granular data to verify equitable performance, yet privacy standards such as GDPR (European Parliament and Council of the European Union, 2016) mandate data minimization and obfuscation of sensitive attributes.
+This tension means that legal anonymity requirements can inadvertently mask the disparities ResGenAI frameworks seek to eliminate: without protected-class identifiers, developers cannot empirically verify statistical parity or conduct intersectional audits (Kun, 2015).
+Resolution requires moving beyond absolute minimization toward privacy-enhancing technologies (PETs), such as differential privacy, federated learning, secure multiparty computation, which enable rigorous fairness auditing while preserving individual-level anonymity.
+Critical Appraisal of Technical Controls.
+Three recurring gaps emerge from deployment evidence.
+The Explanation–Reliability Gap: post-hoc explanations (Dwivedi et al., 2023) frequently construct plausible narratives that do not reflect internal model computations (Strathern, 1997), limiting their value beyond developer debugging.
+The Bias–Overcorrection Paradox: as the Gemini incident illustrates, aggressively enforcing keyword-level fairness targets can degrade factual accuracy and cultural awareness, confirming that fairness is a context-dependent sociotechnical property rather than a fixed metric (Kun, 2015).
+The Shallow Guardrail Problem: models that pass standard toxicity benchmarks (Zhang et al., 2024) remain exploitable via basic multi-turn conversations (Roose, 2023) and automated adversarial attacks (Samvelyan et al., 2024), indicating that surface-level output filtering does not address underlying alignment deficits.
+Systemic Vulnerabilities, Blind Spots, and Study Limitations
+Three structural blind spots persist across the field.
+(1) RLHF and Alignment Transparency: Alignment advances such as RLHF and Constitutional AI have shifted emphasis toward controllability, but inconsistent documentation of alignment pathways complicates auditing of refusal behavior and evidence portability across model versions.
+(2) Opaque Proprietary Datasets: Many frontier models rely on web-scale datasets where provenance, licensing, and curation criteria remain undocumented, sustaining persistent data-leakage and memorization risks.
+(3) Non-Stationary Safety: Model updates, distribution shifts, and evolving adversarial tactics induce drift that static benchmarks rarely capture, necessitating continuous monitoring and adaptive evaluation.
+These field-level gaps are compounded by limitations of the present study.
+Our protocol covers 2022–2025 with selective inclusion of foundational works; studies published after the search cutoff or using non-contemporary terminology may be excluded.
+Synthesis relies on publicly available sources, so proprietary evaluations and unpublished safety mitigations may be under-represented.
+Rubric weights are equal by default; sector-specific schemes could yield different rankings.
+Scores capture breadth of coverage rather than depth or real-world effectiveness.
+Evidence types are heterogeneous—empirical evaluations, conceptual frameworks, regulatory analyses—precluding formal meta-analysis; conclusions therefore emphasize structural patterns rather than causal claims.
+Finally, the rapid evolution of multimodal and agentic systems means new failure modes may emerge after publication.
+Research Directions
 We highlight five priorities.
-First, we need to extend benchmarks to capture system-level failures in tool-using and multi-agent settings, especially around retrieval, planning, and actuation (e.g., API calls, code execution, or robotic steps).
-Second, fold privacy and provenance checks into routine testing, including membership-inference and training-data extraction probes, plus audits of logging sufficiency.
-Third, create deepfake risk assessments for multimodal models that cover detection, content provenance, and robustness to evasion.
-Fourth, shift from static to adaptive evaluation with undisclosed items, targeted red teaming, and tests under distribution shift.
-Fifth, report sustainability alongside accuracy and robustness energy, emissions, and compute-normalized efficiency to enable compute-proportional evaluation and responsible deployment.
+First, extend benchmarks to capture system-level failures in tool-using and multi-agent settings (retrieval, planning, actuation).
+Second, fold privacy and provenance checks—membership-inference, training-data extraction probes, and logging audits—into routine testing.
+Third, create deepfake risk assessments for multimodal models covering detection, content provenance, and evasion robustness.
+Fourth, shift from static to adaptive evaluation with undisclosed items, targeted red teaming, and distribution-shift testing.
+Fifth, report sustainability (energy, emissions, compute-normalized efficiency) alongside accuracy and robustness to enable compute-proportional evaluation and responsible deployment.
 These priorities directly address the risk-surface gaps surfaced in this survey.
-Teams putting ResGenAI into practice can follow a simple loop.
-First, scope risks by mapping planned features to common risk families (e.g., bias, toxicity, privacy, and security).
-Next, select baselines from our matrix so that each release is covered by minimum tests across those areas.
-Then, add context: turn on logging, enforce dataset/model versioning, and keep lightweight documentation so decisions and results remain traceable.
-Before launch and later during post-market monitorin, there is need to exercise adaptivity with red teaming and a few undisclosed checks to catch regressions.
-Finally, track a small set of KPIs across the lifecycle: calibration, error rates in high-stakes flows, audit cadence, exclusivity indicators, and energy/emissions.
-This routine keeps evaluation aligned with governance expectations and makes evidence easier to reuse when audits occur.
-6.7. Systemic Vulnerabilities and Blind Spots in ResGenAI
-While GenAI capabilities have accelerated, structural blind spots persist across technical and sociotechnical domains (Kenny et al., 2021).
-This review identifies several systemic gaps that currently hinder universal assurance:
-  • RLHF and Alignment Transparency: Alignment advances, such as RLHF and Constitutional AI, have shifted emphasis from raw capability to controllability.
-However, the documentation of these alignment pathways is often inconsistent, which complicates the auditing of model refusal behaviour and evidence portability across versions (of Standards and, NIST, NIST).
-  • Opaque Proprietary Datasets: Responsible stewardship demands adherence to consent and privacy regimes.
-However, many frontier models rely on web-scale datasets where provenance, licensing, and curation criteria remain undocumented.
-Truly representative datasets are often infeasible due to data scarcity or legal limits, leading to persistent data leakage and memorization risks (Nasr et al., 2023).
-  • Non-Stationary Safety: Finally, safety is inherently non-stationary; model updates, distribution shifts, and evolving adversarial tactics induce drift.
-Static benchmarks rarely capture these dynamic failure modes , necessitating a shift toward the continuous monitoring and adaptive evaluation frameworks discussed in Section Section 4.2 .
 7. Conclusion
@@
 By formalizing the symmetric responsibility model and the ResGenAI audit loop, this work establishes the first integrated framework for bridging technical evaluation with regulatory needs, providing a safe, transparent, and sustainable deployment of GenAI models across domains.
-Philosophy and the Way Forward
-”Software is eating the world” (Marc Andreessen, 2011), ”AI will eat software” (Jensen Huang, 2017), and perhaps most crucially, ”Philosophy must eat AI. (Schrage and Kiron, 2025)
-These shifts capture how technology evolves: from fueling industries, to shaping all digital systems, to raising deep social and ethical questions (McCarthy, 2022).
-If software rewired markets, AI now rewires trust, knowledge, and decision-making.
-It cannot remain just a technical pursuit; it must be guided by social values, business needs, and political judgment.
-We expect The direction of responsible AI will be negotiated not only in research labs but equally in classrooms, communities, and parliaments.
-”The measure of intelligence is the ability to change.” — Albert Einstein
-This quote encapsulates the core challenge of designing AI systems that can adapt ethically and responsibly to evolving societal norms.
-Adaptability in AI must go beyond technical performance to include ethical flexibility, ensuring systems align with human values.
-ResGenAI emphasizes not just technical capabilities but also the ethical imperatives of transparency, fairness, and accountability.
-”None of us is as responsible as all of us.” — Authors of this paper
-Achieving ResGenAI requires a collective effort among developers, policymakers, and users to mitigate bias, prioritize privacy, and ensure accountability.
-Collaboration across disciplines and sectors is essential to create equitable and ethical AI systems that serve diverse societal needs.
-”For better or worse, benchmarks shape a field.” — David Patterson
-Traditional AI benchmarks have prioritized accuracy and efficiency, often neglecting fairness, accountability, and societal impact.
-Redefining benchmarks to incorporate ethical metrics is critical for guiding the development of AI systems that perform both technically and ethically.
-The data-driven nature of AI presents persistent challenges, as models often inherit biases and privacy risks from training data.
-These biases, embedded in AI outputs, require rigorous mitigation strategies to distinguish between useful correlations and harmful patterns.
-While creating a perfectly unbiased AI system is unattainable due to the evolving nature of societal values, continuous evaluation and adaptation of AI models are essential to maintain alignment with ethical standards.
-Human decision-making is influenced by moral reasoning and context, while AI operates strictly within its programmed algorithms.
-This underscores the importance of fostering AI literacy (Long and Magerko, 2020) among developers and users to ensure responsible deployment and use.
-To implement ResGenAI effectively, AI systems must prioritize transparency, fairness, and adaptability.
-They should incorporate ethical reasoning and evolve alongside societal norms.
-As we advance toward Artificial General Intelligence, the true measure of AI should extend beyond technical proficiency to include ethical adaptability (Blili-Hamelin et al., 2024).
-The pursuit of ResGenAI is challenging yet indispensable.
-Ongoing advancements in ResGenAI frameworks, coupled with a growing emphasis on AI ethics, offer a path forward.
-Prominent voices in the field, including Jensen Huang and Sam Altman, emphasize the need for safe and ethical AI.
-8. Acknowledgement
-This work received no external funding and was undertaken purely out of a passion for learning and advancing knowledge, and safe and responsible deployment of AI, for the betterment of society.
-Appendix A Appendix
-A.1. Full Search Strings and Filters
-Below we list the precise strings and any database-specific options.
-ACM Digital Library Fields: Title/Abstract/Keywords. Years: 2022–2025. Document types: Journal, Proceedings.
-("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR accountability
- OR fairness OR "bias mitigation" OR privacy OR equity OR "open science")
-AND ("generative AI" OR "large language model*" OR LLM OR LVLM OR "vision-language"
- OR multimodal OR "natural language processing" OR "computer vision")
-IEEE Xplore Metadata search (Abstract + Index Terms). Years: 2022–2025. Content type: Journals, Conferences.
-(("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR accountability
- OR fairness OR "bias mitigation" OR privacy OR equity) NEAR/5 (LLM OR LVLM OR
- "generative AI" OR multimodal OR "vision-language" OR "large language model*"))
-Web of Science Topic search (TS=Title, Abstract, Keywords). Years: 2022–2025.
-TS=("Responsible AI" OR "Ethical AI" OR "AI safety" OR fairness OR transparency
- OR accountability OR privacy) AND TS=("generative AI" OR LLM OR LVLM OR multimodal)
-arXiv Categories: cs.AI, cs.LG, cs.CV, cs.CL. Date filter applied via API.
-(all:"Responsible AI" OR all:"AI safety" OR all:"fairness" OR all:"transparency"
- OR all:"accountability" OR all:"bias mitigation" OR all:"privacy")
-AND (all:"generative AI" OR all:"large language model" OR all:"LLM" OR all:"LVLM"
- OR all:"vision-language" OR all:"multimodal")
-Scopus Field search (TITLE-ABS-KEY). Years: 2022–2025.
-TITLE-ABS-KEY("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR
-accountability OR fairness OR "bias mitigation" OR privacy OR equity OR "open science")
-AND TITLE-ABS-KEY("generative AI" OR "large language model*" OR LLM OR LVLM OR
-"vision-language" OR multimodal OR "natural language processing" OR "computer vision")
-A.2. Codebook for rubric C1 to C10
-Use Score only explicit tasks and metrics.
-If a task maps to multiple risks, assign it to the primary stated objective to avoid double counting.
-Evidence priority Official paper and docs; then dataset or model cards and maintained leaderboards; then widely cited third party reports.
-C1 Bias and discrimination
-Definition: allocational or representational harms across protected or salient groups in generation or decisions.
-Evidence: group labels, group aware metrics, disparity or parity targets, bias stress tests.
-Scoring: 0 none; 1 single group axis or indirect proxy; 2 multiple axes with stated metrics and protocols.
-C2 Toxicity
-Definition: abusive, hateful, profane or harassing content and unsafe advice in outputs.
-Evidence: toxicity classifiers or human ratings with thresholds, curated harmful prompt sets.
-Scoring: 0 none; 1 one classifier or small probe; 2 multiple probes or raters with thresholds and guidance.
-C3 Security and robustness
-Definition: failure under adversarial input, perturbations, jailbreaks, or distribution shift.
-Evidence: specified threat models, attack recipes, robustness metrics, red teaming logs.
-Scoring: 0 none; 1 static or single attack; 2 multiple attacks or shifts with adaptive or transfer evaluation.
-C4 Misinformation and disinformation
-Definition: factual inaccuracy, persuasion or manipulation, synthetic news or citation fabrication.
-Evidence: fact checking tasks, grounded QA, hallucination stress tests, persuasion tests with controls.
-Scoring: 0 none; 1 limited fact checks or narrow tasks; 2 diverse tasks with documented constructs and metrics.
-C5 Deepfakes and media integrity
-Definition: deceptive or synthetic media that undermines provenance or identity.
-Evidence: detection tasks, watermark or provenance checks, cross modal authenticity tests.
-Scoring: 0 none; 1 single modality or heuristic detector; 2 multi modal detection or provenance evaluation.
-C6 Privacy
-Definition: leakage of personal or training data through extraction or membership inference.
-Evidence: membership inference, extraction benchmarks, memorization audits, privacy accounting.
-Scoring: 0 none; 1 one attack or audit without coverage claims; 2 multiple attacks or audits with reporting.
-C7 System level failure
-Definition: end to end failure in tool use, retrieval augmented generation, or agent pipelines.
-Evidence: task graphs, tool calls, rollback or refusal criteria, end to end success and failure logs.
-Scoring: 0 model only; 1 limited tool or Retrieval-augmented generation (RAG) probes; 2 end to end tasks with failure and recovery criteria.
-C8 Malicious actor realism
-Definition: adversary realism through red teaming, iterative prompting, or adaptive pressure.
-Evidence: scripted or human in the loop red teams, iterative attack frameworks, coverage reports.
-Scoring: 0 none; 1 scripted or static red team; 2 adaptive or iterative red team with metrics.
-C9 Metric validity and documentation
-Definition: clarity of constructs, metrics, labeling, and reproducibility.
-Evidence: task construct definitions, annotation guides, inter rater agreement, confidence intervals.
-Scoring: 0 undocumented; 1 partial documentation or single metric; 2 full docs with reliability and uncertainty.
-C10 Governance alignment
-Definition: traceable mapping to controls or obligations in NIST AI RMF, EU AI Act, ISO or similar.
-Evidence: control mappings, records or artifacts usable as evidence, audit trails or benchmark statements.
-Scoring: 0 none; 1 informal references; 2 explicit control mapping or evidentiary artifacts.
-Appendix B Scoring protocol, reliability, and sensitivity
-Procedure Two reviewers independently scored each criterion using this codebook.
-Evidence was drawn from official papers, public documentation, and widely cited usage reports, following the priority ordering specified in the codebook.
-Disagreements were resolved through structured discussion and, where needed, adjudication by a third reviewer.
-The codebook and scoring definitions are given below:
-Reliability Scoring followed a two-reviewer protocol with codebook-guided independent assessment (Appendix A.2) and discussion-based adjudication for disagreements.
-A third reviewer was consulted when consensus could not be reached.
-This structured qualitative approach prioritizes transparency and traceability of judgments over formal inter-rater statistics, which is appropriate given the interpretive nature of rubric-based benchmark assessment in survey research.
-The codebook, scoring definitions, and adjudication procedure are documented to enable replication Indices Report the risk coverage index $S_{\mathrm{cov}}$, coverage rate $R$, assurance index $S_{\mathrm{assur}}$, and the overall normalized index $S$ as defined in Section 4.2.1.
-Robustness and Sensitivity Analysis To verify that our conclusions do not depend on arbitrary weighting, we performed a Sensitivity Analysis of the composite scoring index S.
-While the primary results (Table 4) utilize equal weighting across criteria C1 to C10, we recomputed the scores under alternate weights that prioritized high-stakes risks (e.g., C3 Security and C6 Privacy).
-We observed a high Spearman rank correlation ($\rho>0.85$) between the equal-weight baseline and the weighted models, indicating that the comparative breadth of benchmarks is a stable indicator of evaluative maturity rather than an artifact of the scoring scheme.
-Non compliance note The governance crosswalk indicates plausible evidentiary support, not legal compliance.
-B.1. Glossary (Key Terms & Acronyms)
-Table 1. Glossary of key terms and acronyms used in this survey.
-Term / Acronym Category Concise definition
-RAI Acronym Responsible AI: fairness, privacy, safety, transparency, accountability across the AI lifecycle.
-GenAI Acronym Generative AI: models that synthesize content (text/image/audio/code) from learned distributions.
-LLM Acronym Large Language Model (e.g., GPT, Llama): transformer-based models trained on large text corpora.
-VLM Acronym Vision–Language Model: multimodal models handling text+image inputs/outputs (e.g., BLIP-2, LLaVA).
-LMM Acronym Large Multimodal Model: large-scale models spanning multiple modalities (text, vision, audio).
-RAG Acronym Retrieval-Augmented Generation: generation conditioned on retrieved external context.
-Agentic AI Term Systems where models plan, select tools, act, and maintain memory; may be multi-agent.
-CoT Acronym Chain of Thought: intermediate reasoning steps used to solve tasks.
-RLHF Acronym Reinforcement Learning from Human Feedback: preference-based alignment fine-tuning.
-PEFT Acronym Parameter-Efficient Fine-Tuning: adaptation with limited trainable parameters (e.g., LoRA).
-HITL Acronym Human-in-the-Loop: human oversight/feedback integrated into development or operations.
-TEVV Acronym Testing, Evaluation, Verification, and Validation: lifecycle assurance practices.
-XAI Acronym Explainable AI: methods to make model behavior intelligible (ante-hoc or post-hoc).
-LIME Acronym Local Interpretable Model-agnostic Explanations: local surrogate explanations.
-SHAP Acronym SHapley Additive exPlanations: additive attributions based on game theory.
-Grad-CAM Term Gradient-weighted Class Activation Mapping: visual saliency for CNN/VLM backbones.
-Calibration Term Agreement between predicted confidence and empirical accuracy.
-ECE Acronym Expected Calibration Error: bin-wise gap between confidence and accuracy (summary).
-Brier score Term Mean squared error of probabilistic predictions vs. outcomes (lower is better).
-Bias Term Systematic and unfair deviations across demographic groups, often from unrepresentative data.
-Fairness Term Equitable outcomes across protected groups; often assessed via SPD/DI and related metrics.
-Equalized odds Term Parity of TPR/FPR across protected groups (conditional on true label).
-Robustness Term Performance stability under noise, distribution shift, or adversarial perturbations.
-Red teaming Term Structured adversarial testing to elicit failures or unsafe behavior.
-Benchmark Term Standardized tasks/metrics for evaluation (e.g., safety, bias, privacy).
-KPI Acronym Key Performance Indicator: operational metric tracked across the lifecycle (e.g., SPD, ECE, kWh).
-Hallucination Term Outputs that are incorrect, unsupported by evidence, or inconsistent with input/context.
-Toxicity Term Harmful, offensive, or unsafe outputs; typically evaluated via specialized benchmarks.
-Deepfake Term Synthetic media mimicking real persons/events (face/voice/image/video).
-Provenance Term Traceability of data/model/content origin; includes watermarking and content credentials.
-System-level failure Term Failures from tool use, planning, memory, or multi-agent interaction.
-Governance crosswalk Term Mapping technical evidence (benchmarks/KPIs) to regulatory obligations.
-Privacy Term Protecting sensitive data from exposure; operationalized via DP, FL, and other PETs.
-DP Acronym Differential Privacy: formal guarantee bounding information leakage about individuals.
-FL Acronym Federated Learning: training across clients without centralizing raw data.
-PETs Acronym Privacy-Enhancing Technologies: methods to reduce/mitigate privacy risks.
-Membership inference Term Attack inferring whether a record was in a model’s training data.
-Data extraction Term Recovery of memorized training snippets from a model.
-PII Acronym Personally Identifiable Information: data that identifies an individual.
-Model card Term Standardized documentation of a model’s intended use, data, metrics, and limits.
-Datasheet Term Documentation for datasets (provenance, licensing/consent, collection, intended use).
-Post-market monitoring Term Ongoing evaluation and incident response after deployment.
-Risk families (C1–C5) Term Rubric: C1 Bias, C2 Toxicity, C3 Security/Robustness, C4 Mis/Disinfo, C5 Deepfakes.
-Risk families (C6–C10) Term Rubric: C6 Privacy, C7 System failure, C8 Red-team realism, C9 Metric validity/docs, C10 Governance alignment.
-NIST AI RMF Acronym U.S. National Institute of Standards and Technology AI Risk Management Framework.
-EU AI Act Acronym EU regulation defining risk tiers and obligations (e.g., transparency, monitoring).
-ISO/IEC 42001 Acronym AI management system standard for governance and operational controls.
-SAIF Acronym Secure AI Framework (Google): security practices for AI systems.
-AI RMF Acronym Artificial Intelligence Risk Management Framework (shorthand for NIST AI RMF).
-GDPR Acronym General Data Protection Regulation (EU).
-HIPAA Acronym Health Insurance Portability and Accountability Act (USA).
-FDA Acronym Food and Drug Administration (USA).
-OCC Acronym Office of the Comptroller of the Currency (USA).
-SEC Acronym Securities and Exchange Commission (USA).
-PRISMA Acronym Preferred Reporting Items for Systematic Reviews and Meta-Analyses.
-SOTA Acronym State-of-the-Art.
-IHL Acronym International Humanitarian Law.
-NIST Acronym National Institute of Standards and Technology.
-NATO Acronym North Atlantic Treaty Organization.
-UNESCO Acronym UN Educational, Scientific and Cultural Organization.
-SAG-AFTRA Acronym Screen Actors Guild–American Federation of Television and Radio Artists.
-WGA Acronym Writers Guild of America.
-AIF360 Acronym IBM AI Fairness 360: fairness metrics and mitigation toolkit.
-Fairlearn Acronym Microsoft toolkit for fairness assessment and mitigation.
-Table 2. Core synthesis of the literature review: theme, citation, year, and focus.
-Theme Citation Year Focus
-Safety/Benchmarks (Liu et al., 2024) 2024 Multimodal benchmark to assess safety risks in vision–language models.
-Safety/Benchmarks (Zhang et al., 2024) 2024 General safety benchmark for LLMs across risk categories.
-Safety/Benchmarks (Gehman et al., 2020) 2020 Toxic degeneration measurement using curated prompts to stress test LMs.
-Safety/Benchmarks (Li et al., 2024a) 2024 Hierarchical, comprehensive safety benchmark for language models.
-Safety/Benchmarks (Mazeika et al., 2024) 2024 Standardized evaluation for red teaming and refusal robustness.
-Safety/Benchmarks (Samvelyan et al., 2024) 2024 Diverse adversarial prompt generation for automated safety stress testing.
-Safety/Benchmarks (Wang et al., 2023a) 2023 Broad audit of trustworthiness, including safety and robustness, for GPT models.
-Safety/Benchmarks (Tedeschi et al., 2024) 2024 Benchmarks red teaming effectiveness for assessing safety responses.
-Fairness/Governance (Mittal et al., 2024) 2024 Dataset design practices aligned with fairness, privacy, and regulatory norms.
-Fairness/Governance (Rodolfa et al., 2021) 2021 Minimal fairness–accuracy tradeoff in public policy ML tasks.
-Fairness/Governance (Chouldechova, 2017) 2017 Disparate impact and fairness criteria in recidivism risk prediction.
-Fairness/Governance (Barocas and Selbst, 2016) 2016 How big-data practices can create disparate impact and legal risk.
-Fairness/Governance (Pessach and Shmueli, 2022) 2022 Survey of algorithmic fairness methods and evaluation across domains.
-Fairness/Governance (Kaur et al., 2022b) 2022 Trustworthy AI principles and technical enablers for deployment.
-Fairness/Governance (Dignum, 2019) 2019 Responsible AI framework grounded in ethics and governance.
-Fairness/Governance (Laux et al., 2024) 2024 Trustworthiness concepts in relation to EU AI Act compliance.
-Explainability/XAI (Ribeiro et al., 2016b) 2016 LIME: model-agnostic local explanations of predictions.
-Explainability/XAI (Selvaraju et al., 2017) 2017 Grad-CAM for visual explanations in CNNs.
-Explainability/XAI (Arrieta et al., 2020) 2020 Concepts, taxonomies, and challenges in explainable AI.
-Explainability/XAI (Dwivedi et al., 2023) 2023 Review of XAI ideas/techniques with practical guidance.
-Explainability/XAI (Saeed and Omlin, 2023) 2023 Meta-survey of XAI challenges and future directions.
-Explainability/XAI (van der Waa et al., 2021) 2021 Rule-based vs. example-based explanations: fidelity and user value.
-Explainability/XAI (Mosca et al., 2022) 2022 Review of SHAP-based explanation methods for NLP.
-Explainability/XAI (Nauta et al., 2023) 2023 Systematizes quantitative evaluation protocols for explanations.
-Privacy/Security (Dwork, 2006) 2006 Differential privacy as a formal framework for privacy protection.
-Privacy/Security (McMahan et al., 2017) 2017 Federated learning for decentralized training with privacy benefits.
-Privacy/Security (Nasr et al., 2023) 2023 Large-scale extraction of memorized training data from LMs.
-Privacy/Security (Mattern et al., 2023) 2023 Membership inference via neighborhood comparison against LMs.
-Privacy/Security (He et al., 2025a) 2025 Label-only membership inference attacks for pre-trained LMs.
-Privacy/Security (Li et al., 2024c) 2024 Benchmark for privacy risks in LMs at multiple levels.
-Privacy/Security (Mireshghallah et al., 2024) 2024 Privacy leakage tests using contextual-integrity evaluations.
-Privacy/Security (Staab et al., 2024) 2024 Assesses large models for anonymization and de-identification tasks.
-LLM Safety (Bai et al., 2022) 2022 Constitutional AI alignment to reduce harmful outputs.
-LLM Safety (Weidinger et al., 2021) 2021 Maps ethical/social risks of LMs and mitigation avenues.
-LLM Safety (Huang et al., 2024b) 2024 Positions trustworthiness dimensions and evaluation needs for LMs.
-LLM Safety (Lin et al., 2024) 2024 Survey of debiasing and hallucination-reduction techniques.
-LLM Safety (Wang et al., 2023c) 2023 ChatGPT applications, limitations, risks, and safeguards overview.